A Complete Bibliography of Publications in Advances in Computational Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 July 2024 Version 1.02

Title word cross-reference

(3, 1) [DOZ94]. (m, l, s) [BL99]. −1 [Cum96]. 1 [FF05, GT14, Han05, JRS15, NV10, NSW16, TZ07]. 1 < p < ∞ [LWY06]. 2 [Boc04, DL98, GM14, HKR04, Lee03, Möß10, Pot04]. 2.1933 [Pla00]. 3 [CK11, GL13, Hem95, Hem96, HS13, MU14, VW02, vdHS00]. 4 [Han98, Jia11]. 4n [SC08]. 6 [DLUS17]. [−1, 1] [CT05, CKK12]. 2 [WZJ18]. ^{3} [MRS22]. Ax = b [Cul96]. c [Rip99, Sch11b]. C^{0} [CFZ21, SS14, WY20]. C^{1} [KST21, KJ09, LL04, LPSSP00, NZ04, NSZ04, Sab04, SS04, DL98]. C^{2} [AF96, GSS22, Rem12]. C^{3} [CGM01, CM03]. C^{k} [MSS02, PR99]. C^{r} [LLS96, LL98]. C^{ρ} [LLS94]. C^{s} [KV21]. d [DQS20, HX02, Rob98]. E [MP96, Pré94]. ℓ^{1} [GN08, GS13, MZ13]. ℓ_{1} [SHTS14, ZYY16]. ℓ_{1}, 2 [Fli18]. F (z + 1) = b^{F (z)} [PC17]. F_{A} [BW15]. F_{D} [BW15, TZ05]. GC_{n} [HV19a, HV19b, BH17]. H [Dai16, MS00, Peñ11, GH06, Xu04a, ZQ13, ZG16]. H(div) [DMS22]. H^{−1} [She09]. H^{1} [HS20, ZJZ21]. H_{F}^{p} [HLT11]. H_{1} [MSS02]. H_{2}(\overline{Ω}, ∂Ω; Z) [RBGS18]. hp [MW01]. L^{2} [CLZ02, LL10, LL12b, MS16, WQP23, ZW23a]. L^{2}(R) [CHL17]. L^{2}(R^{d}) [CR08]. L^{∞} [Noa14]. L^{p} [War13, RS99, BR23]. L^{p}(R) [LWY06]. L^{p}(R^{n}) [PS23]. l^{q} [BR23]. L_{0} [TGA96]. L_{1} [LW04, DSS09]. L_{2} [KSUW23, DSS09]. L_{2}(R) [Atr12]. L_{p} [Jia95, LWZ19]. L_{p}(R^{d}) [DL04]. l_{q} [LWZ19]. L_{r} [CDP09]. M [CTZ02, Jia00, Tem98]. R^{3} [BST24]. R^{d} [War13]. R^{n} [HNN23, Sor18]. S^{3} [Grä12]. H [AFM23, FMP22]. H_{2} [BGH15, BCS19, BB24]. H [KM19]. N [HCF^{+}21, ZTHW22, BH17, DKLT93, SC08, UG19]. O(N) [GM14, HCF^{+}21, ZTHW22]. P [Dai16, Peñ11, Tyg23, CGG^{+}20, CMM21, CS03, Don19, GH06, Jan98, SL12]. P_{0} [ZT15]. P_{1} [ZT15]. π(x) [Kot08]. q [Rob98]. Q_{1} [HW18]. QR [CM99, LMV11]. R^{2} [KL00]. R^{3} [KL00]. s [Hes09, KL04, KL07]. S^{2} [HL08, Hes09]. σ [SX07]. SO(3) [Grä12]. t^{i}e^{λt} [MPR22]. τ [gTpM02]. θ [MP18]. V [ZXC12]. ε [GH16]. φ [BEG22, WZtX16]. P_{1} ⊕ RT_{0} [LLR24]. W [CGH^{+}96, IB17].

-adaptive [ZQ13, ZG16]. -algorithm [MP96, Pré94]. -analysis [ZYY16]. -approximation [KSUW23]. -bases [MS00]. -channel [Jia00]. -conforming [DMS22]. -continuous [WY20]. -convergence [CDP09]. -d [NV10, TZ07, DL98, FF05, VW02, JRS15]. -dimensional [DQS20, HCF^{+}21, ZTHW22]. -energy [Hes09]. -error [WZJ18, BGH15]. -estimate [CLZ02]. -finite [HW18, LLS96]. -fold [Jia11]. -frame [CGG^{+}20]. -Frames [CS03]. -function [BEG22, IB17]. -functions [WZtX16]. -fundamental [LL98]. -geometrically [ZXC12]. -harmonic [Xu04a]. -inverses [AFM23]. -isometry [SL12]. -matrices [Dai16, Peñ11]. -matrix [FMP22, KM19]. -method [MP18]. -methods [gTpM02]. -minimization [Fli18]. -monotone [KL04, KL07]. -orthogonal [SX07]. -penalty [GS13]. -piecewise [PR99]. -point [DLUS17, SC08]. -poised [BH17]. -refinable [CTZ02]. -regularization [MZ13, SHTS14]. -smooth [KV21]. -sphere [HKR04, Pot04]. -splines [BL99, MSS02, NZ04]. -stable [TGA96]. -subgradient [GH16]. -subspaces [HLT11]. -superconvergence [ZSB01a]. -synthesis [ZYY16]. -term [Tem98]. -values [Tyg23]. -version [Don19, Jan98]. -wave [HX02].

-weighted [MS16]. -widths

[DKLT93, UG19].

/fourth [FZLZ21]. /volume [LLLH23].

1 [BF94]. 19.07.1998 [Sil99]. 1D

[ATS19, BSK19, CHS21].

2 [DMBH24]. 2-chains [RBGS18]. 29

[Fus08a]. 2D

[LRY23, LGQ24, MH22, GQL24]. 2D/ [GQL24]. 2D/3D [LGQ24].

3D [ANGV19, BCC22, CGHK23, GQL24, GSS22, LGQ24, ZD18]. 3D-VAR [ANGV19].

4 [Hem96]. 409b [CN96].

96c [Hem96].

a-posteriori [MRH15, SWH20]. Abel [VNS18]. Abel-type [VNS18]. ability [’Sw19]. absence [GM14]. absorption [NKL^{+}19]. accelerate [KKL^{+}22]. Accelerated [HQHV20, GM22]. Acceleration [Pré94, XY15, CV23, DMMS11, LZY98, MP96, WMB13, ZCY20, ZTHW22]. accelerations [Noa14]. Accuracy [FF05, ZCY20]. Accurate [HX21, MPR22, WWX18, Del17, GM22, GJW20, HBMY14, HW19a, KCCV23, Qiu23, SJJY21]. acoustic [ARPR01, BBR02, GM22, HSSS18, HR02, LMRV23a, LMRV23b, ZP06, ZP24]. across [ZMQ23]. action [Las16]. actions [WW19]. active [KPY09, PDB24]. active-set [KPY09]. Adams [Wil98]. adapted [AL13, Che06, CZS16, GK19, GHLU19, LW23, Str13]. Adaptive [Ano01b, BDD06, BMP^{+}19, BDMR10, BLS02, CHP19, DFR07, EOR18, FYL08, GDMS21, HMX22, Hol01, Kac18, KKL^{+}22, KHM16, LC20, MSZ06, MJM23, PP97, QW11, RTV21, RWT13, WZM21, XZ03, ZW23b, AM24, ASU17, AU20, APV14, ACPV21, BKN11, BMK15, BV24, BCE^{+}09, BF94, CCE^{+}12, DMMS11, Dũn11, Dũn12, FaKT22, Gia20, HBH24, KCW17, KESR22, KC16, LC18, LZL23, LJL20, MSL21, RS19, RSZ11, SW96, SvVW22, Wan22, WBY24, WLPV15, XYZ21, ZL16, ZQ13, ZG16, ZW23c]. adaptively [GJS14]. adaptivity [AK01, CR23]. adding [RY13]. additional [CaL06]. additive [Bre95, PTC12]. ADI [FZLZ21, QX21]. Adini [MS19]. adjacent [Khl18]. Adjoint [LSLS14, BST24, FGO14, FMPS13, YZ24]. ADMM [BC19, TY18]. Adv. [Fus08a, Hem96]. Advances [Ano02d, BOP^{+}15, Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano16]. advection [CLZW13, CtTB21, LZCW21, Zen19]. aerial [ZB99]. aeroacoustic [CEL15]. Affine [CSS98, Jia03, CHSS03, Mai06, PW94]. affine-invariant [Mai06]. age [Kim06]. age-structured [Kim06]. agent [PSW24]. agent-controlled [PSW24]. agglomerated [CH22a]. aggregated [YB21]. aggregation [AL21]. Airy [dLL19]. algebra [BN19, BN21, BKN11, BP11, CLM02, FH05]. Algebraic [JF02, RSA14, RM19, CHS20, CGR13, CCLM21, DMMS11, FÁG19, Hau97, HF19, LMV11, Pan95, SXYY06]. algebras [Kni24]. algorithm [AKKN22, BMB14, BPK96, Bre19, CP99, Che02, CX04, zCpST09, CZS16, DM19b, DES21, DLK12, DGK15, DHWZ18, FKS05, Fou23, GL23, GM14, GR19, GIKV21, Han05, HHX18a, HHX18b, HCH20, HZHL12, HFH15, HCF^{+}21, JZ10, KBL22, KPY09, KC16, LY21, LD07, LWZ19, LWWZ22, Maz07, MU14, MP96, Mur08, Pré94, QW11, Rip99, Rob98, RHZ17, Sid20, Ven94, nCV13, WW12, WLW16, XWL13, Yao19, ZSHZ15, ZTHW22, ZS24, ZXC12, Zwi94, dH94, Yao19]. Algorithmic [Cai02, GH99]. Algorithms [Ano02d, FMR00, GN01, ACM^{+}22, BGM93, BEPS96, CHM18, CDPS22, CW19, DT96, DM23, DT12, FJ99, FRT09, Gon93, GN08, GH16, GS19b, LLS09, LT06, Li04, LSXZ15, LKT18, LT12, MSZ06, MSXZ13, Sab04, SMK22, Tem98, Tem00, Tem01, Tem02, VM17, VK96, WLPV15, XZ01, Yin07, ZQS22, ZS21]. all-electron [SL15]. Allen [RWT13, GWC24, HS20, JlLZ20, KL22, LZL23, QZX23]. almost [CJKK18, GP95, yGmW98]. along [FM18]. alternate [MP96]. Alternating [SYCJ18, TSV21, YF23, YKZ19]. Alternative [LPY10, Sch11b, SLT21]. Ambient [Mai20]. American [lZmCRX03]. Ampère [Awa15, CFZ21]. analog [LPY10]. analyses [KKL03, MNC16]. Analysis [ADL11, DSW20, GS18, GPS21, JS05, KS17, KvL95, KL16, Lee03, LLR24, LJL20, MO19, VW02, WZ19, WZ20a, dFGAJN18, AGN23, Ang00, AK15, BBB22, BSK19, BMSR^{+}16, BBdD21, BC19, BSSM22, BEPS96, CDP03, CCSS09, CR23, CCG10, Che06, CGH^{+}13, CS19, CV23, CGHK23, CDM^{+}24, CKW10, CsL18, DM22, DMBH24, DG13, DLZ16, Fin19, Flo96a, FM04, GAN20, Gat21, GRdSA19, GJMY19, GGRBRG21, yGW07, GL17, HMM21, HL20, JS21, JK08, KK05, KA97, Kum23, LP19, Lee16, LC21, LA22, Lev99, LKT18, LZ20b, LY21, LSY22, LMV11, LY08, Lut04, MH22, MS19, MWW13, Nie98, OS09, OBS15, Onc14, PM17, PLRGVR22, PS98, RMCCG^{+}19, RAB^{+}09, RHZ17, SXZ06, SLN14, TY18, TZ03, TS05, VW05, WYW11, WXWL19, WZ10, WOBL17, XB05, Xu14, YDF97, YL24a, Yao16, YZLL24, Yem22, YXM23, Yin07]. analysis [ZSB01b, ZGGW12, ZYY16, ZW23a, ZL20, ZL21, ZW20, Zwi94]. Analytic [CCK20, CSW96, MXO13, Kot08, Pet00, RAB^{+}09, Sto96]. analyze [FÁG19]. analyzing [KL15, LCC19]. anchoring [SJJY21]. Anderson [CV23]. angled [MB96]. angled-derivative [MB96]. angular [SF14]. Anisotropic [GKM^{+}20, BlH23, BSSM22, CtTB21, CGRS15, DKMT20, GO95, GQ22, HS03, HW18, Kun01, Pei05, WQP23, ZC14]. Announcement [Ano01c]. anomalous [ZD17]. ANOVA [PW24]. antenna [SFDE15]. antennas [MKS02, Ryn00]. Antisymmetric [Jia03]. any [KXZ04, Mon09]. Appell [BCZ05]. Appending [Kun95]. Application [FE15, LL12a, Nee96, lZmCRX03, AGN23, BEG17, BDD06, CLL08, CZS16, CGS93, FRT09, yGSlX03, KF19, Li10a, LZ21, LD07, MSW^{+}22, MU14, PZ16, Pré94, RF19, SWH20, SL12, Svá19, Wal95, Yao19, YZLL24, ZGV22, ZL16, BBF^{+}24]. applications [ABI15, ByLl01, CdGDN20, DST04, DNS21, DQS20, DS13, yGS08, HKR04, HAS05, HX17, IY22, JGW15, JW95, LP19, Low05, MRS18, RT14, RZ10, RSZ11, SW06, Sto96, TLG^{+}22, TX19, WZtX16, YTL^{+}20, vdHS00]. applied [AC05, BBB21, BBB22, BP11, CEL15, Smo07, ZB12]. Applying [MGSS22]. Approach [Zoz15b, AHK^{+}19, ANGV19, AR20, BBF^{+}24, BCS19, BPJ02, BKP20, CCSS09, CCNT21, CC24, CHRX06, CS18, DGMES13, DKK22, ES23, FÁG19, FBDH19, FY13, GHKL18, Grä12, HT12, HQR23, JKK^{+}12, KvD01, LSW17, Leo02, MMNM23, Pan95, PPH^{+}21, RLGGAV19, SBZR19, SLT21, Wal95, WM21, ZD14]. approaches [ZD17, Zoz15a]. approximability [FMP22, Fou23, KM19]. approximants [BCZ05, CDTV99, VK96]. Approximate [CLMR10, FPR^{+}12, FGO14, GN01, FOW14, IMS99, JLZ08, KK05, Kim06, NBL11, Pan95]. approximated [GLS15]. Approximately [CJKK18, DM15, GGRBRG21]. approximating [CGR11, DM22]. Approximation [AMM20, AB11, BDKY02, Bei23, BGN17, BR15, CP04, KP04, Le 05, Li09, Lut04, MZ22, Mha93, MNW99, NSW07, PR19, Pot04, SS16, SS08, SSK24, Xu04a, ZJ06, AS21, AK01, ASS10, Ara20, BX94, BN21, BC99, BMK15, BS05, BW20, BEG22, BNPP20, BB24, BA15, BFCIV22, BW15, CLLS06, Cas07, CL07, CLM96b, CH22b, CP07b, CG17, DPS93, Dar03, DY01, EGL13, FJ99, Flo96a, FHN94, FW15, GHdN12, GZW18, GLX23, Grä12, GS10, yGmW98, Han12, HMX22, HW00, IY22, JS21, KS20, KORS17, KL04, KL07, KSUW23, KvD01, KP18, LP21, LS98, LW04, LW14, LT06, Li94, Li03, Li10a, Li10b, LT98, LB93, LL07b, LP14, Mai20, MM00, Mai05, MN15, MRH15, MNW96, MWW13, MS04, MS16, NW04, NS04, Nee96, NG99, NZ04, ON18b, PWL20]. approximation [PW24, PS98, QZX23, SMK22, Spe17, TLG^{+}22, TX19, Tem98, TYY11, TDG16, WZ20a, War13, WH07, WZ20b, YZB20, YF23, YCQZ24, YZ08, Yoo01, ZP19, ZCL20, ZW23b, ZW23c]. approximations [ALRÁY22, ByLl01, CCG10, CHS21, CW19, Che06, CGH^{+}13, CJN18, DG13, DG16, ER21, FF05, GJU22, GHJZ23, yGSlX03, yGS08, HL05, HR01, HX02, IB17, IMS99, JGW15, Kot08, MRS03, Ryd19, Sid20, SU22, WXWL19, YLBL12, Ye00, ZZP23]. arbitrarily [Han10]. Arbitrary [ZJ23, CMP07a, CHR00, GH18, HW00, LW23, SL05, Ven94]. arc [Leo02]. architecture [Jan98]. Area [HR17, LLLH23]. area- [LLLH23]. area-/volume-preserving [LLLH23]. arguments [Bre19, MS22]. arising [AM14, BNP14, JM00, LN14, LL99, LSXZ15, MV01, TBD94]. Arnoldi [WZtX16]. array [AH96]. arrays [Cal05, SFDE15]. arriving [AGGAV^{+}19]. arrowhead [RDEG24]. art [FRT09]. article [BS23, Zoz15b]. Artificial [PRS24, An20b, Mha93]. aspects [BZA24a, BZA24b, Le 00, Sau95, Svá19, Yoo01, dB00]. Assessment [PJ20]. asset [PGB15]. assimilation [GAN20, KKP19]. Associated [GH03, ADL11, CW09, GRV22, HLT11, KA97, LLLH23, LPSSP00, NSW07, SM00]. association [New96]. astrodynamics [ABI15]. asymmetric [WSL^{+}19]. Asymptotic [AV15, CL07, CV23, FL00, HWZ13, JLZ08, RL07, DHO16, DGS18, LL07a, MWW13, WG19]. asymptotic/numerical [WG19]. Asymptotically [TX19, TDG16, Kac18]. Asymptotics [IS13, DS01]. asynchronous [Tal19]. Atmospheric [CLM96a]. Atomistic [APK18]. Atomistic-continuum [APK18]. atoms [FGO14, Füh16]. ATPH [BCHS22]. attractors [CTZ02]. augmentation [CWX06]. augmented [WLW16, XWL13]. Author [Ano00a, Ano01a, Ano02a, Ano02b, Ano03a, Ano03b, Ano04a, Ano04b]. auto [GLM23]. auto- [GLM23]. autocorrelation [BD10, Plo95]. Automated [Nes16]. Automatic [Hag98, GLM23]. automorphic [Kni24]. autonomous [MO16]. avalanche [IL16]. average [ALRÁY22, ABD08, Atr12, Li03, LD07, RJ00]. averaging [MSZ06]. AWGM [AU20]. axi [Nit22]. axi-symmetric [Nit22]. axial [Jia11]. axis [CCH^{+}08]. axisymmetric [Ara20, BBF^{+}24, BH02, FKS05, KNV01, LD22].

B [AMK19, CHRX06, CM12, FGS16, Flo94, JKLY13, KK99, LM06, MXZ08, MH17, NSS04, Plo95, Str95, Wal95]. B-spline [CHRX06, FGS16, Flo94, KK99, MXZ08, MH17, Wal95]. B-splines [CM12, JKLY13, LM06, NSS04, Plo95, Str95]. background [ACM^{+}22]. backpropagation [KKP24]. backward [AK15, BCZ23, HT11, WLW23, ZL20]. backward-Euler [AK15]. balance [AL21]. Balanced [BDYY20, GPAF18, Kür18, CT17b, PWL20, RDEG24, Tha19, TTC21]. balancing [GKS19]. ball [Li08b, Xu04b, DP06]. balls [DQS20, Noa96]. Banach [CS03, DST04, Tem01, Tem07]. band [MNW04]. band-dominant [MNW04]. banded [AH96]. bandlimited [FPR^{+}12, GP14, Lem09]. bank [CMX07]. banks [EGL13, Jia00, Jia09b]. Barycentric [War96, WSHD07, FHK06, GRB12, GGW24]. basal [IL16]. baseballs [ALDHHZ^{+}19]. based [AS02, AV15, AB97, An20b, ABD08, BN21, BCM99, Bei23, Ben97, BH19, BS97b, BS97a, BS00, CZS16, CFZ21, CJKK18, DL98, DS10, DGMES13, DAP13, DLK12, DHWZ18, DQS20, Dũn09, FÁG19, FM14, FKT20, GZW18, GGW24, GWC24, GS13, HL05, Han05, HH15, HBH24, ILHH22, KKL^{+}22, Khl18, KA97, LHH23, Mac94, MSZ06, MW01, MC10, Nai12, Nes16, NBL11, NKL^{+}19, NP24, PM19, PRS24, SMK22, SS16, SS15, TQ17, TSZ98, VK96, Wat96, XZ01, ZMQ23, Zha01, ZLPX23, ZS21, PP23]. Bases [LW03, LP04, Maz04, CMR22, DS00, DP06, GJS14, GLT93, HHS04, Hua06, JL06, Li09, LWY06, MPR22, MS00, Pau19, QWXZ10, Sab04, TSY10, TYY11]. basic [BZA24a, BZA24b]. Basis [HORU19, ASU17, AZW15, ANGV19, AM14, AS05, BCM99, BGH15, BV24, BFMNP24, BDM20, Cai02, CP93, CEL15, CL16, DSW05, DP13, Fas99a, Fas99b, FF05, FS98, FW15, HK21, Hub12, IUV17, Jüt98, Le 05, LST11, LW14, LM14, LLY06, LL99, LR07, Li05a, Li10a, Li10b, Lic22, LK05, Low05, Mai05, MRH15, MMNM23, PPH^{+}21, PDB24, RS99, Rip99, RBGS18, Sch95, SL15, SS08, SRFH12, WH07, WK20, Wie15, Yan19]. Bayesian [SZJ21]. BDF [LZ20b, LYF^{+}21, LHH23]. BDF2 [WMW21]. be [Kle21, WS01]. Beamlets [She09]. Bedrosian [TSY10]. behavior [LL07b, RL07]. behaviors [YX24b]. behaviour [Mur08, Smo07, WW19]. Bell [Pit16]. Bell-shaped [Pit16]. below [YLBL12]. Beltrami [KN14, KNQ16, O’N18a, PTC12]. BEM [ADG17, CP15, EEK21, HS03, MS98, NT23, Tau19]. Bending [Alb15]. Bernstein [Sab04, ACM20, CMP07a, CP93, CCLM21, Jüt98, LKW17, Sor18, WSM13, WZ20b]. Bernstein-type [Sab04]. Berzolari [CG15]. Bessel [Bre19, KXL24, LYY13, Lóp18, Mac94]. Best [BDKY02, Rau05, CC24, Li94, Tem98]. better [DM23]. between [ATS19, BWX97, BCE^{+}09, FK10, HR17, PZ16, SM00, She13]. Beyond [GN08, Buh06, Maz01]. Bézier [CC94, DP13, Flo94, LKW17, Sor18]. Bezout [GRV22]. Bezout-like [GRV22]. bi [JS99, LN19, TMH21, dCFCF20]. bi-degree [TMH21]. bi-Laplacian [LN19]. bi-objective [dCFCF20]. BiCG [Cul96]. bicubic [KP04]. bifurcation [PPH^{+}21]. bifurcations [CGS93, ER99]. biharmonic [CL07, CLL08, FKS05, HSY16, Jeo98, JM03, Li10a, WXWL19]. bilateral [CW09, ILHH22]. bilaterally [KPY09]. bilayers [Alb15]. bilevel [MHR21]. bilinear [BCS19, BGW21, BGW24, BA15, DHGA23, KV21, Rei93, RGB18]. binary [GZW18, Khl18, LLY06, Lev99, SJJY21]. binder [SEE96]. Binomial [BMSZ01, And20]. binomial-Kumaraswamy [And20]. binormal [KKV15]. biological [’Sw19]. Biorthogonal [Jia11, CHR00, CL02, Kei95, KKL03, KKLY10, PST08]. bit [GKKS23]. bits [MNW04]. Bivariate [MNW96, Rem12, She08, CG04, CGR13, DL98, DS00, DZ04, Flo96b, LS98, LKW17, MH17, NZ04]. Black [PGB15]. blended [LHH23]. blind [Fli18, HR13]. block [AB97, Bai99, sCLC13, CSWP99, HH16, HHX18a, HHX18b, MC23, NBL11, TY18, TZL16, YL24b, dCB07]. block-factorized [NBL11]. block-randomized [YL24b]. blossoming [Maz98]. Blossoms [Maz04, Maz99, Maz07]. body [GM14, HKRS14]. Boltzmann [HXC10]. Boor [JS95, Maz07]. Born [CdGDN20]. both [ZYY16, ZD17]. bottom [TTC21]. bound [Bar03, GWC24, RDEG24, ZW22]. boundaries [BK96, CCLM21, She13, Sou22, Tau19]. Boundary [dLL19, GJMY19, IMS99, MS99, ACK19a, ACK19b, ABI15, AR20, BBB22, BSK19, BBdD21, Boc04, BLV20, BR22, CDS02, CWLH95, CP99, CWX12, CT17a, CHS17, CG17, Del17, DKK22, DQS20, EX20, FLM96, FK98, FZ12, FM18, FY13, FH18, GMM09, GM14, Gil17, GRdSA19, yGmW98, GH06, GPG15, HH15, HJH12, HR02, HZHL12, JSSE97, Jeo98, JM03, JS05, Jia09a, KP95, Kno09, KS17, KP18, Kun95, LD22, Lee03, LR19, LC21, LLR24, LS19, LQW18, MH22, MR96, MB96, Ock96, QC12, RZ98, SH23, SW98, TZ07, TSZ98, TGA96, VC00, Wan22, WLB22, ZCL20, ZGV22, ZS24]. boundary-layer [ZCL20]. boundary-partition-based [DQS20]. bounded [AS21, BBR02, CL16, CH22b, GM17, GKM^{+}20, LWW21, PST95, San15, TDG16]. boundedly [CET21]. Bounds [AS08, Dam07, Dav23, LB93, Mai05, MRS22, NG99, NSW16, SWH20, TMH21, XHC15, ZXC12]. Boussinesq [Lee03]. box [BS05, CGR13, ZW19, Gér19]. box-spline [BS05]. Bracket [Chi03]. branch [IB17]. branched [KC16]. breaking [GH08]. bridging [ESTW19]. Brinkman [AMS22]. Brunel [New96]. bulk [EX20]. bulk-surface [EX20]. Burgers [KP18, WZzS21]. Burridge [XKY15]. Butzer [BS23]. BVMs [AB97]. BVODEs [Mui99]. bypass [HF19, KRP19].

cache [CM99]. cache- [CM99]. Cahn [LGC^{+}20, CZ20, GWC24, HS20, JlLZ20, KL22, LYF^{+}21, LZL23, QZX23, RWT13, SXWG22, WaZL24, ZY20, ZZL^{+}24, ZL21]. Calculation [VC00, WWX18]. calculations [CLR19, SL15, Tad19]. calculus [Lic22, TH24]. Calibration [Tyg23]. camera [FGS13]. can [Kle21, WS01]. canonical [AS24, CL22]. cap [HW12]. capabilities [CLLS06, CLM96b]. capacitance [TW98]. capillary [BSSM22]. Caputo [Gu20]. cardiac [GLS15]. Cardinal [HL16a, HL18, BS05, Buh06, Led15, RS99]. cardioid [ALDHHZ^{+}19]. cardioid-parametric [ALDHHZ^{+}19]. Carleman [CHS21]. Carlo [ESTW19, KR19, PM19, Zhe23]. Carlo-bridging [ESTW19]. carpet [LSY22]. Cartesian [CH22b]. Cas [GGAVGG^{+}19]. cascade [CX04, Li04]. case [AS24, CM03, Han05, KS20, RH12]. Cauchy [Bra01, BR07, CKP13, Cum96, KXL24, WZ10, ZW12]. Cauchy-type [Cum96]. Causality [GGL07]. cavities [QC12, VW02]. Cayley [DLP98, ZS21]. Cayley-transform-based [ZS21]. CDG [YZ24]. cell [ALRÁY22, ABD08, BEPS96, Dad17, JS05, KHM16, LR19, ZQ13, ZG16]. cell-average [ALRÁY22, ABD08]. cell-centered [KHM16]. center [EHV17]. centered [BEPS96, KHM16]. central [TQ17]. centroids [MFB^{+}11]. ceramics [SEE96]. certain [Sto96]. certified [AHK^{+}19, ANGV19, BV24, HK21, Pra23, SFDE15, ZMQ23]. CG [ZT14]. CG-like [ZT14]. chain [PM19, PR15]. chains [DMMS11, RBGS18, ZXC12]. challenges [Onc14]. channel [Jia00, YTL^{+}20]. characteristic [CCE^{+}12, HCH20, LZCW21]. Characterization [CHS20, LLS95, Li04, SZ09, FGG^{+}17, FH20, Fus08a, Fus08b, GP95, LWY06, Mon09]. Characterizations [LY07, Far10]. charged [LW23]. charged-particle [LW23]. Chebyshev [BR13, Cop98, Fou23, MV99, Maz99, Maz01, MC23, Nee96, SC08, TZL16, WK93, Zhu15]. chemorepulsion [GGRBRG21]. chemotaxis [CEHK18, PLRGVR22]. chiral [HPP^{+}19, ZGGW12]. choice [GPT17, JJLR14, Kno09, LLY12, RH12]. Christoffel [LP19]. Ciarlet [CL07]. circle [CJ07, GVSLN96, GKKS23]. circle-preserving [CJ07]. circuit [SS15]. circular [HO17, MKS02, PW08]. circulation [SI15]. clamped [LQW18, MM20]. class [ADL11, ABY13, Atr12, Bai99, BHT11, BDM20, CJKK18, DG13, GMS12, GKS21, Hak00, HR17, HCYY20, HX21, Hub12, IRT23, LLS94, LSXZ15, MSS02, MN00, ZW20]. classes [CKP13, Dai16, KL07, Pei05, Peñ11]. classical [Zoz15a]. Classification [CG04, XP10]. classifiers [ZJ06]. Clenshaw [XHC15]. Clifford [CLM02]. cloak [LSY22]. Closed [Hub12, Dav23, GQL24, LGQ24, Noa96]. closed-loop [GQL24, LGQ24]. closest [GS19a]. closures [SM18]. Clothoid [RW21]. Clough [LL98]. Clustering [CS18]. CMOS [Tal19]. CNNs [YZ23]. co [PR15]. coarse [CH22a, KS18, KXZ04]. coarse-grid [KS18]. coated [CH15]. cocoercive [nCV13]. codes [CHMR97, Hal95, Kir98, NKL^{+}19]. coefficient [FL23, GS13, QZX23, SZJ21, Spe23, SJJL15]. coefficient-based [GS13]. coefficients [AS21, BKP20, CJN18, ÇY23, Ehr00, EGL13, Gia20, Kum23, LS18, Pet02, Wan23a, Wan23b, ZB12, dB00]. Coercivity [HW18]. cohomology [PR15]. Cokriging [BZ18]. Cole [HW19a]. collections [SB22]. Collocation [AHHR16, LS19, AB23, BJX09, CCK20, CCNT21, CDM^{+}24, Cum96, FS98, GK19, Gu17, Gu19, Gu20, yGqW09, GW14, GW21, Han05, Hau97, HS13, HR02, HS16, LLS09, LK05, Liu99, LBCZ18, MWZZ23, MH22, MR96, Nai12, QX21, Ram95, RWT13, Sch09, XH13, YZLL24, ZP24]. colonoscopy [MU14]. colorization [LN19, YKZ19]. combination [FPT06, PW24, TSV21]. combinations [DGMES13, MPS96, MPS97]. combinatorial [TMH21]. Combined [DSS09, KK20, BBB^{+}16, CLCH19, HO15, HCYY20, HY18]. coming [BDM20]. commensurate [Ryd19]. common [Len94]. Commutation [KMMI19]. commuting [GLT93]. compact [Gna07, HWZ22, PS23, Str13, WZzS21, XY20, ZW22]. Compactly [CHSS03, JJK13, JS99, Wu95, CL16, Hub12, Lai06, LLS95, Len96, Wen95]. comparative [BSK19, DM23]. comparing [Sch15]. Comparison [HKR21, BBB^{+}16, Zoz15a]. compass [GMB97]. compatible [TDG16]. compensating [AS24]. competition [Ble07, DLK12]. complements [Dai16, Peñ11]. complete [CtTB21]. completely [Dou94, Xu14]. completion [CMX07, DDLL23, GS19a, GS20, MHR21, WLW16]. completions [MRS14, MRS18]. Complex [NW22, BCC22, BFH^{+}18, DOZ94, FMR00, FH05, FKT20, Han10, IL16, MKS02, PC17, Pau19, SM99, ZHzSR21]. Complex-scaled [NW22]. Complexity [NUWZ17, Tom96, Cho95, GKS19, Pei05]. complicated [Gia20]. component [Gat21, JLW20, LKT18]. components [RST^{+}14]. Composite [yGjW10, CV23, Gia20, HXC10, Pla12]. comprehensive [FGG^{+}17]. compressible [An20b, BBB^{+}16, HCH20, ZCL20]. compressing [MST05]. compression [DPS93, DKLT93, DES21, LW14]. compressive [FBDH19, SL12]. compromise [BCE^{+}09]. Comput [Fus08a, Hem96]. computable [CZ19]. Computation [Ano02c, BJMR24, PR15, ZA10, And20, ANSZ17, BPJ02, BH02, BK24, BPK96, CX04, ER99, GH99, KNV23, LKT18, LD07, MFB^{+}11, Miy19, NP18, NT23, PPH^{+}21, dCFCF20]. Computational [Ano01b, Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano16, Pei05, Sau95, Yoo01, dB00, BZA24a, BZA24b, BOP^{+}15, CGDHRÁ^{+}19, CC24, FÁG19, GRdSA19, Lee03, Sch15, NW04]. Computationally [TL18]. computations [ASU17, GGAVGG^{+}19, GDMS21, KP98, MPR22]. computers [EJ97, vdH93]. Computing [BEG22, CKP13, GH08, GS19a, GKS21, HMS19, HBM03, Jon17, Kni24, KL15, AB23, AR20, BI11, HCF^{+}21, Kir98, LMV11, MXY13, SC08, ZTHW22]. concentrated [GG06, PS23, RST^{+}14]. Concentration [GZ13]. concentrator [HL20]. concepts [ESTW19, WLPV15]. condensation [HH19a]. condensed [ASS10]. condition [Awa13, CZ19, DKK22, Flo94, QC12, Sch95, SS96, Wan22, WL05, ZYY16]. Conditional [Xia13]. conditionally [FM21]. conditioned [ČF11, LWWZ22]. Conditioning [ZP24, Ant18, LG13]. conditions [ACK19a, ACK19b, BBB22, BR22, CDP09, Del17, EX20, EHV17, FLM96, FM18, FS07, Füh16, GW17, Gil17, HXC10, JM03, Jia09a, KK99, Kun95, LLR24, MB96, SLT21, TGA96, ZS24, ZhXpZ16]. conducting [ATS19]. conduction [BDMR10, CN96, HT11]. conduit [HYHH21]. Cone [Leo02]. confidence [KR19]. configurations [BEG17, CG04, CGG^{+}20]. confinement [LQW18]. confluent [DGS18, NP18]. conformal [CWLH95, CL18, HL16b, LLLH23, MKS02, NRR^{+}22, TB16]. Conforming [MN16, DMS22, YZB20, ZMQ23, ZC14]. conic [Flo96a]. conical [EHS02, PW08]. conjecture [Kle21]. conjugate [DDLL23, ZS21]. conjunction [DSS09]. connected [CWLH95, CL18, HL16b, RCJ^{+}23, San15, TB16]. connection [BGM93, NSW07, SM00]. connectivity [NS21, PŠWX03]. conservation [CFC23, KK05, MV01, Yan19, ZQ13]. conservative [HWZ22, LR19, LZ20a, LZ22, WZJ18, XY20, ZW22]. conserving [ABI15, MWZZ23]. consistency [CX06, HFH^{+}07, MNPR06]. consistent [KHM16, MSW^{+}22, Sou22]. constant [KPY09, LY08, Spe23, ZB12]. constant-coefficient [Spe23]. constitutive [JS21]. constrained [BHT11, IUV17, KPY09, SW06, Yao19, KKL^{+}22]. constraint [KXZ04, LC18, MHR21]. constraints [AHK^{+}19, GHKL18, WOBL17, ZW19, ZLPX23]. Constructing [CGM01, FMPS13, GW17, GU02, Lem09, GM14, GIKV21, WM21]. Construction [AF96, BD10, ČF11, CHR00, CaL06, DHO16, FOW14, GS10, JMO00, Lai06, OBS15, Rei93, SX07, WOBL17, ABC14, BLP19, BCZ05, CCH^{+}11, CP14, CMX07, CL02, FHK06, Hua06, KKLY10, LH13, MSS02, PST95, RBGS18, SW98, Str13, ZW23a]. constructions [CHL17, Lic22]. constructive [CGR13, DG16]. contact [HKRS14]. context [HSV09, Wie15]. continuation [Awa15, BSK19, Mur08, She13, Tim21]. continuity [LM04]. Continuous [CS94, DM93, LW23, LWX19, PT21, ALADH^{+}19, ALZ02, BP93, DY10, DAP13, GGRBRG21, JL97, Len96, LA22, Mui99, WY17, WY20, ZT14]. Continuous-stage [LW23]. continuum [APK18]. contraction [PR15]. Control [GS22, KLNS95, AM24, AV15, AHS20, AMPT22, BGW21, CBK01, Cas07, ÇY23, CCE^{+}12, EH24, GN13, HR01, KP18, LSY24, Lee03, LC18, LC21, LZ21, LY01, MRS03, Per22, Wil98, Yan03, ZW23b, ZW23c, dCFCF20, dLdDSM07]. controllability [CHS21]. controlled [PSW24]. controllers [CS18]. controls [MS16]. convection [ANN22, ÇY23, CCE^{+}12, DFOS23, HO17, KY01, KXZ04, LC20, LRS12, Mat09, MG19, OS09, PP97]. convection-dominated [CCE^{+}12, KY01]. converge [CSW14, Gon93]. Convergence [AK15, BQRB13, BBdD21, Bru23, CCSS09, CET21, FS98, Gom95, HX02, HW19b, JL19, KB15, LC18, Li08a, LG13, MH22, MS19, MNC16, Noe95, PK94, Ryn00, Sou22, TY18, WYW11, WZ10, Yao16, Yin07, ACM20, AAFV19, Ara20, Bad19, BC19, CLC16, CDP09, Che06, CS19, CV23, CZ19, CKW10, Don19, FJ99, FP99, FL00, GH06, GS19b, HL18, HS20, KKP19, Led15, LR07, Li04, Li05b, Li08b, LZ20b, LY21, PS98, Tem02, Wan19, Wan23a, Wan23b, WZ20b, Yao19, YXM23, ZJZ21, ZQS22]. Convergent [HPP^{+}19, Lóp18, GQL24, KXZ04, LLS09, LWZ19, WMB13, ZYB18]. conversion [LPY10]. Convex [FPT06, TSV21, sCLC13, CZS16, CCH^{+}15, FBDH19, FHK06, FK10, HH16, HHX18a, HHX18b, Kva14, LD07, LY01, MR10, MRS03, Par22, Pet96, RGB13, SYCJ18, TY18, War96, WSHD07, ZLPX23]. Convexity [CC94, KK99, SS14, Car95, Flo94, LPSSP00]. convolution [Cop98, HJJV97, Li03, YZ13]. convolution-type [YZ13]. convolutional [MZ22]. Convolutions [CK03b]. Conway [AVGGEG^{+}19]. cooperativity [CGDHRÁ^{+}19]. Coorbit [DST04]. coordinate [YWC12, Zwi94]. coordinates [FHK06, KB15, RGB13, War96, WSHD07]. corner [Noa98, Sab04]. corner-cutting [Noa98]. Corrected [Nit22, IRT23, WM21]. Correction [ACK19a, BZA24a, BD21a, HHX18a, LMRV23a, Wan23a, ZW23c, CS19, HV19a, KCW17, LGQ24, WaZL24, WM21, XH13, XYZ21, ZS24]. corrections [FXZ96]. corrector [CSWP99, ZD14]. correlation [GLM23]. Corrigendum [Hem96]. corrupted [XWL13]. Coulomb [HL08]. Counterexamples [LL07a]. Counting [TD21, Kot08]. CountSketch [MB20]. coupled [AMPT22, DFOS23, GLO10, HFH^{+}07, JRS15, LZ20a, LGC^{+}20, MRH15, MS98, OS09, XY20, ZMQ23]. Coupling [HF19, SH23, ADG17, DGMM16, GMM09]. couplings [EEK21]. Courant [CBK01]. CP [YL24b]. Crank [CLCH19, HCYY20, How96, New96, Ock96, SXWG22, aZWL23]. creates [Ito96]. criss [SXYY06]. criss-cross [SXYY06]. criteria [ZV08]. criterion [FBCR13, Tem02]. Critical [Sto96]. Cross [BH19, GLM23, HL96, Ryd19, SXYY06]. cross-correlation [GLM23]. Cross-Gramian-based [BH19]. cross-validation [HL96]. Crouch [JMO00]. Crouzeix [HMM21, YZB20, ZHY19]. cryptographic [Ren09]. crystal [HY14, LMYL19, LS20, RH15, RMCCG^{+}19, Zhu15, aZWL23]. crystals [KK02, SJJY21]. CSOR [HJJV97]. cubature [DG13, GS10, IMS99, Pet00, Xu98, Xu00]. cubic [ANSZ17, BCHS22, CG06, CHS20, ČF11, Flo96a, JL06, NSZ04, PWL20, Pet96, Rem12]. cubic-like [BCHS22]. cubical [Loh23, PR15]. cubics [Noa06, PFMS05]. cuboid [HSY16]. CUR [DM23, VM17]. curl [BKK17]. Current [MM99, AGN23, AMPT22, BMSR^{+}16, Cai02]. Curtis [XHC15]. Curvature [RS01, CC94, HW19b]. Curve [Pot95, Bei23, CM03, Khl18, MU14, Pet96]. curve-based [Bei23]. curved [KB15, SH23]. Curves [CL22, MP04, ABMM19, BX94, BWX97, BGN17, BC94, CHS20, CLM02, CGM01, Der04, FS94, FaKS02, Far10, FGS16, FGG^{+}17, FM18, Flo96b, HBMY14, JKK^{+}12, KLNS95, KKV15, Mon09, Pot95, PS98, RSA14, RM19, Sab04, ZN18]. curvilinear [LQ14]. cut [SLK20]. cuts [DT12]. cutting [Noa98, Sab04]. Cycloidal [MP04]. cylinder [MKS02]. cylindrical [ALADH^{+}19, BGMS07, YDF97].

D [Hem96, HS13, Lee03, CK11, GT14, Han05, NV10, TZ07, Boc04, DL98, FF05, GM14, GL13, Hem95, MU14, VW02, vdHS00, JRS15]. D-problems [Hem96, Hem95]. DAE [GZA20]. damped [ADG17, JSW20]. Damping [TBG18, PDB24]. Darcy [WaZL24, CHM18, MRH15]. Data [CP04, HO15, LW14, WBGG22, ALRÁY22, BCV23, BSK19, Bes03, BE00, CCSS09, CP15, CGW13, CMSS23, DZ04, DSW05, DSS09, ECS18, FHN94, FM04, GAN20, GSS22, GGW24, Grä12, GL13, HQR23, KKP19, KMO^{+}14, KESR22, Kun09, LW04, LP19, LNW02, LP14, MS22, MST05, MJM23, MNW99, MHR21, NSW98, NG99, Pot04, PJ20, RF19, RS99, WL05, WSM13, Yoo01, ZMQ23]. Data-driven [HO15, WBGG22, GGW24, HQR23, PJ20]. data-independent [DSW05]. datasets [CC24]. date [ALADH^{+}19]. Daubechies [DKB99, EGL13]. Davidson [MC23, TZL16]. DCA [HWZ23]. DDFV [DSW20]. deblurring [CZS16, KPY09]. decay [NV10, RCJ^{+}23]. decaying [LP10]. December [Ano16]. decision [AGGAV^{+}19]. decoder [HD20]. Decomposition [BRR18, BHM18, BFFM24, CM99, DT12, ES23, FKS05, GLM23, GL13, Hag98, HKRS14, HFH^{+}07, HQR23, HS03, HYHH21, Hu07, KESR22, Las16, LH13, LKT18, LHY08, MB20, MMNM23, Par20, PS05, PSNP11, SI15, Sto17, YL24b]. Decompositions [DS13, AL13, CCN^{+}16, CW19, CHRX06, DM23, VM17]. deconvolution [Fli18, HR13, WZM21]. decoupled [CHM18, DSW20]. Deep [YZ23, BFFM24, GHJZ23, Kni24, MZ22]. defect [ZS24]. defect-correction [ZS24]. Deficits [BCHL03]. defined [FMR00]. definite [Bai99, FM21, GS10, LS05, Pin04, SM99, WMY13, WMB13, Wen95, Wu95]. definiteness [Luo99]. definition [DQS20]. deflated [PPH^{+}21]. deformation [HSSS18]. degenerate [AGN23, Gna07, LB93]. degeneration [Bes03]. Degree [PR99, ABC14, BR13, DOZ94, FL00, Mar95, Mon09, PK94, SM00, TMH21, Wen95]. degree-raising [FL00]. degrees [Loh23, ZRR23]. DEIM [ZMQ23]. DEIM-based [ZMQ23]. delay [BP93, BPW95, ER99, GH08, WY17, Wan19, ZZL^{+}24]. delay-free [ZZL^{+}24]. delay-integro-differential [WY17]. delays [Dad17, XH13]. delta [CDS02, DRS08, LPY10, Wan08, GKKS23]. demand [GS22]. demixing [Fli18]. Democracy [GHdN12]. denoising [BCE^{+}09, JZ10, MSXZ13, Zhu19]. densely [She09]. densities [DMT03, WBY24]. density [An20b, CJN18, JJK13, WZM21]. depend [MGSS22]. dependence [KL22]. dependent [AKS23, BFMNP24, BKM^{+}21, DFOS23, FaKT22, FL23, GS18, GQL24, HQR23, LZZ09, SW10, TBG18, TB16, TGA96, dFGAJN18]. derangements [IS13]. Derivative [FXZ96, FF05, FW15, KS17, MB96]. derivatives [BRSV15, BS23, LWW21, XL10]. derive [ZRR23]. described [FKT^{+}19]. describing [Ara20]. descriptor [BW20]. Design [Yem22, Pot95]. designing [FGS13]. designs [BMRS20, HL08, Hes09]. desired [Lem09]. Detecting [BLP02, DGMES13]. detection [CGS93, CGW13, ER99, Lee21, MP00, SPS21]. determinants [Maz98]. Determination [LK20, FL23]. determined [BGM93]. Determining [SF14, LAG^{+}24]. deterministic [BZI19, ÇY23, GIKV21, WSM13]. detonation [ZG16]. Developing [LCC19]. Development [CDP03, LSY22]. deviation [Tyg23]. devices [BMRS20]. DG [GM24]. Diagonal [Peñ11, Dai16, FMPS13, Kos15]. diagonalized [LWL17]. Diagonally [vdHS00, EJ97]. diagram [DQS20]. diagrams [CL22, PPH^{+}21]. Dictionary [HBH24, NP24, BN21, Mai06]. Dictionary-based [HBH24, NP24, BN21]. Difference [EX20, BEPS96, Che06, CLL08, CEHK18, Dav23, ES23, FZLZ21, GK19, HWZ22, JRS15, KY01, LCC19, LZ20a, MB96, NV10, PL05, SYCJ18, TZ07, WZJ18, XY20, Xu14, YXM23, ZHzSR21, ZW22, ZLPX23, ZD16]. difference-of-convex [ZLPX23]. different [FKT^{+}19, LQ14, WBY24]. Differentiability [Sau06]. Differential [KK02, ANGV19, BP93, BPW95, BCZ23, ByLl01, BV24, BFCIV22, CJX17, CWX06, CN96, DLP98, EOR18, ESTW19, ER99, FZ12, Fas99b, FY13, GHJZ23, Gu20, GH08, yGS08, yGqW09, GW21, Gut21, Hau97, HK21, HBM03, HS08, JS21, JLZ08, KA97, LMV11, Liu99, Loh23, Mai20, MS18, MMNM23, PM17, QX21, SW96, She13, lWyG06, WY17, WZ19, Wan19, Wat96, WK20, ZRR23, ZL24, ZV08, ZD14, vdHv97]. differential-algebraic [LMV11]. differentiation [KF19, WH07, XL10]. Diffraction [BBR02, EHS02, Wan22]. diffusion [ANN22, BlH23, BCE^{+}09, CC22, CK11, CLM96a, CLZW13, CtTB21, ÇY23, CHS17, CCE^{+}12, DFOS23, DKMT20, DSW20, FZLZ21, GLO10, HO17, HW18, JLW20, KY01, KXZ04, Kim06, KHM16, KC16, Kun01, LC20, LD16, LRS12, LZCW21, MR15, Mat09, MXO13, MG19, MNC16, OS09, PP97, QZX23, SZJ21, Sou22, TDG16, YZLL24, YXM23, ZD17, ZJZ21, ZC14, ZCY20, ZW12]. diffusive [GQ22]. digital [LPY10]. digital-to-analog [LPY10]. dihedral [BJMR24]. dilation [Gro13, Han98, Han06]. dilations [CHSS03]. Dimension [CL12, TQ17, KORS17, Mai06, PŠWX03, TD21, TMH21, ZD18]. Dimension-by-dimension [TQ17]. dimensional [AU20, BNR00, BJMR24, BR15, CMP07b, CCK20, CC24, CGH^{+}13, CCH^{+}08, CLM18b, CLM18a, DLS14, DQS20, EH24, ESTW19, FZLZ21, GDMS21, GJW20, HSY16, HCF^{+}21, JN18, KKP19, KF03, LG13, LZ20a, LYF^{+}21, LZZ09, LL07b, LK20, MR15, Nov00, Onc14, PW24, QX21, SM00, TH19, WHS18, XKY15, ZHzSR21, ZTHW22, ZQ13, ZG16]. Dimensions [ZWDD16, ACK19a, ACK19b, AS08, Bei23, Che02, FP99, Füh16, GL23, HW00, HH19b, JSW20, MS19, O’N18a, WZJ18, WG19, YZLL24]. diminishing [CGP95]. DIMSEMs [EJ97].

DIMSEMs-diagonally [EJ97]. Dirac [CDS02, DRS08, yGSlX03]. Direct [AM14, LL99, Che02, CGHK23, DZ04, GM22, LWWZ22, MS16, SH23, TY18, YX24a, ZGV22]. direction [BS05, LLS96, LL98, SYCJ18, YF23]. Directional [ANZ23, CCH^{+}15, CGRS15, GL17, MSS02, NSS04, RF19]. directly [TQ17]. Dirichlet [AK01, ACH10, BBB22, GLX23, JM03, Lee03, Li05b, LLR24, WL12, ZS24]. Dirichlet-to-Neumann [WL12]. discharge [FKT^{+}19]. discontinuities [AL13, BR15]. discontinuity [ZW19]. Discontinuous [LGC^{+}20, Yan19, Ye04, AK15, BQRB13, BDFS19, BDM20, ÇY23, CHP19, CH22a, Don20, Gia20, KF03, LC20, Lep23, LZL23, Pet02, SXZ06, SH23, WXWL19, WaZL24, ZZP23]. Discrepancy [Leo13, Pla00, RJ00]. Discrete [CMX02, JSSE97, dLdDSM07, AAPCC24, AGGAV^{+}19, BF11, Ble07, Bru23, CJX17, CHS21, CKM99, Dad17, GAN20, HS20, Kir98, LC13, LL13, MZ13, Mui99, Nee96, Rei93, Ryd19, TZ03, VNS18, ZV08]. discrete-time [AGGAV^{+}19, BF11, Ryd19]. discretisation [DNS21]. discretisations [Mat09, WW19]. discretization [AS05, Bra01, CR23, FW96, HBMY14, HL06, HH15, JRS15, MV01, NV10, Ram18, RSZ11, TZ07, ZST19]. discretizations [Ang00, GH18, HL21, KM19, PGB15, XY15, XZ01, ZGV22, ZC14]. Discretized [HT11, CGS93, SXWG22]. discriminating [DGMES13]. disk [CWLH95, CL18]. dispersion [CLMR15]. dispersive [dLL19, HW19a, PM17]. displacement [HCH20]. dissipation [CLMR15]. dissipative [AMM20, BD18]. distance [PZ16]. Distances [BWX97]. distancing [DHSW19]. distinct [VK96]. distributed [BGG24, CS18, FYL08, HH16, LSY24, LKT18, LY01, ZBT^{+}22]. distributed-memory [ZBT^{+}22]. distribution [AR23, HCF^{+}21, LK20, YZ23, ZD17, ZTHW22]. distributions [CDS02, CdGDN20, IY22, Li03]. div [BKK17, JJLR14, dFGAJN18]. divergence [Sor18, Urb95, WWH21]. divergence-free [Urb95, WWH21]. Divergent [Zoz15b, Zoz15a]. Dixon [Chi03]. DL [BFFM24]. DL-ROMs [BFFM24]. DNNs [LP21]. does [MGSS22, Spe23]. Domain [DT12, AS05, BMSR^{+}16, CWLH95, DAP13, EHV17, GL23, GM17, Hag98, HKRS14, HFH^{+}07, HO17, HS03, HYHH21, Hu07, KNV23, KS23, LSY22, LHY08, MMNM23, Par20, PS05, PSNP11, Rem12, Sto17, SL05, VW02, ZhXpZ16]. domains [BB24, BK96, CCNT21, CLL08, CL16, CH22b, DM19a, FKT20, Gia20, GLX23, GKM^{+}20, GH06, yGS08, HL16b, IMS99, KB15, LG13, LHY08, NRR^{+}22, QWS24, Str13, TDG16, YX24a]. dominance [Dai16, Kos15, Peñ11]. Dominant [GDB24, BH19, MNW04]. dominated [CCE^{+}12, GQ22, KY01]. dot [LS05]. double [BST24]. doubly [BCZ23, DX20, Tor16, TB16]. doubly-connected [TB16]. DOUG [Hag98]. Douglas [HH16, TSV21]. Dr [Sil99]. Drift [CCDL20]. Drift-preserving [CCDL20]. driven [DX20, GGW24, HQR23, HO15, PJ20, WSL^{+}19, WBGG22]. driver [ALADH^{+}19]. DtN [Wan22]. Dual [AS24, BBB21, Boc04, CP03, zCpST09, CKK12, CKK16, CJKK18, Chr15, DM15, FGO14, FOW14, GRV22, HYHH21, JJK13, Jüt98, KPY09, LPY10, Lem09, LKW17, LYY13, LY21, Par20, Tad19, nCV13]. dual-porosity-conduit [HYHH21]. Duality [Noa06]. duals [CG10, CSS98]. due [KKP24]. Dyadic [DM99, CS94]. Dynamic [dCB07, AHS20, BHM18, EX20, HQR23, JS21]. dynamical [BH23, BFH^{+}18, DHGA23, GDB24, RGB18, Tan17, UG19, WOBL17]. Dynamics [Dad17, APK18, HPP^{+}19, Kim06, LW23, PSW24, ZST19]. Dyson [KS23].

EARSM [HF19]. earthquakes [XKY15]. ecology [MP18]. eddy [AGN23, BMSR^{+}16]. Edge [CGW13, PGB10, SPS21, GL17, YKZ19]. edge-weighted [YKZ19]. edges [Var96]. Editor [Ano95, MM99]. Editorial [GS12, MM93]. editors [Ano05a, Ano05b, Ano03c, Ano04d, Ano04e, Ano04f]. effect [ALDHHZ^{+}19, Spe23]. effective [HKR04, SWH20]. efficiency [DL04]. Efficient [CH22b, FZLZ21, JGW15, LP21, LMYL19, MV01, PS07, PPH^{+}21, RBGS18, SMK22, VM17, YKZ19, ZMQ23, AZW15, CM99, CCE^{+}12, DGMES13, GS20, HFH15, IP15, LGQ24, LD07, NT23, QZX23, SW98, TL18]. efficiently [LBCZ18]. eigen [Yao19]. eigenfunctions [SS16]. eigenpairs [GKS21]. eigenproblem [CR99]. eigenproblems [JM00]. Eigenvalue [YLBL12, Ara20, BPK96, CL07, EJ97, Gna07, HNT22, HR01, HL21, Kos15, LLZ10, fLLfDX21, MSZ06, Mar94, MC23, MTH21, RSZ11, TZL16, TBD94, WXWL19, WBY24, XY15, XYZ21, YXX22]. Eigenvalues [CGS93, AR20, BJMR24, Jon17, YZB20]. eigenvector [BPK96]. eigenvector/eigenvalue [BPK96]. eighth [TBD94]. eighth- [TBD94]. EIM [RH15]. Elastic [BC94, Ara20, Ble07, BGN17, CGM18, FE15, LD22, SLK20, WBY24]. elasticity [AAQ15, Awa13, BCV23, BQRB13, BBB22, DL23, HZZZ24, LLS03, Lee16, LLR24, MC08, NT23, PGM09]. elastodynamic [BCC22]. elastomers [Zhu15]. elastostatic [GG16]. electric [DD94]. electrical [HLN12]. electrically [ATS19, SFDE15]. Electromagnetic [CH15, BBF^{+}24, BH02, Cai02, DMBH24, HL20, VW05]. Electromagnetics [Ano02c, CCP22]. electron [SL15]. electrophysiology [GLS15]. electrostatic [Tor16]. element [AN20a, AK01, ARPR01, AB11, APV14, AL97, ANN22, BNP14, BBdD21, BDFS19, BK96, BKM^{+}21, CLLS06, CCG10, CZZ17, CBK01, CDS02, CLZ02, CLZW13, CLCH19, CFZ21, CT17a, CKW10, CLR19, CH22a, CG17, DHWZ18, DY01, DLZ16, DL23, EHS02, FF95, FYL08, FH18, GQL24, GAN20, GMM09, GLX23, GH18, GH06, HMX22, HZZ20, HR01, HW18, HCYY20, HL21, HCH20, HX02, HL20, HS20, HZZZ24, JS21, Jan98, Jeo98, JM03, JS05, JLZ08, JK08, KS18, KORS17, KL00, Kum23, KP18, LL04, LSY24, LLS03, LC18, LMRV23a, LMRV23b, LGC^{+}20, LZ20b, LSY22, LFL22, LZL23, LLR24, LZ21, Lic22, LB93, LLC16, LC19, LZCW21, LGN22, LWX19, LL10, LL12b, MR15, MH22, MN16, MSZ06, MC08, MS18, MS99, MS19, MM20, MY17, NBL11, PPH^{+}21, PP97, SW96, SXYY06, SH23, SBMS22, SW98, SvVW22, Svá19, TH24, VW05]. element [WY16, WWX18, WHTZ18, WXWL19, WWH21, WLB22, WLPV15, Xie08, XY15, XZ01, XYZ21, YZB20, YS23, YX24a, YL24a, Ye00, YZ21, ZRR23, ZP06, ZZP23, ZSB01a, Zha14, ZPY15, ZT15, ZHY19, ZZC20, ZW23a, ZSHZ15, ZW20, ZW23b, ZW23c, aZWL23]. element-in-space [SvVW22]. Elementary [Gér19, FH05, Lóp18]. elements [AS02, ASS10, AAQ15, Awa13, BKK17, BK24, Bra01, CMM21, FF95, HN18, HS03, HSY16, HY18, HMM21, KST21, LLS94, LLS96, fLL05, LL07a, LPSSP00, MHR21, MTH21, NW22, SL15, SPS18, SH23, SLK20, TH19, YDF97, YLBL12, Ye04, YZ24, dFGAJN18]. elevation [PK94]. elimination [LQW18, WW12]. Ellipsoidal [LLLH23]. elliptic [AU20, AC05, ACH10, AHC11, CDP03, Cas07, CWX12, CP15, CH22a, DFR07, DNS21, FK98, FZ12, FYL08, Gia20, GM14, GLX23, GH06, HR01, Hol01, HH19b, Jan98, Khl18, Li05a, LMT01, MTH21, OP11, Per22, Pet02, PSNP11, SXZ06, WHTZ18, WHS18, WQP23, XZ03, Yan01, YZ24, ZZP23, ZB12, Zha14, ZYC19, ZHY19, ZCB21, ZL16]. elliptic-type [HH19b]. embedded [BR15, Mai20, She09]. embeddings [GHdN12]. empirical [AAPCC24, CDP09, CHRX06, LP19, MNPR06, Tad19, Yan19]. end [PJ20]. end-to-end [PJ20]. Energetic [ADG17]. Energy [ABI15, GWC24, KU18, Ble07, CZ20, CFC23, EOR18, GZW18, GGRBRG21, HL08, Hes09, KL22, KF19, LSY24, Leo13, LMYL19, LZ22, MWZZ23, MS22, SJJY21, WW19, WZzS21, ZYB18, ZY20]. Energy-conserving [ABI15, MWZZ23]. energy-minimizing [KF19]. energy-preserving [WW19, WZzS21]. energy-stable [GGRBRG21]. Enforcing [HSV09]. engineering [SEE96]. enhanced [BFFM24, KKL^{+}22]. Enhancing [GGAVGG^{+}19, GJU22]. enriched [HT12]. ensemble [JY22]. Entropy [CDP09, BD18, Mai06]. environment [ZGGW12]. epidemic [Kim06]. Equal [CK03a, Möß10]. Equal-Norm [CK03a]. equality [Yao19]. equation [ADG17, AKS23, ANN22, AR20, AC00, BSK19, BBdD21, Bru23, CLC16, CC22, CBK01, Che02, CFZ21, CJW22, CGHK23, CHS17, CGS93, DFOS23, DD94, DLS14, FZ12, FYL08, dLL19, FKT20, FL23, GT14, Gil17, GM24, GM22, GWC24, yGSlX03, GW14, Han05, HS13, HR02, HY18, HS16, HS20, Jeo98, JM03, JlLZ20, JSW20, JLW20, KU18, KS23, KP95, KL22, KA97, KN14, KNQ16, KP18, Kun01, LRY23, LM14, Li05b, Li08a, Li09, Li10a, LMYL19, LS20, LZ22, LZL23, LQW18, LJ18, LL07b, LWX19, MWZZ23, MGSS22, MS99, MS16, NV10, NSD24, PM17, PGB15, PTC12, PP23, QZX23, QX21, Ryn00, Sch09, SW96, Spe23, SLK20, Sto17, SXWG22, Tau19, TZ07, TGA96, WL12, WZJ18, WZ19, WSL^{+}19, WWH21, WZzS21, WZ10, Xu14, YZLL24, YXM23, ZT14, ZSB01b, ZHY19, ZHzSR21, ZZL^{+}24, ZW12, dCFCF20]. equation-based [FKT20]. equations [AL21, AG98, AB02, An20b, AB23, Ang00, APV14, ACPV21, ANGV19, ACH10, AHC11, Awa15, Bai99, BNP14, BP93, BPW95, BCZ23, ByLl01, BF11, BV24, BK16, Boc04, Bra01, Bre95, BFCIV22, CDP03, CR23, CCG10, CJX17, CHS21, CMX97, CP99, CWX06, CLL08, CLZW13, CS19, CLCH19, CZ20, CtTB21, CS94, ÇY23, CLM18b, CLM18a, CN96, Cum96, CT17b, DFR07, DM93, DPS93, DX20, DM19a, DM19b, DKK22, DS01, DKMT20, DNS21, DL23, DX10, EOR18, ESTW19, ER99, FH05, Fas99b, FMP22, FR10, dLL19, FZLZ21, GS18, GAN20, GRV22, GH99, GLO10, GH18, GHJZ23, Gu17, Gu19, GH08, GQ22, yGS08, yGqW09, GW21, Gut21, Han06, HBMY14, HMX22, Hau97, HL06, HWZ22, HKR21, HQR23, HK21, HBM03, HS08, HCYY20, HSY16, HY18, HH19b, Hua06, HZHL12, HW19a, HJYZ23, Jan98, JJLR14]. equations [JSSE97, JLZ08, JY22, JK08, JRS15, KNP98, KY01, KKP19, KL17, KHM16, KC16, Kum23, LP21, LD22, LN14, Lee03, Li04, Li08b, LC09, LC13, LD16, LCC19, LZ20a, LZY98, LMV11, Liu99, LHY08, LY08, LLC16, LC19, LGN22, LWWZ22, LK20, LLY12, Lut04, MS22, Mai20, MN16, MS18, MXO13, MG19, MMNM23, MY17, MS98, Nai12, O’N18a, Pen98, Pet02, Pla12, PS98, PF23, Qiu23, RZ98, She13, SY19, SS15, SZJ21, SvVW22, Sun14, TH19, gTpM02, TQ17, TTC21, TSZ98, VW02, VNS18, lWyG06, WY16, WY17, WW19, Wan19, WMW21, WLW23, WK20, WQP23, XH13, Xie08, XY20, XZ03, YTL^{+}20, YL24a, YZ13, YZ21, YZ24, ZVW95, Zen19, ZL24, ZV08, ZT15, ZYB18, ZCL20, ZJZ21, ZGV22, ZD14, ZCY20, ZQS22, ZhXpZ16, ZWZ19, dH94, dFGAJN18, vdHv97]. equilibria [dCFCF20]. equilibrium [AKKN22, BSSM22, sCLC13, HQHV20, HMQ22, HH19a, KS23]. Equivalence [SC08]. equivariant [Kni24]. Erasures [CK03a, AS24]. erasures-a [AS24]. ergodic [ZXC12]. ERM [ZXC12]. Errata [Ano04c, Ano08]. Erratum [Ano16, CLM18b, Dũn12, Fus08a]. Error [Alz10, Ang00, Ano01b, BFFM24, CCG10, CDM^{+}24, Dav23, DM22, DMS22, GAN20, GRB12, yGW07, KF03, LB93, Per22, Sch95, YDF97, ZL21, aKT17, AK01, ARPR01, BN19, Bar03, BBB21, BGH15, Bra01, CLC16, CR23, CC22, CBK01, CK11, CLZW13, CCE^{+}12, DFOS23, DY01, DLZ16, ER21, GMBS23, GLX23, GHJZ23, GN13, HORU19, HR01, HWZ13, JS21, JK08, Kac18, KY01, KK05, KLS19, KMMI19, KL17, KNV01, KK20, Kun01, LC21, LS20, LMV11, LMT01, LY01, LWW21, LGN22, LL10, LL12b, MS22, MSZ06, MRH15, MW01, NSW16, Pet00, Pet02, PS05, PP23, Qiu23, RGB13, RDEG24, Sch15, SWH20, SY19, TZ03, WZJ18, WXWL19, WZzS21, WMW21, War13, WSM13, XHC15, Xu14, Yan01, Yan03, YS23, Zha01, ZSB01a, ZHzSR21, ZW22, ZC14, ZL20]. error-optimal [Sch15]. errors [CLMR15, DM15, Nes16, Sch23, ST23, Wan08]. ESDIRK [BBB^{+}16]. essence [JAA22]. Essentially [MTH21]. estimate [BBB21, CC22, CK11, CLZ02, DY01, LL10, MHR21, Qiu23, ZHzSR21]. estimates [ARPR01, AL97, BFFM24, CLC16, CHS21, CLZW13, DFOS23, DMS22, DG16, FS98, Füh08, Fus08a, Fus08b, GMBS23, GRB12, Gom95, GLX23, GHJZ23, GZ13, GN13, KY01, KL17, KNV01, KK20, LS20, LMT01, LY01, LWW21, LGN22, Low05, LL12b, MS22, MSZ06, Per22, PS05, PP23, PR99, RGB13, Sch95, SY19, Spe17, WZJ18, WZzS21, WMW21, War13, WSM13, YS23, Zha01, ZC14]. Estimating [ER21]. Estimation [Ano01b, ST23, BN19, BL09, BMP^{+}19, Bra01, Ehr00, KL19, KF03, Kun01, MRH15, MW01, NP24, WK93, aKT17]. estimations [WZM21]. estimator [HORU19]. estimators [AK01, Pet02, Yan01, Yan03, ZSB01b, ZSB01a]. Euclidean [Bün11, Kun09, NSW98]. Euler [AK15, Boc23, CLC16, JL19, KvL95, XKY15, ZL24]. evaluated [ST23]. evaluating [WZtX16]. Evaluation [CWLH95, GP13, Bar03, Bre19, CN96, DP13, GG16, GR19, GJW20, IP15, Joh13, Lee21, Pan95, Sch23, SB22, WG19, aKT17]. evolution [GH99, GH18, MU14, RS19, SvVW22]. evolutionary [BPW95, LWWZ22, Xu14]. evolving [ABMM19, WK20]. Ewald [Tor16]. Exact [And20, AR23, ATS19, SLN14, ZV08]. exactness [HW12]. example [PW94]. examples [Kle21]. exceptional [Boc23]. Excesses [BCHL03]. Existence [MSS02, CCH^{+}11, DX20, Far10, LM06]. exotic [KXL24, lZmCRX03]. Expanded [SPS18]. expansion [CL07, HLN12, LL07a, NP18, aKT17]. expansions [Atr12, CI13, Dam07, DHO16, DGS18, GJU22, HWZ13, JLZ08, LWW21, LT13, Lóp18, RTV21, Tem07, WZL17, WWX18, YX24b, vdMNS03]. expectation [HCF^{+}21, ZTHW22]. experimental [CBK01]. Experiments [Wil98, KMO^{+}14]. expiry [ALADH^{+}19]. Explicit [ALRÁY22, BCZ05, CHL17, CHMR97, CMR07, CLMR10, Hal95, HH15, Nie98, PIZ01, ZP06]. exploitation [AG98]. exploiting [AD08]. Exploring [CDTV99]. exponent [Don19, Pla00]. Exponential [BD21b, CM12, JKLY13, KSUW23, ZPR03, ZCL20, CDPS22, CsL18, Don19, LW23, LM06, Miy19, NV10, RCJ^{+}23, WZtX16, ZL17, ZP19, BD21a]. exponential-polynomial [CDPS22]. Exponentially [CFC23]. exponentials [CCS16, JGW15]. expression [CGDHRÁ^{+}19]. extended [SSK24, Yem22]. extension [AVGGEG^{+}19, CMR22, DZ04, TY18, YX24a]. Extensions [CG19, HR13, CKK16, GGAVGG^{+}19]. extensive [CHS20, SO20]. exterior [GMM09, yGS08, yGjW10, LST11, LG13, Lic22, SH23, TH24]. extraction [MSW^{+}22]. extragradient [AKKN22]. extraordinary [Pra98]. extrapolated [aZWL23]. Extrapolation [LLZ10, fLL05, Xie08, BRZ94, CL07, HZHL12, JLZ08, RZ98, Sid17]. Extremal [SW04].

factor [Han98]. factorisation [Gér19]. Factorization [CCS16, AH96, GKS19, GMRS97, Sau06]. factorized [NBL11]. factors [MN15]. fairness [GU02, KP04]. false [Kle21]. Families [MH17, CL02, Gro13, Led15]. family [DLUS17, KST21, Wat94, Xu98]. Fan [Hof06]. far [SFDE15]. Fast [AZW15, BCM99, GS19b, HPS19, IP15, KXL24, Khl18, KP98, KCCV23, KN14, MB20, MFB^{+}11, Par22, AC00, BMB14, BA15, Bra01, Cai02, CJX17, Che02, CLL08, DPS93, DX20, DHSW19, DES21, DLK12, GG16, Gut21, Han05, JZ10, Joh13, KS23, LD22, LT98, LLY12, QX21, SFDE15, WLB22, YB21, YZ13, ZGV22, SF14, Ste98]. fast-multipole [WLB22]. Fatou [Cho95]. FBPINNs [MMNM23]. FDM [ATS19]. feasibility [sCLC13]. feasible [ZLPX23]. feature [AAPCC24, MSW^{+}22]. Feedback [AHS20, AKS23, KP18]. feedforward [Mha93]. FEM [ADG17, AMK19, Don19, EEK21, FMP22, MW01, MS98, Per22, SPS18, SY19, Spe23, WBY24, Yan03, ZSB01b]. FEMs [GS18]. Fer [Ram18]. ferromagnetic [APK20]. FES [ZRR23]. Feynman [NSD24]. FFT [BZI19, BKP20, DG16]. FFTs [HKR04, Nes16]. fidelity [BZ18, GDB24]. field [BK24, BKM^{+}21, DD94, GHKL18, LMYL19, LGC^{+}20, LS20, MSW^{+}22, QZX23, SFDE15, ZY20]. fields [BH02, GG16, LW23, Sor18, TS05]. Filon [XHC15]. filter [CMX07, CW09, EGL13, Jia09b, TL18]. filtering [MP23]. filters [CL02, Jia09b, MWW13, dVR19]. financial [ALZ02]. Finding [NZ22, Yao16, Yao19]. Fine [JN18]. Finite [BF03, CG17, EHS02, FKT^{+}19, JK08, KP18, LMWX13, MMNM23, NBL11, VW05, AK01, ARPR01, AB11, APV14, AL97, AAQ15, BNP14, BDYY20, BBdD21, BDFS19, BKK17, BK24, BEPS96, BK96, Bra01, CL12, CDP03, CLLS06, CCG10, CZZ17, CMP23, CBK01, CCN^{+}16, CLZ02, Che06, CLL08, CWX12, CGH^{+}13, CLZW13, CLCH19, CFZ21, CEHK18, CKW10, CLR19, CLM18b, CLM18a, CH22a, Dav23, DHWZ18, DY01, DLZ16, EBS22, FF95, FWW06, FYL08, FZLZ21, GQL24, GAN20, GMM09, GK19, GLX23, GH18, Grö03, GQ22, HMX22, HZZ20, Hem95, Hem96, HR01, HW18, HCYY20, HL21, HCH20, HX02, HL20, HS20, HZZZ24, JS21, Jan98, JLZ08, JRS15, KY01, KST21, KS18, KORS17, KH05, KL00, KHM16, Kum23, LLS94, LLS96, LSY24, LLS03, LC18, Leo13, LC09, LC13, LGC^{+}20, LZ20b, LSY22, LZL23, LLR24, LZ21]. finite [Lic22, LL07a, LB93, LLC16, LZCW21, LGN22, LPSSP00, LWX19, LL10, LL12b, MR15, MN16, MSZ06, MC08, MY17, MTH21, NV10, Noe95, PSW24, PL05, PP97, PP23, SL15, SW96, SXYY06, SRFH12, SvVW22, SLK20, Sun08, Svá19, TZ05, TH19, TZ07, TH24, WY16, WZL17, WZJ18, WXWL19, Wan22, WLPV15, WQP23, Xie08, XY15, XY20, XZ01, XYZ21, YDF97, YS23, YX24a, YL24a, Ye00, YZ21, YZ24, YXM23, ZRR23, ZST19, ZZP23, ZSB01a, Zha14, ZPY15, ZT15, ZYC19, ZCB21, ZW22, ZW23a, ZD16, ZSHZ15, ZW20, ZW23b, ZW23c, dFGAJN18]. finite-difference [JRS15, KY01]. Finite-element [NBL11, KS18]. finite-frequency [BDYY20]. finite-part [WZL17]. finite-volume [Hem95, Hem96]. finite-volume-finite-difference [CEHK18]. finitely [MNW04]. fire [ABMM19, FPR^{+}12]. first [CLMR10, HR02, HZHL12, LWWZ22, Nai12, Pla12, ZLPX23, Zhu19]. first- [LWWZ22]. first-order [ZLPX23, Zhu19]. fitted [CFC23, KL16, ZL16]. Fitting [JF02, BCM99, DZ04, DSS09, ECS18, MJM23, RW21, WK93]. Fix [GW17, WL05]. Fixed [CZS16, LSXZ15, MHR21]. fixed-point [LSXZ15, MHR21]. flat [KS20, SRFH12]. flatness [GL17]. Flexible [HL05, MC23]. flights [ZD17]. Flory [ZY20]. flow [ACM^{+}22, ATS19, BDFS19, BSSM22, CLL08, CT17a, DHWZ18, dLL19, GHKL18, HFH15, HJYZ23, KS18, KH05, KCCV23, KESR22, LG13, Nit22, WBGG22, lXyG02, Ye04, ZPY15, ZGV22]. flows [BBB^{+}16, CD15, Don20, FZ19, GHLU19, GJMY19, KBL22, KMMI19, KvL95, PIZ01, PCM^{+}96, PJ20, RHZ17, SBZR19, TL18, Tha19, TH24, ZS24, aZWL23]. fluctuations [KK20]. fluid [AMK19, KMMI19, PJ20]. fluids [BDFS19, DGMM16, KRP19, SJJY21]. flux [CtTB21, DLZ16, LK20]. focal [PW08]. focusing [LRY23]. Fokker [WSL^{+}19]. fold [Jia11]. FOM [Cul96]. forced [FKT20]. Foreword [Buh99, HN96, How96]. form [AM14, CCLM21, FGS16, Hub12, Khl18, LL99, LKW17, Rei93]. Formal [Sal98]. format [SSK24]. formed [Mai06, SEE96]. forming [RTV21]. forms [GGW24, Loh23, Mar95, SM00, ZRR23]. formula [CG15, Fra99, SS13]. formulae [CG04, GS10, MC10, Xu98, Xu00]. formulas [BB07, FXZ96, Jeo98, SX07, Wal95]. formulation [BKM^{+}21, CL18, DGMM16, HJH12, Lep23, MS16, Par20, RS19]. formulations [TW98]. Fortran [Kir98]. forward [NSD24]. FOSLS [RS19]. four [GMS12, NSS04]. four-directional [NSS04]. four-point [GMS12]. Fourier [SF14, Ste98, BZI19, CGW13, FQ10, JSW20, LS20, LL21, LL07b, LP14, QW11, SL05, SS13, TZ03, YX24b, YCQZ24, ZL21]. Fourier-spectral [LS20, ZL21]. fourth [CFZ21, DNS21, HWZ22, ILHH22, Li05a, LCC19, MS19, WHTZ18, WZzS21, ZZC20, ZW22]. fourth-order [CFZ21, DNS21, FZLZ21, HWZ22, ILHH22, Li05a, LCC19, MS19, WZzS21, ZZC20, ZW22]. Fractional [MP18, AMK19, BlH23, CC22, CDM^{+}24, DKMT20, FZLZ21, GM17, Gu20, GW21, HB24, HX17, HS20, JS21, JlLZ20, JLW20, KM19, KS17, LD16, LJ18, LZCW21, LWW21, LBCZ18, MS18, MNC16, NSD24, PM17, PF23, QX21, Ryd19, Sou22, SU12, TDG16, WZ19, WLW23, YCQZ24, YZLL24, YXM23, ZJZ21, ZHzSR21, ZD14, ZD16, ZW12, ZW23b, ZW23c]. fractional-order [Ryd19]. Fractional-step [MP18]. fractional-time [BlH23]. Frame [AS21, CR08, MRS18, BlH23, BMRS20, CP14, CGG^{+}20, DFR07, JAA22, Jia09b, KKL03, LL11, LS18, MR10, MRS14, PŠWX03, PWL20]. framelet [CCSS09]. framelets [Lem09]. Frames [BCHL03, BC20, BF03, CK03a, CC03, GH03, HM03, HLT11, Jia03, Sun08, AS24, BS06, BH15, BLP19, CL12, CCH^{+}11, CK07, CCN^{+}16, CKK16, CJKK18, CGKK21, Chr15, CSS98, CHSS03, DST04, DM15, Far10, FGG^{+}17, FM18, FS07, FOW14, FWW06, Füh08, Grö03, JJK13, LPY10, LMO14, LH13, LA22, LL13, MNPW00, NS21, Onc14, ON18b, PR19, Søn07, Zha05, CS03]. framework [AM24, ABC14, BFFM24, BFCIV22, CL22, DAP13, GZA20, Par20, Pra23, WZtX16, ZRR23]. Franklin [NP18]. Fredholm [DM19b, FR10]. Free [KL04, PCM^{+}96, BGT20, CG17, DAP13, JW95, KL22, Kum23, KF19, LAG^{+}24, LLS03, LR19, Ock96, Urb95, WWH21, YZ21, ZY20, ZZL^{+}24, ZL16]. Free-knot [KL04]. freedom [Loh23, ZRR23]. Freeknot [KL07]. frequency [BDYY20, GG06, KCW17, KL19, LA22, LHY08, MGSS22, Onc14, RST^{+}14, SF14]. frequency-concentrated [GG06]. friction [DKK22, Don20, HKRS14, XKY15]. Friedman [NP18]. Friedrichs [BDM20]. Froissart [BC99]. Full [DS01, Fou23, ASS10, Spe23, Tyg23]. full-space [Spe23]. fully [CHS21, GAN20, HS20, LWL17, VNS18, Wan19]. Function [HM03, Alz10, AM14, Boc04, BEG22, BCZ05, CTZ02, CH22b, CGH^{+}96, DSW05, EH24, EGST13, ER21, FF05, GR19, IB17, JW95, Kot08, KSUW23, LP19, LW14, LLY06, LL99, LT13, Mac94, Mai10, MNW99, MP00, NG99, Rip99, SMK22, Sch95]. functional [Gat21, Liu99, RLGGAV19, SSK24, Sun14, ZD17]. functional-differential [Liu99]. functionals [MR10, SM00, Tad19]. Functions [Zoz15b, BCM99, Bei23, BGH15, BGW24, Bes03, BMB14, BZI19, BRR18, Bre19, BDM20, BW15, Cai02, Car95, Cas07, CP03, CX04, CSW14, CL16, CTZ02, CG10, CJKK18, Chr15, CSW96, De 06, DG16, DLU19, DX10, DGS18, Fas99a, Fas99b, FPR^{+}12, FHN94, FS98, Füh08, FW15, GP13, GHdN12, GG06, GGL07, GJU22, Gro13, GH16, GS10, yGSlX03, GYZ17, Han98, Hub12, HJ23, JS99, Jüt98, KKLY10, KL04, KL07, KR19, Le 05, LST11, LM14, Len96, Li04, Li05a, LK05, LWL17, LWW21, Lóp18, Low05, LP14, MM00, MZ22, MST05, MNW04, MS04, NP18, NSW16, PŠWX03, PRS24, Pet00, Pin04, PT21, RZ10, RS99, RY13, Sau06, SW06, Sch11a, SS08, SRFH12, Sto96, SM99, SYCJ18, TZ05, War13, WH07, Wen95, WK20, Wu95, WFL02, WZtX16, Xu98, YX24b, Zoz15a]. fundamental [AC05, Ant18, FK98, LL98, Li05b, Li08a, Li08b, SB22, WZ10]. fundamentals [HCF^{+}21]. furnaces [BGMS07, BGG24]. fusion [CCH^{+}11, CP14, HB24, LH13]. fuzzy [DLK12, Tal19].

G [DDP14]. G-symplectic [DDP14]. Gabor [CL12, CC03, CKK12, CJKK18, CGKK21, Chr15, DM15, FS07, FGO14, Füh08, JJK13, LL13, LL11, LW03, ON18b, RST^{+}14, Søn07]. Galerkin [AMK19, AU20, AK15, AS05, BN19, BQRB13, BDFS19, BDM20, CMX97, CMX02, CP15, CT17a, ÇY23, CHP19, CH22a, Cop98, CCE^{+}12, DHK98, DS01, FLM96, GS18, Gia20, Gom95, HL06, Hua06, JL19, KNP98, KF03, Kum23, LN14, LC20, Lep23, LZ20b, LLC16, LGN22, OBS15, Ryn00, SXZ06, SH23, VNS18, WY16, WY17, WXWL19, WaZL24, WY20, YS23, Yan19, ZZP23, ZB12]. Galerkin-characteristic [CCE^{+}12]. Galerkin-mixed [GS18]. Game [AVGGEG^{+}19]. gamma [GR19]. Gasca [CP07a]. Gauss [Dou94, yGqW09, LT13, MWZZ23, SC08, WZL17]. Gaussian [CC94, CSW14, KS20, MC10, RS99, SX07, SHTS14]. Gaussians [Xia13, YZ08]. GC [CG19]. Gegenbauer [ESM13, WZ20a]. gene [CGDHRÁ^{+}19]. general [AN20a, Alb15, ABC14, BB24, BCZ05, BRZ94, CH22b, FHK06, Gut21, HL16a, HL18, Han06, HKR21, Kos15, LS19, RL07]. generalisation [SU22]. Generalised [HJ23, CtTB21]. Generalization [NG99, ZXC12, CGP95, CK07, CKBP11, JS95, MNPR06, ’Sw19]. Generalized [BRS13, CdGDN20, Dar03, SL05, WL05, gYyG12, Zoz15b, AB11, AKKN22, BR23, CLC16, CG15, sCLC13, CHL17, DP06, GW17, GP13, Gat21, GRB12, GH99, GW14, IS13, KRP19, fLLfDX21, LWL17, MC23, MN00, Pen98, SF14, ZL20, Zoz15a]. generalized-functions [Zoz15a]. generally [Gon93]. generated [CG06, Chr15, CSW96, JS99]. generating [CK07, NSS04]. Generation [FWW06, KF19, SR02, CGR11, Cop98, RLGGAV19]. Generators [Jia03]. Gennes [ZY20]. geodesic [BMK15, SLN14]. geodesics [NZ22, SLN14]. Geometric [GPG15, HNN23, Mai10, MO16, PFMS05, BS06, BI11, BWX97, FH05, JKK^{+}12, Kni24, LM04, Leo02, LO08, MS18, RW21, Wan19, ZLPX23]. Geometrical [Ano02d]. geometrically [ZXC12]. geometries [GJW20]. geometry [BZA24a, BZA24b, BBF^{+}24, FÁG19, Fin19, NS21, PW08, PGM09, PGB10, RAB^{+}09]. geothermal [GQL24, LGQ24]. Gersgorin [Hof06, CKBP11]. Gibbs [RY13]. Ginzburg [GS18, HX02, LL07b, ZHzSR21]. given [BSK19, NZ22]. Givens [CM99]. Global [Kim06, KP18, LSW17, Len94, Yao19]. globally [BF94, LWZ19]. GMRES [BCM99, Cul96, MGSS22, MS98, Smo07, WLB22, dH94, dCB07]. GMRES/FOM [Cul96]. good [Rip99, Sch11b, WS01]. Gordon [LZ22, GW14, WZJ18]. Gordon-Schrödinger [WZJ18]. GPU [ZBT^{+}22]. grad [JJLR14, dFGAJN18]. graded [CDM^{+}24, QX21]. Gradient [DY01, BMK15, DSS09, DDLL23, HJYZ23, Par22, TTC21, Wan12, Yan01, Yan03, YTL^{+}20, Zha01, ZS21]. gradients [YWC12]. Graduated [BCE^{+}09]. Gram [CK07]. Gramian [BH19, Ryd19]. graph [DGMES13, DT12, LN19, MSL21]. graphs [KK02]. Grassmann [BZA24a, BZA24b, ZS21]. grating [BBR02, LN14, Wan22]. Greedy [GN01, Tem01, Tem07, TYY11, CDPS22, DT96, GN08, LT06, LT12, Pra23, QW11, Tem98, Tem00, Tem02]. green [FY13, Cai02]. Gregory [Ano93]. Greville [CHY11]. grid [BNP14, BMK15, Che06, CLCH19, CH22a, CH22b, DG16, Hem95, HCYY20, HCH20, KS18, KHM16, MR15, PGB15, XZ01, ZZP23, ZTHW22, Hem96]. grids [ANSZ17, BNR00, FF05, Hag98, KXZ04, LQ14, Noe95, Pla00, PS05, SXYY06, Ste98, WWX18, Wie97, Zha01]. Gross [LZ20a]. Grossman [JMO00]. groups [MP23]. growth [RH15, RS99, RMCCG^{+}19, ZST19]. Grünwald [CC22]. Guaranteed [HFH^{+}07]. guarantees [HD20, Kac18]. Gudkov [Hof06]. guesses [CV23]. Guest [Ano95]. guided [BPJ02]. Gupta [Ock96].

H [WZJ18, WY17]. H2Opus [ZBT^{+}22]. Haar [ZhXpZ16]. habil [Sil99]. Hadamard [WZL17]. Half [YF23, GG16, Li94, LWL17, lWyG06, gYyG12]. half-plane [Li94]. Half-quadratic [YF23]. half-space [GG16]. halfband [MWW13]. Hamilton [TQ17]. Hamiltonian [AMM20, ABI15, CCDL20, DDP14, HBH24, JM00, Mar94, MO16]. Hammerstein [KNP98]. hand [FR10]. handbook [BZA24a, BZA24b]. Hankel [AM14, BW20, Li94]. Hankel-norm [BW20]. Hanrahan [Han05]. hardening [GN13]. Hardy [CS06, HN18]. harmonic [CCP22, FMP22, HS13, HY18, KP98, LHY08, LWX19, Sch09, WWX18, WZtX16, Xu04a, YX24a]. harmonics [FM14]. HCT [LLS94]. HDG [LC21, MNC16]. heat [BDMR10, Bru23, CN96, FKT20, HT11, HXC10, JGW15, LM14, LK20, MS16, PCM^{+}96, Tau19, TGA96, WG19]. heat-conduction [CN96]. heating [BGMS07]. Heaviside [RY13]. Hele [LYF^{+}21, ZL21]. helical [Mon09]. Helix [PW94]. Helmholtz [BNPP20, CHP19, DRS08, DLS14, GMM09, GPS21, HH15, HJH12, HY18, Kir98, Li08b, Li09, MGSS22, Spe23, Sto17]. hemivariational [Yao16]. Hermite [ABD08, ABY13, BCHS22, CCMM21, CCS16, CMSS23, DM99, EH24, ECS18, FaKS02, Fas99a, Fra99, Füh08, yGSlX03, Hak00, HNN23, JL06, KJ09, Luo99, MST05, MS12, Moo17, PFMS05, RW21, RSA14, SX95, TQ17]. Hermitian [WMY13, WMB13]. Hessenberg [EGST13]. Hessian [DNS21]. Heston [Zhe23]. heterogeneous [NW22, TS05]. Heuristic [MN15, DGMES13]. Hexagonal [Jia09b]. hidden [CLM96b]. Hierarchical [Spe17, AU20, BCC22, HORU19, Han05, SO20]. High [BNR00, CEHK18, Gat21, HBMY14, IRT23, LD16, LZ22, NP18, Nov00, SLK20, TH19, ZD16, ZZ09, AU20, Bei23, CLM03, CCK20, CC24, EH24, ESTW19, GL23, GM22, GDB24, Han10, HZZ20, HKR04, HS03, JMO00, Jia09b, JN18, LL12a, LWX19, MS04, NKL^{+}19, PW24, Ram18, TL18, TTC21, WZJ18, ZJ23, ZZL^{+}24, Yem22]. high-dimensional [AU20, CCK20, EH24, ESTW19, PW24]. high-fidelity [GDB24]. High-order [CEHK18, HBMY14, IRT23, LZ22, SLK20, GL23, GM22, HZZ20, HKR04, JN18, NKL^{+}19, Ram18, TTC21, ZJ23, ZZL^{+}24, Yem22]. high-pass [Jia09b]. High-precision [NP18]. High-resolution [Gat21]. Higher [CWX12, Fin19, MRS03, ACK19a, ACK19b, CDP03, Füh16, GW17, KL16, Mat09, Noe95, OBS15, WOBL17, Wie97]. Higher-order [CWX12, MRS03, GW17]. highest [ZD18]. highly [DGMES13, KXL24, XHC15]. Hilbert [BMB14, BMP^{+}19, Chr15, Dam07, HT11, MXY13, RY16, SM99, Yao16, vdMNS03]. Hilliard [CZ20, LYF^{+}21, WaZL24, ZL21, LGC^{+}20, SXWG22, ZY20, ZZL^{+}24]. Hilliard-Darcy [WaZL24]. histogram [CP07b]. Hodograph [RM19, AF96, FS94, FaKS02, FGS16, JKK^{+}12, KJ09, PFMS05, Pot95, RSA14]. Hölder [AAFV19]. holonomic [WOBL17]. homoclinic [CS93]. homogeneous [AK01, BC99, CHP19, DST04, DMBH24, GJMY19, Jia09a, Kni24, LR07, ZGGW12, ZN18, ZS24]. homogenized [Str95]. homological [PR15]. homology [PR15]. Homotopy [PM17, AF96, Tim21]. Hood [fLL05]. Hopf [CGS93, ER99, Far10]. hot [Kle21]. Householder [FWW06]. hp} [CMM21, CHP19, CH22a, GW21, Spe23, WY20]. hp}-FEM [Spe23]. hp}-Version [CHP19, CH22a, GW21, WY20]. HT [AU20]. HT-AWGM [AU20]. Huggins [ZY20]. Hunting [CS93]. Huygens [IKS10].

Hybrid [CM96, LQ14, WG19, Yem22, ACPV21, BPJ02, CCNT21, sCLC13, CEHK18, HQHV20, LZ20b, Per22, PR19, Zen19, ZYC19]. hybrid-mixed [ACPV21]. hybrid-variable [Zen19]. hybridizable [LC20]. hydrodynamic [GZW18, NKL^{+}19]. Hyper [DM19a, AZW15]. Hyper-reduced [DM19a]. hyper-reduction [AZW15]. Hyperbolic [Len96, CLM18b, CLM18a, CKM99, DLU19, KK05, KF03, MV01, ZQ13]. hypercube [WZ20a]. hypergeometric [CDTV99, DGS18, LT13, NP18]. hyperinterpolation [Li08b]. hypersingular [DLS14, HWZ13, MS99].

ideal [She08]. idealized [Jia09b]. ideals [MS04, Xu00]. identical [XP10]. identification [BH23, BJX09, CJW22, FYL08, GZA20, SJJL15, Tan17]. identity [GMB97, TSY10]. II [BN21, CMX97, DPS93, HL06, Hof06, MPS97, SS13, ZTHW22, zCFX07]. ill [Ant18, BL09, BE00, LLY12, XY12]. ill-posed [BL09, BE00, LLY12, XY12]. Image [LN19, BlH23, BCE^{+}09, CZS16, Del17, DLK12, GLM23, HB24, ILHH22, JZ10, LSXZ15, MSXZ13, RWT13, SX09, Zhu19]. images [HMS19, RY16, SF14, ZW19]. imaging [BKN11, DGK15, Par20, KMO^{+}14]. immersed [CLLS06, CZZ17, CKW10, GPG15]. immiscible [BDFS19, SJJY21]. immobile [LZCW21]. impedance [BR22, HLN12, QC12]. implementation [DHO16, Jan98, Svá19, dH94]. Implicit [EBS22, Wie15, XKY15, BB07, CLC16, CLM03, EJ97, GH08, GPG15, HBM03, JSW20, JL19, Mui99, Wan23a, Wan23b, vdH93, vdHv97, vdHS00]. implicit-time [GPG15]. implicitly [FMR00]. implicitly-defined [FMR00]. imply [Sch23]. importance [YL24b]. imposed [Awa13]. improper [GGAVGG^{+}19]. Improved [Fus08a, Fus08b, KL17, FJ99, Sto17]. Improvement [BCC22, ZL24]. imputation [RF19]. inclusion [CGM18]. inclusions [nCV13]. incomplete [GR19]. incompressible [An20b, Ang00, BQRB13, BBB22, CLL08, DHWZ18, DL23, GHLU19, GJMY19, HQR23, HFH15, KS18, KH05, LHH23, TH24, lXyG02, Ye04, ZVW95, ZS24]. inconsistent [Sou22]. increasing [CSW14]. increasingly [KS20, SRFH12]. indefinite [YZ24]. independence [Ito96]. independent [CTZ02, DS00, DSW05, SB22]. Index [Ano02a, Ano02b, Ano03a, Ano03b, Ano04a, Ano04b, Ano00a, Ano01a, Cum96]. indexing [dB00]. indicator [ZQ13, ZG16]. indicators [KLS19, LQ14, SBMS22]. induced [CGR13]. induction [BGMS07]. Inequalities [GG06, Alz10, Bad19, BR07, RZ10]. inequality [JRS15, LB93, Yao16, ZLPX23]. inertial [AKKN22]. Inexact [WLB22]. inf [dFGAJN18]. infiltration [ZST19]. infinite [BD10, Buh06, FF05, HN18, KNV23, NW22, lXyG02, Zhe06]. infinitely [RZ10]. influence [BS23, CV23, EHV17]. informed [DM22, MMNM23]. inhomogeneous [LK20]. initial [Alb15, CJW22, CV23, Kac18, WY20, vdH93]. initial-value [vdH93]. injectivity [FK10]. inkjet [ALADH^{+}19]. inner [Joh13, Pin04]. innovation [Sun08]. input [BHM18, CP15]. input-output [BHM18]. insertion [BGM93, KLNS95, Wal95]. instabilities [Sch23]. instability [KKP24, TBD94]. integer [CHSS03, Ryd19]. integer-order [Ryd19]. integrability [RS01]. Integral [KNQ16, AG98, AR20, Boc04, BLV20, CWLH95, CMX97, CP99, Che02, CHS17, Cum96, DX20, Dar03, DD94, DES21, FR10, dLL19, FKT20, GHKL18, Gil17, Gna07, GJMY19, Gu17, Gu19, HBMY14, HKR21, HJH12, HR02, Hua06, HZHL12, HS16, JSSE97, Kir98, KN14, LD22, LN14, LC18, Li09, LZY98, LQW18, LLY12, MS99, Nai12, O’N18a, Pla12, PF23, RZ98, Ryn00, TW98, TSZ98, VC00, VNS18, WL12, WM21, XH13, Xu14, YZ13, ZGV22, ZW23b, ZW23c]. Integrals [Zoz15b, CLMR10, CKP13, DL98, DG13, GGAVGG^{+}19, HWZ13, IRT23, MV99, Nit22, NUWZ17, TX19, VC00, WZL17, XHC15, Zoz15a]. integrand [ACH21]. integrate [FPR^{+}12]. integrated [BPJ02]. Integration [HW00, DDP14, DL98, ESM13, HW12, KXL24, KR19, MO16, Nov00, NSW16, Pei05, SW04, ZB12]. integrators [CCDL20, HPP^{+}19, KL16, KL19, OBS15, WOBL17, WZtX16, ZZL^{+}24, dLdDSM07]. integro [Gut21, JS21, JLZ08, QX21, WY17, ZV08]. integro-differential [Gut21, JS21, JLZ08, QX21]. integrodifferential [Qiu23]. Interaction [LRY23]. interactions [APK20, SMK22]. interface [BK96, CZZ17, CKW10, DLZ16, FZ12, HH19b, KP95, SH23, WHS18, ZL16]. interface-fitted [ZL16]. interfaces [BK96, ZMQ23]. Interior [AL97, Lep23, LWX19, NV10]. interpolant [Dũn09, SW08]. interpolants [ABC14, ABY13, GMB97, MS07, Rem12, Spe17]. Interpolating [CT05, AF96, BL99, CGM01, DL98, GRV22]. Interpolation [BLS02, CG06, DG13, GSS22, GY99, Noa02, Pet96, RGB13, dB07, AAPCC24, AFM23, ABD08, AM14, BNR00, BCHS22, BS05, ByLl01, BGW21, BGW24, Buh06, Cal05, CDPS22, Car95, CG04, CS06, CM03, DSW05, DS10, DAP13, DM99, FM21, FaKS02, FGS13, Fas99a, FP99, FE15, FF05, Fra99, GS00, GRB12, GO05, GGW24, Hak00, HL16a, HL18, HQR23, HNN23, JKK^{+}12, JLM19, KJ09, KF19, Kun09, Kva14, LL98, LW04, LM04, Le 00, LLY06, LL99, LR07, LZY98, LL07a, Luo99, LNW02, MS00, NSW98, NZ04, NSZ04, Osw04, PFMS05, RZ10, RS99, Rip99, RSA14, Sau95, SX95, Sch95, Sch11b, SS96, SS15, SRFH12, TBG18, Ven94, WS01, WL05, Xu00, Xu04b, dB00, dB06, vdMNS03, zCFX07]. interpolation-based [GGW24]. interpolation-free [DAP13]. interpolative [DM23, MB20, VM17]. interpolators [Led15]. Interpolatory [DHGA23, RGB18, Zhe06, CGR11, CGR13, DG16, DLUS17, JS99, KKLY10, ZA10]. intersecting [LD07]. intersection [BX94]. intersections [HFH^{+}07]. Intertwining [Cal05]. interval [ČF11, JL06, Jia09a, MS99, PST95, gTpM02]. intervals [Kür18, LMV11]. Introduction [AX98, Bar94, HAS05, Mar95]. intrusive [PJ20]. invariance [CL12]. Invariant [MT98, BDYY20, CHL17, FaKS02, GU02, HHS04, Mai06, MXZ08, NSW16, Sun10, Wat94, vdMNS03]. invariants [DGMES13]. invasion [PLRGVR22]. Inverse [LAG^{+}24, BDMR10, BR22, CR99, CGM18, CH15, CHY10, HO15, HXC10, IKS10, NT23, QC12, RSZ11, San15, TLG^{+}22, WL12, YZB20]. inverses [AFM23, FMP22, KM19, NBL11, SU12]. Inversion [SS13, GPT17, Li03, LL12a, SZJ21]. Inverted [BKK17, BK24]. invertible [JAA22]. investigation [JY22, KL22, PSW24]. investigations [BZ18]. involving [JS21, Sto96, Tom96, nCV13]. ion [YTL^{+}20]. Irrational [yGS08]. Irregular [CC03, CLL08, FS07, Noe95, Pla95, Zha01, ZhXpZ16]. isochronous [HHX18a, HHX18b]. isogeometric [EEK21, KV21, ZP24]. isometry [SL12]. isomorphism [DGMES13]. isospectral [PIZ01]. isotropic [BCE^{+}09, PGM09, RAB^{+}09, ZW19]. issue [Ano16, BOP^{+}15, CHY10, XY12]. Issues [BPW95, Cai02]. iterated [BS97b, BS97a, BS00, CTZ02]. iterates [CCK20]. Iteration [An20b, BCM99, MSW^{+}22, RSZ11]. iterations [MGSS22, WLB22]. Iterative [Cul96, ZZP23, ZCY20, BBB^{+}16, FP99, GO95, HPS19, IKS10, LNW02, WW12, WMY13, ZT14, ZY20, ZQS22, dH94, dCB07].

Jacobi

[ByLl01, CR99, DHO16, TQ17, gYyG12]. Jacobi-type [DHO16]. Jacobian [ZD14]. Jacobian-predictor [ZD14]. Janus [KCCV23]. John [Ano93, How96, New96]. joining [NZ22]. Joint [CCMM21, FRT09, Möß10]. Jordan [CdGDN20]. Julia [Cho95]. jump [ZhXpZ16]. junction [WHS18]. Justen [HR13].

Kac [NSD24]. Kaczmarz [JWJ23]. Kantorovich [BR07]. Kármán [DNS21, MN16, Mur08]. KdV [XY20]. Kelvin [BNP14]. Kernel [YB21, BMP^{+}19, BJX09, DS10, DL04, EH24, FY13, FaKT22, Gna07, HH15, HS08, HT11, HJJV97, JSSE97, JGW15, KNP98, PS23, RY16, SMK22, SS16, SB22, Xu14, YZ13, vdMNS03]. kernel-based [DS10, SS16]. kernel-independent [SB22]. kernel-split [FaKT22, HH15]. kernels [AL21, BRS13, BRSV15, CMP21, Cop98, DX20, FM21, FJ99, FY13, GW17, Gut21, HS13, KS20, LMWX13, LS05, Opf06, Qiu23, SW06, Sch09, SHTS14, ZJ06]. Kerr [CGHK23]. Kerr-type [CGHK23]. kind [Bad19, CMX97, CP99, CHS17, DM19b, HR02, HZHL12, HS16, LL04, Nai12, O’N18a, Pla12, TW98]. Kinematic [SR02]. Kirchhoff [MM20]. Klein [GW14, LZ22, WZJ18]. Knopoff [XKY15]. Knot [BGM93, KLNS95, KL04, LY07, Wal95]. knots [Plo95, Str95]. Kohn [CGH^{+}13, GP14]. Kolmogorov [DM22, UG19]. Korobov [MZ22]. Korteweg [dLL19]. Kozakiewicz [BS23]. Krylov [BR23, ZVW95]. KSR [BF94]. KSR-1 [BF94]. Kumaraswamy [And20]. Kutta [CSWP99, SLT21, BP93, Boc23, BS97b, BS97a, BS00, CHMR97, CLM03, CMR07, CLMR10, CLMR15, Hal95, JMO00, Liu99, Mui99, vdHdS97, vdHS00]. Kutta-collocation [Liu99].

L1 [MSXZ13]. L1/TV [MSXZ13]. labels [ALADH^{+}19]. lacunary [Ven94]. lag [BP93]. lagging [MG19]. Lagrange [Cal05, CG04, Hu07, Kun95, LL98, NZ04, NSZ04, WLW16]. Lagrangian [BBF^{+}24, TH24, XWL13]. Laguerre [DGS18, yGjW10, LWL17, PGM09, PGB10, lWyG06, lXyG02]. Lambert [CGH^{+}96, IB17]. laminar [ZCL20]. Landau [GS18, HX02, LL07b, ZHzSR21]. Laplace [BBdD21, BW15, HS13, KP95, KN14, LM14, LL12a, O’N18a, PTC12, Sch09, WZ10]. Laplacian [BJMR24, CDM^{+}24, Jon17, KM19, LN19, LBCZ18, RF19, SU12, TDG16, ZW23b, ZW23c]. Laplacians [CW16]. Large [HSSS18, AMM20, BF11, BW20, JWJ23, KESR22, Kür18, LY21, SI15, Sch23, SFDE15, SB22, ZPY15, ZY20]. large-scale [AMM20, BF11, BW20, JWJ23, KESR22]. laser [GN13, NKL^{+}19]. Latin [FÁG19]. lattice [FGO14, GH99, NSW16]. Lattices [CC03, CG06, GIKV21]. Laurent [GMRS97, Lev99, MS04]. Lauricella [BW15, TZ05]. law [CFC23, JS21, KP18]. laws [KK05, MV01, XKY15, Yan19, ZQ13]. Lax [GLS15]. layer [BST24, CLM96b, FaKT22, GJW20, IMS99, NT23, ST23, SB22, WG19, ZCL20, aKT17]. layered [Cai02, Wat96]. layers [Kno09, LZZ09, ZL17]. Learning [CMP21, GS13, MNPR06, XP10, YZ08, YWC12, BFFM24, ESTW19, GZ13, GYZ17, GS19b, KKL^{+}22, Kni24, LMWX13, LS05, PJ20, RZ10, SW06, YZ23]. learning-based [KKL^{+}22]. Least [BNPP20, FHN94, JF02, WK93, CHY11, Dav23, LK05, Pla00, Sto96, SW10, Tom96, ZPR03]. Least-Squares [BNPP20, JF02, LK05, Sto96]. Lebesgue [Las16]. Leer [CT17b, TTC21]. Leer-type [CT17b, TTC21]. Leffler [GP13]. Left [SU12, ZN18]. Left-inverses [SU12]. leg [Wan19]. Legendre [yGqW09, yGjW10, GW21, HKR04, LWW21, gTpM02, lXyG02]. lemma [HHS04]. Letnikov [CC22]. level [BDFS19, Bre95, DMMS11, HL06, HT12]. level-set [BDFS19]. Levenberg [AAFV19]. Lévy [DX20, WSL^{+}19]. Lévy-driven [DX20]. Lie [Kni24, ZN18]. life [FRT09, AVGGEG^{+}19]. like [BCHS22, GRV22, ZT14]. likelihood [BL09, ZP19]. Limitations [CLM96b, KKP24]. Limited [BR23, Kür18, YX24b]. limiting [EBS22]. line [Dam07, GP13, LWL17, MV99, NUWZ17, VC00, lWyG06, gYyG12]. Linear [GPS10, GZW18, HL96, HJYZ23, TLG^{+}22, Zen19, AH96, ABY13, Bai99, BN19, BN21, BCV23, BKN11, BCS19, BDYY20, BP11, BD21a, BD21b, BDM20, BKM^{+}21, Cas07, CC24, CJ07, CD15, CFZ21, CL18, DM19b, DNS21, DL23, Dũn09, FL23, GLX23, GPAF18, GJU22, GKS21, Han10, HZZZ24, Ito96, JWJ23, KA97, KL17, KvD01, LP21, LLS03, LY07, LLR24, LMV11, LY08, MC08, MPS96, MPS97, MO16, MV01, MS16, NT23, PGM09, PP97, Sch15, SXWG22, TZL16, UG19, VK96, WMB13, WZJ18, WW19, WMW21, Wil98, ZJZ21, dH94, dCB07, vdHdS97, GJU22]. linear-constant [LY08]. linear-phase [Han10]. Linear/Ridge [GJU22]. linearization [BR22, HX02]. linearized [GS18, HWZ22, LGN22, YS23, ZYB18]. linearizing [XWL13]. Linearly [SW06, CTZ02, DS00, JSW20]. lines [BEG17, HV19b]. link [Len94]. Liouville [GM17, KS17, LS19, Ram18]. Lippmann [Che02, GM22]. Lipschitz [GH16, LWZ19, Wan23a, Wan23b, ZL24]. liquid [HY14, SJJY21, Zhu15, aZWL23]. Live [BCV23]. Ljusternik [Yao19]. LLT [zCpST09]. load [BCV23]. loads [FE15, GM14]. Local [AAFV19, ANSZ17, DHWZ18, FGS16, HZZ20, KK05, Mat09, Sun10, XZ01, YX24b, ACK19a, ACK19b, ARPR01, AZW15, ABC14, APV14, AS05, BCE^{+}09, CFC23, DL98, DZ04, GY00, Han12, HHS04, HT12, HSV09, HY18, Kac18, KK20, LSW17, Len94, LFL22, LWY06, MSZ06, MRS18, MJM23, Spe17, SZ09, ZBT^{+}22]. Localization [CP03, Lem09, MSST14]. localizations [Kos15]. Localized [Grö03, Sun14, PS23, Rau05, RST^{+}14]. Locally [DS00, CTZ02, GH16, KC16, ZGV22]. locations [DSW05]. locking [LLS03]. locking-free [LLS03]. Loewner [GZA20, Pra23]. logarithmic [JSSE97, KNP98]. logarithmic-kernel [JSSE97]. logic [Tal19]. Long [WW19, APK20, LL07b, Xu14]. long-range [APK20]. Long-time [WW19, LL07b]. loop [GQL24, LGQ24]. Lorentz [DKB99]. loss [GYZ17, PRS24]. Lotka [SC08]. Low [BD18, AAPCC24, AL21, AL97, BGT20, BH15, CLMR15, DDLL23, GS20, GKS21, HZZZ24, KvL95, LHY08, SMK22]. low-frequency [LHY08]. low-order [HZZZ24]. low-rank [AL21, BGT20, DDLL23, GS20, GKS21, SMK22]. low-redundancy [BH15]. low-subsonic [KvL95]. lower [Mai05]. lowest [SPS18, ZT15]. lumped [BGG24]. Lyapunov [LMV11, Pen98].

machine [ESTW19, KKL^{+}22, RZ10, SW06]. Machines [EPP00, CX06]. macro [AS02, ASS10, LL04]. macro-element [LL04]. macro-elements [AS02, ASS10]. macroscopic [WBGG22]. magnetic [HPP^{+}19, HW19a, LW23]. magnetisation [APK18]. magnetization [AMPT22]. magnetohydrodynamic [DMS22, RHZ17]. magnetohydrodynamics [GDMS21]. magnetostatic [CJ24]. Magnus [ALDHHZ^{+}19]. MALDI [KMO^{+}14]. manifold [BZA24a, BZA24b, CMSS23, HQR23, Mai10]. manifold-valued [CMSS23]. manifolds [Dav23, GH16, Hol01, HW19b, LR07, Mai06, Moo17, Noa14, YZ08, ZS21]. Manning [Don20]. many [MNW04]. map [Far10, San15, WL12]. mapping [CWLH95, EGST13, TB16]. mappings [FK10, NRR^{+}22]. maps [BNPP20, FPT06, GMBS23, HMS19, HR17]. marching [Awa15, DD94, DHSW19]. Mariano [CP07a]. market [ALZ02]. Markov [DMMS11, PM19, ZXC12]. Marquardt [AAFV19]. Maruyama [ZL24]. masked [LL21]. Masks [Jia03, CZ19, De 06, Gon93, Han10, LP10, Zhe06]. Mass [LR19, MWZZ23, LLLH23, WBY24]. Mass- [MWZZ23]. massive [CC24]. matched [LZZ09]. Matching [QWS24]. Material [XB05]. materials [APK20, HXC10]. Matérn [BRS13]. Math. [Fus08a, Hem96]. Mathematica [CKP13]. Mathematical [RMCCG^{+}19, SEE96, CLM96a, SX09]. mathematics [BOP^{+}15, CC24, Ano16, Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b]. Matrices [BMSZ01, Chi03, AD08, AH96, BCC22, BGM93, CR99, CG19, Dai16, EGL13, FMP22, FMPS13, GP95, HX21, Kni24, LH13, MPR22, Möß10, Peñ11, SO20, Tom96, ZP24]. Matrix [Low05, AC00, BD93, BCV23, BI11, BR07, CMX07, Cop98, DY10, DES21, ESM13, EGST13, ER21, FKS05, FMP22, FE15, Fus08a, Fus08b, GM17, GK19, GS19a, GKS21, Han06, HHS04, JL97, KM19, KA97, fLLfDX21, MB20, Miy19, NT23, VM17, WLW16, dVR19, DPS93]. Matrix-valued [Low05, Fus08a, Fus08b, GKS21, HHS04]. Matzinger [DKB99]. Max [MP23]. MAXBET [MSW^{+}22]. Maximal [BS05]. maximizing [Bün11]. Maximum [CK11, CHSS03, GWC24, JlLZ20, KL17, ZP19, ZZL^{+}24]. maximum-norm [KL17]. Maxwell [JL19, AB02, CGHK23, FMP22, HKR21, HY18, HW19a, KCW17, LCC19, LHY08, LWX19, VW02, Xie08, YX24a]. Mean [Wan23a, Wan23b, BD21a, BD21b, FK10, RGB13]. Mean-square [Wan23a, Wan23b, BD21a, BD21b]. means [BI11, Xu04a]. Measurable [GH03]. measure [CMP07a]. measurement [FZ12]. measurements [JR23, LL21, Zwi94]. measures [GU02, MSST14, MRS03, MT98]. measuring [DL04]. mechanical [HZHL12, TBG18]. media [BSSM22, Cai02, CGHK23, DGMM16, HW19a, TS05]. medial [CCH^{+}08]. median [Osw04]. median-interpolation [Osw04]. medians [GY99]. melting [PCM^{+}96]. memoriam [Ano93]. memory [BR23, DHSW19, IP15, Jan98, MS22, ZBT^{+}22]. menus [RLGGAV19]. Merrien [Sab04]. mesh [BS05, DHSW19, DM99, Kac18, KF19, LLS96, LL98, MS18, MSS02, MS19, NSS04, SBMS22, Wan19, ZL16]. mesh-free [KF19]. meshes [AN20a, AAQ15, ANN22, CDM^{+}24, CH22a, DSW20, DY01, FPT06, GH06, HSY16, KP98, Kun01, LFL22, Loh23, LL10, LL12b, Mat09, Pla95, PW08, PGB10, QX21, TMH21, WWH21, WQP23, ZWDD16, ZD18, ZL17, ZYC19, ZW23a, ZC14, ZD16, ZW20]. Meshfree [TS05]. meshless [AS05, Dav23, ECS18, FS98, HS08, HS13, KH05, Sch09, Sch15, ZB12]. metamaterial [LCC19]. Metamorphosis [RY16]. method [AAPCC24, AAFV19, AF96, AU20, ARPR01, AC05, An20b, AB11, Ant18, APV14, ACPV21, AL97, ANN22, AC00, ACH10, AHC11, AS05, Bad19, BQRB13, BBB21, BH23, BV24, BZ18, BBdD21, BFMNP24, BK16, Boc04, BK24, BK96, BS97a, CJX17, CEL15, CDS02, sCLC13, CMX97, CLZ02, CLL08, CZS16, CFZ21, CGHK23, CDM^{+}24, CT17a, CKW10, CL02, CN96, CHY11, CCE^{+}12, DX20, DHK98, Dav23, DHSW19, DMS22, DLS14, DHWZ18, DNS21, DL23, DDLL23, EBS22, EX20, ECS18, EHV17, FK98, FGO14, FKT20, FaKT22, FH18, GM17, GK19, Gia20, GG16, Gna07, Gom95, GM24, Gu17, Gu19, Gu20, GS20, GQ22, GH06, yGW07, yGjW10, GW14, GW21, Gut21, GPG15, Han05, HKRS14, HL06, HH16, HR13, HO17, HMQ22, HK21, HAS05, HX17, Hu07, HXC10, HY14, HY18, HWZ23, HS16]. method [HW19a, HL20, HS20, IUV17, Jan98, JM03, JS05, JWJ23, JK08, KNP98, KY01, KS18, KH05, KHM16, Kum23, LN14, LLS03, LC18, LMRV23a, LMRV23b, Li05b, Li08a, Li08b, LC09, LG13, LC13, LZ20a, LGC^{+}20, LS20, fLLfDX21, LFL22, LZL23, LLR24, LZ21, LZY98, LQW18, LLC16, LWL17, LJ18, LC19, LZCW21, LGN22, LBCZ18, MR15, MS18, MR96, MV99, MST05, MSL21, MM20, MP18, MG19, MC23, MY17, MS98, Mur08, NKL^{+}19, OP11, PM17, PPH^{+}21, Pla12, PDB24, QX21, Ryn00, SL15, SXYY06, SH23, SBMS22, SW98, SvVW22, SYCJ18, Svá19, TLG^{+}22, Tau19, TZL16, VNS18, lWyG06, WY16, WY17, WHS18, WWH21, WMW21, Wan22, WLB22, WZ10, WY20, WLW23, Wie15, WLPV15, WZtX16, Xie08, lXyG02, Xu14, XYZ21, YB21, YF23, YS23, YX24a, YL24a, Yao16, YZ21, YZLL24, Yem22, gYyG12, YL24b, YXX22, ZT14, ZP06, ZL24, Zha14]. method [ZPY15, ZT15, ZYC19, ZP19, ZY20, ZZC20, ZJZ21, ZCB21, ZW23a, ZW12, ZL16, ZL21, ZhXpZ16, lZmCRX03, ZQ13, Zhu15, ZG16, dCB07, HORU19]. Methods [Ano01b, AN20a, ASU17, AMS22, ABI15, AK15, Awa15, BP93, BN19, BN21, BCM99, BSK19, BDFS19, BBB^{+}16, BDMR10, BLV20, BEPS96, BJX09, BR23, BD21a, BD21b, BKM^{+}21, BS97b, Bur97, BS00, But97, CDP03, CLM03, CLMR10, CLMR15, CZZ17, CMX02, CWX06, CWX12, CS19, CLCH19, CL16, CEHK18, CJ24, ÇY23, CCP22, CSWP99, CHP19, CH22a, CFC23, Cul96, Cum96, DDP14, DFR07, DPS93, DMMS11, DS01, DRS08, DLZ16, DT12, EOR18, EJ97, FLM96, Fas99b, FYL08, FR10, FKT^{+}19, dLL19, FS98, FZLZ21, GQL24, GMM09, GO95, GWC24, yGqW09, HBMY14, Hau97, HQHV20, HS08, HCYY20, HYHH21, HR02, Hua06, HZHL12, HJYZ23, HZZZ24, IKS10, JMO00, JSSE97, JlLZ20, JLZ08, JY22, Khl18, KXZ04, KLS19, KN14, KNQ16, LW04, LSY24, Lee03, Lee16, LC20, LC21, Lep23, Li05a, LCC19, LW23]. methods [LS19, LK05, LMV11, Liu99, LWX19, LLY12, LL10, LL12b, LNW02, MWZZ23, MH22, MN16, MNC16, Nai12, NSD24, Nie98, Par22, PIZ01, PP23, PS98, Ram95, RWT13, RZ98, Sch15, SWH20, SXZ06, SX09, Sid17, SLT21, SXWG22, gTpM02, TX19, TBD94, TGA96, VC00, Wan12, WL12, WMY13, WMB13, WHTZ18, WG19, WW19, Wan19, Wan23a, Wan23b, Wie15, Wil98, XB05, XZ03, XY12, YKZ19, YXM23, ZVW95, Zen19, ZZP23, ZB12, ZJ23, ZLPX23, ZD16, ZCY20, ZWZ19, vdH93, vdHdS97, vdHv97, vdHS00]. metric [AAFV19, CZS16, LD07]. MFS [FKS05]. MHD [ATS19, CZ20, DHWZ18, LZ20b, LHH23, ZYB18]. Microlocal [GL17]. micromagnetic [BK24]. midpoint [SC08]. Milstein [Wan23a, Wan23b]. Mindlin [AL97, Ye00]. Mini [KL00]. Minimal [Sid17, Xu98, BN21, CsL18, FM18, PGM09, Pra23, Wen95]. minimax [Yao16, Yao19, Zwi94]. Minimization [MR10, BR23, CM96, zCpST09, CZS16, DT12, Fli18, FRT09, GN08, HH16, HHX18a, HHX18b, HWZ23, LWZ19, MHR21, MNPR06, TY18, WFL02, YKZ19]. minimizations [ZYY16]. minimizers [MRS18]. Minimizing [Bün11, Far10, FGS13, FGG^{+}17, KF19]. Minimum [Noa14, ABC14]. Minkowski [FMR00, FH05, FGMS19, KJ09]. MIRKs [Ben97]. miscible [HCH20]. misclassification [CM96]. missing [CCSS09, MHR21, Sch11a]. Mittag [GP13]. Mittag-Leffler [GP13]. Mixed [AAQ15, Li05a, Tim21, lXyG02, ACPV21, AK15, Awa13, BBB21, BBB22, CHM18, DGMM16, DLZ16, DL23, FF95, GS18, GQL24, GAN20, GMM09, Gil17, HCYY20, HZZZ24, JLZ08, Lee16, LLR24, LMT01, MM20, MS16, Per22, SPS18, SY19, TD21, Xie08, YDF97]. mixture [SJJY21]. mixtures [GZW18]. MLPG [AS05]. mobile [LZCW21]. mobile/immobile [LZCW21]. modal [RTV21]. mode [BHM18, CHRX06, HQR23, HW19a, LQW18]. Model [BOP^{+}15, BFH^{+}18, SS15, ALDHHZ^{+}19, AV15, AHK^{+}19, AK19, AHS20, And20, AL97, BNP14, BN19, BN21, BGT20, BGH15, BCS19, CHM18, zCpST09, CZ20, CS18, DHGA23, FE15, GLS15, GZA20, GDB24, GHLU19, GGRBRG21, HB24, HBH24, HF19, HY14, HX02, HL20, JZ10, KU18, KORS17, Kim06, KCW17, Kür18, LAG^{+}24, LGC^{+}20, LSY22, LZCW21, MSXZ13, NP24, Ock96, PWL20, PLRGVR22, RH15, RGB18, SI15, SM18, ’Sw19, Tha19, XKY15, Yan03, Yan19, ZPR03, ZY20, Zhe23, Zhu19]. Modeling [Ano02c, IL16, ABMM19, BZ18, BWX97, BFFM24, DAP13, GGW24, GH18, HQR23, KKL^{+}22, KMMI19, LR19, PJ20, TS05, ZPR03]. Modelling [APK20, APK18, KRP19, SEE96]. models [AGN23, ARAK09, AR23, BGG24, CLM96a, CGH^{+}13, CW09, DM19a, DNS21, EX20, FKT^{+}19, GZW18, MO19, PGB15, RWT13, Wat96, WBGG22, YKZ19, ZW19, ZST19]. modes [BPJ02]. modification [FGS16, ZP19]. Modified [KL00, WMY13, Bai99, dLL19, HW18, Li08b, LMYL19, RJ00, WLW16]. modular [TL18]. modulation [GKKS23]. moduli [NRR^{+}22]. Mollification [BB07]. mollifiers [CGW13]. moment [BR07, Füh16, TQ17]. moment-based [TQ17]. moments [CHR00, CHSS03, CaL06, CCH^{+}15, Han10, ZTHW22]. momentum [WW19]. Monge [Awa15, CFZ21]. mono [Mui99]. mono-implicit [Mui99]. monomials [dB07]. Monotone [Kva14, ABY13, BC19, yGmW98, KL04, KL07, nCV13, Wan12]. monotonic [Xu14]. monotonicity [FP98, FBCR13]. monotonicity-preserving [FBCR13]. Monte [ESTW19, KR19, PM19, Zhe23]. Moreau [Bad19]. Moreau-regularized [Bad19]. MoRePaS [BOP^{+}15]. Morley [HSY16, HMM21, WHTZ18, YLBL12]. morphing [LD07]. mortar [FH18]. mortaring [LRS12]. motion [BNP14, FKT^{+}19, WSL^{+}19]. motions [FGS13]. motivated [GPS21, Wie97]. moving [BMSR^{+}16, FE15, GRdSA19, GJW20, Ock96, TZ03, Tau19]. MR0019410 [CN96]. MR1338896 [Hem96]. MR2438345 [Fus08a]. MR2783300 [Dũn12]. MR3428571 [Ano16]. MR3428572 [Ano16]. MR3428573 [Ano16]. MR3428574 [Ano16]. MR3428575 [Ano16]. MR3428576 [Ano16]. MR3428577 [Ano16]. MR3428578 [Ano16]. MR3428579 [Ano16]. MR3428580 [Ano16]. MRA [PŠWX03]. Multi [HL06, LSXZ15, RZ98, BlH23, BR15, CMR07, CC22, CaL06, CHS17, DHSW19, FZLZ21, GW21, HH16, HD20, HCF^{+}21, HJYZ23, KV21, KCW17, LY07, LZ20a, LMWX13, MSW^{+}22, Onc14, PGB15, Qiu23, gTpM02, ZBT^{+}22, ZTHW22, dVR19]. multi-asset [PGB15]. multi-block [HH16]. multi-dimensional [BR15, FZLZ21, LZ20a, Onc14]. multi-frame [BlH23]. multi-frequency [KCW17]. multi-GPU [ZBT^{+}22]. multi-interval [gTpM02]. multi-knot [LY07]. Multi-level [HL06]. multi-mesh [DHSW19]. multi-order [GW21]. Multi-parameter [RZ98]. multi-patch [KV21]. multi-revolution [CMR07]. multi-space [HD20]. Multi-step [LSXZ15, HJYZ23]. multi-subdomain [CHS17]. multi-task [LMWX13]. multi-term [CC22, Qiu23]. multi-variate [HCF^{+}21, ZTHW22]. multi-view [MSW^{+}22]. multi-wavelet [dVR19]. multi-wavelets [CaL06]. multibody [FE15]. multicategory [CX06]. multidelay [ZV08]. multidelay-integro-differential [ZV08]. multidimensional [BRR18]. Multidomain [BK16]. multifidelity [PM19]. multifilter [Jia00]. multiframelets [LYY13]. Multifrequency [BR22]. Multigrid [BH02, DKMT20, Bad19, BA15, BEPS96, GPG15, Hem95, Hem96, KXZ04, KvL95, SXYY06, ZVW95]. multiharmonic [LZ21]. multilayer [WFL02]. multilayered [Mha93]. Multilevel [BMK15, CWX06, KP95, Kun95, XH13, XYZ21, Zhe23, CHM18, DMMS11, Fas99b, GJS14, GO95, HX02, Nie98, YXX22]. multimedia [Ren09]. Multiobjective [IUV17]. Multiparameter [BJX09]. multiphase [ZGV22]. Multiple [HJH12, CDTV99, GPT17, GIKV21, MXO13, Plo95, TY18, Yao16]. multiple-block [TY18]. multiplication [AC00, BD93, Khl18]. multiplicative [GQL24, QZX23]. multiplier [Hu07, WLW16]. multipliers [JAA22, Kun95, SYCJ18, YF23]. multiply [HL16b, San15]. multiply-connected [HL16b]. multipoint [DLZ16]. multipole [GG16, WLB22, YB21]. multiquadratic [FHN94]. Multiquadrics [BLS02, Buh06, HL16a, HL18]. multiresolution [ADL11, AL13, ABD08, CGRS15, GDMS21, Jia11, KKL03, RAB^{+}09]. Multiscale [CL16, CLR19, Opf06, ACPV21, APK18, APK20, BJX09, GKM^{+}20, GL17, Hua06, JK08, KORS17, LW14, LSW17, Li03, LLY12, MR15, MTH21, SBZR19, TH19, Var96]. multisplitting [WMY13]. Multistep [FJ99, BD21a, BD21b, BS97b, BS97a, BS00, Wil98]. multitasking [BMRS20]. Multivariate [ACH21, BL99, Car95, CsL18, LP04, MS04, SRFH12, ALRÁY22, BC99, BGW24, BCZ05, Cal05, CHR00, CMX07, CDTV99, DS13, FH20, Lai06, Le 00, Li04, NSW16, Sau95, SX95, Sau06, SS96, SW08, SU22, WZ20a, dB00, Sch04]. Multiwavelet [HM03, PS98, LYY13, CMX97]. multiwavelets [CMSS23, MS97]. multiwindow [JJK13]. Mumford [KPY09].

Nash [dCFCF20]. nat [Sil99]. native [Fus08a, Fus08b]. natural [Dou94, SH23]. nature [LO08]. Navier [An20b, Ang00, APV14, CHM18, CR23, CS19, CGS93, DFOS23, DM19a, DNS21, GAN20, GHKL18, HL06, HQR23, JY22, JK08, LC13, LGC^{+}20, LC19, PCM^{+}96, WWH21, ZVW95, ZCL20, ZQS22, dFGAJN18]. Near [BDKY02, DSW05, GIKV21, MM00, Nit22, SFDE15, ST23]. near-field-to-far-field [SFDE15]. Near-optimal [DSW05, GIKV21]. near-singular [Nit22]. nearly [DL23]. necessary [MNPR06]. Nédélec [Xie08]. negative [GS20, HX21]. nematic [SJJY21, aZWL23]. Nernst [HMX22, YTL^{+}20, YL24a]. Nested [ZT14, KP98]. nets [Sto96]. network [CD15, GHJZ23, KKP24, Len96, Mha93, SM18, ’Sw19]. Networks [EPP00, Noa96, BE00, CS18, CLM96b, DM22, EBS22, GS22, Len94, MM00, MZ22, MSL21, MNW99, MMNM23, Nee96, PRS24, Wat96]. Neumann [AHC11, BST24, KP18, LST11, LC21, WL12, ZB12]. Neural [SM18, AM24, BE00, CLM96b, DM22, GHJZ23, KKP24, Len94, Len96, MM00, MZ22, Mha93, MMNM23, Nee96, PRS24, ’Sw19, Wat96]. neutral [Wan19, ZV08]. neutron [YXM23]. Newton [CG15, CCK20, EHV17, HY14, HWZ23, fLLfDX21, YXX22]. Newton-penalty [HY14]. Newtonian [DGMM16, KRP19, Li10b]. NFFT [PT21]. Nicolson [CLCH19, HCYY20, SXWG22, aZWL23]. Nirenberg [GP14]. Nitsche [LRS12]. Nitsche-mortaring [LRS12]. no [Fus08a, Hem96]. node [CDPS22, Gom95]. noise [GQL24, LO08, QZX23]. noisy [BE00, MHR21, NG99, RH12]. Non [Dũn09, EEK21, LW03, YZB20, ZY20, AK01, AS24, ABY13, APK18, Atr12, ANSZ17, BBB21, BSSM22, Bru23, CJ07, CGR11, DMBH24, DLUS17, DY01, FBDH19, Fra99, GJMY19, GS20, HH19a, JKLY13, JM00, Khl18, Lev99, LHY08, LWZ19, LP14, MO16, MG19, PSNP11, PJ20, Ryd19, SO20, TYY11, TMH21, Wan12, WMY13, WMB13, Wan23a, Wan23b, XP10, YZ24, ZBT^{+}22, ZMQ23, ZWDD16, ZL24, ZB12, ZJZ21, ZD16, ZS24]. non-adjacent [Khl18]. non-autonomous [MO16]. non-canonical [AS24]. non-commensurate [Ryd19]. Non-conforming [YZB20, ZMQ23]. non-constant [ZB12]. non-convex [FBDH19]. non-equilibrium [BSSM22, HH19a]. non-extensive [SO20]. non-greedy [TYY11]. non-Hermitian [WMY13, WMB13]. non-homogeneous [AK01, DMBH24, GJMY19, ZS24]. non-identical [XP10]. non-intrusive [PJ20]. Non-iterative [ZY20]. Non-linear [Dũn09, ABY13, CJ07, ZJZ21]. non-Lipschitz [LWZ19, Wan23a, Wan23b, ZL24]. non-local [ZBT^{+}22]. non-monotone [Wan12]. non-negative [GS20]. non-null [BBB21]. non-oscillatory [DLUS17]. non-overlapping [LHY08, PSNP11]. non-periodic [LP14]. non-polynomial [Fra99]. non-rectangular [ZWDD16]. non-self [YZ24]. non-selfadjoint [JM00]. non-stationary [Bru23, CGR11, JKLY13]. non-steady [MG19]. Non-symmetric [EEK21]. Non-Uniform [LW03, Atr12, ANSZ17, DY01, Lev99, TMH21, ZD16]. non-zero [APK18]. nonconforming [LLS03, LY08, LC19, MC08, ZZC20]. nonconjugate [AR23]. nonconservative [Tha19]. nonconvex [SL12]. noncovex [ZW19]. nonequispaced [Nes16, Ste98]. nonhomogeneous [AC05, AMS22, HY18, Li04]. nonintrusive [CEL15]. Nonlinear [AL13, DLU19, Noa98, ON18b, AGN23, AKS23, ADL11, AZW15, BQRB13, BMP^{+}19, BFFM24, CLC16, CLCH19, CGHK23, DKLT93, EH24, GH99, GRdSA19, Gu20, GMS12, GQ22, yGmW98, GW14, Gut21, HCYY20, HL21, JSW20, KU18, KNV01, LRY23, LT98, MWZZ23, MN15, MG19, Osw04, QWXZ10, SM18, SWH20, SPS18, SY19, Sun14, TL18, WZ19, WY20, XZ01, XZ03, YTL^{+}20, YS23, Yan19, YZLL24, ZPR03, ZHzSR21, ZCB21, ZCY20]. Nonlinearity [Ito96, MG19]. Nonlocal [ACK19b, CCP22, AN20a, FLM96, SPS18, SZJ21, TDG16, ACK19a]. nonmatching [PS05]. nonnegative [CZ19, Gon93]. nonpositive [HW19b]. nonregular [LT98]. Nonsmooth [MS22, BV24, ZW19]. nonsquare [fLLfDX21]. Nonstandard [MC10]. Nonstationary [DM15, DMS22, LC20, ON18b, Pit16]. nonsymmetric [XYZ21, dH94]. Norm [CK03a, BW20, Bün11, CK11, KL17, Nie98, PRS24, Tad19, ZPR03]. norm-based [PRS24]. normal [GG16, GS19a, HCF^{+}21, ZTHW22]. Normalized [BF03, JKLY13]. norms [MRS18, Wat94]. note [BI11, CLZ02, Dai16, Del17, Kun09, Lar18, Ste98]. novel [BlH23]. nozzles [HH19a]. Null [KCW17, AKKN22, BBB21]. number [CsL18, LWX19, MGSS22, RJ00, TL18]. number-theoretic [RJ00]. numbers [Sch95]. Numerical [AGN23, AG98, AMS22, ACM^{+}22, ABMM19, BBB22, BGG24, BDFS19, BSSM22, CGH^{+}13, CGHK23, CCH^{+}15, Cum96, DDP14, DL98, ER99, FR10, FQ10, FZ19, GT14, Hau97, HW12, HXC10, HL20, JLW20, JY22, KNV23, Kei95, KRP19, KL22, KMO^{+}14, Mar94, NSD24, OS09, Pen98, PLRGVR22, PSW24, PST08, RCJ^{+}23, RHZ17, She13, SJJL15, TBD94, WH07, XY12, YZLL24, ZVW95, ZGGW12, ZB12, ZD17, ADG17, BPW95, BSK19, BF11, BGMS07, Bes03, Ble07, Bre19, BP11, Cas07, CX04, CJN18, CCDL20, CN96, CT17b, DHK98, DD94, DLP98, EJ97, FM14, FKT^{+}19, FKT20, GVSLN96, HO17, Hol01, HH19a, HJYZ23, KXL24, KKP24, Kle21, KS17, LLS09, LMYL19, LL12a, LJ18, Lut04, Mur08, NRR^{+}22, OP11, PF23, SW04, SJJY21, Tha19, WHS18, WZ19, WG19, WLPV15, XL10, ZWZ19, BMSR^{+}16]. numerically [AD08, CJ24]. NURBS [BX94]. Nyström [CSWP99, DLS14, HBMY14, HH15, LG13, SLT21, Tau19]. Nyström-type [CSWP99].

objective [dCFCF20]. objects

[LD22, SF14, SB22]. oblique [CG10].

oblivious [DES21]. Observability [JRS15]. observable [LSLS14]. Observations [DHK98, GPT17, KKP19, XWL13]. observers [GT14]. obstacle [BBR02, BR22, IKS10, LRY23, LMT01, Yan01]. obstacles [ACM^{+}22, CH15, ZGGW12]. ocean [SI15]. ODE [AB97, Ben97, BKP20]. ODEs [AM24, Bur97, EJ97, JM00, WK93]. offline [Tad19]. offline/online [Tad19]. offset [PGB10]. offsets [CCH^{+}08]. oil [KKL^{+}22]. Oldroyd [AMK19]. Oldroyd-B [AMK19]. One [GKKS23, HMQ22, ZYY16, CLM96b, CLM18b, CLM18a, HT12, KORS17, KKP19, LZZ09, MS12, Wan19, XKY15, ZMQ23, ZG16]. One-bit [GKKS23]. one-dimensional [CLM18b, CLM18a, KKP19, LZZ09, XKY15, ZG16]. one-leg [Wan19]. one-level [HT12]. One-step [HMQ22]. one-way [ZMQ23]. Online [GYZ17, GS19b, HBH24, Tad19, Yin07]. online-adaptive [HBH24]. onset [IL16]. onto [CWLH95]. open [CL18, NW22]. operational [ESM13]. operations [PR15, RLGGAV19]. operator [AR20, BMK15, DFR07, FQ10, GM14, Lut04, MWZZ23, MC10, RH12, RSZ11, XZ03]. operators [ACK19a, ACK19b, BDD06, BC20, BC19, CMP23, CS06, CCS16, DL04, DY10, DES21, FY13, Gna07, GLT93, GP14, GKS21, JL97, KP98, Kir98, KK02, Len96, Li94, LT98, MR10, MXZ08, PL05, TSV21, nCV13, WM21, ZBT^{+}22, ZD16]. operon [CGDHRÁ^{+}19]. Optics [Ano02c, BKN11, BPJ02]. Optimal [AMPT22, BEG17, BMRS20, CS06, CKW10, Dũn11, GL13, HMM21, HS20, KR19, LO08, LWW21, MRS14, Maz04, Mui99, Qiu23, Wan19, ZP19, AM24, BB24, CLC16, CLZ02, CGG^{+}20, DSW05, DP13, EH24, Fou23, GHdN12, GIKV21, GH06, GN13, HPS19, HJJV97, KS20, KP95, LSY24, Lee03, LC18, LZ20b, LZ21, LLLH23, LY01, LGN22, MS22, MRS03, MTH21, NZ04, Per22, RH12, RJ00, RS19, Sch15, SY19, TX19, WZJ18, WZM21, WQP23, Yan03, YZLL24, ZW23b, ZW23c, dLdDSM07, Dũn12]. optimality [CP93, LC18, MM00]. Optimally [GQL24, LP04, ČF11, RST^{+}14]. Optimization [SLN14, AHK^{+}19, BHT11, BLV20, BA15, CD15, ESM13, GHKL18, HMQ22, IUV17, KKL^{+}22, LSXZ15, LL21, PDB24, TBG18]. optimize [RMCCG^{+}19]. Optimized [Onc14, SXWG22, AAPCC24, CCP22]. optimizing [KLS19]. option [lZmCRX03]. orbital [KL19]. orbits [CS93]. Order [But97, FW15, YZ24, AK19, AHS20, AL97, BS05, BOP^{+}15, BGH15, BCS19, BD18, BA15, BFFM24, CDP03, CLM03, CWX12, CZ20, CFZ21, CS18, CEHK18, CLM18b, CLM18a, DM19a, Don20, DNS21, Fin19, FS98, FZLZ21, GL23, GW17, GLS15, GZW18, GM22, GGW24, GH18, GHLU19, GW21, HBMY14, HZZ20, HWZ22, HKR04, HBH24, HS03, HLN12, HZZZ24, ILHH22, IRT23, JMO00, JS21, JY22, JN18, KU18, KMMI19, KCW17, KL16, Kür18, LR19, Li05a, LD16, LCC19, LZ20b, LYF^{+}21, LZ22, LL12a, LWWZ22, LHH23, MRS03, Mat09, MS19, MS04, Möß10, NS04, NKL^{+}19, Noe95, NZ04, OBS15, PM17, PSNP11, Qiu23, Ram18, RHZ17, Ryd19, SI15, SM18, SPS18, SS15, SBZR19, SLK20, SJJY21, SLT21, TX19, TTC21, TBD94, TGA96, WHTZ18, WZ19, WZzS21, WY20, WOBL17, Wie97, WZ20b, XL10, Xu14, ZP06, ZD18, ZPR03]. order [ZT15, ZYB18, ZZC20, ZW22, ZJ23, ZLPX23, ZZL^{+}24, ZD16, ZCY20, ZZ09, Zhu19, Yem22]. Order-preserving [FW15]. ordering [CM99, Dou94]. orders [Bre19, Wan19]. ordinary [yGqW09]. oriented [BWX97]. originated [MH17, Zhe06]. originating [ANZ23]. Orthogonal [Bes03, CMP07a, FM04, LH13, AB23, BFFM24, CI13, Dam07, DHO16, ES23, JS99, Jia00, KESR22, MPS96, MPS97, PF23, QX21, Rei93, Sal98, SI15, SX07, TSY10, Wat94, YF23]. Orthogonality [MPS96, MPS97]. orthogonalization [CK07]. orthogonally [Wat94]. orthomaps [ZB99]. Orthonormal [QWXZ10, CaL06, Han98, Han10, HHS04, Lai06]. oscillators [KXL24]. oscillatory [DLUS17, KXL24, NUWZ17, XHC15]. Oseen [ACPV21, LFL22, LLC16]. osmosis [HB24]. osmotic [LR19]. outflow [Kno09]. output [BHM18]. outputs [DHGA23]. overlapping [HS03, LHY08, Par20, PTC12, PSNP11]. oversampled [MH22].

p [WY17]. package [BD93, ZBT^{+}22]. packet [CR08]. packets [ANZ23]. Padé [BC99, BNPP20, BCZ05, CDTV99, Dar03]. Padé-type [Dar03]. pages [Ano16]. pair [ZT15]. Pairs [CG10, Chr15, NSS04, CKK16, GLS15, Lem09]. pairwise [GYZ17]. panel [HH15]. panel-based [HH15]. parabolic [AGN23, AN20a, AV15, BSK19, BV24, BA15, CHS21, CJW22, FLM96, GLO10, HK21, HCYY20, KL17, LSY24, Le 05, Li05a, LZ21, LB93, LGN22, RS19, SPS18, SvVW22, TH19, gTpM02, VK96, WMW21]. Parallel [AB97, AH96, BP93, BF94, BS97b, BS97a, Bur97, BS00, Chi03, Jan98, Tal19, vdHdS97, ATS19, Ben97, But97, CSWP99, DHSW19, DHWZ18, EJ97, Moo17, RCJ^{+}23, WMY13, WLPV15, XZ01, ZSHZ15, ZQS22, dH94, vdH93]. Parameter [HO17, AHK^{+}19, BCS19, BFMNP24, FaKT22, GPT17, GLM23, HO15, JJLR14, Kno09, LJL20, LLY12, Nes16, OP11, QWS24, RH12, Rip99, RZ98, Sch11b, She13, TBG18, Wie15, WK93]. parameter-dependent [BFMNP24, FaKT22]. Parameter-uniform [HO17, OP11]. parameter-varying [BCS19]. Parameterization [Jia00, BGT20, CLM02, CL18]. Parameterizations [Jia03, HL16b, KV21, LLLH23]. parameterized [ANGV19, BOP^{+}15, CM99, HX21, fLLfDX21, RGB18, SS15, Tad19]. parameters [BRR18, EHV17, FYL08, KLS19, LAG^{+}24, RDEG24]. Parametric [BGH15, KKV15, ALDHHZ^{+}19, BNPP20, CC94, FE15, GPS10, GDB24, HBH24, KK99, LP21, MO19, SFDE15]. parametrized [AHS20, BFFM24, DM19a, MN15, Yan19, ZMQ23]. Pareto [AR23, dCFCF20]. Parseval [BLP19, CK07]. part [BSK19, WZL17, Yao19, BN19, BN21, DMBH24]. partial [ANGV19, Atr12, BV24, CJX17, CN96, EOR18, ESTW19, yGS08, HK21, HS08, LS18, Mai20, PM17, SW96, She13, WK20]. partially [HHX18a, HHX18b]. particle [KH05, KCCV23, LW23, XB05]. particular [AMS22]. partition [CM12, DHWZ18, DQS20, HL05, NSS04, ZSHZ15]. partitioning [Wie15]. partitions [CGKK21, MS07, SS04]. Parzen [ZZ09]. pass [Jia09b]. Past [Ano03c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b]. patch [KV21]. Patches [Kra02, BLP02, GPS10]. path [Tim21]. pattern [RTV21]. pattern-forming [RTV21]. PCA [WW12]. PDE [AV15, AHK^{+}19, BHT11, CW09, DAP13, IUV17, ILHH22, KKL^{+}22, Sch15]. PDE-constrained [KKL^{+}22, BHT11, IUV17]. PDEs [AHS20, AKS23, BFFM24, CCNT21, CH22a, DM22, Le 05, LBCZ18, MN15, MSL21, VK96]. peakiness [GKS19]. pedestrian [PSW24]. penalizations [Per22]. penalized [BL09]. penalty [An20b, GS13, HY14, Lep23, LWX19, Mai20, PWL20, WHTZ18]. pencils [CG06, fLLfDX21]. perceptrons [WFL02]. perfect [CJKK18, EGL13]. perfectly [LZZ09]. Performance [HD20]. period [SC08]. Periodic [CDS02, Zha05, AB02, AB23, CP03, CG10, DS01, FW15, GY00, LL13, LZ21, LP14, MP00, She13, Tor16]. periodization [Søn07]. peristaltic [BLV20]. permittivities [HKR21]. permutation [NSW16]. permutation-invariant [NSW16]. Persistence [CL22, PZ16]. personalized [RLGGAV19]. perturbation [CCNT21, Che06, WHTZ18, ZZC20]. perturbations [She08]. Perturbed [Ram95, RSZ11, AB02, BBR02, CK11, CFZ21, GLO10, HO17, Kun01, LRS12, MR15, MR96, MXO13, MS19, MP18, OS09, OP11, ZC14]. Petrov [AS05, CMX97, CMX02, DS01, Hua06, WY17, WY20]. PH [CHS20, GSS22]. Phase [BH23, BDFS19, BH15, BSSM22, GHKL18, Han10, JR23, KL16, LMYL19, LGC^{+}20, LS20, LL21, LS18, Tha19, TSV21, ZY20]. phase-field [ZY20]. phases [QWXZ10]. phenomenon [BC99, RY13]. photoacoustic [DGK15]. photographs [ZB99]. photonic [KK02]. physical [FKT^{+}19, TS05, ZRR23]. physics [DM22, HL21, MMNM23]. physics-informed [DM22, MMNM23]. pi [Len94, Len96]. Piecewise [AB23, Gu17, Wen95, ABC14, BGM93, Cas07, HL05, JSSE97, KPY09, LY07, LM06, PR99]. piecewise-polynomial [ABC14]. PINNs [DM22]. PinT [LWWZ22]. pipeline [CM99]. pipeline-efficient [CM99]. Pitaevskii [LZ20a]. pivoting [DM23]. planar [BCHS22, BLP02, BGN17, CG19, CHS20, CCLM21, DOZ94, KV21, LLS03, Zwi94]. Planck [HMX22, WSL^{+}19, YTL^{+}20, YL24a]. plane [GMM09, GM22, HBMY14, HPP^{+}19, HY18, Li94, PC17]. plasma [CG17, LAG^{+}24]. plastically [SEE96]. plastically-formed [SEE96]. plate [AL97, FP99, Joh13, MM20, Ye00, Yoo01, ZB99]. plates [ATS19, LQW18]. plus [SS08]. POD [AV15, BFFM24, GH18, GHLU19, RH15]. Poincaré [KP98]. point [AKKN22, CZS16, DSW05, DLUS17, Dou94, Fou23, GMS12, yGmW98, KXL24, Khl18, KS17, KF19, Leo13, LSXZ15, LS19, MRS22, MR96, MHR21, SC08, ZL20]. points [BLP02, Gom95, GH08, IP15, LQW18, NZ22, Pan95, Pra98, SW04, Sto96, Ven94, WHS18, Yao19]. Pointwise [Dam07, GLX23, ZHzSR21, NV10, PP23, WZzS21, WZM21]. poised [BH17]. Poisson [ALZ02, AK15, BL09, CBK01, HMX22, JL19, Li05b, Li08a, PP23, YTL^{+}20, YL24a, Yem22, ZSB01b, ZhXpZ16]. polarized [AMPT22]. poles [VK96]. pollution [Spe23]. Polycircular [NRR^{+}22]. polygonal [ANN22, DSW20, GH06, LFL22, WWH21]. polygons [FHK06, FK10, Jon17, KB15, LD07, RGB13]. polyharmonic [BRS13]. polyhedrons [HR17]. Polynomial [GS00, JZ04, MNPW00, TMH21, WZ20b, Xu00, Xu04b, zCFX07, ABC14, ANZ23, Bar03, BGM93, BNR00, BR13, BFCIV22, CDPS22, CT05, Cho95, DS00, Der04, FGG^{+}17, Fra99, GRV22, GDB24, HMS19, HW12, JKLY13, JLM19, Le 00, Lev99, MV99, Maz99, MWW13, MS00, Mon09, Nie07, Pan95, PS07, PR99, Sau95, SPS21, Sid17, Sor18, Tim21, Wen95, WS01, ZJ06, dB00]. polynomially [LP10]. polynomials [ACM20, Bün11, CMP07a, CI13, CCS16, DZ04, DHO16, DKB99, DGS18, DS13, Flo96b, GMRS97, IP15, Jüt98, LKW17, MPS96, MPS97, MS04, MN00, Pot04, PF23, Rau05, Rei93, Sal98, SX07, SS08, WSM13, Xu04a, dB06]. polytopes [CCH^{+}15, War96]. population [AL21, Tyg23]. poroelasticity [BKM^{+}21]. porosity [HYHH21]. porous [BSSM22, DGMM16]. posed [BL09, BE00, LLY12, XY12]. Positive [LS05, Bai99, De 06, FM21, GP95, GS10, GQ22, GGRBRG21, Luo99, Pin04, SM99, WMY13, WMB13, Wen95, Wu95]. positivity [CEHK18, DSW20, LM06, Maz01]. positivity-preserving [CEHK18, DSW20]. post [HZZ20, LZY98]. post-processed [HZZ20]. post-processing [LZY98]. postbuckling [Mur08]. postconditioning [FJ99]. posterior [And20, AR23]. Posteriori [Ano01b, BBB21, Bra01, CR23, CBK01, CK11, DFOS23, DY01, EBS22, GMBS23, HORU19, KY01, KL17, LC21, LMRV23a, LMRV23b, MSZ06, MRH15, MW01, SWH20, WXWL19, ZC14, Kun01, CLM18b, CLM18a, AK01, ARPR01, DLZ16, GN13, HR01, KNV01, LY01, LLY12, Pet02, PS05, Yan01, Yan03, Zha01, ZSB01b]. potential [BST24, FL23, GM24, Li09, SB22, aKT17]. potentials [CGG^{+}20, EX20, FaKT22, GJW20, IMS99, Li10b, ST23, SJJY21, WG19]. Poussin [SU22]. Powell [AS02, LLS96, LL98, MS07, Rem12, SW08]. power [ASS10, CS18, GKS19, LS98, NSW07, RS99]. power-growth [RS99]. powers [PF23]. Practical [WMB13, CN96, WZL17]. precise [Jon17]. precision [GPS10, NP18, Tim21]. Preconditioned [LNW02, BCM99, MS98, RSZ11]. preconditioner [Bre95, FH18, HS03, HH19b, LK05, PM19, PTC12, PL05, Sto17, Zhu15]. preconditioners [Bai99, DKMT20, MS99, NBL11, SW98, TSZ98]. Preconditioning [GK19, Wan12, vdH93, BHT11, BPK96, CHY11, GPS21, JS95, KP95, KvL95, Kun95, LZ21, MV01, ZT14]. prediction [CC24]. predictive [AV15]. predictor [CSWP99, ZD14]. Preface [AS03, Ano00b, Ano02d, Ano04g, Ano05c, Ano06, BP11, CHY10, Dah95, GOZ03, NW04, Sch04, SX09, Twi96, XY12, BOP^{+}15]. preimage [KNV23]. prescribed [CL02, FMPS13, Kac18, MRS18]. presence [AL13]. Preservation [CK03b, CLMR10, CMP07b, FP98]. preserve [MXZ08]. Preserving [CP04, GZA20, BGW21, CP93, Car95, CJ07, CCDL20, CEHK18, CCS16, CGM01, CM03, CP07b, Der04, DSW20, EOR18, FBCR13, FW15, GO05, GWC24, HBH24, HR17, JlLZ20, JSW20, KNP98, KU18, KBL22, KvD01, LLLH23, PSW24, SS14, WW19, WZzS21, YL24a, ZJ23]. pressure [BSSM22, BR15, Lep23, LGQ24, WaZL24, YZ21]. pressure-correction [LGQ24, WaZL24]. pressure-robust [YZ21]. pricing [lZmCRX03]. primal [DGMM16, KPY09, LY21]. primal-dual [KPY09, LY21]. primal-mixed [DGMM16]. Prime [LMO14, Kot08]. prime-counting [Kot08]. principal [Gat21, IB17, LKT18, SLN14]. principle [AHS20, GWC24, IKS10, JlLZ20, ZZL^{+}24]. principles [Kos15, LSLS14, MSST14, RT14]. printer [ALADH^{+}19]. prior [Fou23]. priori [JS21, WXWL19]. probabilistic [AVGGEG^{+}19, BFMNP24]. probability [ACM20]. problem [AK01, AMK19, Ara20, AGGAV^{+}19, ACH10, AHC11, BST24, BSK19, BBF^{+}24, BGMS07, BMSR^{+}16, BE00, CK11, sCLC13, CGM18, CH15, CL07, CMX07, CZS16, CJW22, CDM^{+}24, CL16, CMM21, CHP19, CG17, FLM96, FF95, FRT09, GMM09, GRdSA19, GPS21, GGRBRG21, yGmW98, GN13, HO17, HKR21, HPS19, HT11, HXC10, HCH20, HFH15, KL00, KS17, KCW17, LAG^{+}24, LST11, LR19, LMRV23a, Lep23, LMRV23b, Li05b, LLZ10, fLLfDX21, LFL22, fLL05, MR15, MC08, MS19, MM20, MP18, Mur08, NT23, O’N18a, OP11, Pei05, QC12, TW98, Wan08, WHTZ18, Wan22, WBY24, WZ10, WLW23, XY15, Yan03, YTL^{+}20, YS23, Yem22, YXX22, ZZP23, ZB12, ZPY15, ZZC20, ZC14, ZW12, vdH93]. problems [AGN23, AN20a, AU20, ARPR01, AC05, ABI15, AKKN22, BCC22, BL09, BKN11, BBB22, BHT11, Bes03, BFMNP24, BDMR10, BA15, BK96, BS97a, BS00, But97, CZZ17, Cas07, sCLC13, Che06, CWX12, CHY10, CP15, CHS17, CHY11, CCE^{+}12, DDP14, DRS08, DSW20, DLZ16, EH24, EHS02, EX20, FK98, FKS05, GM17, Gia20, GM14, Gil17, Gna07, GLX23, GO95, GH06, yGjW10, Hak00, HN18, HNT22, HKRS14, HH16, HHX18a, HHX18b, HH15, Hem95, Hem96, HR01, HQHV20, HMQ22, HO15, HJH12, HW18, HX17, HL21, JS21, JN18, Kac18, KORS17, KXZ04, KL19, KPY09, KF03, KNV01, LSY24, Lee03, LC18, LC20, LC21, Li05a, LSXZ15, LS19, LZ21, LRS12, LMT01, LY01, MSZ06, Mar94, MO16, MR96, Mat09, MC23, MTH21, MNC16, NW22, OS09, Ock96, Par20, Par22, Per22, PP97]. problems [PSNP11, Ram18, RS19, RSZ11, SWH20, SXZ06, SPS18, SH23, Sou22, Sto96, Sun14, TLG^{+}22, TZL16, TDG16, Tom96, TBD94, VW05, WL12, WHS18, WXWL19, Wat94, WY20, XZ01, XY12, XYZ21, Yan01, YX24a, Yao19, ZP24, ZMQ23, Zha14, ZL17, ZYC19, ZCB21, ZL16, ZL20, ZW23b, ZW23c, dCFCF20]. problems-application [SWH20]. procedure [BRZ94, PSNP11, Sch11b, SU22, Tad19, Yan19]. process [ALZ02, DAP13]. processed [HZZ20]. processes [BBF^{+}24, CDP09, KBL22, SU12]. processing [Jia11, LSXZ15, LZY98, PZ16, RT14, SX09, Sun14]. processor [AH96]. Procrustes [Wat94]. product [Bün11, CP15, Flo94, FP98, GO95, KK99, LS05, MJM23, Pin04, RS19]. production [GGRBRG21]. products [FMR00, FGMS19, GY00]. Professor [Sil99]. programming [AHS20]. programs [ZLPX23]. projection [APV14, CLMR15, HT12, KS18, KH05, LH13, LFL22, Mat09, TSV21, ZMQ23]. projections [BLP19]. projectors [She08]. prolate [LST11]. Prony [BH23, JLM19, Sid20, ZP19]. Proof [FP99, Spe23]. propagation [ABMM19]. proper [BFFM24, ES23, KESR22, SI15]. properties [AB11, BS06, Boc23, BD21a, BD21b, CR08, CHL17, CGKK21, CGR11, CL02, DQS20, GGL07, JL97, Mai10, MPS96, MPS97, Mha93, MNW99, PŠWX03, PR19, Pla12, RAB^{+}09, ZP24]. property [BH17, CGP95, CM12, HV19a, JKLY13, MP96, Pré94, SL12]. proportional [XH13]. proportionality [CHMR97]. Prössdorf [Sil99]. prototype [RLGGAV19]. Proximity [MSXZ13, LSXZ15]. Pseudo [Awa15, Mai06, FÁG19, JSW20, LZ20a]. Pseudo-dimension [Mai06]. pseudo-random [FÁG19]. pseudo-spectral [JSW20]. pseudo-spectral-difference [LZ20a]. Pseudoaffinity [Maz07]. pseudodifferential [DPS93, DS01, FQ10, GP14, PS98]. pseudoframes [LO08]. pseudomonotone [HQHV20]. pseudoparabolic [FL23]. pseudospectral [GZW18, yGSlX03, lWyG06]. Publisher [Ano01c]. Puiseux [WZL17]. pumps [BLV20]. pursuit [QWS24]. pyramids [RF19]. Pythagorean [AF96, FS94, FaKS02, FGS16, JKK^{+}12, KJ09, KKV15, PFMS05, Pot95, RSA14, RM19]. Pythagorean-hodograph [AF96, FS94, FaKS02, FGS16, JKK^{+}12, PFMS05, Pot95, RSA14]. Python [CLR19].

QCMC [HL16b]. QMR [Cul96].

QMR/BiCG [Cul96]. QTT [KORS17].

QTT-finite-element [KORS17].

quadrangulations [NSZ04]. Quadratic [CLLS06, MP04, MS07, Noa02, BGW24, CLMR10, DHGA23, Kva14, WQP23, YF23, YZLL24, ZCB21, ZW20]. quadratic-bilinear [BGW24, DHGA23]. Quadratic-Cycloidal [MP04]. quadratically [MN15]. quadratization [LZ22]. Quadrature [Nai12, SB22, ACH21, BF94, FXZ96, FM14, FaKT22, GVSLN96, HZHL12, MC10, SX07, ST23, Tad19, WM21, XHC15, Yan19, aKT17]. quadratures [LL12a]. quadrilateral [AAQ15, FF95, Jia11, KST21, LY08, LPSSP00, LL10, LL12b, MC08, ZYC19]. quadrilaterals [NRR^{+}22]. quadtree [KHM16]. quality [SBMS22]. qualocation [JSSE97]. quantification [GPS21, HSSS18, WBGG22]. Quantile [SHTS14]. quantiles [Xia13]. quantization [LPY10, Wan08]. quantum [HL21, ZJ23]. quartic [HWZ23, MH17]. Quasi [KKL03, KKLY10, NV10, CP99, CSS98, DL98, DGMM16, Dũn09, HL16b, LM04, LC18, MS07, RH12, Rem12, Spe17, WLW23, WL05]. quasi-affine [CSS98]. Quasi-biorthogonal [KKL03]. quasi-conformal [HL16b]. quasi-interpolant [Dũn09]. quasi-interpolants [MS07, Rem12, Spe17]. quasi-interpolating [DL98]. quasi-interpolation [LM04, WL05]. Quasi-interpolatory [KKLY10]. quasi-Newtonian [DGMM16]. quasi-optimal [RH12]. quasi-optimality [LC18]. quasi-reversibility [WLW23]. quasi-wavelet [CP99]. quasigeostrophic [SI15]. quasilinear [ANN22, CH22a, HK21, LB93]. quasiuniform [GH06]. Quaternion [Far10, FGMS19]. queueing [AGGAV^{+}19]. quintic [CHS20, Far10, FGS16, GSS22, SS04]. quintics [KJ09]. quotient [MS04]. quotients [ES23].

R [MRS22]. Rachford [HH16, TSV21]. Radau [BS97b, BS97a, BS00]. Radial [BRSV15, LR07, Li10a, Li10b, BCM99, BGH15, BRR18, CL16, CSW96, DSW05, Fas99a, Fas99b, FF05, FS98, FZ19, FW15, GW17, Hub12, LST11, LW14, LM14, LLY06, Li05a, Li09, LK05, Low05, Mai05, RS99, Rip99, Sch95, SS08, SRFH12, WH07, Wen95, WK20, Wu95, Xu98, dB06]. radial-basis [RS99]. Radiation [MKS02]. radii [AMM20]. radiosity [AC00, Han05]. radius [CCMM21, DY10, Möß10]. Radon [CG15]. raising [FL00]. Ramlau [HR13]. Random [PS23, BKP20, CCNT21, CP15, FÁG19, JWJ23, QZX23]. Randomized [AK19, BN19, BN21, CW19, LA22, LKT18, Boc23, DM23, JWJ23, MB20, SMK22, YL24b, ZZ09]. range [APK20]. ranges [BDYY20]. Rank [NSW16, AAPCC24, AL21, BGT20, CMP23, DDLL23, GIKV21, GS20, GKS21, LMWX13, MRS22, SMK22, Sid17]. Rank- [NSW16]. rank-1 [GIKV21]. Rapid [Lee21, Bre19]. Rate [Li08b, Ara20, Sun08]. rates [ACM20, BC19, CET21, FJ99, HL18, LR07, Wan23a, Wan23b]. Rational [FM18, TSY10, BEG22, BB24, CLM02, DP13, DOZ94, EGL13, FM21, Far10, FGS13, FGG^{+}17, GMB97, JJK13, JW95, Pot95, Sid20, TBG18, VK96, gYyG12]. Raviart [CL07, HCYY20, HMM21, YZB20, ZHY19]. ray [GL13]. RB [SWH20]. RBF [AFM23, Bru23, FM21, GK19, Han12, LSW17, Sch11b]. RBF-interpolation [Sch11b]. RBFs [Fus08a, Fus08b, NSW07]. re [DHSW19]. re-distancing [DHSW19]. reaction [ANN22, CK11, DFOS23, GLO10, KC16, Kun01, MR15, MXO13, SZJ21, ZC14]. Real [LS18, ANSZ17, BMB14, Bre19, Dam07, FRT09, GP13, GS19a, IB17, NUWZ17, Pan95, Pin04, VK96]. real-life [FRT09]. real-time [ANSZ17]. real-valued [IB17]. realizations [RDEG24]. realized [Tal19]. reciprocal [BEG22]. Reconstruction [AL21, DMT03, Var96, ARAK09, Atr12, BDMR10, BRR18, CJKK18, CCH^{+}15, Don20, DM15, EGL13, FPR^{+}12, GP14, JLW20, KL15, SW06, Sun10]. reconstructions [BD18]. recoverable [KL15]. recovered [ZW19]. Recovering [WL12, GN08]. Recovery [FZ12, AAPCC24, BCV23, CCSS09, CS06, CFZ21, CCLM21, DY01, Dũn09, Dũn11, Fou23, Gat21, GL13, JLM19, KKL^{+}22, XWL13, Yan01, Yan03, YTL^{+}20, Zha01, Dũn12]. recovery-based [CFZ21]. rectangles [CDP03]. Rectangular [Awa13, HNT22, DM99, Rem12, ZWDD16]. rectilinear [KNV23]. recurrence [EHV17, MXZ08, RL07]. recursion [Wal95]. recursions [Wal95]. Recursive [Cop98, BR22, TZ03]. Reduced [ASU17, GLS15, MRH15, PGB15, AZW15, ANGV19, BHM18, BV24, BFMNP24, BFFM24, CEL15, DM19a, GGW24, GH18, HK21, IUV17, KMMI19, LAG^{+}24, LR19, PPH^{+}21, PDB24, RH15, SI15, SL15, Sid17, SBZR19, Wie15, Yan19, HORU19]. reduced-basis [IUV17]. reduced-order [GH18, SI15]. Reducing [Ant18]. reduction [AHK^{+}19, AK19, AHS20, AZW15, BN19, BN21, BGT20, BOP^{+}15, BGH15, BFH^{+}18, BCS19, CS18, DHGA23, FE15, GZA20, GDB24, GHLU19, HBH24, HO15, KU18, KCW17, Kür18, NP24, PS07, RGB18, SM18, SS15, Tom96, Yan19, ZN18, Zhu19]. redundancy [BH15]. Refinable [Gro13, HM03, CX04, CTZ02, De 06, GGL07, JS99, KKLY10, LLS95, Li04, MS04, Pit16, Sau06, dVR19]. refined [GJS14, KP95, KP98, ZGV22]. Refinement [DX10, ARPR01, CHP19, DM93, DY10, Han06, JL97, Li04]. refinements [HNN23]. reflection [MP23, Wan12]. reflectivity [DMT03]. regard [TYY11]. regime [ZW22]. region [DLK12]. regions [CMP07b, CCLM21, JW95, KvL95, Pet96, San15, TB16]. register [CM99]. register- [CM99]. regression [SHTS14, SW10, YZ23]. regular [CET21, Led15]. regularities [YX24b]. Regularity [MS97, CCK20, MXO13]. Regularization [EPP00, ’Sw19, Zoz15b, Zoz15a, BJX09, DSS09, FBDH19, FQ10, GS13, HT11, LSY24, MJM23, MZ13, RH12, SHTS14, WLW23, ZW12]. Regularized [SW10, Bad19, DGK15, GYZ17, Hu07, WZ10, ZW19]. Regularizers [MMP13]. regularizing [Pla12]. Reguralizers [Lar18]. Reissner [AL97, Ye00]. related [DGS18, HH19b, MPR22, Sid17, TDG16]. Relation [PZ16]. relations [EHV17, RL07]. relaxation [HJJV97, WLB22, vdHv97]. relaxations [Las16]. Reliable [CCE^{+}12]. ReLU [LP21]. Remarks [HJJV97, DT96, Dou94]. removal [RY13, SEE96]. renewal [AB23]. representation [BDD06, CMP23, CWLH95, CLM02, Lev99, MSST14, MNW04, Nie98]. Representations [Ano02d, CP93, Dar03, Dũn09, GN08, GL17, Hub12, TZ05]. representer [Sch21]. representing [RBGS18]. reprint [CN96]. Reproducing [FY13, BMP^{+}19, DLU19, HT11, JKLY13, PS23, RY16, vdMNS03]. Reproduction [JZ04]. rer [Sil99]. Residual [LC21, APV14, BN21, Mai20, MW01]. residual-based [MW01]. Residual-type [LC21]. residue [GGAVGG^{+}19]. resistance [BGG24]. resistive [GDMS21]. resolution [BlH23, Gat21, HFH^{+}07, ILHH22, JLM19, NT23]. resonance [NW22]. respect [CMP07a, Rei93]. response [TZL16]. restarted [BR23, dH94]. restoration [Del17, FRT09]. Restricted [CMP03, SL12]. Restriction [AD08]. restrictions [BMRS20]. result [JS95]. results [CHS21, GVSLN96, HW18, Kun09, MN00, NSW98]. retarded [Hau97]. RETRACTED [Zoz15b]. Retraction [ZLPX23]. Retraction-based [ZLPX23]. retrieval [BH23, BH15, JR23, LL21, LS18, TSV21, Wan12]. reversibility [WLW23]. reversible [CS93]. revisited [BRZ94]. revolution [Ara20, CMR07]. Reynolds [TL18]. Riccati [AKS23, BF11, KA97]. Richardson [JLZ08]. ridge [GJU22, Mai10, WFL02, GJU22]. Riemann [Bra01, BS23, EGST13, JN18, KS17, LS19]. Riemannian [DDLL23, GH16, HW19b, LL21, Noa06, Noa14, WYW11, YZ08]. Riesz [Grö03, HHS04, Leo13]. right [FR10]. right-hand [FR10]. rightmost [GKS21]. Rigorous [SWH20]. ripplets [Pit16]. risk [MNPR06]. RKDG [ZQ13, ZG16]. RKN [Ram95]. RMS [Nes16]. Robin [HYHH21, PSNP11, SJJL15]. Robin-Robin [HYHH21]. Robin-type [PSNP11]. Robust [AAPCC24, Kun01, LSY24, LZ21, ZC14, AHK^{+}19, BlH23, ÇY23, Don20, YF23, YZ21]. robustness [ZSB01b, ZYY16]. role [Joh13, Ock96, SBMS22]. ROMs [BFFM24]. Rosenbrock [BBB^{+}16]. rotating [SF14]. rotation [FaKS02, Far10, FGS13, FGG^{+}17]. rotation-invariant [FaKS02]. rotation-minimizing [Far10, FGS13, FGG^{+}17]. rotational [CS19, LGQ24, WaZL24]. rounding [Bar03]. Roundoff [TZ03]. RT [SPS18]. Rubenchik [ZWZ19]. rule [Nit22, SC08]. Ruled [SR02]. rules [Han10, IRT23, NSW16, RH12, RJ00, WYW11, WZL17, WM21]. Runge [BP93, Boc23, BS97b, BS97a, BS00, CHMR97, CLM03, CMR07, CLMR10, CLMR15, CSWP99, Hal95, JMO00, Liu99, Mui99, SLT21, vdHdS97, vdHS00].

Sabin [AS02, LLS96, LL98, MS07, Rem12, SW08]. Sacker [LMV11]. saddle [Yao19, ZL20]. safe [HKR04]. salesman [Wan08]. Salzman [Han05]. same [CGR11]. sample [GKS19]. sampler [FPR^{+}12]. samples [CsL18, SW10]. Sampling [BGT20, RZ10, WSM13, vdMNS03, ARAK09, Atr12, BH23, Dũn09, Dũn11, Dũn12, Gat21, GZ13, JJK13, JWJ23, PS23, Pra23, RTV21, Søn07, SZ09, Sun10, Sun14, Tan17, XP10, YL24b, ZZ09]. Sampling-free [BGT20]. sampling-reconstruction [ARAK09]. satellite [FM04]. satisfying [TSY10]. Saul’yev [GWC24]. SAV [JY22, LS20, ZL21]. SBFs [NSW07]. scalability [JAA22]. scalable [MMNM23, WLPV15]. scalar [BR07, HN18, Khl18]. scale [AMM20, BF11, BW20, BRSV15, CS94, HN18, HL21, JWJ23, KESR22, Plo95, RH15, SI15]. Scaled [BLS02, BBdD21, CMP21, NW22]. scales [MXO13]. scaling [CP03, Han98]. Scattered [DZ04, ECS18, FHN94, Grä12, IP15, Kun09, LW04, LNW02, MNW99, NSW98, NG99, Pot04, WL05, WSM13, Yoo01]. scattered-data [NSW98]. scattering [BCV23, BR22, Cai02, CGM18, GL23, Gil17, GM22, HSSS18, HR02, IKS10, LD22, LN14, MKS02, QC12, VW05, WL12, YZB20, ZGGW12]. scheme [BNP14, BCZ23, BBB22, BS05, CLC16, CC22, CJ07, CL07, CLCH19, CtTB21, CLM18b, CLM18a, CGRS15, CT17b, FGS13, HFH^{+}07, HWZ22, HH15, HW18, HCYY20, JSW20, JL19, LZ20b, LYF^{+}21, LY08, LHH23, MB96, PSW24, Qiu23, SJJY21, SXWG22, Tha19, TTC21, WaZL24, Wie97, WQP23, XY20, ZYB18, ZW22, aZWL23]. schemes [ADL11, BD18, Ble07, Boc23, Bru23, CMR07, CD15, CMX97, CZ20, CZ19, CGR11, CGR13, DD94, DLUS17, DLU19, DSW20, DS13, Flo96a, GAN20, GRV22, GZW18, GWC24, GMS12, GGRBRG21, HW19b, JKLY13, Jia95, JN18, LLY06, Lev99, LP10, LQ14, LD16, LMYL19, LSY22, LZ22, LGQ24, MS12, Mui99, MH17, Nie07, Noe95, Ram18, SW96, Sou22, TQ17, TDG16, WZJ18, WZzS21, Zhe06, ZW20]. Schmidt [CK07, Chr15]. Schnirelman [Yao19]. Schoenberg [SS96]. Scholes [PGB15]. Schrödinger [BK16, CLC16, GRdSA19, GM24, JSW20, KU18, LRY23, LZ22, LJ18, MWZZ23, SY19, WZJ18, XY20]. Schur [Dai16, Peñ11]. Schwarz [Bre95, BR07, CCP22, PTC12, SXWG22]. Schwinger [Che02, GM22]. SDEs [CJN18, IY22, Wan23a, Wan23b]. SDFEM [ZL17]. Second [BA15, CZ20, CHS17, HLN12, O’N18a, SJJY21, TW98, TGA96, ZYB18, Bad19, Bra01, CMX97, CP99, DM19b, Don20, FZLZ21, GZW18, GGW24, HS16, JY22, LZ20b, LYF^{+}21, LWWZ22, LHH23, Mar95, PSNP11, Qiu23, RHZ17, TX19, WY20, XL10, Xu14, ZP06, ZCY20]. second- [FZLZ21]. second-/fourth-order [FZLZ21]. Second-kind [O’N18a, TW98]. Second-order [BA15, CZ20, HLN12, SJJY21, TGA96, Don20, GGW24, JY22, LZ20b, LYF^{+}21, LWWZ22, LHH23, Qiu23, TX19, WY20, ZCY20]. sectional [HW19b]. sections [Flo96a]. sectorial [FM04]. segment [MNW96]. segmentation [DLK12, KPY09, RWT13]. segments [Pet96]. Seidel [Dou94]. seismic [Wan12]. selecting [Rip99, Sch11b]. selection [AAPCC24, CHMR97, CDPS22, GLM23, Kac18]. self [FMPS13, MSW^{+}22, SM00, YZ24]. self-adjoint [FMPS13]. self-associated [SM00]. self-consistent-field [MSW^{+}22]. selfadjoint [JM00]. Sell [LMV11]. Semi [Buh06, PIZ01, PDB24, TH24, BS05, Bru23, CLC16, DNS21, JWJ23, JRS15, LC13, PP97, WW19]. Semi-active [PDB24]. semi-cardinal [BS05]. semi-discrete [Bru23, LC13]. semi-discretisations [WW19]. semi-discretization [JRS15]. Semi-explicit [PIZ01]. semi-implicit [CLC16]. Semi-infinite [Buh06]. Semi-Lagrangian [TH24]. semi-linear [DNS21, PP97, WW19]. semi-randomized [JWJ23]. semidefinite [Las16]. semidiscrete [CW16]. semidiscretized [VK96]. semilinear [AV15, Cas07, CK11, FLM96, KL17, LGN22, WMW21, ZZP23]. seminorms [AM14, LL99]. sensing [FBDH19, SL12]. Sensitivity [KA97]. separable [CS03, HHX18a, HHX18b, TY18]. separating [lZmCRX03]. separation [Leo13]. sequences [BD10, BR07, FÁG19, GGL07, Grö03, LY07, LYY13, MP96, SZ09]. sequential [PJ20]. series [CDTV99, LT13, Pré94, QW11]. set [BLP02, BDFS19, Boc23, FGG^{+}17, GLT93, KPY09, Pan95, PS23, RLGGAV19, XKY15]. set-valued [XKY15]. sets [BH17, BS06, CG15, CG19, Cho95, FMR00, FH05, FGMS19, HV19a, HV19b, KF19, Leo13, LL13, LL11, WSHD07, Wie15]. setting [ZHzSR21]. several [GS00, MS12, Xu00]. SfePy [CLR19]. Shah [KPY09]. shallow [CT17b, Don20, IL16, KKP19, TTC21, vdHS00]. Sham [CGH^{+}13]. Shape [BLV20, CP93, CMP07b, CM03, CP04, CP07b, Der04, GO05, CH15, CGM01, DM19a, GHKL18, HSSS18, HSV09, KvD01]. Shape-Preserving [CP04, CM03, GO05]. shaped [Pit16]. shapes [Alb15, BJMR24]. shared [DHSW19, Jan98]. shared-memory [Jan98]. Sharp [CC22]. Shaw [LYF^{+}21, ZL21]. shearlet [GL13, PR19]. shearlet-wavelet [PR19]. shearlets [BLP19, SPS21]. shells [Ara20, YDF97]. shift [CHL17, HHS04, Sun10]. shift-invariant [CHL17, HHS04, Sun10]. shifted [Yoo01]. Shishkin [Mat09, MS19, ZL17]. shooting [CS93]. Shortest [dVR19]. sided [MS18]. sides [FR10]. sideways [LK20]. Siegfried [Sil99]. sifting [DAP13]. Sigma [Wan08, LPY10, Len94, Len96, GKKS23]. sigmoidal [Sto96]. sign [RDEG24]. Signal [MSST14, Gat21, Lee21, PZ16, RT14, Sun14]. signals [BRR18, PS23, Ren09, Var96, Wan12]. signatures [CC94]. Simple [JlLZ20, SO20, CG04, CKBP11, JWJ23, MST05, MY17, Spe23, SB22, ZLPX23]. Simpler [DM23]. simplices [Wie97]. simplified [HY14, RHZ17]. simply [CWLH95, CL18]. simply-connected [CWLH95, CL18]. simulating [KCCV23]. Simulation [Zhu15, BGMS07, BGG24, BDFS19, FE15, FKT^{+}19, FZ19, HPP^{+}19, XKY15, Zhe23]. simulations [APK20, BK24, CEL15, KS18, ZGV22, ZG16]. Simultaneous [CJW22, Füh08, LT06, PCM^{+}96]. sine [LZ20a]. Single [gTpM02, EJ97, NT23]. single-eigenvalue [EJ97]. singly [Tor16]. singular [ACH21, ByLl01, Bes03, Boc04, Che06, CKP13, DX20, FR10, Gu19, HX17, HS16, HX21, IRT23, KXL24, LKT18, LWW21, Nit22, Pla12, Smo07, WHTZ18, WM21, ZZC20]. singularities [Fin19, HZHL12, MRS22, MV99, MP00]. Singularity [KNP98, lZmCRX03]. singularity-separating [lZmCRX03]. singularly [CK11, CFZ21, GLO10, HO17, Kun01, LRS12, MR15, MR96, MXO13, MS19, MP18, OS09, OP11, ZC14]. six [CMP07b]. six-dimensional [CMP07b]. sixth [CLM18b, CLM18a]. sixth-order [CLM18b, CLM18a]. size [BS00, GIKV21, WMW21]. skyrmion [HPP^{+}19]. sleeve [Bei23]. slice [AAPCC24]. slice/feature [AAPCC24]. slipping [ATS19]. slits [KNV23, San15]. Small [Sch23, BGH15, CKK12, NSS04]. SMMP [BD93]. Smolyak [Pet00]. Smooth [AS02, BLP19, BST24, BK96, Car95, GLX23, HBMY14, JSSE97, KV21, KR19, LP14, MM00, Par22, RZ10, YX24a, ZLPX23]. smoothing [BB07, FJ99, Fas99b, MST05]. Smoothness [Osw04, Pra98, CSW14, FBCR13, Li04, MT98, SRFH12, TD21, WYW11, ZD18]. snapshot [ASU17]. snapshots [GHLU19]. Sobolev [CLCH19, DG16, FY13, Han06, KL07, LYY13, PR19, War13]. Sobolev-type [KL07]. Software [WLPV15, CKP13, Mac94, ZBT^{+}22]. solidification [PCM^{+}96]. solidification/melting [PCM^{+}96]. Solution [FH05, LST11, ADG17, BPW95, BF11, BA15, Bra01, Cum96, DPS93, DMMS11, DLP98, EHS02, EJ97, GM14, HH19a, KS17, KC16, Mar94, Pen98, PF23, WZ19, Wat94, ZMQ23, ZVW95, ZYY16]. Solutions [Han06, AC05, AMS22, AB23, Ant18, ATS19, CLZW13, CJN18, CN96, ER99, FK98, HZZ20, HBM03, JLZ08, JN18, KK05, Kim06, KF03, LP21, Li05b, Li08a, Li08b, Li09, Mur08, RL07, Ryn00, She13, SB22, WZ10, XH13, Yao16, ZSB01a, ZV08]. Solvability [HS08]. Solvable [Wat96]. solve [ARPR01]. solver [Ben97, Cai02, GM22, KS23, LD22, RS19, YZ13, ZGV22]. solvers [AB97, BBB^{+}16, HPS19, HH19b, KP95, KCCV23, Sch15, SW96, vdHdS97]. Solving [Fas99b, HS13, JM00, LM14, PC17, Sch09, WLW23, WK20, AC00, BSK19, BBF^{+}24, sCLC13, CJ24, Cul96, DM19b, HQHV20, HJYZ23, Li05b, LSXZ15, LJ18, LBCZ18, LLY12, MP18, MC23, MS00, MMNM23, TLG^{+}22, TQ17, WHS18, XYZ21, Yao19, ZT14, ZCL20, ZhXpZ16]. Some [CHM18, DT96, Dou94, GVSLN96, KY01, MN00, CLM96a, CKP13, Dai16, DS01, HX17, Kle21, Le 00, LT98, LZZ09, LPSSP00, Peñ11, RT14, SW98, TX19, ZQS22, dCFCF20]. source [BBB21, CJW22, JLW20]. Space [GHJZ23, BLP02, BV24, BKK17, CGM01, DKMT20, FS94, Far10, FM18, Fus08a, Fus08b, GG16, GM24, GHLU19, HN18, HD20, HK21, HT11, KCW17, KJ09, LSY24, NV10, Pin04, RS19, RM19, She13, SZJ21, Spe23, SvVW22, TH19, TZ07, Yao16, YZLL24, ZCL20, ZD16]. space-adapted [GHLU19]. space-fractional [DKMT20]. Space-time [GHJZ23, BV24, GM24, HK21, LSY24, RS19, SZJ21]. spaced [LP14]. Spaces [GH03, AS08, BGM93, BMP^{+}19, CLLS06, CMP07b, CS06, CS03, CG10, DST04, DS00, Don19, FY13, GHdN12, GJS14, Han06, HHS04, Jia95, KV21, Kni24, LYY13, LT98, Lut04, MZ22, Maz99, NS21, NSW16, PR19, RY16, SS16, Spe17, Sun08, Sun10, Tad19, Tem01, Tem07, TMH21, WYW11, ZRR23, ZWDD16, ZD18, ZN18, dB07, vdMNS03]. spanned [dB07]. Sparse [BD93, CCH^{+}11, CP15, Fli18, GKS19, Hem95, JLM19, WW12, BNR00, BF11, BZI19, BKP20, Gut21, KKP19, KL15, MR15, NBL11, PWL20, PGB15, Pla00, SU12, SJJL15, YCQZ24, ZTHW22, Hem96]. Sparse-grid [Hem95, Hem96]. sparsification [SO20]. sparsity [BZI19, BDMR10, FRT09, GN08, Lar18, MMP13]. Spatial [CC24, CP04, FaKS02, FZLZ21, FL23, HS20, JLW20, Leo02, PFMS05, WW19]. spatial-dependent [FL23]. spatially [GZW18]. SPDEs [KK20]. special [BOP^{+}15, CHY10, Tom96, XY12]. specified [FM18]. Spectral [GMRS97, Gu19, Gu20, yGSlX03, JL97, MS18, Smo07, ACH10, AHC11, BK16, CJX17, CP14, CCMM21, CP15, CGW13, DY10, Gu17, yGW07, yGjW10, GW21, Gut21, HL06, HX17, HY18, HS16, HW19a, JSW20, Lee21, Lep23, LY07, LZ20a, LS20, LMV11, LWL17, LL07b, LBCZ18, MFB^{+}11, MSL21, Möß10, Mur08, PPH^{+}21, WWX18, WW19, lXyG02, gYyG12, ZP06, ZP24, ZJZ21, ZL21, Zhu15]. spectrum [FMPS13]. sphere [BC94, DST04, DG13, FM14, FM04, HKR04, HR17, HWZ23, HJ23, KN14, KNQ16, LW14, LM14, Leo13, Leo02, Li08a, MNW99, MNPW00, MNW04, NW04, Pot04, PP23, SW04, SS08, SM99, WS01, Xu04a, Xu04b, zCFX07]. Spheres [Noa02, Fas99a, GMBS23, KF19, Kun09, Le 05, Luo99, NSW98]. Spherical [FW96, BMK15, DG16, FM14, HL08, Hes09, HW12, IP15, Le 05, NS04, PTC12, ST23, WWX18]. spheroid [CP15, LST11]. spin [AMPT22, CLM02]. spin-polarized [AMPT22]. Spline [Jia09a, JF02, AS08, BS05, ČF11, CHRX06, CMR22, DS00, DS01, DX10, FGS16, FP99, Flo94, Gér19, GJS14, KV21, KK99, KvD01, LW04, Leo02, MS07, MJM23, MXZ08, MXY13, MH17, PWL20, QX21, Rem12, SS96, Sor18, Spe17, SZ09, TMH21, Wal95, YZLL24, ZWDD16, ZD18]. splines [AF96, AHHR16, ANSZ17, ANZ23, BR13, BL99, CDPS22, CM12, CGR13, CH22b, CKM99, DL98, DZ04, DOZ94, FH20, GSS22, GO05, JKLY13, JL06, Joh13, KLNS95, KL04, KL07, KP04, Kva14, LL98, LS98, LLS95, Leo02, LM06, MR96, MSS02, Maz01, MNW96, NS04, NSS04, NZ04, NSZ04, PTC12, Plo95, PW94, Sch04, SS04, SS14, Str95, TD21, War13, Yoo01, ZB99, Zhe06, ZA10, dVR19]. split [AKKN22, FaKT22, HH15, TLG^{+}22]. splits [AS02]. Splitting [BCZ23, HZHL12, XWL13, HH16, HHX18a, HHX18b, nCV13, ZT14, ZPY15]. splittings [GO95, WMB13]. spots [Kle21]. square [BD21a, BD21b, SW10, Wan23a, Wan23b]. Squares [BNPP20, JF02, CHY11, Dav23, FÁG19, FHN94, LK05, Sto96, Tom96, WK93]. SRLW [HWZ22]. SSOR [Bai99]. stabilisation [Mat09]. stabilities [LMYL19]. Stability [CK03b, CS19, DS10, DMBH24, LS20, NSW98, WMW21, ZV08, ZQS22, ARAK09, AMM20, AV15, BD21a, BD21b, But97, CMP23, DHK98, DP13, ER99, Fus08a, Fus08b, HJYZ23, JS95, Kei95, Kun09, Low05, MN15, MB96, MNPR06, PST08, XB05, YXM23, PS98]. Stabilization [CMR22, AKS23, HT12, JJLR14, KP18, LFL22, NV10, TL18, TZ07, YX24a, dFGAJN18]. Stabilized [Ye00, BQRB13, BBB22, CLM18b, CLM18a, JY22, KLS19, LC09, LC13, MU14, SI15, SXWG22, XY15, ZT15, ZQS22, ZS24]. stabilizer [Kum23, YZ21]. stabilizer-free [Kum23, YZ21]. Stable [BH15, CDPS22, CCLM21, LP04, XL10, BD18, CZ20, GJS14, GZW18, GWC24, GGRBRG21, LYF^{+}21, She13, SJJY21, TGA96, WQP23, Ye04, ZYB18, ZY20, dFGAJN18]. stage [LW23]. Stair [lWyG06]. staircase [Zhu19]. standard [BDD06]. star [GMB97]. starting [CLM03]. State [AKS23, Ang00, CLM18b, CLM18a, DMMS11, ER99, GHKL18, HMX22, HO15, NP24]. State-dependent [AKS23]. Stationary [LLY06, Bre95, Bru23, CGR11, DLU19, JKLY13, LC09, MS12]. statistical [BEG17]. statistics [HSV09]. steady [Ang00, CLM18b, CLM18a, DMMS11, ER99, HMX22, MG19, ZQS22]. steady-state [Ang00, CLM18b, CLM18a, DMMS11, HMX22]. steam [HH19a]. steel [GN13]. Stefan [HXC10, JL19]. Stekloff [YZB20]. Steklov [KP98, LLZ10, YXX22]. step [BS00, CLM03, HMQ22, HFH15, HJYZ23, LSXZ15, MP18, WMW21, ZS24]. step-size [BS00, WMW21]. stepping [JlLZ20, KC16, VK96, WY20, ZPY15, ZY20]. Stepsize [CHMR97, Hal95, Wil98]. stepsizes [LY21]. Stieltjes [Lee21]. stiff [Ben97, BS97a, BS00, But97, EJ97]. Stochastic [ACM20, ÇY23, BCZ23, BNPP20, BD21a, BD21b, CCG10, CJX17, CCK20, CCDL20, DX20, DM19b, GQL24, MM00, QZX23, SU12, TX19, WSM13, WZ20b, YL24b, ZL24, Zhe23]. Stokes [LGC^{+}20, AMS22, An20b, Ang00, APV14, AR20, BBB21, Bre95, BR15, CHM18, CR23, CCG10, CLL08, CS19, CL16, CMM21, CT17a, CGS93, DFOS23, DM19a, DKK22, DNS21, FF95, GAN20, GHKL18, GJMY19, GJW20, HL06, HQR23, JJLR14, JY22, JK08, KL00, KCCV23, Lep23, LC09, LG13, LC13, fLL05, LY08, LC19, MRH15, MY17, Nit22, PCM^{+}96, WY16, WWH21, WaZL24, XY15, YZ21, ZVW95, ZT15, ZCL20, ZQS22, dFGAJN18]. Stokes/Darcy [CHM18]. Stokesian [BLV20]. Straight [San15, DD94]. Strang [GW17, WL05]. strategies [GS22, MN15, WLB22]. strategy [ARPR01, CGR13, GPT17, Hal95, LLY12]. Stray [BK24]. streamers [Ram18]. streamline [KY01]. streams [GSS22]. strength [CJW22]. stress [GG16, Lee16]. Strictly [Luo99, Pin04, SM99, GP95]. strip [KNV23, lXyG02]. strong [DGMM16, GS19b, MGSS22, SJJY21, TX19]. Strongly [GJS14, OS09]. structural [ARPR01, CCN^{+}16]. structural-acoustic [ARPR01]. Structure [BGW21, AB02, GZA20, HBH24, JSW20, KBL22, PSW24, Wat96, YL24a, ZJ23]. Structure-preserving [BGW21, HBH24, JSW20, KBL22, PSW24, YL24a, ZJ23]. Structured [BKN11, BGW24, GGW24, ZL20, BZI19, GDB24, GN08, Kim06, Lar18, MMP13, ZPR03]. structures [JN18, KC16]. Studies [LZZ09]. study [CGDHRÁ^{+}19, CHS20, DM23, EHV17, ILHH22, ’Sw19]. Sturm [GM17, Ram18]. subdiffusion [MS22]. Subdivision [CD15, Gon93, JZ04, Jia95, LP10, ADL11, CJ07, CCMM21, CZ19, CGR11, CGR13, CCS16, CGRS15, DM93, De 06, DLUS17, DLU19, DS13, FBCR13, GRV22, Gér19, GO05, GMS12, HW19b, JKLY13, LLY06, Lev99, MS12, Moo17, MH17, Nie07, Osw04, PK94, Pra98, PR99, RS01, RW21, WYW11, Zhe06, ZA10]. subdomain [CHS17]. subgradient [GH16]. subgrid [ZQS22]. submanifolds [Mai20]. subpopulation [Tyg23]. subregularity [AAFV19]. subsonic [KvL95, ZW22]. subspace [BPK96, GO95, PS23, XWL13, ZVW95]. subspaces [Atr12, BH19, BR23, GDB24, HLT11, LO08, LL11, SZ09]. substructuring [HH19b]. suffer [Spe23]. Sufficient [FS07, CZ19, MNPR06]. sum [CMP07a, Han10, JGW15, TZ05, WFL02]. sum-of-exponentials [JGW15]. summaries [Joh13]. summarizing [CL22]. summation [Pré94]. sums [Tor16, ZP19]. Sumudu [PM17]. sup [dFGAJN18]. Super [LL13, LT12, BlH23, ILHH22, JLM19]. super-resolution [BlH23, ILHH22, JLM19]. Supercloseness [Kum23, ZL17]. superconductivity [GS18]. superconductors [HX02]. Superconvergence [CZZ17, HSY16, LY08, Zha14, ZHY19, Gom95, HZZ20, HMM21, LLZ10, LL12b, MNC16, YL24a, YZ24, ZSB01a]. Superconvergent [YTL^{+}20, YS23]. superlinear [GGRBRG21]. superposition [Nee96]. supersmoothness [FH20]. SUPG [Kno09]. Support [EPP00, CX06, CKK12, NSS04]. supported [CL16, CKK12, CHSS03, Hub12, JJK13, JS99, Lai06, LLS95, Len96, Wen95, Wu95]. Supports [CTZ02, KL15]. suppression [LO08]. Surface [Kra02, ABMM19, BX94, Ble07, Don20, EX20, FM14, FHN94, GN13, HFH^{+}07, Jia11, KBL22, PCM^{+}96, ZL16]. surface/surface [BX94]. Surfaces [JF02, SR02, ALADH^{+}19, BST24, CW16, CC94, CLM02, CL18, DP13, Flo94, GU02, HFH^{+}07, KK99, KP04, Leo02, LPSSP00, MT98, MJM23, O’N18a, PGM09, Pra98, PR99, RS01, ST23, WK20]. surrogate [BZ18, HQR23, KKL^{+}22]. surrogates [EH24]. survey [CBK01, RT14]. suspensions [KCCV23]. SVM [ZJ06]. sweeping [Sto17]. swelling [LR19]. switched [GPAF18]. Sylvester [ZT14]. symbol [GP14, Plo95]. symbols [Nie07]. Symmetric [Han98, Han10, Jia03, Bai99, EEK21, JKLY13, JS95, Lee16, MC23, Nit22, WYW11]. Symmetric/Antisymmetric [Jia03]. symmetries [GH99]. symmetry [AD08, AG98, Awa13, BJMR24, CGKK21, Jia11]. Symplectic [NS21, DDP14, Ram95]. synchrosqueezing [LJL20]. synthesis [ZYY16]. System [Tan17, AMPT22, AK15, AGGAV^{+}19, BBB21, BH23, BR15, DFOS23, DMS22, GQL24, GLO10, Gu17, Gu19, Gu20, HYHH21, KA97, LZ20b, LYF^{+}21, LGQ24, LHH23, MRH15, MXO13, MS00, OS09, SC08, WaZL24, ZW22, ZJ23, ZL21, vdHdS97]. Systems [CMP03, ALADH^{+}19, AMM20, AMS22, AZW15, Bai99, BGT20, BOP^{+}15, BFH^{+}18, BCS19, BDYY20, BW20, BGW21, BGW24, Bes03, BMP^{+}19, BKK17, BDM20, CS93, CCDL20, CEHK18, CTZ02, CR08, CKK16, CHL17, CJ24, CET21, Dad17, DHGA23, DLP98, FE15, GPAF18, GGW24, GDB24, GKM^{+}20, HBH24, Hol01, JWJ23, JL19, Kac18, KKLY10, Kür18, LZZ09, Liu99, Mar94, MV01, MS98, NW22, PDB24, RTV21, RGB18, Ryd19, SXYY06, SW04, Smo07, Tal19, TBG18, Tor16, UG19, WMY13, WMB13, WOBL17, dH94, dCB07]. systolic [AH96]. syzygy [CG19]. Szego [CMP07a].

T [ZD18, TMH21, ZWDD16]. T-meshes [ZD18, TMH21, ZWDD16]. Table [Mac94]. Table-based [Mac94]. tangent [HPP^{+}19]. tangentially [MU14]. task [LMWX13]. Taylor [Del17, fLL05]. Taylorlet [Fin19]. Tchebycheffian [PW94]. technique [BHT11, CS93, FP99, Gil17, KH05, Ren09, SFDE15]. techniques [LZ21, Sor18]. technology [Tal19]. temperature [APK18, LK20, TTC21]. tempered [IY22, LD16, Qiu23]. tension [BRR18, CKM99, MR96, ZL16]. Tensioned [LM04]. Tensor [FP98, GO95, MRS22, MHR21, AAPCC24, CW19, CP15, DDLL23, Flo94, ILHH22, KK99, MB20, MJM23, RS19, SMK22, SSK24, YF23, YL24b]. tensor-based [ILHH22, SMK22]. Tensor-product [FP98, Flo94, KK99, MJM23]. tensorial [ZW23a]. tensors [WBY24]. tenth [TBD94]. tenth- [TBD94]. term [CC22, ILHH22, JLW20, Qiu23, RL07, Tem98]. terms [BBB21, CC94, Lóp18, MG19, Tha19, XZ03]. Ternary [GLM23, CCMM21]. tessellations [GS10]. tesseral [FM04]. test [Tad19]. tests [Mac94]. tetrahedral [FPT06, Kun01, SS04]. Tetration [Pau19]. tetris [CP14]. their [ByLl01, CLLS06, CSS98, yGS08, HLT11, HJYZ23, JGW15, KF19, NSW07, PŠWX03, RSA14, She08, Wat96]. themes [Hof06]. theorem [CKBP11, GGAVGG^{+}19, Sch21]. theorems [GP14]. theoretic [RJ00]. Theoretical [BZ18, GRdSA19, ILHH22]. theoretically [Wie97]. theory [AM14, LL99, LS05, MNPR06, Ock96, ’Sw19, YZ23, dLdDSM07]. there [Sch21]. thermal [LK20, TBD94]. thermistor [YS23]. thermo [BBF^{+}24]. thermo-electromagnetic [BBF^{+}24]. thermoelastic [RCJ^{+}23]. thermoelectrical [BGMS07]. thin [DD94, FP99, Joh13, LQW18, Ryn00, Yoo01, ZB99]. thin-plate [Joh13, Yoo01]. third [BD18, PM17, SM00]. Thomas [HCYY20]. three [BS05, BKM^{+}21, GL23, GDMS21, HHX18a, HHX18b, HSY16, HH19b, JR23, JSW20, LLS96, LL98, LG13, LYF^{+}21, LL07b, MSS02, O’N18a, RL07, WHS18, ZS24]. three-block [HHX18a, HHX18b]. three-dimensional [GDMS21, LG13, LYF^{+}21, WHS18]. three-direction [BS05]. three-field [BKM^{+}21]. three-step [ZS24]. threshold [Lee21]. thresholding [AL13]. Tight [BF03, BS06, CK03a, Jia03, PŠWX03, CCSS09, CHSS03, FWW06, Jia09b, LMO14, PWL20]. Tikhonov [HT11]. tilings [CTZ02, FPT06].

Time [DD94, TB16, VK96, AMK19, ALZ02, AGGAV^{+}19, ANSZ17, Awa15, BlH23, BBF^{+}24, BF11, BDYY20, BV24, BKM^{+}21, CC22, CCP22, DFOS23, FMP22, FZLZ21, GL23, GS18, GQL24, GG06, GM24, GZW18, GHJZ23, GPG15, HQR23, HK21, HY18, HS20, JlLZ20, JLW20, KS23, KC16, Kür18, LSY24, Lem09, LA22, LSY22, LZ21, LZZ09, LHY08, LJ18, LZCW21, LL07b, LWX19, LJL20, MNC16, Onc14, PM17, Rau05, RS19, RST^{+}14, Ryd19, SZJ21, SvVW22, TH19, gTpM02, TGA96, VW02, WW19, WY20, WLW23, Xu14, YX24a, YZLL24, ZPY15, ZY20, ZJZ21, ZCY20, ZW12, dFGAJN18, HL06]. time-concentrated [GG06]. Time-dependent [TB16, BKM^{+}21, DFOS23, GS18, GQL24, HQR23, LZZ09, TGA96, dFGAJN18]. time-domain [GL23, LSY22, VW02]. time-fractional [AMK19, FZLZ21, HS20, JlLZ20, JLW20, LZCW21, PM17, WLW23, YZLL24, ZJZ21]. time-frequency [LA22, Onc14, RST^{+}14]. time-harmonic [CCP22, FMP22, HY18, LHY08, LWX19, YX24a]. time-invariant [BDYY20]. Time-marching [DD94]. time-periodic [LZ21]. time-space [TH19]. Time-stepping [VK96, JlLZ20, ZY20]. time-varying [BBF^{+}24, LJL20]. timer [ALZ02]. TM [LN14]. Tocher [LL98]. Toeplitz [WLW16]. tokamak [LAG^{+}24]. tolerance [CHMR97]. tomographic [GL13]. tomography [HLN12]. tool [Sch15]. Top [DMMS11]. Top-level [DMMS11]. topography [ACM^{+}22, CT17b, Don20, IL16, TTC21]. topological [Cho95, HLN12]. topology [HFH^{+}07, ST23]. Toric [Kra02]. torus [Pot04]. Total [BL09, LM06, BCE^{+}09, DGK15, DT12, FRT09, JZ10, KPY09, Maz01, PZ16, ZPR03]. totally [GP95]. traces [HJH12]. tracking [Tim21]. tractability [KSUW23, NSW16]. traffic [WBGG22]. train [CW19, SSK24].

Training [BE00, WFL02, KKP24]. transaction [ALZ02]. transfer [BGW24, HXC10, PCM^{+}96]. transform [ANSZ17, BMB14, CCH^{+}08, DLP98, Fin19, FW96, Lee21, LL12a, LJL20, MXY13, PM17, RST^{+}14, SL05, SS13, TZ03, YCQZ24, ZS21, SF14]. transformation [SFDE15]. transformations [FWW06, PW24]. transforms [ADL11, DMT03, Dam07, HKR04, KXL24, Kei95, LM14, PST08, Ste98]. Transient [BGMS07, Awa15, BMSR^{+}16, CT17a, LC13, VW05]. transition [HF19]. translates [AM14, CL12, LL99]. Translation [CGKK21, LSLS14, MXZ08, vdMNS03]. transmission [CJW22, DRS08, HKR21, HJH12, WBY24]. transport [BDM20, EBS22, GS22, HPS19, KBL22, LP21, LAG^{+}24, Moo17, NKL^{+}19, PM19]. transport-based [PM19]. transport-hydrodynamic [NKL^{+}19]. transportations [LLLH23]. transverse [HW19a]. trapezoidal [IRT23, Nit22, Pla12, WM21]. trapping [MGSS22]. traps [ZD17]. traveling [Wan08]. travelling [She13]. treatment [Bes03, Hau97, Hol01, Tha19]. Tree [BDKY02]. Trefftz [BDM20, CHP19]. Tresca [DKK22, HKRS14]. triadic [ZA10]. triangle [AS02, yGW07]. Triangular [LLS94, CLZW13, DP13, SL05, WQP23, ZL17, ZW20]. triangulated [NSZ04]. triangulations [FPT06, NS04, Rem12]. Tribute [Sil99]. Trigonometric [RM19, CHS20, DLU19, DS13, KLNS95, LN14, LWY06, Pot04, Rau05, RSA14, SPS21]. trimmed [HFH^{+}07]. triple [WHS18]. triply [Tor16]. trivariate [AS08, LL04, ZD18]. troubled [ZQ13, ZG16]. troubled-cell [ZQ13, ZG16]. truncated [CGW13, DKK22, HCF^{+}21, Wan22, ZL24, ZTHW22]. truncation [BDYY20, GPAF18, Kür18, RDEG24]. tryptophan [CGDHRÁ^{+}19]. tubal [AAPCC24]. tube [KRP19]. Tucker [AU20, CW19]. Tucker-adaptive [AU20]. tumor [PLRGVR22, ZST19]. tuning [Nes16]. turbine [FZ19]. turbulence [HF19]. turbulent [LAG^{+}24, SBZR19]. TV [MSXZ13, RWT13]. TVD [BD18]. twelfth [TBD94]. twelfth-order [TBD94]. twist [FM18]. Two [ASS10, CLM03, CLCH19, CCH^{+}08, CH22a, HN18, HCYY20, HL21, Jia03, Plo95, RF19, ZWZ19, BNP14, BJMR24, BDFS19, BSSM22, Bre95, Bün11, CGDHRÁ^{+}19, Che02, CS94, Dad17, DLS14, FP99, GDMS21, GJW20, GGRBRG21, yGmW98, HKRS14, HT12, HCH20, HFH15, JY22, JN18, KS17, KF03, LSY22, LS19, LD07, LK20, MR15, MS18, MR96, MS19, OS09, OP11, QX21, RH15, SM00, Sun14, Tha19, WZJ18, WG19, WZzS21, XZ01, YZLL24, YZ24, ZZP23, ZHzSR21, ZQ13]. two- [GDMS21]. two-body [HKRS14]. two-cell [Dad17]. Two-dimensional [CCH^{+}08, BJMR24, GJW20, JN18, KF03, LK20, MR15, QX21, SM00, ZHzSR21, ZQ13]. Two-grid [CLCH19, CH22a, HCYY20, BNP14, HCH20, XZ01, ZZP23]. two-level [Bre95, HT12]. two-phase [BDFS19, BSSM22, Tha19]. two-point [yGmW98, KS17, LS19, MR96]. Two-scale [HL21, Plo95, CS94, RH15]. two-sided [MS18]. Two-step [CLM03, HFH15]. type [AN20a, Awa15, BR07, BS97b, BS97a, BS00, CGHK23, Chr15, CKP13, CSWP99, CN96, Cum96, CT17b, Dar03, DHO16, DOZ94, DY01, ECS18, Fra99, GO95, HH19b, JMO00, KXL24, KL07, LLS94, LLS96, LC21, LT12, LL07b, MSZ06, Maz98, PSNP11, RWT13, Sab04, SS04, SS96, TTC21, VNS18, WZL17, Wan23a, Wan23b, Xu14, Yan01, Yan03, YZ13]. type-4 [SS04]. types [CGDHRÁ^{+}19, LL98].

ultra [Jon17]. ultra-precise [Jon17]. unbiased [CC24, DM19b]. unbounded [BC20, CJN18, GZ13, LHY08, QWS24]. uncertain [ÇY23]. Uncertainty [GY00, GS22, GPS21, HSSS18, LSLS14, MSST14, RT14, WBGG22]. Unconditional [CLC16, WZJ18, YL24a, LMYL19]. Unconditionally [LZ20b, LGN22, SY19, YS23, WQP23, ZYB18, ZY20]. unevenly [LP14]. unfitted [YX24a]. Unified [ZW23a, Bar03, Grä12, Lee16, Wal95, ZW20]. Uniform [Che06, CLZW13, DGS18, LW03, Ram18, TZ07, ZW22, ACM20, Atr12, ANSZ17, DG16, DY01, HO17, Lev99, OP11, PRS24, Rem12, TMH21, ZD16]. uniform-grid [DG16]. Uniformity [LW03]. Uniformly [KXZ04, Par22]. unimodular [BD10]. unions [Grö03]. Uniqueness [GP14, JR23, DX20, ZYY16]. unisolvent [Cal05]. unit [CWLH95, FGMS19, GVSLN96, LM14, Leo13, Li08a, Li08b, Xu04a, Xu04b, zCFX07]. unitarily [vdMNS03]. unitary [DLP98, GLT93]. unity [CM12, CGKK21, DHWZ18, HL05, NSS04, ZSHZ15]. univariate [AL21, Fra99, GRV22, MH17]. Universal [CGG^{+}20]. University [New96]. unknown [BR22, Wie15, ZPR03]. unlearning [HL96]. unordered [LS18]. unorganised [BLP02]. unstable [She13]. unsteady [BBB^{+}16, DM19a, FZ19, GJW20]. unstructured [Hag98, Wie97]. unsymmetric [LLS09]. updates [AZW15]. upwind [Ang00, Wie97]. upwind-discretizations [Ang00]. usage [HV19b]. use [RSA14]. Using [Cas07, Lev99, ZRR23, Zoz15b, AAPCC24, AH96, BGH15, BGG24, BD18, BMP^{+}19, BLV20, BR22, BR23, CLR19, CGW13, CH22b, Dou94, EHV17, FF95, FGS16, Fin19, FHN94, FS98, GGAVGG^{+}19, GO05, yGSlX03, HFH^{+}07, IUV17, KL19, Le 05, LW14, LLY06, Li05a, LN19, LZY98, LWL17, MB20, MNW99, MNW04, PTC12, PSW24, PDB24, RLGGAV19, RWT13, Ryd19, Sch09, SS08, WL12, WZL17, YLBL12, Yoo01, ZB99, ZPR03, ZT15, Zhu15]. utilizing [GH18].

validated [HBM03]. validation [HL96]. Vallée [SU22]. value [ABI15, CWX12, CJW22, EH24, FK98, FK10, GM14, yGmW98, HH15, Kac18, KS17, LKT18, LS19, MR96, RGB13, Rip99, Sch11b, WY20, vdH93]. valued [CMSS23, Füh08, Fus08a, Fus08b, GMB97, GKS21, HHS04, IB17, Low05, Sid20, XKY15]. values [CLM03, Fou23, HX21, KSUW23, MC10, Tyg23]. Vandermonde [Maz98, Tom96]. Vanishing [Füh16, BP93, CHR00, CHSS03, CaL06, GT14, Wan19]. VAR [ANGV19]. Variable [BR13, PW24, An20b, BZ18, Bra01, BS00, CT17b, DM19a, FZLZ21, GM24, Kum23, TTC21, WZ19, Wan19, WMW21, Zen19]. variable-fidelity [BZ18]. variable-order [FZLZ21, WZ19]. variables [GS00, MS12, Xu00]. variably [CMP21]. Variants [HH15]. variate [HCF^{+}21, ZTHW22]. variation [BL09, BCE^{+}09, CGP95, DGK15, DT12, FRT09, JZ10, KPY09, LWW21, PZ16, QW11]. variation-penalized [BL09]. Variational [CW16, CJ24, ESTW19, KL19, SZJ21, Bad19, HD20, JK08, KL16, LB93, OBS15, Par20, SBZR19, WOBL17, YKZ19, dLdDSM07]. variations [Hof06]. varying [BBF^{+}24, BCS19, LJL20, Xia13]. Vector [ABD08, EPP00, Sid20, AC00, CX06, Füh08, GO05, GMB97, Han06, Rob98, SM00, Sal98, Sor18, WWX18, dVR19]. vector-valued [Füh08, GMB97, Sid20]. vectorizing [Dou94]. Vectors [HM03]. velocity [CS19, Lep23]. velocity-pressure [Lep23]. VEM [CR23]. Ventcel [SXWG22]. Verified [Miy19]. Version [CHP19, CH22a, Don19, GH06, GW21, Jan98, WY17, WY20]. versions [Sid20]. versus [Cul96]. vertex [HH19b, RJ00]. vertex-modified [RJ00]. vertex-related [HH19b]. via [AHS20, BH23, BRSV15, BJX09, CDPS22, CMP21, CP14, Dam07, FY13, FGO14, Gat21, GLM23, GS10, GL13, HS13, Kos15, LM04, Las16, LM14, MHR21, MS04, Moo17, NP18, PR15, PF23, Ram18, SF14, WWX18, WW19, WFL02, ZA10]. vibration [ARPR01, LMRV23a, LMRV23b, LQW18, MM20]. vibrational [PDB24]. vibrations [Ara20]. view [MSW^{+}22]. Virtual [AN20a, ANN22, BKM^{+}21, CMM21, DL23, LMRV23a, LMRV23b, LFL22, LC19, MM20, MU14, SBMS22, WWH21, ZZC20, aZWL23]. viscoelastic [Kum23, ZPY15]. viscoelasticity [JS21]. viscosity [GT14, ZPY15]. viscosity-splitting [ZPY15]. viscous [SJJY21, WZzS21, ZZL^{+}24]. Vlasov [AK15]. Voigt [BNP14]. volatility [Zhe23]. Volterra [BJX09, DX20, DES21, Gu17, Gu19, Gut21, HS16, Pla12, Qiu23, SW96, SC08, WY17, XH13]. volume [Ano16, CDP03, CLZ02, CWX12, CEHK18, CLM18b, CLM18a, EBS22, FKT^{+}19, GQ22, HZZ20, Hem95, Hem96, HR17, HW18, KH05, KHM16, LN14, LC09, LC13, LGC^{+}20, LZL23, LL10, LL12b, Noe95, PSW24, PP23, Wan22, WQP23, ZST19, Zha14, ZT15, ZYC19, ZCB21, ZW23a, ZW20]. volume-preserving [LLLH23]. Voronoi [DQS20, GS10, PP23]. Voronoï-based [PP23]. Vries [dLL19].

Wachspress [DOZ94, FK10, KB15].

Warping [ZB99]. water [ACM^{+}22, CT17b, Don20, KBL22, KKP19, TTC21, vdHS00]. watermarking [Ren09]. Wave [NKL^{+}19, ADG17, Bru23, CR08, GT14, GMBS23, HY18, HX02, JRS15, Kum23, LWX19, MWZZ23, NV10, She13, SLK20, TZ07, WW19, WLW23, ZP24, ZG16, dCFCF20]. Wave-based [NKL^{+}19]. Waveform [vdHv97, HJJV97]. waveguide [HN18]. waveguides [DMBH24]. Wavelet [DPS93, DKLT93, GLT93, HR02, JL06, KESR22, TSZ98, AU20, ANSZ17, ANZ23, BS06, BDMR10, CP99, CMX02, CP03, CKK16, DMT03, Dũn09, Ehr00, EGL13, FOW14, FW96, Füh16, Gom95, Kei95, KKLY10, LT98, MSL21, PR19, PL05, PST08, RS19, RSZ11, RWT13, SvVW22, XZ03, Zha05, dVR19]. Wavelet-based [TSZ98]. wavelet-Galerkin [Gom95]. wavelet-in-time [SvVW22]. Wavelets [Pla95, CT05, ČF11, CHR00, CS94, CSW96, CaL06, DKB99, FM04, GY00, Han98, Han10, JS99, Jia09a, Jia11, KKL03, Lai06, PŠWX03, PST95, PW24, Rau05, Str13, Urb95, WZM21, ZhXpZ16]. wavenumber [BCC22]. waves [ACM^{+}22, RCJ^{+}23, ZP06]. wavetrain [She13]. way [ZMQ23]. Weak [GN01, IY22, Tem00, BC20, Flo94, Kum23, LLC16, LGN22, Tem02, WY16, WWH21]. weakly [Awa13, Gu19, HS16, Lee16, Par22, Pla12]. WEB [AHHR16, HAS05]. Web-method [HAS05]. WEB-splines [AHHR16]. weight [BMRS20, Xu98]. Weighted [CP14, GQ22, Dam07, ECS18, Kva14, MS16, YKZ19]. well [CT17b, LWWZ22, Tha19, TTC21]. well-balanced [CT17b, Tha19, TTC21]. well-conditioned [LWWZ22]. Wendland [CSW14, HJ23, Sch11a, War13]. WENO [BD18, JN18, LQ14, TQ17]. which [MXZ08]. while [Dou94]. Whitney [SS96]. Whittle [BRS13]. whole [BKK17]. wideband [DMT03, DG13]. widths [DKLT93, UG19]. wildland [ABMM19]. Wilson [ZYC19]. window [CJKK18, PT21, TZ03]. windowed [SS13]. windows [CKK12, ZZ09]. wire [DD94, Ryn00]. wirebasket [FH18]. within [Jon17]. without [KXZ04, Par20]. work [CP07a]. Worst [KS20]. Worst-case [KS20].

X [GL13]. X-ray [GL13]. Xu [WHTZ18].

Young [MRS03]. Yukawa [KNQ16].

Zakharov [ZW22, ZJ23, ZWZ19]. Zero [JW95, APK18, BD10, Flo96b, GG16]. Zerofree [JW95]. Zeta [WM21]. zonal [FM04, MNW99]. Zooming [LSW17].

References

Ahookhosh:2019:LCL

[AAFV19] Masoud Ahookhosh, Francisco J. Aragón Artacho, Ronan M. T. Fleming, and Phan T. Vuong. Local convergence of the Levenberg– Marquardt method under Hölder metric subregularity. Advances in Computational Mathematics, 45(5-6):2771– 2806, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                                      tronic). URL https://link.
                                                      springer.com/article/10.
                                                      1007/s10444-019-09708-7.

Ahmadi-Asl:2024:RLT

[AAPCC24] Salman Ahmadi-Asl, Anh-Huy Phan, Cesar F. Caiafa, and Andrzej Cichocki. Robust low tubal rank tensor recovery using discrete empirical interpolation method with optimized slice/feature selection. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                                      tronic). URL https://link.
                                                      springer.com/article/10.
                                                      1007/s10444-024-10117-8.

Arnold:2015:MFE

[AAQ15] Douglas N. Arnold, Gerard Awanou, and Weifeng Qiu. Mixed finite elements for elasticity on quadrilateral meshes. Advances in Computational Mathematics, 41 (3):553–572, June 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9376-x.

Amodio:1997:POS

[AB97] Pierluigi Amodio and Luigi Brugnano. Parallel ODE solvers based on block BVMs. Advances in Computational Mathematics, 7(1–2):5–26, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018926314954. Par-

allel methods for ODEs.

Ammari:2002:MEP

[AB02] Habib Ammari and Gang Bao. Maxwell’s equations in a perturbed periodic structure. Advances in Computational Mathematics, 16(2– 3):99–112, April 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014402300371. Mod-

eling and computation in optics and electromagnetics.

Anitescu:2011:APG

[AB11] C. Anitescu and U. Banerjee.

Approximation properties of the generalized finite element method. Advances in Computational Mathematics, 34 (4):369–390, May 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9159-y.

Ando:2023:POC

[AB23] Alessia Andò and Dimitri Breda. Piecewise orthogonal collocation for computing periodic solutions of renewal equations. Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10094-4.

Antonelli:2014:GFC

[ABC14] Michele Antonelli, Carolina Vittoria Beccari, and Giulio Casciola. A general framework for the construction of piecewisepolynomial local interpolants of minimum degree. Advances in Computational Mathematics, 40(4):945–976, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9335-y.

Arandiga:2008:VCA

[ABD08] Francesc Aràndiga, Antonio Baeza, and Rosa Donat. Vector cell-average multiresolution based on Hermite interpolation. Advances in Computational Mathematics, 28 (1):1–22, January 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9007-7.

Amodio:2015:ECM

[ABI15] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. Energy-conserving methods for Hamiltonian boundary value problems and applications in astrodynamics. Advances in Computational Mathematics, 41(4): 881–905, August 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9390-z.

Ambroz:2019:NMW

[ABMM19] Martin Ambroz, Martin Balažovjech, Matej Medl’a, and Karol Mikula. Numerical modeling of wildland surface fire propagation by evolving surface curves. Advances in Computational Mathematics, 45(2): 1067–1103, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9650-4.

Arandiga:2013:NCN

[ABY13] Francesc Aràndiga, Antonio Baeza, and Dionisio F. Yáñez. A new class of non-linear monotone Hermite interpolants. Advances in Computational Mathematics, 39(2):289–309, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-012-9280-1.

Atkinson:2000:FMV

[AC00] Kendall Atkinson and David Da-Kwun Chien. A fast matrix–vector multiplication method for solving the radiosity equation. Advances in Computational Mathematics, 12(2–3):151–174, February 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018961001209.

Alves:2005:NMF

[AC05] C. J. S. Alves and C. S. Chen. A new method of fundamental solutions applied to nonhomogeneous elliptic problems. Advances in Computational Mathematics, 23(1– 2):125–142, July 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1833-5.

Atkinson:2010:SME

[ACH10] Kendall Atkinson, David Chien, and Olaf Hansen. A spectral method for elliptic equations: the Dirichlet problem. Advances in Computational Mathematics, 33(2): 169–189, August 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9125-8.

Atkinson:2021:MQS

[ACH21] Kendall Atkinson, David Chien, and Olaf Hansen. Multivariate quadrature of a singular integrand. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09869-4.

Aksoylu:2019:CNO

[ACK19a] Burak Aksoylu, Fatih Celiker, and Orsan Kilicer. Correction to: Nonlocal operators with local boundary conditions in higher dimensions. Advances in Computational Mathematics, 45(1): 493, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9632-6.

See [ACK19b].

Aksoylu:2019:NOL

[ACK19b] Burak Aksoylu, Fatih Celiker, and Orsan Kilicer. Nonlocal operators with local boundary conditions in higher dimensions. Advances in Computational Mathematics, 45(1):453–492, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9624-6.

See correction [ACK19a].

Adell:2020:SBP

[ACM20] José A. Adell and Daniel Cárdenas-Morales. Stochastic Bernstein polynomials: uniform convergence in probability with rates. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09770-6.

Ambrose:2022:NAW

[ACM^{+}22] David M. Ambrose, Roberto Camassa, Jeremy L. Marzuola, Richard M. McLaughlin, Quentin Robinson, and Jon Wilkening. Numerical algorithms for water waves with background flow over obstacles and topography. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09957-z.

Araya:2021:AMH

[ACPV21] Rodolfo Araya, Cristian Cárcamo, Abner H. Poza, and Frédéric Valentin. An adaptive multiscale hybridmixed method for the Oseen equations. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09833-8.

Aimi:2008:RMN

[AD08] A. Aimi and M. Diligenti.

Restriction matrices for numerically exploiting symmetry. Advances in Computational Mathematics, 28(3): 201–235, April 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9019-y.

Aimi:2017:EBF

[ADG17] A. Aimi, M. Diligenti, and C. Guardasoni. Energetic BEM–FEM coupling for the numerical solution of the damped wave equation. Advances in Computational Mathematics, 43(3): 627–651, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9500-1.

Amat:2011:ACN

[ADL11] S. Amat, K. Dadourian, and J. Liandrat. Analysis of a class of nonlinear subdivision schemes and associated multiresolution transforms. Advances in Computational Mathematics, 34(3): 253–277, April 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9151-6.

Albrecht:1996:CPH

[AF96] Gudrun Albrecht and Rida T. Farouki. Construction of C^{2} Pythagorean-hodograph interpolating splines by the homotopy method. Advances in Computational Mathematics, 5(1):417–442, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124754.

Angleitner:2023:MHI

[AFM23] Niklas Angleitner, Markus Faustmann, and Jens Markus Melenk. H-inverses for RBF interpolation. Advances in Computational Mathematics, 49(6):??, December 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10069-5.

Allgower:1998:NES

[AG98] Eugene L. Allgower and Kurt Georg. Numerical exploitation of symmetry in integral equations. Advances in Computational Mathematics, 9(1–2):1–20, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018929204901. Nu-

merical treatment of boundary integral equations.

Atencia:2019:ADP

[AGGAV^{+}19] Ivan Atencia, José L. Galán-García, Gabriel Aguilera-Venegas, Pedro Rodríguez-Cielos, and M. Ángeles Galán-García. An arriving decision problem in a discrete-time queueing system. Advances in Computational Mathematics, 45(4): 1863–1879, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09663-3.

Acevedo:2023:NAN

[AGN23] Ramiro Acevedo, Christian Gómez, and Paulo Navia. Numerical analysis of nonlinear degenerate parabolic problems with application to eddy current models. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10052-0.

Anderson:1996:PFB

[AH96] I. J. Anderson and S. K. Harbour. Parallel factorization of banded linear matrices using a systolic array processor. Advances in Computational Mathematics, 5(1): 1–14, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124732.

Atkinson:2011:SME

[AHC11] Kendall Atkinson, Olaf Hansen, and David Chien. A spectral method for elliptic equations: the Neumann problem. Advances in Computational Mathematics, 34(3): 295–317, April 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9154-3.

Apprich:2016:CWS

[AHHR16] Christian Apprich, Klaus Höllig, Jörg Hörner, and Ulrich Reif. Collocation with WEB-splines. Advances in Computational Mathematics, 42(4):823–842, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9444-x.

Alla:2019:CMR

[AHK^{+}19] Alessandro Alla, Michael Hinze, Philip Kolvenbach, Oliver Lass, and Stefan Ulbrich. A certified model reduction approach for robust parameter optimization with PDE constraints. Advances in Computational Mathematics, 45(3):1221– 1250, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9653-1.

Alla:2020:FCP

[AHS20] Alessandro Alla, Bernard Haasdonk, and Andreas Schmidt. Feedback control of parametrized PDEs via model order reduction and dynamic programming principle. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09744-8.

Ainsworth:2001:PEE

[AK01] Mark Ainsworth and Donald W. Kelly. A posteriori error estimators and adaptivity for finite element approximation of the nonhomogeneous Dirichlet problem. Advances in Computational Mathematics, 15(1–4): 3–23, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014240508621.

Asadzadeh:2015:CAB

[AK15] Mohammad Asadzadeh and Piotr Kowalczyk. Convergence analysis for backward-Euler and mixed discontinuous Galerkin methods for the Vlasov–Poisson system. Advances in Computational Mathematics, 41(4): 833–852, August 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9388-6.

Alla:2019:RMO

[AK19] Alessandro Alla and J. Nathan Kutz. Randomized model order reduction. Advances in Computational Mathematics, 45(3):1251– 1271, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10444-018-09655-9.

Arfat:2022:IEA

[AKKN22] Yasir Arfat, Poom Kumam, Muhammad Aqeel Ahmad Khan, and Parinya Sa Ngiamsunthorn. An inertial extragradient algorithm for equilibrium and generalized split null point problems. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09920-4.

Alla:2023:SDR

[AKS23] Alessandro Alla, Dante Kalise, and Valeria Simoncini. State-dependent Riccati equation feedback stabilization for nonlinear PDEs. Advances in Computational Mathematics, 49 (1):??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09998-4.

afKlinteberg:2017:EEQ

[aKT17] Ludvig af Klinteberg and Anna-Karin Tornberg. Error estimation for quadrature by expansion in layer potential evaluation. Advances in Computational Mathematics, 43(1):195–234, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9484-x;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        016-9484-x.pdf.

Arnold:1997:IEL

[AL97] Douglas N. Arnold and Xiaobo Liu. Interior estimates for a low order finite element method for the Reissner–Mindlin plate model. Advances in Computational Mathematics, 7(3): 337–360, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018907205385.

Amat:2013:NTM

[AL13] Sergio Amat and Jacques Liandrat. Nonlinear thresholding of multiresolution decompositions adapted to the presence of discontinuities. Advances in Computational Mathematics, 38(1):133–146, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9231-2.

Ahrens:2021:RLR

[AL21] Robin Ahrens and Sabine Le Borne. Reconstruction of low-rank aggregation kernels in univariate population balance equations. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09871-w.

Aguirre-Lopez:2019:CIS

[ALADH^{+}19] Mario A. Aguirre-López, F-Javier Almaguer, O. Díaz-Hernández, Gerardo J. Escalera Santos, and Javier Morales-Castillo. On continuous inkjet systems: a printer driver for expiry date labels on cylindrical surfaces. Advances in Computational Mathematics, 45(4): 2019–2028, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09682-0.

Alben:2015:BBG

[Alb15] Silas Alben. Bending of bilayers with general initial shapes. Advances in Computational Mathematics, 41(1): 1–22, February 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9347-2.

Aguirre-Lopez:2019:CPM

[ALDHHZ^{+}19] Mario A. Aguirre-López, O. Díaz-Hernández, Filiberto Hueyotl-Zahuantitla, Javier Morales-Castillo, F.- Javier Almaguer, and Gerardo J. Escalera Santos. A cardioid-parametric model for the Magnus effect in baseballs. Advances in Computational Mathematics, 45(4): 2097–2109, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09694-w.

Amat:2022:EMA

[ALRÁY22] Sergio Amat, David Levin, Juan Ruiz-Álvarez, and Dionisio F. Yáñez. Explicit multivariate approximations from cell-average data. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09997-5.

Albeverio:2002:CTF

[ALZ02] Sergio Albeverio, Lan-Jun Lao, and Xue-Lei Zhao. A continuous time financial market with a Poisson process as transaction timer. Advances in Computational Mathematics, 16(4): 305–330, May 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014566730969.

Alzer:2010:EFI

[Alz10] Horst Alzer. Error function inequalities. Advances in Computational Mathematics, 33(3):349–379, October 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9139-2.

Arteaga:2014:DFS

[AM14] Cristian Arteaga and Isabel Marrero. Direct form seminorms arising in the theory of interpolation by Hankel translates of a basis function. Advances in Computational Mathematics, 40(1): 167–183, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9303-6.

Aghili:2024:OCF

[AM24] Joubine Aghili and Olga Mula. An optimal control framework for adaptive neural ODEs. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10149-0.

Al-Maskari:2019:GFT

[AMK19] Mariam Al-Maskari and Samir Karaa. Galerkin FEM for a time-fractional Oldroyd-B fluid problem. Advances in Computational Mathematics, 45(2):1005– 1029, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9649-x.

Aliyev:2020:ASR

[AMM20] Nicat Aliyev, Volker Mehrmann, and Emre Mengi. Approximation of stability radii for large-scale dissipative Hamiltonian systems. Advances in Computational Mathematics, 46(1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09763-5.

An:2022:OCC

[AMPT22] Xin An, Ananta K. Majee, Andreas Prohl, and Thanh Tran. Optimal control for a coupled spin-polarized current and magnetization system. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09947-1.

Alves:2022:NMP

[AMS22] Carlos J. S. Alves, Nuno F. M. Martins, and Ana L.

Silvestre. Numerical methods with particular solutions for nonhomogeneous Stokes and Brinkman systems. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09937-3.

Adak:2020:VEM

[AN20a] D. Adak and S. Natarajan. Virtual element methods for nonlocal parabolic problems on general type of meshes. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09811-0.

An:2020:IPM

[An20b] Rong An. Iteration penalty method for the incompressible Navier–Stokes equations with variable density based on the artificial compressible method. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09757-3.

Andrade:2020:EPC

[And20] J. A. A. Andrade. Exact posterior computation for the binomial-Kumaraswamy model. Advances in Computational Mathematics, 46 (6):??, December 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09821-y.

Angermann:2000:EAU

[Ang00] Lutz Angermann. Error analysis of upwind-discretizations for the steady-state incompressible Navier–Stokes equations. Advances in Computational Mathematics, 13(2): 167–198, May 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018902210295.

Aretz-Nellesen:2019:VPP

[ANGV19] Nicole Aretz-Nellesen, Martin A. Grepl, and Karen Veroy. 3D-VAR for parameterized partial differential equations: a certified reduced basis approach. Advances in Computational Mathematics, 45(5-6):2369– 2400, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09713-w.

Arrutselvi:2022:VEM

[ANN22] M. Arrutselvi, E. Natarajan, and S. Natarajan. Virtual element method for the quasilinear convection– diffusion–reaction equation on polygonal meshes. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09990-y.

Anonymous:1993:MJG

[Ano93] Anonymous. In memoriam John Gregory. Advances in Computational Mathematics, 1(1):iii, February 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02070818.

Anonymous:1995:GE

[Ano95] Anonymous. The Guest Editor. Advances in Computational Mathematics, 4(1– 2):i, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123469.

Anonymous:2000:AI

[Ano00a] Anonymous. Author index.

Advances in Computational Mathematics, 13(4):417–418, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1023/A:1016615818384.

Anonymous:2000:P

[Ano00b] Anonymous. Preface. Advances in Computational Mathematics, 12(4):i–ii, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1018925420773.

Anonymous:2001:AI

[Ano01a]  Anonymous.  Author in-

dex. Advances in Computational Mathematics, 14(4): 393–394, May 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016819732496.

Anonymous:2001:PEE

[Ano01b] Anonymous. A posteriori error estimation and adaptive computational methods. Advances in Computational Mathematics, 15(1–4): 1–2, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016858615238.

Anonymous:2001:PA

[Ano01c] Anonymous. Publisher’s announcement. Advances in Computational Mathematics, 14(1):1–6, January 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016633820248.

Anonymous:2002:AIa

[Ano02a]  Anonymous.  Author in-

dex. Advances in Computational Mathematics, 16(4): 389–390, May 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016822023144.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index.

Advances in Computational Mathematics, 17(4):397–398, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016232931523.

Anonymous:2002:MCO

[Ano02c]  Anonymous.  Modeling

and computation in optics and electromagnetics. Advances in Computational Mathematics, 16(2–3):97– 98, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016809007214.

Anonymous:2002:PAG

[Ano02d] Anonymous. Preface; advances in geometrical algorithms and representations. Advances in Computational Mathematics, 17 (1–2):1–3, July 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016806617055.

Anonymous:2003:AIa

[Ano03a] Anonymous. Author index. Advances in Computational Mathematics, 18(2–4):431– 432, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021307330201.

Anonymous:2003:AIb

[Ano03b] Anonymous. Author index. Advances in Computational Mathematics, 19(4):413–414, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1024388324172.

Anonymous:2003:PEA

[Ano03c] Anonymous. Past Editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 19(1–3):337, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022848030422.

Anonymous:2004:AIa

[Ano04a]  Anonymous.  Author in-

dex. Advances in Computational Mathematics, 20(4): 427–428, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1027303623367.

Anonymous:2004:AIb

[Ano04b] Anonymous. Author index.

Advances in Computational Mathematics, 21(3–4):449– 450, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032143.
19990.2b.

Anonymous:2004:E

[Ano04c] Anonymous. Errata. Advances in Computational Mathematics, 21(3–4):445, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032141.
51652.c6;   http://link.
springer.com/content/pdf/
10.1023/B:ACOM.0000032141.
51652.c6.pdf.

Anonymous:2004:PEAa

[Ano04d] Anonymous. Past Editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 20(4):425, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1027315116562.

Anonymous:2004:PEAb

[Ano04e] Anonymous. Past Editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 21(1–2):219, July 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016693.
57844.49.

Anonymous:2004:PEAc

[Ano04f] Anonymous. Past Editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 21(3–4):447, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032142.
27408.87.

Anonymous:2004:P

[Ano04g] Anonymous. Preface. Advances in Computational Mathematics, 20(1–3):1–3, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025814926647.

Anonymous:2005:PEAa

[Ano05a] Anonymous. Past editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 22(1):101, January 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-5323-1.

Anonymous:2005:PEAb

[Ano05b] Anonymous. Past editors of Advances in Computational Mathematics. Advances in Computational Mathematics, 22(2):199, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-6475-
0.

Anonymous:2005:P

[Ano05c] Anonymous. Preface. Advances in Computational Mathematics, 23(1–2):1–3, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-004-2232-
7.

Anonymous:2006:P

[Ano06] Anonymous. Preface. Advances in Computational Mathematics, 24(1–4):1–3, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7205-3.

Anonymous:2008:E

[Ano08] Anonymous. Errata. Advances in Computational Mathematics, 29(1):91–100, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-008-9090-
7.

Anonymous:2016:EAC

[Ano16] Anonymous. Erratum to: Advances in Computational Mathematics, volume 41 December 2015, issue 6, pages 1391–1641 [MR3428571; MR3428572; MR3428573; MR3428574; MR3428575; MR3428576; MR3428577; MR3428578; MR3428579; MR3428580]. Advances in Computational Mathematics, 42(4):1013– 1014, August 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9464-1;
 http://link.springer.com/
content/pdf/10.1007/s10444-
016-9464-1.pdf. Due to an

oversight during issue building, the page numbers of issue 6 overlaps issue 5. Issue 5 start and end page is 955– 1389, while issue 6 start and end page is 955–1205. Corrected page numbers are supplied here.

Averbuch:2017:LCS

[ANSZ17] Amir Averbuch, Pekka Neittaanmäki, Etay Shefi, and Valery Zheludev. Local cubic splines on non-uniform grids and real-time computation of wavelet transform. Advances in Computational Mathematics, 43(4): 733–758, August 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9504-x.

Antunes:2018:RIC

[Ant18] Pedro R. S. Antunes. Reducing the ill conditioning in the method of fundamental solutions. Advances in Computational Mathematics, 44(1): 351–365, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9548-6.

Averbuch:2023:DWP

[ANZ23] Amir Averbuch, Pekka Neittaanmäki, and Valery Zheludev. Directional wavelet packets originating from polynomial splines. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10024-4.

Arjmand:2018:ACM

[APK18] Doghonay Arjmand, Mikhail Poluektov, and Gunilla Kreiss. Atomistic-continuum multiscale modelling of magnetisation dynamics at nonzero temperature. Advances in Computational Mathematics, 44(4):1119– 1151, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9575-3.

Arjmand:2020:MLR

[APK20] Doghonay Arjmand, Mikhail Poluektov, and Gunilla Kreiss. Modelling long-range interactions in multiscale simulations of ferromagnetic materials. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09747-5.

Araya:2014:ARL

[APV14] Rodolfo Araya, Abner H. Poza, and Frédéric Valentin. An adaptive residual local projection finite element method for the Navier– Stokes equations. Advances in Computational Mathematics, 40(5–6):1093–1119, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9343-6.

Askham:2020:BIE

[AR20] Travis Askham and Manas Rachh. A boundary integral equation approach to computing eigenvalues of the Stokes operator. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09774-2.

Andrade:2023:EPD

[AR23] J. A. A. Andrade and P. N.

Rathie. Exact posterior distribution for nonconjugate Pareto models. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10030-6.

Arabyan:2020:RCA

[Ara20] Mariam Arabyan. On the rate of convergence for approximation of an eigenvalue problem describing vibrations of axisymmetric revolution elastic shells. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-020-09775-1.

Acosta-Reyes:2009:SSR

[ARAK09] Ernesto Acosta-Reyes, Akram Aldroubi, and Ilya Krishtal. On stability of samplingreconstruction models. Advances in Computational Mathematics, 31(1–3):5–34, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9083-6.

Alonso:2001:PEE

[ARPR01] Ana Alonso, Anahí Dello Russo, Claudio Padra, and Rodolfo Rodríguez. A posteriori error estimates and a local refinement strategy for a finite element method to solve structuralacoustic vibration problems. Advances in Computational Mathematics, 15(1–4):25–59, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014243118190. A

posteriori error estimation and adaptive computational methods.

Alfeld:2002:SME

[AS02] Peter Alfeld and Larry L. Schumaker. Smooth macroelements based on Powell– Sabin triangle splits. Advances in Computational Mathematics, 16(1):29–46, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014299228104.

Aldroubi:2003:P

[AS03] Akram Aldroubi and Qiyu Sun. Preface. Advances in Computational Mathematics, 18(2–4):91–92, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021340203871.

Atluri:2005:BMD

[AS05] Satya N. Atluri and Shengping Shen. The basis of meshless domain discretization: the meshless local Petrov– Galerkin (MLPG) method. Advances in Computational Mathematics, 23(1–2):73– 93, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1813-9.

Alfeld:2008:BDT

[AS08] Peter Alfeld and Larry L.

Schumaker. Bounds on the dimensions of trivariate spline spaces. Advances in Computational Mathematics, 29(4):315–335, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9051-6.

Adcock:2021:FAB

[AS21] Ben Adcock and Mohsen Seifi. Frame approximation with bounded coefficients. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09820-z.

Arambasic:2024:DFC

[AS24] Ljiljana Arambasić and Diana Stoeva. Dual frames compensating for erasures-a noncanonical case. Advances in Computational Mathematics, 50(1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10104-5.

Alfeld:2010:TCM

[ASS10] Peter Alfeld, Larry L. Schumaker, and Tatyana Sorokina. Two condensed macro-elements with full approximation power. Advances in Computational Mathematics, 32(4):381–391, May 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/

10.1007/s10444-008-9110-
7.

Ali:2017:RBM

[ASU17] Mazen Ali, Kristina Steih, and Karsten Urban. Reduced basis methods with adaptive snapshot computations. Advances in Computational Mathematics, 43(2): 257–294, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9485-9.

Atreas:2012:CNU

[Atr12] Nikolaos D. Atreas. On a class of non-uniform average sampling expansions and partial reconstruction in subspaces of L_{2}(R). Advances in Computational Mathematics, 36(1):21–38, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9177-4.

Arslan:2019:EFS

[ATS19] Sinem Arslan and Münevver Tezer-Sezgin. Exact and FDM solutions of 1D MHD flow between parallel electrically conducting and slipping plates. Advances in Computational Mathematics, 45(4): 1923–1938, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-019-09669-x.

Ali:2020:HAH

[AU20] Mazen Ali and Karsten Urban. HT-AWGM: a hierarchical Tucker-adaptive wavelet Galerkin method for highdimensional elliptic problems. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09797-9.

Alla:2015:ASP

[AV15] Alessandro Alla and Stefan Volkwein. Asymptotic stability of POD based model predictive control for a semilinear parabolic PDE. Advances in Computational Mathematics, 41(5):1073– 1102, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9381-0.

Aguilera-Venegas:2019:PEC

[AVGGEG^{+}19] Gabriel Aguilera-Venegas, José Luis Galán-García, Rocío Egea-Guerrero, María Á. Galán-García, Pedro Rodríguez-Cielos, Yolanda Padilla-Domínguez, and María Galán-Luque. A probabilistic extension to Conway’s Game of Life. Advances in Computational Mathematics, 45(4):

2111–2121, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09696-8.

Awanou:2013:RME

[Awa13] Gerard Awanou. Rectangular mixed elements for elasticity with weakly imposed symmetry condition. Advances in Computational Mathematics, 38(2):351–367, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9240-1.

Awanou:2015:PTC

[Awa15] Gerard Awanou. Pseudo transient continuation and time marching methods for Monge–Ampère type equations. Advances in Computational Mathematics, 41(4): 907–935, August 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9391-y.

Atkinson:1998:I

[AX98] Kendall Atkinson and Yuesheng Xu. Introduction. Advances in Computational Mathematics, 9(1–2):0, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/A:1018968919014.

Amsallem:2015:FLR

[AZW15] David Amsallem, Matthew J. Zahr, and Kyle Washabaugh. Fast local reduced basis updates for the efficient reduction of nonlinear systems with hyper-reduction. Advances in Computational Mathematics, 41(5):1187– 1230, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9409-0.

Zou:2023:ECN

[aZWL23] Guang an Zou, Xuyang Wang, and Jian Li. An extrapolated Crank–Nicolson virtual element scheme for the nematic liquid crystal flows. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10028-0.

Borzi:2015:SOA

[BA15] Alfio Borzì and Sergio González Andrade. Second-order approximation and fast multigrid solution of parabolic bilinear optimization problems. Advances in Computational Mathematics, 41(2): 457–488, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9369-9.

Badea:2019:CMM

[Bad19] Lori Badea. On the convergence of a multigrid method for Moreau-regularized variational inequalities of the second kind. Advances in Computational Mathematics, 45(5-6):2807–2832, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09709-6.

Bai:1999:CMB

[Bai99] Zhong-Zhi Bai. A class of modified block SSOR preconditioners for symmetric positive definite systems of linear equations. Advances in Computational Mathematics, 10(2):169–186, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018974514896.

Barry:1994:I

[Bar94]  Phillip Barry.  Introduc-

tion. Advances in Computational Mathematics, 2(1): ix–x, January 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF02519032.

Barrio:2003:URE

[Bar03]  Roberto Barrio.  A uni-

fied rounding error bound for polynomial evaluation. Advances in Computational Mathematics, 19(4):385–399, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1024203520270.

Beatson:2007:MFI

[BB07] R. K. Beatson and H.-Q. Q. Bui. Mollification formulas and implicit smoothing. Advances in Computational Mathematics, 27(2): 125–149, August 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7512-3.

Borghi:2024:MHO

[BB24] Alessandro Borghi and Tobias Breiten. H_{2} optimal rational approximation on general domains. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10125-8.

Blom:2016:CRE

[BBB^{+}16] David S. Blom, Philipp Birken, Hester Bijl, Fleur Kessels, Andreas Meister, and Alexander H. van Zuijlen. A comparison of Rosenbrock and ESDIRK methods combined with iterative solvers for unsteady compressible flows. Advances in Computational Mathematics, 42(6):1401–1426, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9468-x;
http://link.springer.com/
content/pdf/10.1007/s10444-
016-9468-x.pdf.

Barrios:2021:PEE

[BBB21] Tomás P. Barrios, Edwin M. Behrens, and Rommel Bustinza. An a posteriori error estimate for a dual mixed method applied to Stokes system with non-null source terms. Advances in Computational Mathematics, 47(5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09906-2.

Barrios:2022:NAS

[BBB22] Tomás P. Barrios, Edwin M. Behrens, and Rommel Bustinza. Numerical analysis of a stabilized scheme applied to incompressible elasticity problems with Dirichlet and with mixed boundary conditions. Advances in Computational Mathematics, 48(4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09956-0.

Bertrand:2021:CAS

[BBdD21] Fleurianne Bertrand, Daniele Boffi, and Gonzalo G. de Diego. Convergence analysis of the scaled boundary finite element method for the Laplace equation. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09852-z.

Benitez:2024:LAS

[BBF^{+}24] Marta Benítez, Alfredo Bermúdez, Pedro Fontán, Iván Martínez, and Pilar Salgado. A Lagrangian approach for solving an axisymmetric thermoelectromagnetic problem. Application to time-varying geometry processes. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10121-y.

Bonnet-Bendhia:2002:DAG

[BBR02] Anne-Sophie Bonnet-Bendhia and Karim Ramdani. Diffraction by an acoustic grating perturbed by a bounded obstacle. Advances in Computational Mathematics, 16(2– 3):113–138, April 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014437813575. Mod-

eling and computation in optics and electromagnetics.

Brunnett:1994:ECS

[BC94] Guido Brunnett and Peter E.

Crouch. Elastic curves on the sphere. Advances in Computational Mathematics, 2(1): 23–40, January 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02519034.

Becuwe:1999:FPM

[BC99] Stefan Becuwe and Annie Cuyt. On the Froissart phenomenon in multivariate homogeneous Padé approximation. Advances in Computational Mathematics, 11 (1):21–40, July 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018911623074.

Bot:2019:AMO

[BC19] Radu Ioan Bot and Ernö Robert Csetnek. ADMM for monotone operators: convergence analysis and rates. Advances in Computational Mathematics, 45(1):327–359, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9619-3.

Bellomonte:2020:FWF

[BC20] Giorgia Bellomonte and Rosario Corso. Frames and weak frames for unbounded operators. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09773-3.

Bagur:2022:IHM

[BCC22] Laura Bagur, Stéphanie Chaillat, and Patrick Ciarlet, Jr. Improvement of hierarchical matrices for 3D elastodynamic problems with a complex wavenumber. Advances in Computational Mathematics, 48(2):??, April 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09921-3.

Bollt:2009:GAI

[BCE^{+}09] Erik M. Bollt, Rick Chartrand, Selim Esedoğlu, Pete Schultz, and Kevin R. Vixie. Graduated adaptive image denoising: local compromise between total variation and isotropic diffusion. Advances in Computational Mathematics, 31(1–3): 61–85, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9082-7.

Balan:2003:DEF

[BCHL03] Radu Balan, Peter G.

Casazza, Christopher Heil, and Zeph Landau. Deficits and excesses of frames. Advances in Computational Mathematics, 18(2–4):93– 116, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021360227672. Frames.

Bay:2022:HIP

[BCHS22] Thierry Bay, Isabelle Cattiaux-Huillard, and Laura Saini. Hermite interpolation by planar cubic-like ATPH. Advances in Computational Mathematics, 48(5):??, October 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-022-09978-8.

Beatson:1999:FFR

[BCM99] R. K. Beatson, J. B. Cherrie, and C. T. Mouat. Fast fitting of radial basis functions: methods based on preconditioned GMRES iteration. Advances in Computational Mathematics, 11(2–3):253– 270, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018932227617. Ra-

dial basis functions and their applications.

Benner:2019:BMH

[BCS19] Peter Benner, Xingang Cao, and Wil Schilders. A bilinear H_{2} model order reduction approach to linear parameter-varying systems. Advances in Computational Mathematics, 45(5-6):2241– 2271, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09695-9.

Barcelo:2023:LLM

[BCV23] Juan Antonio Barceló, Carlos Castro, and Mari Cruz Vilela. Live load matrix recovery from scattering data in linear elasticity. Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10087-3.

Borwein:2005:ECG

[BCZ05] P. B. Borwein, A. Cuyt, and P. Zhou. Explicit construction of general multivariate Padé approximants to an Appell function. Advances in Computational Mathematics, 22(3):249–273, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-2600-
8.

Bao:2023:SSB

[BCZ23] Feng Bao, Yanzhao Cao, and He Zhang. Splitting scheme for backward doubly stochastic differential equations. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10053-z.

Bank:1993:SMM

[BD93] Randolph E. Bank and Craig C. Douglas. Sparse matrix multiplication package (SMMP). Advances in Computational Mathematics, 1(1): 127–137, February 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02070824.

Benedetto:2010:CIU

[BD10] John J. Benedetto and Somantika Datta. Construction of infinite unimodular sequences with zero autocorrelation. Advances in Computational Mathematics, 32(2): 191–207, February 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9100-9.

Biswas:2018:LDE

[BD18] Biswarup Biswas and Ritesh Kumar Dubey. Low dissipative entropy stable schemes using third order WENO and TVD reconstructions. Advances in Computational Mathematics, 44(4):1153– 1181, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9576-2.

Buckwar:2021:CEM

[BD21a] Evelyn Buckwar and Raffaele D’Ambrosio. Correction to: Exponential meansquare stability properties of stochastic linear multistep methods. Advances in Computational Mathematics, 47 (6):??, December 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09901-7.

See [BD21b].

Buckwar:2021:EMS

[BD21b] Evelyn Buckwar and Raffaele D’Ambrosio. Exponential mean-square stability properties of stochastic linear multistep methods. Advances in Computational Mathematics, 47(4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09879-2.

See correction [BD21a].

Barinka:2006:AAO

[BDD06] Arne Barinka, Stephan Dahlke, and Wolfgang Dahmen. Adaptive application of operators in standard representation. Advances in Computational Mathematics, 24(1– 4):5–34, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7631-2.

Bezchlebova:2019:NST

[BDFS19] Eva Bezchlebová, Vít Dolejší, Miloslav Feistauer, and Petr Sváček. Numerical simulation of two-phase flow of immiscible fluids by the finite element, discontinuous Galerkin and level-set methods. Advances in Computational Mathematics, 45(4): 1993–2018, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09681-1.

Baraniuk:2002:NBT

[BDKY02] R. G. Baraniuk, R. A. De-Vore, G. Kyriazis, and X. M. Yu. Near best tree approximation. Advances in Computational Mathematics, 16 (4):357–373, May 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014554317692.

Buet:2020:TDG

[BDM20] Christophe Buet, Bruno Despres, and Guillaume Morel. Trefftz discontinuous Galerkin basis functions for a class of Friedrichs systems coming from linear transport. Advances in Computational Mathematics, 46 (3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09755-5.

Bonesky:2010:AWM

[BDMR10] Thomas Bonesky, Stephan Dahlke, Peter Maass, and Thorsten Raasch. Adaptive wavelet methods and sparsity reconstruction for inverse heat conduction problems.

Advances in Computational Mathematics, 33(4):385–411, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9147-2.

Benner:2020:BTL

[BDYY20] Peter Benner, Xin Du, Guanghong Yang, and Dan Ye. Balanced truncation of linear time-invariant systems over finite-frequency ranges. Advances in Computational Mathematics, 46 (6):??, December 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09823-w.

Burger:2000:TNN

[BE00] Martin Burger and Heinz W.

Engl. Training neural networks with noisy data as an ill-posed problem. Advances in Computational Mathematics, 13(4):335–354, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016641629556.

Bachoc:2017:OCL

[BEG17] François Bachoc, Martin Ehler, and Manuel Gräf. Optimal configurations of lines and a statistical application. Advances in Computational Mathematics, 43(1): 113–126, February 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9478-8;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        016-9478-8.pdf.

Boito:2022:CRE

[BEG22] Paola Boito, Yuli Eidelman, and Luca Gemignani. Computing the reciprocal of a φfunction by rational approximation. Advances in Computational Mathematics, 48 (1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09917-z.

Beinert:2023:ACB

[Bei23] Robert Beinert. Approximation of curve-based sleeve functions in high dimensions. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10088-2.

Bendtsen:1997:PSO

[Ben97] Claus Bendtsen. A parallel stiff ODE solver based on MIRKs. Advances in Computational Mathematics, 7(1– 2):27–36, March 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018978331792. Par-

allel methods for ODEs.

Bramble:1996:AMA

[BEPS96] James H. Bramble, Richard E.

Ewing, Joseph E. Pasciak, and Jian Shen. The analysis of multigrid algorithms for cell centered finite difference methods. Advances in Computational Mathematics, 5(1): 15–29, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124733.

Bespalov:2003:OSS

[Bes03] Alexei Bespalov. Orthogonal systems of singular functions and numerical treatment of problems with degeneration of data. Advances in Computational Mathematics, 19(1– 3):159–182, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022862704316. Chal-

lenges in computational mathematics (Pohang, 2001).

Bull:1994:PGA

[BF94] J. M. Bull and T. L. Freeman.

Parallel globally adaptive quadrature on the KSR-1. Advances in Computational Mathematics, 2(4):357–373, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02521604.

Benedetto:2003:FNT

[BF03] John J. Benedetto and Matthew Fickus. Finite normalized tight frames. Advances in Computational Mathematics, 18(2–4):357– 385, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1021323312367. Frames.

Benner:2011:NSL

[BF11] Peter Benner and Heike Faßbender. On the numerical solution of largescale sparse discrete-time Riccati equations. Advances in Computational Mathematics, 35(2–4):119– 147, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9174-7.

Brugnano:2022:NFP

[BFCIV22] Luigi Brugnano, Gianluca Frasca-Caccia, Felice Iavernaro, and Vincenzo Vespri. A new framework for polynomial approximation to differential equations. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09992-w.

Brivio:2024:EEP

[BFFM24] Simone Brivio, Stefania Fresca, Nicola Rares Franco, and Andrea Manzoni. Error estimates for POD–DL-ROMs: a deep learning framework for reduced order modeling of nonlinear parametrized PDEs enhanced by proper orthogonal decomposition. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10110-1.

Benner:2018:MRC

[BFH^{+}18] Peter Benner, Heike Faßbender, Michael Hinze, Tatjana Stykel, and Ralf Zimmermann. Model reduction of complex dynamical systems. Advances in Computational Mathematics, 44(6):1687– 1691, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9640-6.

Billaud-Friess:2024:PRB

[BFMNP24] Marie Billaud-Friess, Arthur Macherey, Anthony Nouy, and Clémentine Prieur. A probabilistic reduced basis method for parameterdependent problems. Advances in Computational Mathematics, 50(2):??, April 2024. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
024-10114-x.

Bermudez:2024:NSR

[BGG24] A. Bermúdez, D. Gómez, and D. González. Numerical simulation of resistance furnaces by using distributed and lumped models. Advances in Computational Mathematics, 50(2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10120-z.

Benner:2015:PMO

[BGH15] Peter Benner, Sara Grundel, and Nils Hornung. Parametric model order reduction with a small H_{2}-error using radial basis functions. Advances in Computational Mathematics, 41(5):1231– 1253, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9410-7.

Barry:1993:KIA

[BGM93] Phillip J. Barry, Ronald N.

Goldman, and Charles A. Micchelli. Knot insertion algorithms for piecewise polynomial spaces determined by connection matrices. Advances in Computational Mathematics, 1(2): 139–171, June 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02071383.

Bermudez:2007:TNS

[BGMS07] A. Bermúdez, D. Gómez, M. C. Muñiz, and P. Salgado. Transient numerical simulation of a thermoelectrical problem in cylindrical induction heating furnaces. Advances in Computational Mathematics, 26(1– 3):39–62, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7470-9.

Brander:2017:APE

[BGN17] David Brander, Jens Gravesen, and Toke Bjerge Nørbjerg. Approximation by planar elastic curves. Advances in Computational Mathematics, 43(1):25–43, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/

                                        10.1007/s10444-016-9474-
                                        z.

Beattie:2020:SFM

[BGT20] Christopher Beattie, Serkan Gugercin, and Zoran Tomljanović. Sampling-free model reduction of systems with low-rank parameterization. Advances in Computational Mathematics, 46(6):??, December 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09825-8.

Benner:2021:SPI

[BGW21] Peter Benner, Serkan Gugercin, and Steffen W. R. Werner. Structure-preserving interpolation of bilinear control systems. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09863-w.

Benner:2024:SIM

[BGW24] Peter Benner, Serkan Gugercin, and Steffen W. R. Werner. Structured interpolation for multivariate transfer functions of quadratic-bilinear systems. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10444-024-10109-8.

Borm:2002:MCA

[BH02] S. Börm and R. Hiptmair.

Multigrid computation of axisymmetric electromagnetic fields. Advances in Computational Mathematics, 16(4): 331–356, May 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014533409747.

Bodmann:2015:SPR

[BH15] Bernhard G. Bodmann and Nathaniel Hammen. Stable phase retrieval with lowredundancy frames. Advances in Computational Mathematics, 41(2):317–331, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-014-9359-
y.

Bayramyan:2017:NPP

[BH17] Vahagn Bayramyan and Hakop Hakopian. On a new property of n-poised and GC_{n} sets. Advances in Computational Mathematics, 43 (3):607–626, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9499-3.

Benner:2019:CGB

[BH19] Peter Benner and Christian Himpe. Cross-Gramian-based dominant subspaces. Advances in Computational Mathematics, 45(5-6):2533– 2553, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09724-7.

Beinert:2023:PRS

[BH23] Robert Beinert and Marzieh Hasannasab. Phase retrieval and system identification in dynamical sampling via Prony’s method. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10059-7.

Benner:2018:RIO

[BHM18] Peter Benner, Christian Himpe, and Tim Mitchell. On reduced input-output dynamic mode decomposition. Advances in Computational Mathematics, 44(6):1751– 1768, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9592-x.

Benzi:2011:PTC

[BHT11] Michele Benzi, Eldad Haber, and Lauren Taralli. A preconditioning technique for a class of PDE-constrained optimization problems. Advances in Computational Mathematics, 35(2–4):149– 173, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9173-8.

Bini:2011:NCM

[BI11] Dario Andrea Bini and Bruno Iannazzo. A note on computing matrix geometric means. Advances in Computational Mathematics, 35(2–4):175– 192, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9165-0.

Berghaus:2024:CLE

[BJMR24] David Berghaus, Robert Stephen Jones, Hartmut Monien, and Danylo Radchenko. Computation of Laplacian eigenvalues of two-dimensional shapes with dihedral symmetry. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10138-3.

Brenner:2009:MRV

[BJX09] Marty Brenner, Ying Jiang, and Yuesheng Xu. Multiparameter regularization for Volterra kernel identification via multiscale collocation methods. Advances in Computational Mathematics, 31(4):421–455, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9077-4.

Bramble:1996:FEM

[BK96] James H. Bramble and J. Thomas King. A finite element method for interface problems in domains with smooth boundaries and interfaces. Advances in Computational Mathematics, 6(1):109–138, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127700.

Birem:2016:MSM

[BK16] Mira Birem and Christian Klein. Multidomain spectral method for Schrödinger equations. Advances in Computational Mathematics, 42(2): 395–423, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9429-9.

Boulmezaoud:2024:SFC

[BK24] Tahar Z. Boulmezaoud and Keltoum Kaliche. Stray field computation by inverted finite elements: a new method in micromagnetic simulations. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10139-2.

Boulmezaoud:2017:IFE

[BKK17] Tahar Z. Boulmezaoud, Keltoum Kaliche, and Nabil Kerdid. Inverted finite elements for div–curl systems in the whole space. Advances in Computational Mathematics, 43(6):1469–1489, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9532-1.

Burger:2021:VEM

[BKM^{+}21] Raimund Bürger, Sarvesh Kumar, David Mora, Ricardo Ruiz-Baier, and Nitesh Verma. Virtual element methods for the threefield formulation of timedependent linear poroelasticity. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09826-7.

Bardsley:2011:SLA

[BKN11] Johnathan M. Bardsley, Sarah Knepper, and James Nagy. Structured linear algebra problems in adaptive optics imaging. Advances in Computational Mathematics, 35(2–4):103– 117, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9172-9.

Bochmann:2020:SFA

[BKP20] Maximilian Bochmann, Lutz Kämmerer, and Daniel Potts. A sparse FFT approach for ODE with random coefficients. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09807-w.

Bouhamidi:1999:MIS

[BL99] A. Bouhamidi and A. Le Méhauté. Multivariate interpolating (m, l, s)-splines. Advances in Computational Mathematics, 11(4):287–314, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1023/A:1018940429434.

Bardsley:2009:TVP

[BL09] Johnathan M. Bardsley and Aaron Luttman. Total variation-penalized Poisson likelihood estimation for ill-posed problems. Advances in Computational Mathematics, 31(1–3):35–59, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9081-8.

Blesgen:2007:CEE

[Ble07] T. Blesgen. On the competition of elastic energy and surface energy in discrete numerical schemes. Advances in Computational Mathematics, 27(2):179–194, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9028-5.

Ben-loghfyry:2023:NRF

[BlH23] Anouar Ben-loghfyry and Abdelilah Hakim. A novel robust fractional-time anisotropic diffusion for multi-frame image super-resolution. Advances in Computational Mathematics, 49(6):??, December 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-023-10079-3.

Bercovier:2002:DPP

[BLP02] Michel Bercovier, Moshe Luzon, and Elan Pavlov. Detecting planar patches in an unorganised set of points in space. Advances in Computational Mathematics, 17(1– 2):153–166, July 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1015285722950. Ad-

vances in geometrical algorithms and representations.

Bodmann:2019:SPC

[BLP19] Bernhard G. Bodmann, Demetrio Labate, and Basanta R. Pahari. Smooth projections and the construction of smooth Parseval frames of shearlets. Advances in Computational Mathematics, 45(5-6):3241– 3264, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09736-3.

Bozzini:2002:AIS

[BLS02] Mira Bozzini, Licia Lenarduzzi, and Robert Schaback. Adaptive interpolation by scaled multiquadrics. Advances in Computational Mathematics, 16 (4):375–387, May 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014584220418.

Bonnet:2020:SOS

[BLV20] Marc Bonnet, Ruowen Liu, and Shravan Veerapaneni. Shape optimization of Stokesian peristaltic pumps using boundary integral methods. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09761-7.

Bilato:2014:AFH

[BMB14] Roberto Bilato, Omar Maj, and Marco Brambilla. An algorithm for fast Hilbert transform of real functions. Advances in Computational Mathematics, 40(5–6):1159– 1168, December 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9345-4.

Behera:2015:MAG

[BMK15] Ratikanta Behera, Mani Mehra, and N. K.-R. Kevlahan. Multilevel approximation of the gradient operator on an adaptive spherical geodesic grid. Advances in Computational Mathematics, 41(3):663–689, June 2015.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9382-
                                        z.

Bobade:2019:AEN

[BMP^{+}19] Parag Bobade, Suprotim Majumdar, Savio Pereira, Andrew J. Kurdila, and John B. Ferris. Adaptive estimation for nonlinear systems using reproducing kernel Hilbert spaces. Advances in Computational Mathematics, 45(2): 869–896, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9639-z.

Benac:2020:OFD

[BMRS20] María J. Benac, Pedro Massey, Mariano Ruiz, and Demetrio Stojanoff. Optimal frame designs for multitasking devices with weight restrictions. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09762-6.

Bermudez:2016:TEC

[BMSR^{+}16] A. Bermúdez, R. Muñoz-Sola, C. Reales, R. Rodríguez, and P. Salgado. A transient eddy current problem on a moving domain.

Numerical analysis. Advances in Computational Mathematics, 42(4):757–789, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9441-0.

Boyd:2001:BM

[BMSZ01] Geoff Boyd, Charles A. Micchelli, Gilbert Strang, and Ding-Xuan Zhou. Binomial matrices. Advances in Computational Mathematics, 14 (4):379–391, May 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1012207124894.

Balabanov:2019:RLA

[BN19] Oleg Balabanov and Anthony Nouy. Randomized linear algebra for model reduction. Part I: Galerkin methods and error estimation. Advances in Computational Mathematics, 45(5-6):2969– 3019, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09725-6.

Balabanov:2021:RLA

[BN21] Oleg Balabanov and Anthony Nouy. Randomized linear algebra for model reduction — Part II: minimal residual methods and dictionary-based approximation. Advances in Computational Mathematics, 47 (2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09836-5.

Bajpai:2014:TGF

[BNP14] Saumya Bajpai, Neela Nataraj, and Amiya K. Pani. On a two-grid finite element scheme for the equations of motion arising in Kelvin– Voigt model. Advances in Computational Mathematics, 40(5–6):1043–1071, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9340-1.

Bonizzoni:2020:LSP

[BNPP20] Francesca Bonizzoni, Fabio Nobile, Ilaria Perugia, and Davide Pradovera. Leastsquares Padé approximation of parametric and stochastic Helmholtz maps. Advances in Computational Mathematics, 46(3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09749-3.

Barthelmann:2000:HDP

[BNR00] Volker Barthelmann, Erich Novak, and Klaus Ritter.

High dimensional polynomial interpolation on sparse grids. Advances in Computational Mathematics, 12(4): 273–288, March 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018977404843. Mul-

tivariate polynomial interpolation.

Bochniak:2004:DSF

[Boc04] Marius Bochniak. The dual singular function method for 2D boundary integral equations. Advances in Computational Mathematics, 20(4): 293–310, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1027370428226.

Bochacik:2023:PES

[Boc23] Tomasz Bochacik. On the properties of the exceptional set for the randomized Euler and Runge–Kutta schemes. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10012-8.

Benner:2015:MOR

[BOP^{+}15] Peter Benner, Mario Ohlberger, Anthony T. Patera, Gianluigi Rozza, Danny C.

Sorensen, and Karsten Urban. Model order reduction of parameterized systems (MoRePaS): Preface to the special issue of advances in computational mathematics. Advances in Computational Mathematics, 41(5): 955–960, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9443-y;
 http://link.springer.com/
content/pdf/10.1007/s10444-
015-9443-y.pdf.

Baker:1993:PCR

[BP93] Christopher T. H. Baker and Christopher A. H. Paul. Parallel continuous Runge– Kutta methods and vanishing lag delay differential equations. Advances in Computational Mathematics, 1(3): 367–394, October 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02072017.

Bru:2011:PNA

[BP11] Rafael Bru and Juan Manuel Peña. Preface: numerical and applied linear algebra. Advances in Computational Mathematics, 35(2–4): 99–102, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.

1007/s10444-011-9170-y;
http://link.springer.com/
content/pdf/10.1007/s10444-
011-9170-y.pdf.

Bermudez:2002:HAC

[BPJ02] A. Bermúdez, D. Gómez Pedreira, and P. Joly. A hybrid approach for the computation of guided modes in integrated optics. Advances in Computational Mathematics, 16(2– 3):229–261, April 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014454317209. Mod-

eling and computation in optics and electromagnetics.

Bramble:1996:SPA

[BPK96] James H. Bramble, Joseph E.

Pasciak, and Andrew V. Knyazev. A subspace preconditioning algorithm for eigenvector/eigenvalue computation. Advances in Computational Mathematics, 6(1):159–189, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127702.

Baker:1995:INS

[BPW95] C. T. H. Baker, C. A. H.

Paul, and D. R. Willé. Issues in the numerical solution of evolutionary delay differential equations.

Advances in Computational Mathematics, 3(1–2):171– 196, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02988625.

Baroli:2013:CSD

[BQRB13] Davide Baroli, Alfio Quarteroni, and Ricardo Ruiz-Baier. Convergence of a stabilized discontinuous Galerkin method for incompressible nonlinear elasticity. Advances in Computational Mathematics, 39(2): 425–443, August 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9286-8.

Brezinski:2007:CSK

[BR07] C. Brezinski and M. Raydan. Cauchy–Schwarz and Kantorovich type inequalities for scalar and matrix moment sequences. Advances in Computational Mathematics, 26(1– 3):71–80, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7646-8.

Bosner:2013:VDP

[BR13] Tina Bosner and Mladen Rogina. Variable degree polynomial splines are Chebyshev splines. Advances in Computational Mathematics, 38(2): 383–400, February 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9242-z.

Buscaglia:2015:AMD

[BR15] Gustavo Buscaglia and Vitoriano Ruas. Approximation of the multi-dimensional Stokes system with embedded pressure discontinuities. Advances in Computational Mathematics, 41(3): 599–634, June 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9378-8.

Borges:2022:MIO

[BR22] Carlos Borges and Manas Rachh. Multifrequency inverse obstacle scattering with unknown impedance boundary conditions using recursive linearization. Advances in Computational Mathematics, 48(1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09915-1.

Buccini:2023:LMR

[BR23] Alessandro Buccini and Lothar Reichel. Limited memory restarted l^{p}– l^{q} minimization methods using generalized Krylov subspaces. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10020-8.

Brandts:2001:CRE

[Bra01] Jan Brandts. The Cauchy– Riemann equations: discretization by finite elements, fast solution of the second variable, and a posteriori error estimation. Advances in Computational Mathematics, 15(1–4):61–77, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014217225870. A

posteriori error estimation and adaptive computational methods.

Brenner:1995:TLA

[Bre95] Susanne C. Brenner. A two-level additive Schwarz preconditioner for the stationary Stokes equations. Advances in Computational Mathematics, 4(1–2):111– 126, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123475.

Bremer:2019:ARN

[Bre19] James Bremer. An algorithm for the rapid numerical evaluation of Bessel functions of real orders and arguments. Advances in Computational Mathematics, 45(1): 173–211, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9613-9.

Bozzini:2018:DRM

[BRR18] Mira Bozzini, Christophe Rabut, and Milvia Rossini. Decomposition and reconstruction of multidimensional signals by radial functions with tension parameters. Advances in Computational Mathematics, 44(4): 1003–1040, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9571-7.

Bozzini:2013:GWM

[BRS13] Mira Bozzini, Milvia Rossini, and Robert Schaback. Generalized Whittle–Matérn and polyharmonic kernels. Advances in Computational Mathematics, 39(1):129–141, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-012-9277-
9.

Bozzini:2015:RKS

[BRSV15] Mira Bozzini, Milvia Rossini, Robert Schaback, and Elena Volontè. Radial kernels via scale derivatives. Advances in Computational Mathematics, 41(2):277–291, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9366-
                                        z.

Brummelhuis:2023:CNS

[Bru23] Raymond Brummelhuis. Convergence of non-stationary semi-discrete RBF schemes for the heat and wave equation. Advances in Computational Mathematics, 49 (5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10058-8.

Brezinski:1994:GEP

[BRZ94] C. Brezinski and M. Redivo-Zaglia. A general extrapolation procedure revisited. Advances in Computational Mathematics, 2(4):461–477, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02521609.

Burrage:1997:PIMb

[BS97a] K. Burrage and H. Suhartanto. Parallel iterated method based on multistep Runge–Kutta of Radau type for stiff problems. Advances in Computational Mathematics, 7(1–2):59–77, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018982432701. Par-

allel methods for ODEs.

Burrage:1997:PIMa

[BS97b] K. Burrage and H. Suhartanto. Parallel iterated methods based on multistep Runge–Kutta methods of Radau type. Advances in Computational Mathematics, 7(1–2):37–57, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018930415863. Par-

allel methods for ODEs.

Burrage:2000:PIM

[BS00] K. Burrage and H. Suhartanto. Parallel iterated methods based on variable stepsize multistep Runge–Kutta methods of Radau type for stiff problems. Advances in Computational Mathematics, 13(3):257–270, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018906311204.

Bejancu:2005:MAO

[BS05] Aurelian Bejancu and Malcolm A. Sabin. Maximal approximation order for a boxspline semi-cardinal interpolation scheme on the threedirection mesh. Advances in Computational Mathematics, 22(3):275–298, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-003-2601-
                                        7.

Benedetto:2006:TFG

[BS06] John J. Benedetto and Songkiat Sumetkijakan. Tight frames and geometric properties of wavelet sets. Advances in Computational Mathematics, 24(1–4):35–56, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7623-2.

Butzer:2023:BKA

[BS23] P. L. Butzer and R. L. Stens. The Butzer–Kozakiewicz article on Riemann derivatives of 1954 and its influence. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10086-4.

Belonosov:2019:CAN

[BSK19] Andrey Belonosov, Maxim Shishlenin, and Dmitriy Klyuchinskiy. A comparative analysis of numerical methods of solving the continuation problem for 1D parabolic equation with the data given on the part of the boundary. Advances in Computational Mathematics, 45(2): 735–755, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9631-7.

Bouadjila:2022:NAT

[BSSM22] Khaled Bouadjila, Ali Samir Saad, Mazen Saad, and Wissal Mesfar. Numerical analysis for two-phase flow with non-equilibrium capillary pressure in anisotropic porous media. Advances in Computational Mathematics, 48(5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09972-0.

Beale:2024:ADL

[BST24] J. Thomas Beale, Michael Storm, and Svetlana Tlupova. The adjoint double layer potential on smooth surfaces in R^{3} and the Neumann problem. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10111-0.

Buhmann:1999:F

[Buh99]  Martin Buhmann.  Fore-

word. Advances in Computational Mathematics, 11(2– 3):i, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018963607144.

Buhmann:2006:SIC

[Buh06] M. D. Buhmann. Semiinfinite cardinal interpolation with multiquadrics and beyond. Advances in Computational Mathematics, 24(1– 4):57–80, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7612-5.

Bunger:2011:MME

[Bün11]  F. Bünger.  Minimizing

and maximizing the Euclidean norm of the product of two polynomials. Advances in Computational Mathematics, 35(2–4):193– 215, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9158-z.

Burrage:1997:PMO

[Bur97]  Kevin Burrage.  Parallel

methods for ODEs. Advances in Computational Mathematics, 7(1–2):1–3, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018997130884.

Butcher:1997:OSP

[But97] J. C. Butcher. Order and stability of parallel methods for stiff problems. Advances in Computational Mathematics, 7(1–2):79–96, March 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018934516771. Par-

allel methods for ODEs.

Bernreuther:2024:ACS

[BV24] Marco Bernreuther and Stefan Volkwein. An adaptive certified space-time reduced basis method for nonsmooth parabolic partial differential equations. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10137-4.

Butler:2015:LAL

[BW15] Ronald W. Butler and Andrew T. A. Wood. Laplace approximation of Lauricella functions F_{A} and F_{D}. Advances in Computational Mathematics, 41(6):1451– 1473, December 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9397-5.

Benner:2020:HNA

[BW20] Peter Benner and Steffen W. R. Werner. Hankel-norm approximation of large-scale descriptor systems. Advances in Computational Mathematics, 46(3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09750-w.

Bogacki:1997:DBO

[BWX97] Przemysław Bogacki, Stanley Weinstein, and Yuesheng Xu. Distances between oriented curves in geometric modeling. Advances in Computational Mathematics, 7(4): 593–621, August 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018923609019.

Bajaj:1994:NAS

[BX94] Chandrajit L. Bajaj and Guo Liang Xu. NURBS approximation of surface/ surface intersection curves.

Advances in Computational Mathematics, 2(1):1–21, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/BF02519033.

Ben-yu:2001:JIA

[ByLl01] Guo Ben-yu and Wang Lilian. Jacobi interpolation approximations and their applications to singular differential equations. Advances in Computational Mathematics, 14 (3):227–276, April 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016681018268.

Bertram:2018:TIN

[BZ18] Anna Bertram and Ralf Zimmermann. Theoretical investigations of the new cokriging method for variablefidelity surrogate modeling. Advances in Computational Mathematics, 44(6):1693– 1716, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9585-1.

Bendokat:2024:CGM

[BZA24a] Thomas Bendokat, Ralf Zimmermann, and P. A. Absil. Correction: A Grassmann manifold handbook: basic geometry and computational aspects. Advances in Computational Mathematics, 50(2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10115-w.

See [BZA24b].

Bendokat:2024:GMH

[BZA24b] Thomas Bendokat, Ralf Zimmermann, and P. A. Absil. A Grassmann manifold handbook: basic geometry and computational aspects. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10090-8.

See correction [BZA24a].

Bittens:2019:DSF

[BZI19] Sina Bittens, Ruochuan Zhang, and Mark A. Iwen. A deterministic sparse FFT for functions with structured Fourier sparsity. Advances in Computational Mathematics, 45(2):519–561, April 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-018-9626-
                                        4.

Cai:2002:AIE

[Cai02] Wei Cai. Algorithmic issues for electromagnetic scattering in layered media:

Green’s functions, current basis, and fast solver. Advances in Computational Mathematics, 16(2–3):157– 174, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014429611757. Mod-

eling and computation in optics and electromagnetics.

Calvi:2005:IUA

[Cal05] Jean-Paul Calvi. Intertwining unisolvent arrays for multivariate Lagrange interpolation. Advances in Computational Mathematics, 23 (4):393–414, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1840-
6.

Chui:2006:COM

[CaL06] Charles K. Chui and Jian ao Lian. Construction of orthonormal multi-wavelets with additional vanishing moments. Advances in Computational Mathematics, 24 (1–4):239–262, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7610-
7.

Carnicer:1995:MCP

[Car95] J. M. Carnicer. Multivariate convexity preserving interpolation by smooth functions. Advances in Computational Mathematics, 3(4): 395–404, May 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02432005.

Casas:2007:UPL

[Cas07]  Eduardo Casas.  Using

piecewise linear functions in the numerical approximation of semilinear elliptic control problems. Advances in Computational Mathematics, 26(1–3):137– 153, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4142-0.

Carstensen:2001:ESP

[CBK01] Carsten Carstensen, Sören Bartels, and Roland Klose. An experimental survey of a posteriori Courant finite element error control for the Poisson equation. Advances in Computational Mathematics, 15(1–4):79– 106, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014269318616. A

posteriori error estimation and adaptive computational methods.

Cheng:1994:CPB

[CC94] Zheng Xing Cheng and Charles K. Chui. Convexity of parametric Bézier surfaces in terms of Gaussian curvature signatures. Advances in Computational Mathematics, 2(4):437–459, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521608.

Casazza:2003:GFI

[CC03] Peter G. Casazza and Ole Christensen. Gabor frames over irregular lattices. Advances in Computational Mathematics, 18(2–4):329– 344, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021356503075. Frames.

Cao:2022:SEE

[CC22] Dewei Cao and Hu Chen.

Sharp error estimate of Grünwald–Letnikov scheme for a multi-term time fractional diffusion equation. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09999-3.

Castrillon-Candas:2024:SBL

[CC24] Julio Enrique Castrillón-Candás. Spatial best linear unbiased prediction: a computational mathematics approach for high dimensional massive datasets. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10132-9.

Chen:2020:DPN

[CCDL20] Chuchu Chen, David Cohen, Raffaele D’Ambrosio, and Annika Lang. Driftpreserving numerical integrators for stochastic Hamiltonian systems. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09771-5.

Cui:2012:REE

[CCE^{+}12] Ming Cui, Zhangxin Chen, Richard E. Ewing, Guan Qin, and Hongsen Chen. Reliable and efficient error control for an adaptive Galerkincharacteristic method for convection-dominated diffusion problems. Advances in Computational Mathematics, 37(3):319–353, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9210-7.

Cao:2010:EAF

[CCG10] Yanzhao Cao, Zheng Chen, and Max Gunzburger. Error analysis of finite element approximations of the stochastic Stokes equations. Advances in Computational Mathematics, 33(2): 215–230, August 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9127-6.

Choi:2008:TDO

[CCH^{+}08] Hyeong In Choi, Sung Woo Choi, Chang Yong Han, Tae wan Kim, Song-Hwa Kwon, Hwan Pyo Moon, Kyeong Hah Roh, and Nam-Sook Wee. Two-dimensional offsets and medial axis transform. Advances in Computational Mathematics, 28(2): 171–199, February 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9036-5.

Calderbank:2011:SFF

[CCH^{+}11] Robert Calderbank, Peter G.

Casazza, Andreas Heinecke, Gitta Kutyniok, and Ali Pezeshki. Sparse fusion frames: existence and construction. Advances in Computational Mathematics, 35 (1):1–31, July 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9162-3.

Collowald:2015:NRC

[CCH^{+}15] Mathieu Collowald, Annie Cuyt, Evelyne Hubert, Wen-Shin Lee, and Oliver Salazar Celis. Numerical reconstruction of convex polytopes from directional moments. Advances in Computational Mathematics, 41(6):1515– 1535, December 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9401-0.

Castrillon-Candas:2020:ARS

[CCK20] Julio E. Castrillón-Candás and Mark Kon. Analytic regularity and stochastic collocation of high-dimensional Newton iterates. Advances in Computational Mathematics, 46(3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09791-1.

Conti:2021:SRP

[CCLM21] Costanza Conti, Mariantonia Cotronei, Demetrio Labate, and Wilfredo Molina. Stable recovery of planar regions with algebraic boundaries in Bernstein form. Advances in Computational Mathematics, 47(2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09843-0.

Charina:2021:JSR

[CCMM21] M. Charina, C. Conti, T. Mejstrik, and J.-L. Merrien. Joint spectral radius and ternary Hermite subdivision. Advances in Computational Mathematics, 47 (2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09854-x.

Chan:2016:SDF

[CCN^{+}16] Alice Z.-Y. Chan, Martin S.

Copenhaver, Sivaram K. Narayan, Logan Stokols, and Allison Theobold. On structural decompositions of finite frames. Advances in Computational Mathematics, 42 (3):721–756, June 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9440-1.

Castrillon-Candas:2021:HCP

[CCNT21] Julio E. Castrillón-Candás, Fabio Nobile, and Raúl F. Tempone. A hybrid collocation– perturbation approach for PDEs with random domains. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09859-6.

Claeys:2022:NOS

[CCP22] Xavier Claeys, Francis Collino, and Emile Parolin. Nonlocal optimized Schwarz methods for time-harmonic electromagnetics. Advances in Computational Mathematics, 48 (6):??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09987-7.

Conti:2016:FHS

[CCS16] Costanza Conti, Mariantonia Cotronei, and Tomas Sauer. Factorization of Hermite subdivision operators preserving exponentials and polynomials. Advances in Computational Mathematics, 42(5):1055– 1079, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9453-4.

Cai:2009:CAT

[CCSS09] Jian-Feng Cai, Raymond H. Chan, Lixin Shen, and Zuowei Shen. Convergence analysis of tight framelet approach for missing data recovery. Advances in Computational Mathematics, 31(1–3):87– 113, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9084-5.

Charina:2015:SSN

[CD15] Maria Charina and Geir Dahl. Subdivision schemes, network flows and linear optimization. Advances in Computational Mathematics, 41 (3):507–528, June 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9371-2.

Cordero:2020:GBJ

[CdGDN20] Elena Cordero, Maurice de Gosson, Monika Dörfler, and Fabio Nicola. Generalized Born–Jordan distributions and applications. Advances in Computational Mathematics, 46(4):??, August 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
020-09788-w.

Chen:2024:EAC

[CDM^{+}24] Minghua Chen, Weihua Deng, Chao Min, Jiankang Shi, and Martin Stynes. Error analysis of a collocation method on graded meshes for a fractional Laplacian problem. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10146-3.

Cai:2003:DAH

[CDP03] Zhiqiang Cai, Jim Douglas, Jr., and Moongyu Park. Development and analysis of higher order finite volume methods over rectangles for elliptic equations. Advances in Computational Mathematics, 19 (1–3):3–33, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022841012296. Chal-

lenges in computational mathematics (Pohang, 2001).

Caponnetto:2009:ECC

[CDP09] Andrea Caponnetto, Ernesto De Vito, and Massimiliano Pontil. Entropy conditions for L_{r}-convergence of empirical processes. Advances in Computational Mathematics, 30(4):355–373, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/

10.1007/s10444-008-9072-
9.

Campagna:2022:SIE

[CDPS22] R. Campagna, S. De Marchi, E. Perracchione, and G. Santin. Stable interpolation with exponential-polynomial splines and node selection via greedy algorithms. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09986-8.

Celorrio:2002:PDD

[CDS02] Ricardo Celorrio, Víctor Domínguez, and Francisco-Javier Sayas. Periodic Dirac delta distributions in the boundary element method. Advances in Computational Mathematics, 17(3):211–236, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016002302671.

Cuyt:1999:EMP

[CDTV99] Annie Cuyt, Kathy Driver, Jieqing Tan, and Brigitte Verdonk. Exploring multivariate Padé approximants for multiple hypergeometric series. Advances in Computational Mathematics, 10(1): 29–49, January 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018918429917.

Chertock:2018:HOP

[CEHK18] Alina Chertock, Yekaterina Epshteyn, Hengrui Hu, and Alexander Kurganov. Highorder positivity-preserving hybrid finite-volume-finite-difference methods for chemotaxis systems. Advances in Computational Mathematics, 44(1):327–350, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-017-9545-
                                        9.

Casenave:2015:NRB

[CEL15] Fabien Casenave, Alexandre Ern, and Tony Lelièvre. A nonintrusive reduced basis method applied to aeroacoustic simulations. Advances in Computational Mathematics, 41(5):961–986, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9365-0.

Csetnek:2021:CRB

[CET21] Ernö Robert Csetnek, Andrew Eberhard, and Matthew K. Tam. Convergence rates for boundedly regular systems. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09891-6.

Cerna:2011:COC

[ČF11] Dana Černá and Václav Finěk. Construction of optimally conditioned cubic spline wavelets on the interval. Advances in Computational Mathematics, 34(2): 219–252, February 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9152-5.

Conte:2023:EFM

[CFC23] Dajana Conte and Gianluca Frasca-Caccia. Exponentially fitted methods with a local energy conservation law. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10049-9.

Chen:2021:RBL

[CFZ21] Hongtao Chen, Xiaobing Feng, and Zhimin Zhang. A recovery-based linear C^{0} finite element method for a fourth-order singularly perturbed Monge–Ampère equation. Advances in Computational Mathematics, 47 (2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09847-w.

Carnicer:2004:CBC

[CG04] J. M. Carnicer and M. Gasca. Classification of bivariate configurations with simple Lagrange interpolation formulae. Advances in Computational Mathematics, 20(1– 3):5–16, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1025823727706. Ap-

proximation and applications.

Carnicer:2006:ILG

[CG06] J. M. Carnicer and M. Gasca. Interpolation on lattices generated by cubic pencils. Advances in Computational Mathematics, 24(1–4):113– 130, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7638-8.

Christensen:2010:POD

[CG10] Ole Christensen and Say Song Goh. Pairs of oblique duals in spaces of periodic functions. Advances in Computational Mathematics, 32(3): 353–379, April 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9115-x.

Carnicer:2015:NFG

[CG15] J. M. Carnicer and C. Godés.

A Newton formula for generalized Berzolari–Radon sets. Advances in Computational Mathematics, 41(2):373–386, April 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-014-9362-
3.

Cui:2017:FEA

[CG17] Jintao Cui and Thirupathi Gudi. Finite element approximation of a free boundary plasma problem. Advances in Computational Mathematics, 43(3): 517–535, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9495-7.

Carnicer:2019:EPG

[CG19] Jesús M. Carnicer and Carmen Godés. Extensions of planar GC sets and syzygy matrices. Advances in Computational Mathematics, 45 (2):655–673, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9630-8.

Cantu-Gonzalez:2019:CSG

[CGDHRÁ^{+}19] José Roberto Cantú-González, O. Díaz-Hernández, Elizeth Ramírez-Álvarez, C. I. Enríquez Flores, A. Flores Rosas, and Gerardo J. Escalera Santos. A computational study of the gene expression in the tryptophan operon with two types of cooperativity. Advances in Computational Mathematics, 45(4): 1843–1851, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-09661-x.

Chen:2020:UOC

[CGG^{+}20] X. Chen, V. Gonzalez, E. Goodman, S. Kang, and K. A. Okoudjou. Universal optimal configurations for the p-frame potentials. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09745-7.

Corless:1996:LF

[CGH^{+}96] R. M. Corless, G. H. Gonnet, D. E. G. Hare, D. J. Jeffrey, and D. E. Knuth. On the Lambert W function. Advances in Computational Mathematics, 5(1):329–359, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124750;   https:
//nickhigham.wordpress.
com/2016/08/23/corless-
knuth-and-lambert-w/.

Chen:2013:NAF

[CGH^{+}13] Huajie Chen, Xingao Gong, Lianhua He, Zhang Yang, and Aihui Zhou. Numerical analysis of finite dimensional approximations of Kohn–Sham models. Advances in Computational Mathematics, 38(2):225–256, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9235-y.

Chen:2023:NAD

[CGHK23] Meng Chen, Rong Gao, Yan He, and Linghua Kong. Numerical analysis of the direct method for 3D Maxwell’s equation in Kerr-type nonlinear media. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10029-z.

Christensen:2021:TPU

[CGKK21] Ole Christensen, Say Song Goh, Hong Oh Kim, and Rae Young Kim. Translation partitions of unity, symmetry properties, and Gabor frames. Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09851-0.

Costantini:2001:CSP

[CGM01] P. Costantini, T. N. T. Goodman, and C. Manni. Constructing C^{3} shape preserving interpolating space curves. Advances in Computational Mathematics, 14(2): 103–127, February 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016664630563.

Chapko:2018:ISP

[CGM18] Roman Chapko, Drossos Gintides, and Leonidas Mindrinos. The inverse scattering problem by an elastic inclusion. Advances in Computational Mathematics, 44(2): 453–476, April 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9550-z.

Carnicer:1995:GVD

[CGP95] J. M. Carnicer, T. N. T. Goodman, and J. M. Peña. A generalization of the variation diminishing property. Advances in Computational Mathematics, 3(4): 375–394, May 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02432004.

Conti:2011:AIN

[CGR11] Costanza Conti, Luca Gemignani, and Lucia Romani. From approximating to interpolatory non-stationary subdivision schemes with the same generation properties. Advances in Computational Mathematics, 35(2–4):217– 241, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9175-6.

Conti:2013:CAS

[CGR13] Costanza Conti, Luca Gemignani, and Lucia Romani. A constructive algebraic strategy for interpolatory subdivision schemes induced by bivariate box splines. Advances in Computational Mathematics, 39(2):395–424, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9285-9.

Cotronei:2015:ADM

[CGRS15] Mariantonia Cotronei, Daniele Ghisi, Milvia Rossini, and Tomas Sauer. An anisotropic directional multiresolution and subdivision scheme. Advances in Computational Mathematics, 41(3):709–726, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-014-9384-
                                        x.

Cliffe:1993:EDN

[CGS93] K. A. Cliffe, T. J. Garratt, and A. Spence. Eigenvalues of the discretized Navier–Stokes equation with application to the detection of Hopf bifurcations. Advances in Computational Mathematics, 1(3): 337–356, October 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02072015.

Cochran:2013:EDT

[CGW13] Doug Cochran, Anne Gelb, and Yang Wang. Edge detection from truncated Fourier data using spectral mollifiers. Advances in Computational Mathematics, 38(4): 737–762, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9258-4.

Chaulet:2015:EIS

[CH15] N. Chaulet and H. Haddar. Electromagnetic inverse shape problem for coated obstacles. Advances in Computational Mathematics, 41(6):1615–1641, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9406-3.

Congreve:2022:TGE

[CH22a] Scott Congreve and Paul Houston. Two-grid hpversion discontinuous Galerkin finite element methods for quasilinear elliptic PDEs on agglomerated coarse meshes. Advances in Computational Mathematics, 48(5):??, October 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09968-w.

Coppe:2022:EFA

[CH22b] Vincent Coppé and Daan Huybrechs. Efficient function approximation on general bounded domains using splines on a Cartesian grid. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09946-2.

Chen:2002:FDA

[Che02] Yu Chen. A fast, direct algorithm for the Lippmann– Schwinger integral equation in two dimensions. Advances in Computational Mathematics, 16(2–3):175– 190, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014450116300. Mod-

eling and computation in optics and electromagnetics.

Chen:2006:UCA

[Che06] Yanping Chen. Uniform convergence analysis of finite difference approximations for singular perturbation problems on an adapted grid. Advances in Computational Mathematics, 24(1–4):197– 212, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7641-0.

Chionh:2003:PDM

[Chi03] Eng-Wee Chionh. Parallel Dixon matrices by bracket. Advances in Computational Mathematics, 19(4):373–383, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1024264830187.

Christensen:2017:ECP

[CHL17] Ole Christensen, Marzieh Hasannasab, and Jakob Lemvig. Explicit constructions and properties of generalized shift-invariant systems in L^{2}(R). Advances in Computational Mathematics, 43 (2):443–472, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9492-x.

Cai:2018:SMD

[CHM18] Mingchao Cai, Peiqi Huang, and Mo Mu. Some multilevel decoupled algorithms for a mixed Navier–Stokes/Darcy model. Advances in Computational Mathematics, 44(1): 115–145, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9537-9.

Calvo:1997:SST

[CHMR97] M. Calvo, D. J. Higham, J. I.

Montijano, and L. Rández. Stepsize selection for tolerance proportionality in explicit Runge–Kutta codes. Advances in Computational Mathematics, 7(3):361–382, May 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/A:1018959222223.

Chong:1995:PTC

[Cho95] C. T. Chong. The polynomial topological complexity of Fatou–Julia sets. Advances in Computational Mathematics, 3(4):369–374, May 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02432003.

Congreve:2019:ARE

[CHP19] Scott Congreve, Paul Houston, and Ilaria Perugia. Adaptive refinement for hpversion Trefftz discontinuous Galerkin methods for the homogeneous Helmholtz problem. Advances in Computational Mathematics, 45(1): 361–393, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9621-9.

Chen:2000:CMB

[CHR00] Di-Rong Chen, Bin Han, and Sherman D. Riemenschneider. Construction of multivariate biorthogonal wavelets with arbitrary vanishing moments. Advances in Computational Mathematics, 13 (2):131–165, May 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018950126225.

Christiansen:2015:PDG

[Chr15] Lasse Hjuler Christiansen.

Pairs of dual Gabor frames generated by functions of Hilbert–Schmidt type. Advances in Computational Mathematics, 41(6):1537– 1554, December 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9402-7.

Chen:2006:BSA

[CHRX06] Qiuhui Chen, Norden Huang, Sherman Riemenschneider, and Yuesheng Xu. A Bspline approach for empirical mode decompositions. Advances in Computational Mathematics, 24(1–4):171– 195, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7614-3.

Claeys:2017:SKB

[CHS17] X. Claeys, R. Hiptmair, and E. Spindler. Second kind boundary integral equation for multi-subdomain diffusion problems. Advances in Computational Mathematics, 43(5):1075– 1101, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9517-0.

Cattiaux-Huillard:2020:CES

[CHS20] Isabelle Cattiaux-Huillard and Laura Saini. Characterization and extensive study of cubic and quintic algebraic trigonometric planar PH curves. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09772-4.

Casanova:2021:CEC

[CHS21] Pedro González Casanova and Víctor Hernández-Santamaría. Carleman estimates and controllability results for fully discrete approximations of 1D parabolic equations. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09885-4.

Chui:2003:CST

[CHSS03] Charles K. Chui, Wenjie He, Joachim Stöckler, and Qiyu Sun. Compactly supported tight affine frames with integer dilations and maximum vanishing moments. Advances in Computational Mathematics, 18(2–4):159– 187, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021318804341. Frames.

Cheng:2010:PSI

[CHY10] Jin Cheng, Benny Yiu Chung Hon, and Masahiro Yamamoto. Preface to the special issue on inverse problems. Advances in Computational Mathematics, 33(4):381–383, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9166-z;
http://link.springer.com/
content/pdf/10.1007/s10444-
010-9166-z.pdf.

Cui:2011:GMP

[CHY11] Xiaoke Cui, Ken Hayami, and Jun-Feng Yin. Greville’s method for preconditioning least squares problems. Advances in Computational Mathematics, 35(2–4):243– 269, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9171-x.

Cantero:2013:EOP

[CI13] María José Cantero and Arieh Iserles. On expansions in orthogonal polynomials. Advances in Computational Mathematics, 38(1): 35–61, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9225-0.

Chalmoviansky:2007:NLC

[CJ07] Pavel Chalmovianský and Bert Jüttler. A nonlinear circle-preserving subdivision scheme. Advances in Computational Mathematics, 27(4):375–400, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9011-y.

Ciarlet:2024:VMS

[CJ24] Patrick Ciarlet, Jr. and Erell Jamelot. Variational methods for solving numerically magnetostatic systems. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10089-1.

Christensen:2018:ADG

[CJKK18] Ole Christensen, Augustus J. E. M. Janssen, Hong Oh Kim, and Rae Young Kim. Approximately dual Gabor frames and almost perfect reconstruction based on a class of window functions. Advances in Computational Mathematics, 44(5):1519– 1535, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9595-7.

Chen:2018:NDA

[CJN18] Linghua Chen, Espen Robstad Jakobsen, and Arvid Naess. On numerical density approximations of solutions of SDEs with unbounded coefficients. Advances in Computational Mathematics, 44 (3):693–721, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9558-4.

Chen:2022:SII

[CJW22] Shuli Chen, Daijun Jiang, and Haibing Wang. Simultaneous identification of initial value and source strength in a transmission problem for a parabolic equation. Advances in Computational Mathematics, 48(6): ??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09983-x.

Cao:2017:FDS

[CJX17] Yanzhao Cao, Ying Jiang, and Yuesheng Xu. A fast discrete spectral method for stochastic partial differential equations. Advances in Computational Mathematics, 43(5):973–998, October 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9513-4.

Casazza:2003:ENT

[CK03a] Peter G. Casazza and Jelena Kovačević. Equal-norm tight frames with erasures. Advances in Computational Mathematics, 18(2–4):387– 430, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1021349819855. Frames.

Chae:2003:PSU

[CK03b] Gee Ju Chae and Hong Oh Kim. On the preservation of stability under convolutions. Advances in Computational Mathematics, 19(1– 3):183–193, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1022811229621. Chal-

lenges in computational mathematics (Pohang, 2001).

Casazza:2007:GGS

[CK07] Peter G. Casazza and Gitta Kutyniok. A generalization of Gram–Schmidt orthogonalization generating all Parseval frames. Advances in Computational Mathematics, 27(1):65–78, July 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7478-1.

Chadha:2011:MNP

[CK11] Naresh M. Chadha and Natalia Kopteva. Maximum norm a posteriori error estimate for a 3 d singularly perturbed semilinear reaction–diffusion problem. Advances in Computational Mathematics, 35 (1):33–55, July 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9163-2.

Cvetkovic:2011:SGG

[CKBP11] Ljiljana Cvetković, Vladimir Kostić, Rafael Bru, and Francisco Pedroche. A simple generalization of Geršgorin’s theorem. Advances in Computational Mathematics, 35(2–4):271– 280, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9143-6.

Christensen:2012:GWS

[CKK12] Ole Christensen, Hong Oh Kim, and Rae Young Kim. Gabor windows supported on [−1, 1] and dual windows with small support. Advances in Computational Mathematics, 36(4):525–545, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9189-
                                        0.

Christensen:2016:EWS

[CKK16] Ole Christensen, Hong Oh Kim, and Rae Young Kim. On extensions of wavelet systems to dual pairs of frames. Advances in Computational Mathematics, 42(2): 489–503, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9432-1.

Costantini:1999:DHT

[CKM99] Paolo Costantini, Boris I. Kvasov, and Carla Manni. On discrete hyperbolic tension splines. Advances in Computational Mathematics, 11(4):331–354, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018988312596.

Conceicao:2013:CSC

[CKP13] Ana C. Conceição, Viktor G. Kravchenko, and José C. Pereira. Computing some classes of Cauchy type singular integrals with Mathematica software. Advances in Computational Mathematics, 39(2):273–288, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9279-7.

Chou:2010:OCA

[CKW10] So-Hsiang Chou, Do Y.

Kwak, and K. T. Wee. Optimal convergence analysis of an immersed interface finite element method. Advances in Computational Mathematics, 33(2):149–168, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9122-y.

Cotronei:2002:MCF

[CL02] Mariantonia Cotronei and Maria Laura Lo Cascio. A method for the construction of families of biorthogonal filters with prescribed properties. Advances in Computational Mathematics, 17(3): 199–210, October 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016040516242.

Chen:2007:AEE

[CL07] Wei Chen and Qun Lin.

Asymptotic expansion and extrapolation for the eigenvalue approximation of the biharmonic eigenvalue problem by Ciarlet–Raviart scheme. Advances in Computational Mathematics, 27(1):95–106, July 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-007-9031-
x.

Cahill:2012:DIF

[CL12] Jameson Cahill and Shidong Li. Dimension invariance of finite frames of translates and Gabor frames. Advances in Computational Mathematics, 37(4):505–520, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9222-3.

Chernih:2016:MMC

[CL16] A. Chernih and Q. T. Le Gia. Multiscale methods with compactly supported radial basis functions for the Stokes problem on bounded domains. Advances in Computational Mathematics, 42(5):1187– 1208, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9458-z.

Choi:2018:LFD

[CL18] Gary Pui-Tung Choi and Lok Ming Lui. A linear formulation for disk conformal parameterization of simply-connected open surfaces. Advances in Computational Mathematics, 44(1): 87–114, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9536-x.

Chung:2022:PCC

[CL22] Yu-Min Chung and Austin Lawson. Persistence curves: a canonical framework for summarizing persistence diagrams. Advances in Computational Mathematics, 48 (1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09893-4.

Cai:2016:UCO

[CLC16] Wentao Cai, Jian Li, and Zhangxin Chen. Unconditional convergence and optimal error estimates of the Euler semi-implicit scheme for a generalized nonlinear Schrödinger equation. Advances in Computational Mathematics, 42(6):1311– 1330, December 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-016-9463-2.

Chen:2019:TGF

[CLCH19] Chuanjun Chen, Kang Li, Yanping Chen, and Yunqing Huang. Two-grid finite element methods combined with Crank–Nicolson scheme for nonlinear Sobolev equations. Advances in Computational Mathematics, 45(2): 611–630, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9628-2.

Chen:2008:FFD

[CLL08] Guo Chen, Zhilin Li, and Ping Lin. A fast finite difference method for biharmonic equations on irregular domains and its application to an incompressible Stokes flow. Advances in Computational Mathematics, 29(2): 113–133, August 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9043-6.

Camp:2006:QIF

[CLLS06] Brian Camp, Tao Lin, Yanping Lin, and Weiwei Sun. Quadratic immersed finite element spaces and their approximation capabilities. Advances in Computational Mathematics, 24(1–4):81– 112, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4139-8.

Chatwin:1996:ADS

[CLM96a] P. C. Chatwin, D. M. Lewis, and N. Mole. Atmospheric diffusion: some new mathematical models. Advances in Computational Mathematics, 6(1):227–242, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127705.  John

Crank 80th birthday special issue.

Chui:1996:LAC

[CLM96b] C. K. Chui, Xin Li, and H. N. Mhaskar. Limitations of the approximation capabilities of neural networks with one hidden layer. Advances in Computational Mathematics, 5(1):233–243, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124745.

Choi:2002:CAS

[CLM02] Hyeong In Choi, Doo Seok Lee, and Hwan Pyo Moon. Clifford algebra, spin representation, and rational parameterization of curves and surfaces. Advances in Computational Mathematics, 17 (1–2):5–48, July 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1015294029079. Ad-

vances in geometrical algorithms and representations.

Calvo:2003:TSH

[CLM03] M. Calvo, M. P. Laburta, and J. I. Montijano. Twostep high order starting values for implicit Runge–Kutta methods. Advances in Computational Mathematics, 19 (4):401–412, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1024281802353.

Clain:2018:EPS

[CLM18a] Stéphane Clain, Raphaël Loubère, and Gaspar J. Machado. A posteriori stabilized sixth-order finite volume scheme for one-dimensional steady-state hyperbolic equations. Advances in Computational Mathematics, 44(2): 571–607, April 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9556-6.

See erratum [CLM18b].

Clain:2018:EEP

[CLM18b] Stéphane Clain, Raphaël Loubère, and Gaspar J. Machado. Erratum to: A posteriori stabilized sixth-order finite volume scheme for onedimensional steady-state hyperbolic equations. Advances in Computational Mathematics, 44(2):609, April 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-017-9563-

7. See [CLM18a].

Calvo:2010:APQ

[CLMR10] M. Calvo, M. P. Laburta, J. I.

Montijano, and L. Rández. Approximate preservation of quadratic first integrals by explicit Runge–Kutta methods. Advances in Computational Mathematics, 32(3): 255–274, April 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9105-4.

Calvo:2015:RKP

[CLMR15] M. Calvo, M. P. Laburta, J. I.

Montijano, and L. Rández. Runge–Kutta projection methods with low dispersion and dissipation errors. Advances in Computational Mathematics, 41(1):231–251, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9355-2.

Cimrman:2019:MFE

[CLR19] Robert Cimrman, Vladimír Lukes, and Eduard Rohan. Multiscale finite element calculations in Python using SfePy. Advances in Computational Mathematics, 45(4): 1897–1921, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09666-0.

Chen:2002:NOE

[CLZ02] Zhongying Chen, Ronghua Li, and Aihui Zhou. A note on the optimal L^{2}-estimate of the finite volume element method. Advances in Computational Mathematics, 16 (4):291–303, May 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014577215948.

Chen:2013:UEE

[CLZW13] Hongtao Chen, Qun Lin, Junming Zhou, and Hong Wang. Uniform error estimates for triangular finite element solutions of advection– diffusion equations. Advances in Computational Mathematics, 38(1):83–100, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9228-x.

Chen:1996:HMM

[CM96] Chunhui Chen and O. L.

Mangasarian. Hybrid misclassification minimization. Advances in Computational Mathematics, 5(1):127–136, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124738.

Carrig:1999:POC

[CM99] James J. Carrig, Jr. and Gerard G. L. Meyer. A parameterized ordering for cache-, register- and pipelineefficient Givens QR decomposition. Advances in Computational Mathematics, 10(1): 97–113, January 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018970413988.

Costantini:2003:SPI

[CM03] P. Costantini and C. Manni.

Shape-preserving C^{3} interpolation: the curve case. Advances in Computational Mathematics, 18(1):41–63, January 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021270530342.

Christensen:2012:EBS

[CM12] Ole Christensen and Peter Massopust. Exponential B-splines and the partition of unity property. Advances in Computational Mathematics, 37(3):301–318, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9208-1.

Chernov:2021:EHV

[CMM21] A. Chernov, C. Marcati, and L. Mascotto. p — and hp — virtual elements for the Stokes problem. Advances in Computational Mathematics, 47(2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09831-w.

Carnicer:2003:RS

[CMP03] J. M. Carnicer, E. Mainar, and J. M. Peña. Restricted systems. Advances in Computational Mathematics, 18(1): 79–90, January 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1021267126479.

Cachafeiro:2007:OPR

[CMP07a] A. Cachafeiro, F. Marcellán, and C. Pérez. Orthogonal polynomials with respect to the sum of an arbitrary measure and a Bernstein–Szegő measure. Advances in Computational Mathematics, 26(1–3):81– 104, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7644-x.

Carnicer:2007:SPR

[CMP07b] J. M. Carnicer, E. Mainar, and J. M. Peña. Shape preservation regions for sixdimensional spaces. Advances in Computational Mathematics, 26(1–3):121– 136, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7505-2.

Campi:2021:LVS

[CMP21] C. Campi, F. Marchetti, and E. Perracchione. Learning via variably scaled kernels. Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09875-6.

Carnicer:2023:SRF

[CMP23] J. M. Carnicer, E. Mainar, and J. M. Peña. On the stability of the representation of finite rank operators. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10057-9.

Calvo:2007:EMR

[CMR07] M. Calvo, J. I. Montijano, and L. Rández. On explicit multi-revolution Runge–Kutta schemes. Advances in Computational Mathematics, 26(1–3):105– 120, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7209-z.

Chu:2022:SSB

[CMR22] Ba-Duong Chu, Florian Martin, and Ulrich Reif. Stabilization of spline bases by extension. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09945-3.

Cotronei:2023:HMM

[CMSS23] Mariantonia Cotronei, Caroline Moosmüller, Tomas Sauer, and Nada Sissouno. Hermite multiwavelets for manifold-valued data. Advances in Computational Mathematics, 49(3):??, June 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
023-10042-2.

Chen:1997:PGM

[CMX97] Zhongying Chen, Charles A.

Micchelli, and Yuesheng Xu. The Petrov–Galerkin method for second kind integral equations. II. Multiwavelet schemes. Advances in Computational Mathematics, 7(3): 199–233, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018994802659.

Chen:2002:DWP

[CMX02] Zhongying Chen, Charles A.

Micchelli, and Yuesheng Xu. Discrete wavelet Petrov– Galerkin methods. Advances in Computational Mathematics, 16(1):1–28, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014273420351.

Chen:2007:MCP

[CMX07] Qiuhui Chen, Charles A.

Micchelli, and Yuesheng Xu. On the matrix completion problem for multivariate filter bank construction. Advances in Computational Mathematics, 26(1–3):

173–204, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7535-9.

Crank:1996:PMN

[CN96] J. Crank and P. Nicolson. A practical method for numerical evaluation of solutions of partial differential equations of the heatconduction type [reprint of MR0019410 (8,409b)]. Advances in Computational Mathematics, 6(1):207–226, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127704.  John

Crank 80th birthday special issue.

Cope:1998:RGG

[Cop98]  Davis K. Cope.  Re-

cursive generation of the Galerkin–Chebyshev matrix for convolution kernels. Advances in Computational Mathematics, 9(1–2):21–35, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018933305810. Nu-

merical treatment of boundary integral equations.

Carnicer:1993:SPR

[CP93] J. M. Carnicer and J. M.

Peña. Shape preserving representations and optimality of the Bernstein basis. Advances in Computational Mathematics, 1(2): 173–196, June 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02071384.

Chen:1999:QWA

[CP99] Han-Lin Chen and Si-Long Peng. A quasi-wavelet algorithm for second kind boundary integral equations. Advances in Computational Mathematics, 11(4):355–375, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018992413504.

Chen:2003:LDP

[CP03] Han-Lin Chen and Si-Long Peng. Localization of dual periodic scaling and wavelet functions. Advances in Computational Mathematics, 19(1–3):195– 210, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022882423107. Chal-

lenges in computational mathematics (Pohang, 2001).

Costantini:2004:SPA

[CP04] Paolo Costantini and Francesca Pelosi. Shape-preserving approximation of spatial data. Advances in Computational Mathematics, 20(1–3):25–51, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1025803122254. Ap-

proximation and applications.

Carnicer:2007:WMG

[CP07a] J. M. Carnicer and J. M. Peña. The work of Mariano Gasca. Advances in Computational Mathematics, 26(1– 3):1–8, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-006-9024-1.

Costantini:2007:SPH

[CP07b] Paolo Costantini and Francesca Pelosi. Shape preserving histogram approximation. Advances in Computational Mathematics, 26(1–3):205– 230, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-8008-2.

Casazza:2014:WFF

[CP14] Peter G. Casazza and Jesse Peterson. Weighted fusion frame construction via spectral tetris. Advances in Computational Mathematics, 40 (2):335–351, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9310-7.

Chernov:2015:STP

[CP15] Alexey Chernov and Duong Pham. Sparse tensor product spectral Galerkin BEM for elliptic problems with random input data on a spheroid. Advances in Computational Mathematics, 41(1):77–104, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9350-7.

Calvetti:1999:IEJ

[CR99] D. Calvetti and L. Reichel. On an inverse eigenproblem for Jacobi matrices. Advances in Computational Mathematics, 11(1): 11–20, July 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018951304418.

Christensen:2008:FPW

[CR08] Ole Christensen and Asghar Rahimi. Frame properties of wave packet systems in L^{2}(R^{d}). Advances in Computational Mathematics, 29(2): 101–111, August 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9038-3.

Canuto:2023:PEA

[CR23] Claudio Canuto and Davide Rosso. A posteriori error analysis and adaptivity for a VEM discretization of the Navier–Stokes equations. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10081-9.

Champneys:1993:HHO

[CS93] Alan R. Champneys and Alastair Spence. Hunting for homoclinic orbits in reversible systems: a shooting technique. Advances in Computational Mathematics, 1(1): 81–108, February 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02070822.

Chui:1994:CTS

[CS94] Charles K. Chui and Xian Liang Shi. Continuous twoscale equations and dyadic wavelets. Advances in Computational Mathematics, 2(2): 185–213, March 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF02521107.

Christensen:2003:FSB

[CS03] Ole Christensen and Diana T. Stoeva. p-Frames in separable Banach spaces. Advances in Computational Mathematics, 18(2–4):117– 126, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021364413257. Frames.

Cavaretta:2006:ORI

[CS06] A. S. Cavaretta and A. Sharma.

Optimal recovery of interpolation operators in Hardy spaces. Advances in Computational Mathematics, 24(1–4):131–141, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4134-0.

Cheng:2018:CAM

[CS18] Xiaodong Cheng and Jacquelien M. A. Scherpen. Clustering approach to model order reduction of power networks with distributed controllers. Advances in Computational Mathematics, 44(6):1917– 1939, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9617-5.

Chen:2019:SCA

[CS19] Feng Chen and Jie Shen. Stability and convergence analysis of rotational velocity correction methods for the Navier–Stokes equations. Advances in Computational Mathematics, 45(5-6):3123– 3136, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09729-2.

Cuyt:2018:MEA

[CsL18] Annie Cuyt and Wen shin Lee. Multivariate exponential analysis from the minimal number of samples. Advances in Computational Mathematics, 44(4): 987–1002, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9570-8.

Chui:1998:AFQ

[CSS98] Charles K. Chui, Xianliang Shi, and Joachim Stöckler. Affine frames, quasi-affine frames, and their duals. Advances in Computational Mathematics, 8(1–2):1–17, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018975725857.

Chui:1996:AWG

[CSW96] C. K. Chui, J. Stöckler, and J. D. Ward. Analytic wavelets generated by radial functions. Advances in Computational Mathematics, 5(1): 95–123, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124736.

Chernih:2014:WFI

[CSW14] A. Chernih, I. H. Sloan, and R. S. Womersley. Wendland functions with increasing smoothness converge to a Gaussian. Advances in Computational Mathematics, 40(1):185–200, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9304-5.

Cong:1999:RKN

[CSWP99] Nguyen Huu Cong, Karl Strehmel, Rüdiger Weiner, and Helmut Podhaisky. Runge– Kutta–Nyström-type parallel block predictor–corrector methods. Advances in Computational Mathematics, 10 (2):115–133, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018930732643.

Capobianco:2005:IPW

[CT05] M. R. Capobianco and W. Themistoclakis. Interpolating polynomial wavelets on [−1, 1]. Advances in Computational Mathematics, 23 (4):353–374, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-1828-
                                        2.

Choi:2017:GBE

[CT17a] Young Ok Choi and Johannes Tausch. The Galerkin boundary element method for transient Stokes flow. Advances in Computational Mathematics, 43(3):473–493, June 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-016-9493-
                                        9.

Cuong:2017:WBV

[CT17b] Dao Huy Cuong and Mai Duc Thanh. A well-balanced van Leer-type numerical scheme for shallow water equations with variable topography. Advances in Computational Mathematics, 43(5):1197– 1225, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9521-4.

Cheng:2021:GCF

[CtTB21] Hanz Martin Cheng and Jan ten Thije Boonkkamp. A generalised complete flux scheme for anisotropic advection– diffusion equations. Advances in Computational Mathematics, 47(2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09846-x.

Cheung:2002:SLL

[CTZ02] Hoi Ling Cheung, Canqin Tang, and Ding-Xuan Zhou. Supports of locally linearly independent M-refinable functions, attractors of iterated function systems and tilings. Advances in Computational Mathematics, 17(3):257–268, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1011906532250.

Cullum:1996:IMS

[Cul96] Jane Cullum. Iterative methods for solving Ax = b, GMRES/FOM versus QMR/ BiCG. Advances in Computational Mathematics, 6(1): 1–24, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127693.

Cuminato:1996:NSC

[Cum96] Jose A. Cuminato. Numerical solution of Cauchytype integral equations of index −1 by collocation methods. Advances in Computational Mathematics, 6(1): 47–64, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127695.

Chen:2023:ACA

[CV23] Kewang Chen and Cornelis Vuik. Asymptotic convergence analysis and influence of initial guesses on composite Anderson acceleration. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10095-3.

Chui:2009:PMA

[CW09] Charles K. Chui and Jianzhong Wang. PDE models associated with the bilateral filter. Advances in Computational Mathematics, 31(1–3): 131–156, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9095-2.

Carl:2016:VLS

[CW16] Wolfgang Carl and Johannes Wallner. Variational Laplacians for semidiscrete surfaces. Advances in Computational Mathematics, 42(6):1491–1509, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9472-1;
http://link.springer.com/
content/pdf/10.1007/s10444-
016-9472-1.pdf.

Che:2019:RAA

[CW19] Maolin Che and Yimin Wei. Randomized algorithms for the approximations of Tucker and the tensor train decompositions. Advances in Computational Mathematics, 45(1):395–428, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9622-8.

Chandler-Wilde:1995:EBI

[CWLH95] S. N. Chandler-Wilde, J. Levesley, and D. M. Hough. Evaluation of a boundary integral representation for the conformal mapping of the unit disk onto a simply-connected domain. Advances in Computational Mathematics, 3(1–2): 115–135, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03028363.

Chen:2006:MAM

[CWX06] Zhongying Chen, Bin Wu, and Yuesheng Xu. Multilevel augmentation methods for differential equations. Advances in Computational Mathematics, 24(1–4):213– 238, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4092-6.

Chen:2012:HOF

[CWX12] Zhongying Chen, Junfeng Wu, and Yuesheng Xu. Higher-order finite volume methods for elliptic boundary value problems. Advances in Computational Mathematics, 37(2):191–253, August 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9201-8.

Chen:2004:CAN

[CX04] Qiuhui Chen and Yuesheng Xu. The cascade algorithm for the numerical computation of refinable functions. Advances in Computational Mathematics, 20(1– 3):17–24, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025844410775. Ap-

proximation and applications.

Chen:2006:CMS

[CX06] Di-Rong Chen and Dao-Hong Xiang. The consistency of multicategory support vector machines. Advances in Computational Mathematics, 24(1–4):155– 169, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7207-1.

Ciloglu:2023:SDG

[ÇY23] Pelin Çiloglu and Hamdullah Yücel. Stochastic discontinuous Galerkin methods for robust deterministic control of convection–diffusion equations with uncertain coefficients. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10015-5.

Cheng:2019:NCS

[CZ19] Li Cheng and Xinlong Zhou.

A new computable sufficient condition for the convergence of subdivision schemes with nonnegative masks. Advances in Computational Mathematics, 45(3):

1273–1290, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-09656-8.

Chen:2020:SOE

[CZ20] Rui Chen and Hui Zhang. Second-order energy stable schemes for the new model of the Cahn–Hilliard–MHD equations. Advances in Computational Mathematics, 46 (6):??, December 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09822-x.

Chen:2016:FPA

[CZS16] Dai-Qiang Chen, Yan Zhou, and Li-Juan Song. Fixed point algorithm based on adapted metric method for convex minimization problem with application to image deblurring. Advances in Computational Mathematics, 42(6):1287–1310, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9462-3.

Cao:2017:SIF

[CZZ17] Waixiang Cao, Xu Zhang, and Zhimin Zhang. Superconvergence of immersed finite element methods for interface problems. Advances in Computational Mathematics, 43(4):795–821, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9507-7.

Dadi:2017:DTC

[Dad17] Z. Dadi. Dynamics of twocell systems with discrete delays. Advances in Computational Mathematics, 43(3): 653–676, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9501-0.

Dahmen:1995:P

[Dah95] Wolfgang Dahmen. Preface. Advances in Computational Mathematics, 4(1–2): iii–iv, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123470.

Dai:2016:NDD

[Dai16]  Ping-Fan Dai.  A note

on diagonal dominance, Schur complements and some classes of H-matrices and P - matrices. Advances in Computational Mathematics, 42 (1):1–4, February 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-014-9375-y.

Damelin:2007:PBO

[Dam07] S. B. Damelin. Pointwise bounds of orthogonal expansions on the real line via weighted Hilbert transforms. Advances in Computational Mathematics, 26(4): 451–471, May 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7472-7.

Diop:2013:PBI

[DAP13] El Hadji S. Diop, Radjesvarane Alexandre, and Valérie Perrier. A PDE based and interpolation-free framework for modeling the sifting process in a continuous domain. Advances in Computational Mathematics, 38 (4):801–835, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9260-x.

Daras:2003:GPT

[Dar03] Nicholas J. Daras. Generalized Padé-type approximation and integral representations. Advances in Computational Mathematics, 18(1): 1–39, January 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1023/A:1021273702443.

Davydov:2023:EBL

[Dav23] Oleg Davydov. Error bounds for a least squares meshless finite difference method on closed manifolds. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10044-0.

deBoor:2000:CAM

[dB00]  Carl de Boor.  Compu-

tational aspects of multivariate polynomial interpolation: indexing the coefficients. Advances in Computational Mathematics, 12(4): 289–301, March 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018929521682. Mul-

tivariate polynomial interpolation.

deBoor:2006:IRP

[dB06] Carl de Boor. On interpolation by radial polynomials. Advances in Computational Mathematics, 24(1–4): 143–153, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7630-3.

deBoor:2007:ISS

[dB07]  C. de Boor.  Interpola-

tion from spaces spanned by monomials. Advances in Computational Mathematics, 26(1–3):63–70, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-005-7538-
6.

daCunha:2007:DBG

[dCB07] R. D. da Cunha and D. Becker. Dynamic block GMRES: an iterative method for block linear systems. Advances in Computational Mathematics, 27(4):423–448, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9012-5.

deCarvalho:2020:CNP

[dCFCF20] Pitágoras Pinheiro de Carvalho, Enrique Fernández-Cara, and Juan Bautista Límaco Ferrel. On the computation of Nash and Pareto equilibria for some bi-objective control problems for the wave equation. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09812-z.

Davies:1994:TMN

[DD94] Penny J. Davies and Dugald B.

Duncan. Time-marching numerical schemes for the electric field integral equation on a straight thin wire. Advances in Computational Mathematics, 2(3): 279–317, April 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521113.

Duan:2023:RCG

[DDLL23] Shan-Qi Duan, Xue-Feng Duan, Chun-Mei Li, and Jiao-Fen Li. Riemannian conjugate gradient method for low-rank tensor completion. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10036-0.

DAmbrosio:2014:NIH

[DDP14] Raffaele D’Ambrosio, Giuseppe De Martino, and Beatrice Paternoster. Numerical integration of Hamiltonian problems by G-symplectic methods. Advances in Computational Mathematics, 40(2): 553–575, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-013-9318-z.

DeVilliers:2006:RFS

[De 06] Johan De Villiers. On refinable functions and subdivision with positive masks. Advances in Computational Mathematics, 24(1–4):281– 295, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7628-x.

DellAcqua:2017:NTB

[Del17] Pietro Dell’Acqua. A note on Taylor boundary conditions for accurate image restoration. Advances in Computational Mathematics, 43(6):1283–1304, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9525-0.

Derriennic:2004:SPP

[Der04] Marie Madeleine Derriennic. Shape preserving polynomial curves. Advances in Computational Mathematics, 20 (4):273–292, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1027303701445.

Dolz:2021:FOM

[DES21] J. Dölz, H. Egger, and V. Shashkov. A fast and oblivious matrix compression algorithm for Volterra integral operators. Advances in Computational Mathematics, 47(6):??, December 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09902-6.

deFrutos:2018:AGD

[dFGAJN18] Javier de Frutos, Bosco García-Archilla, Volker John, and Julia Novo. Analysis of the grad–div stabilization for the time-dependent Navier– Stokes equations with inf– sup stable finite elements. Advances in Computational Mathematics, 44(1):195–225, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9540-1.

Dakroub:2023:PEE

[DFOS23] Jad Dakroub, Joanna Faddoul, Pascal Omnes, and Toni Sayah. A posteriori error estimates for the time-dependent Navier– Stokes system coupled with the convection–diffusion– reaction equation. Advances in Computational Mathematics, 49(4):??, August 2023.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-023-10066-
                                        8.

Dahlke:2007:AFM

[DFR07] Stephan Dahlke, Massimo Fornasier, and Thorsten Raasch. Adaptive frame methods for elliptic operator equations. Advances in Computational Mathematics, 27(1):27–63, July 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7501-6.

Dominguez:2013:ICA

[DG13] V. Domínguez and M. Ganesh. Interpolation and cubature approximations and analysis for a class of wideband integrals on the sphere. Advances in Computational Mathematics, 39(3–4):547– 584, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-012-9293-9.

Dominguez:2016:SEC

[DG16] V. Domínguez and M. Ganesh. Sobolev estimates for constructive uniform-grid FFT interpolatory approximations of spherical functions. Advances in Computational Mathematics, 42(4):843–887, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9445-9.

Dong:2015:ATV

[DGK15] Yiqiu Dong, Torsten Görner, and Stefan Kunis. An algorithm for total variation regularized photoacoustic imaging. Advances in Computational Mathematics, 41(2): 423–438, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9364-1.

Dehmer:2013:EHA

[DGMES13] Matthias Dehmer, Martin Grabner, Abbe Mowshowitz, and Frank Emmert-Streib. An efficient heuristic approach to detecting graph isomorphism based on combinations of highly discriminating invariants. Advances in Computational Mathematics, 39(2):311–325, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9281-0.

Dominguez:2016:PMF

[DGMM16] Sebastián Domínguez, Gabriel N.

Gatica, Antonio Márquez, and Salim Meddahi. A primal-mixed formulation for the strong coupling of quasi-Newtonian fluids with porous media. Advances in Computational Mathematics, 42(3): 675–720, June 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9439-7.

Dunster:2018:UAE

[DGS18] T. M. Dunster, A. Gil, and J. Segura. Uniform asymptotic expansions for Laguerre polynomials and related confluent hypergeometric functions. Advances in Computational Mathematics, 44(5):1441– 1474, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9589-5.

daCunha:1994:PIR

[dH94] Rudnei Dias da Cunha and Tim Hopkins. A parallel implementation of the restarted GMRES iterative algorithm for nonsymmetric systems of linear equations. Advances in Computational Mathematics, 2(3): 261–277, April 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521112.

Diaz:2023:IMR

[DHGA23] Alejandro N. Diaz, Matthias Heinkenschloss, Ion Victor Gosea, and Athanasios C. Antoulas. Interpolatory model reduction of quadraticbilinear dynamical systems with quadratic-bilinear outputs. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10096-2.

Dallas:1998:ONS

[DHK98] Allan G. Dallas, G. C. Hsiao, and R. E. Kleinman. Observations on the numerical stability of the Galerkin method. Advances in Computational Mathematics, 9(1–2): 37–67, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018941607627. Nu-

merical treatment of boundary integral equations.

Deano:2016:CIA

[DHO16] Alfredo Deaño, Daan Huybrechs, and Peter Opsomer. Construction and implementation of asymptotic expansions for Jacobi-type orthogonal polynomials. Advances in Computational Mathematics, 42(4):791–822, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9442-z.

Diamantopoulos:2019:SMP

[DHSW19] Georgios Diamantopoulos, Andreas Hössinger, Siegfried Selberherr, and Josef Weinbub. A shared memory parallel multi-mesh fast marching method for re-distancing. Advances in Computational Mathematics, 45(4):2029– 2045, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09683-z.

Dong:2018:LPF

[DHWZ18] Xiaojing Dong, Yinnian He, Hongbo Wei, and Yuhong Zhang. Local and parallel finite element algorithm based on the partition of unity method for the incompressible MHD flow. Advances in Computational Mathematics, 44(4): 1295–1319, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9582-4.

Drouiche:1999:DMW

[DKB99] Karim Drouiche, Djalil Kateb, and Daniel Boichu. Daubechies– Matzinger wavelets and Lorentz polynomials. Advances in Computational Mathematics, 10(3–4):239– 260, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018934800784.

Djoko:2022:SEU

[DKK22] Jules K. Djoko, Jonas Koko, and Sognia Konlack. Stokes equations under Tresca friction boundary condition: a truncated approach. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09933-7.

DeVore:1993:WCN

[DKLT93] R. A. DeVore, G. Kyriazis, D. Leviatan, and V. M. Tikhomirov. Wavelet compression and nonlinear nwidths. Advances in Computational Mathematics, 1(2): 197–214, June 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02071385.

Donatelli:2020:MPA

[DKMT20] Marco Donatelli, Rolf Krause, Mariarosa Mazza, and Ken Trotti. Multigrid preconditioners for anisotropic spacefractional diffusion equations. Advances in Computational Mathematics, 46 (3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09790-2.

Dagnino:1998:NID

[DL98] C. Dagnino and P. Lamberti. Numerical integration of 2-D integrals based on local bivariate C^{1} quasiinterpolating splines. Advances in Computational Mathematics, 8(1–2):19–31, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018927809928.

Dekel:2004:MEK

[DL04] S. Dekel and D. Leviatan. On measuring the efficiency of kernel operators in L_{p}(R^{d}). Advances in Computational Mathematics, 20(1–3):53–65, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1025838129042. Ap-

proximation and applications.

Duan:2023:MVE

[DL23] Huoyuan Duan and Ziliang Li. A mixed virtual element method for nearly incompressible linear elasticity equations. Advances in Computational Mathematics, 49 (1):??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10013-7.

deLeon:2007:DVI

[dLdDSM07] Manuel de León, David Martín de Diego, and Aitor Santamaría-Merino. Discrete variational integrators and optimal control theory. Advances in Computational Mathematics, 26(1–3):251– 268, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4093-5.

Dong:2012:FAI

[DLK12] Fangfang Dong, Chunxiao Liu, and De-Xing Kong. A fast algorithm for image segmentation based on fuzzy region competition. Advances in Computational Mathematics, 37(4):521–542, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9221-4.

Franco-de-Leon:2019:BIM

[dLL19] Mariano Franco de Leon and John Lowengrub. Boundary integral methods for dispersive equations, Airy flow and the modified Korteweg de Vries equation. Advances in Computational Mathematics, 45(1):99–135, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9607-7.

Diele:1998:CTN

[DLP98] F. Diele, L. Lopez, and R. Peluso. The Cayley transform in the numerical solution of unitary differential systems. Advances in Computational Mathematics, 8(4): 317–334, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018908700358.

Dominguez:2014:NMT

[DLS14] Víctor Domínguez, Sijiang L. Lu, and Francisco-Javier Sayas. A Nyström method for the two dimensional Helmholtz hypersingular equation. Advances in Computational Mathematics, 40(5–6): 1121–1157, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-014-9344-
                                        5.

Donat:2019:NSS

[DLU19] Rosa Donat and Sergio López-Ureña. Nonlinear stationary subdivision schemes reproducing hyperbolic and trigonometric functions. Advances in Computational Mathematics, 45(5-6):3137– 3172, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09731-8.

Donat:2017:FNO

[DLUS17] Rosa Donat, Sergio López-Ureña, and Maria Santágueda. A family of non-oscillatory 6-point interpolatory subdivision schemes. Advances in Computational Mathematics, 43(4):849–883, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9509-5.

Du:2016:PEA

[DLZ16] Shaohong Du, Runchang Lin, and Zhimin Zhang. A posteriori error analysis of multipoint flux mixed finite element methods for interface problems. Advances in Computational Mathematics, 42(4):921–945, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9447-7.

Dahmen:1993:CRE

[DM93] Wolfgang Dahmen and Charles A. Micchelli. Continuous refinement equations and subdivision. Advances in Computational Mathematics, 1(1): 1–37, February 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02070819.

Dubuc:1999:DHI

[DM99] Serge Dubuc and Jean-Louis Merrien. Dyadic Hermite interpolation on a rectangular mesh. Advances in Computational Mathematics, 10(3– 4):343–365, May 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018943002601.

Dorfler:2015:NGF

[DM15] Monika Dörfler and Ewa Matusiak. Nonstationary Gabor frames — approximately dual frames and reconstruction errors. Advances in Computational Mathematics, 41(2): 293–316, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9358-z.

DalSanto:2019:HRO

[DM19a] Niccolò Dal Santo and Andrea Manzoni. Hyperreduced order models for parametrized unsteady Navier– Stokes equations on domains with variable shape. Advances in Computational Mathematics, 45(5-6):2463– 2501, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09722-9.

Dimov:2019:NUS

[DM19b] I. T. Dimov and S. Maire. A new unbiased stochastic algorithm for solving linear Fredholm equations of the second kind. Advances in Computational Mathematics, 45(3): 1499–1519, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09676-y.

DeRyck:2022:EAP

[DM22] Tim De Ryck and Siddhartha Mishra. Error analysis for physics-informed neural networks (PINNs) approximating Kolmogorov PDEs. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09985-9.

Dong:2023:SBC

[DM23] Yijun Dong and Per-Gunnar Martinsson. Simpler is better: a comparative study of randomized pivoting algorithms for CUR and interpolative decompositions. Advances in Computational Mathematics, 49(4):??, August 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
023-10061-z.

Demkowicz:2024:SAE

[DMBH24] Leszek Demkowicz, Jens M. Melenk, Jacob Badger, and Stefan Henneking. Stability analysis for electromagnetic waveguides. Part 2: non-homogeneous waveguides. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10130-x.

DeSterck:2011:TLA

[DMMS11] H. De Sterck, K. Miller, T. Manteuffel, and G. Sanders. Top-level acceleration of adaptive algebraic multilevel methods for steady-state solution to Markov chains. Advances in Computational Mathematics, 35(2–4):375– 403, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-010-9168-x.

Ding:2022:EEH

[DMS22] Qianqian Ding, Shipeng Mao, and Jiaao Sun. Error estimates of H(div)-conforming method for nonstationary magnetohydrodynamic system. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09964-0.

Dahlke:2003:RWR

[DMT03] Stephan Dahlke, Peter Maass, and Gerd Teschke. Reconstruction of wideband reflectivity densities by wavelet transforms. Advances in Computational Mathematics, 18(2–4):189–209, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/A:1021303718373.

Frames.

Droniou:2021:HDM

[DNS21] Jérome Droniou, Neela Nataraj, and Devika Shylaja. Hessian discretisation method for fourth-order semi-linear elliptic equations: applications to the von Kármán and Navier– Stokes models. Advances in Computational Mathematics, 47(2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09837-4.

Dong:2019:EEC

[Don19] Zhaonan Dong. On the exponent of exponential convergence of p-version FEM spaces. Advances in Computational Mathematics, 45(2): 757–785, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9637-1.

Dong:2020:RSO

[Don20] Jian Dong. A robust secondorder surface reconstruction for shallow water flows with a discontinuous topography and a Manning friction. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09783-1.

Douglas:1994:SRC

[Dou94] Craig C. Douglas. Some remarks on completely vectorizing point Gauss–Seidel while using the natural ordering. Advances in Computational Mathematics, 2(2): 215–222, March 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF02521108.

Dikshit:1994:WTR

[DOZ94] H. P. Dikshit, A. Ojha, and R. A. Zalik. Wachspress type rational complex planar splines of degree (3, 1). Advances in Computational Mathematics, 2(2): 235–249, March 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521110.

Delgado:2006:GBB

[DP06] Jorge Delgado and Juan Manuel Peña. On the generalized Ball bases. Advances in Computational Mathematics, 24(1–4): 263–280, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7636-x.

Delgado:2013:ERT

[DP13] Jorge Delgado and Juan Manuel Peña. On the evaluation of rational triangular Bézier surfaces and the optimal stability of the basis. Advances in Computational Mathematics, 38(4): 701–721, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9256-6.

Dahmen:1993:WAM

[DPS93] W. Dahmen, S. Prössdorf, and R. Schneider. Wavelet approximation methods for pseudodifferential equations. II. Matrix compression and fast solution. Advances in Computational Mathematics, 1(3):259–335, October 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02072014.

Duan:2020:BPB

[DQS20] Xianglong Duan, Chaoyu Quan, and Benjamin Stamm. A boundary-partition-based Voronoi diagram of d-dimensional balls: definition, properties, and applications. Advances in Computational Mathematics, 46(3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09765-3.

Dominguez:2008:DDM

[DRS08] V. Domínguez, M.-L. Rapún, and F.-J. Sayas. Dirac delta methods for Helmholtz transmission problems. Advances in Computational Mathematics, 28(2):119–139, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10444-006-9015-2.

Davydov:2000:LLI

[DS00] Oleg Davydov and Larry L.

Schumaker. Locally linearly independent bases for bivariate polynomial spline spaces. Advances in Computational Mathematics, 13(4):355–373, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016626526861.

Dominguez:2001:FAS

[DS01] Víctor Domínguez and Francisco-Javier Sayas. Full asymptotics of spline Petrov– Galerkin methods for some periodic pseudodifferential equations. Advances in Computational Mathematics, 14 (1):75–101, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016620816451.

DeMarchi:2010:SKB

[DS10] Stefano De Marchi and Robert Schaback. Stability of kernel-based interpolation. Advances in Computational Mathematics, 32(2): 155–161, February 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9093-4;

                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        008-9093-4.pdf.

Dyn:2013:DTP

[DS13] Nira Dyn and Maria Skopina. Decompositions of trigonometric polynomials with applications to multivariate subdivision schemes. Advances in Computational Mathematics, 38(2):321–349, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9239-7.

Didas:2009:CDG

[DSS09] Stephan Didas, Simon Setzer, and Gabriele Steidl. Combined l_{2} data and gradient fitting in conjunction with l_{1} regularization. Advances in Computational Mathematics, 30(1):79–99, January 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9061-4.

Dahlke:2004:CSB

[DST04] Stephan Dahlke, Gabriele Steidl, and Gerd Teschke. Coorbit spaces and Banach frames on homogeneous spaces with applications to the sphere. Advances in Computational Mathematics, 21(1–2):147– 180, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016435.
42220.fa.

DeMarchi:2005:NOD

[DSW05] Stefano De Marchi, Robert Schaback, and Holger Wendland. Near-optimal dataindependent point locations for radial basis function interpolation. Advances in Computational Mathematics, 23(3):317–330, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1829-1.

Dong:2020:ADP

[DSW20] Qiannan Dong, Shuai Su, and Jiming Wu. Analysis of the decoupled and positivitypreserving DDFV schemes for diffusion problems on polygonal meshes. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09748-4.

DeVore:1996:SRG

[DT96] R. A. DeVore and V. N.

Temlyakov. Some remarks on greedy algorithms. Advances in Computational Mathematics, 5(1):173–187, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124742.

Duan:2012:DDM

[DT12] Yuping Duan and Xue-Cheng Tai. Domain decomposition methods with graph cuts algorithms for total variation minimization. Advances in Computational Mathematics, 36(2):175–199, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9213-4.

Dung:2009:NLS

[Dũn09] Dinh Dũng. Non-linear sampling recovery based on quasiinterpolant wavelet representations. Advances in Computational Mathematics, 30 (4):375–401, May 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9074-7.

Dung:2011:OAS

[Dũn11]  Dinh Dũng.  Optimal

adaptive sampling recovery. Advances in Computational Mathematics, 34(1): 1–41, January 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10444-009-9140-9.

See erratum [Dũn12].

Dung:2012:EOA

[Dũn12] Dinh Dũng. Erratum to: Optimal adaptive sampling recovery [MR2783300]. Advances in Computational Mathematics, 36(4):605–606, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-011-9198-
z;   http://link.springer.
com/content/pdf/10.1007/
s10444-011-9198-z.pdf. See

[Dũn11].

deVilliers:2019:SMW

[dVR19] Johan de Villiers and Dinna Ranirina. Shortest multiwavelet matrix filters for refinable vector splines. Advances in Computational Mathematics, 45(5-6):3327– 3365, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09740-7.

Dubickas:2010:RES

[DX10] Artūras Dubickas and Zhiqiang Xu. Refinement equations and spline functions. Advances in Computational Mathematics, 32(1):1–23, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9079-2.

Dai:2020:LDS

[DX20] Xinjie Dai and Aiguo Xiao. Lévy-driven stochastic Volterra integral equations with doubly singular kernels: existence, uniqueness, and a fast EM method. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09780-4.

Du:2001:GRT

[DY01] Liu Du and Ningning Yan. Gradient recovery type a posteriori error estimate for finite element approximation on non-uniform meshes. Advances in Computational Mathematics, 14(2):175–193, February 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016676917360.

Didenko:2010:SRM

[DY10] Victor Didenko and Wee Ping Yeo. The spectral radius of matrix continuous refinement operators. Advances in Computational Mathematics, 33 (1):113–127, July 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-009-9124-9.

Davydov:2004:SDF

[DZ04] Oleg Davydov and Frank Zeilfelder. Scattered data fitting by direct extension of local polynomials to bivariate splines. Advances in Computational Mathematics, 21(3–4):223–271, October 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032041.
68678.fa.

Eimer:2022:IFV

[EBS22] Matthias Eimer, Raul Borsche, and Norbert Siedow. Implicit finite volume method with a posteriori limiting for transport networks. Advances in Computational Mathematics, 48(3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09939-1.

Esmaeilbeigi:2018:SDF

[ECS18] M. Esmaeilbeigi, O. Chatrabgoun, and M. Shafa. Scattered data fitting of Hermite type by a weighted meshless method. Advances in Computational Mathematics, 44 (3):673–691, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-017-9555-7.

Elasmi:2021:NSI

[EEK21] Mehdi Elasmi, Christoph Erath, and Stefan Kurz. Non-symmetric isogeometric FEM–BEM couplings. Advances in Computational Mathematics, 47(5):??, October 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09886-3.

Ephremidze:2013:ADW

[EGL13] L. Ephremidze, A. Gamkrelidze, and E. Lagvilava. An approximation of Daubechies wavelet matrices by perfect reconstruction filter banks with rational coefficients. Advances in Computational Mathematics, 38(1):147–158, January 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9232-1.

Escribano:2013:HMR

[EGST13] Carmen Escribano, Antonio Giraldo, M. Asunción Sastre, and Emilio Torrano. The Hessenberg matrix and the Riemann mapping function. Advances in Computational Mathematics, 39(3–4):525– 545, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9291-y.

Ehring:2024:HKS

[EH24] Tobias Ehring and Bernard Haasdonk. Hermite kernel surrogates for the value function of high-dimensional nonlinear optimal control problems. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10128-5.

Ehrich:2000:EWC

[Ehr00] Sven Ehrich. On the estimation of wavelet coefficients. Advances in Computational Mathematics, 13 (2):105–129, May 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018998009387.

Elschner:2002:FES

[EHS02] Johannes Elschner, Rainer Hinder, and Gunther Schmidt. Finite element solution of conical diffraction problems. Advances in Computational Mathematics, 16(2–3):139– 156, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.

1023/A:1014456026778. Mod-

eling and computation in optics and electromagnetics.

Ezquerro:2017:SIC

[EHV17] J. A. Ezquerro and M. A. Hernández-Verón. A study of the influence of center conditions on the domain of parameters of Newton’s method by using recurrence relations. Advances in Computational Mathematics, 43(5):1103– 1129, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9518-z.

Enenkel:1997:DDI

[EJ97] Robert F. Enenkel and Kenneth R. Jackson. DIMSEMsdiagonally implicit singleeigenvalue methods for the numerical solution of stiff ODEs on parallel computers. Advances in Computational Mathematics, 7(1–2): 97–133, March 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018986500842. Par-

allel methods for ODEs.

Eidnes:2018:AEP

[EOR18] Sølve Eidnes, Brynjulf Owren, and Torbjørn Ringholm. Adaptive energy preserving methods for partial differential equations. Advances in Computational Mathematics, 44(3):815–839, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-017-9562-
8.

Evgeniou:2000:RNS

[EPP00] Theodoros Evgeniou, Massimiliano Pontil, and Tomaso Poggio. Regularization networks and support vector machines. Advances in Computational Mathematics, 13 (1):1–50, April 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018946025316.

Engelborghs:1999:NCS

[ER99] Koen Engelborghs and Dirk Roose. Numerical computation of stability and detection of Hopf bifurcations of steady state solutions of delay differential equations. Advances in Computational Mathematics, 10(3– 4):271–289, May 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018986817622.

Eshghi:2021:EEM

[ER21] Nasim Eshghi and Lothar Reichel. Estimating the error in matrix function approximations. Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09882-7.

Eskew:2023:NAP

[ES23] Sarah Locke Eskew and John R. Singler. A new approach to proper orthogonal decomposition with difference quotients. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10011-9.

Elgindy:2013:OGO

[ESM13] Kareem T. Elgindy and Kate A. Smith-Miles. On the optimization of Gegenbauer operational matrix of integration. Advances in Computational Mathematics, 39(3–4): 511–524, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-012-9289-5.

Eigel:2019:VMC

[ESTW19] Martin Eigel, Reinhold Schneider, Philipp Trunschke, and Sebastian Wolf. Variational Monte Carlo-bridging concepts of machine learning and high-dimensional partial differential equations. Advances in Computational Mathematics, 45(5-6):2503– 2532, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09723-8.

Epshteyn:2020:DPM

[EX20] Yekaterina Epshteyn and Qing Xia. Difference potentials method for models with dynamic boundary conditions and bulk-surface problems. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09798-8.

Falcon:2019:CAG

[FÁG19] Raúl M. Falcón, Víctor Álvarez, and Félix Gudiel. A computational algebraic geometry approach to analyze pseudo-random sequences based on Latin squares. Advances in Computational Mathematics, 45(4): 1769–1792, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9654-0.

Farouki:2002:HIR

[FaKS02] Rida T. Farouki, Mohammad al Kandari, and Takis Sakkalis. Hermite interpolation by rotation-invariant spatial Pythagorean-hodograph curves. Advances in Computational Mathematics, 17 (4):369–383, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1016280811626.

Fryklund:2022:AKS

[FaKT22] Fredrik Fryklund, Ludvig af Klinteberg, and Anna-Karin Tornberg. An adaptive kernel-split quadrature method for parameterdependent layer potentials. Advances in Computational Mathematics, 48(2):??, April 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
022-09927-5.

Farouki:2010:QHM

[Far10] Rida T. Farouki. Quaternion and Hopf map characterizations for the existence of rational rotation-minimizing frames on quintic space curves. Advances in Computational Mathematics, 33(3): 331–348, October 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9138-3;
 http://link.springer.com/
content/pdf/10.1007/s10444-
009-9138-3.pdf.

Fasshauer:1999:HIR

[Fas99a] Gregory E. Fasshauer. Hermite interpolation with radial basis functions on spheres. Advances in Computational Mathematics, 10(1):81–96, January 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018914229009.

Fasshauer:1999:SDE

[Fas99b] Gregory E. Fasshauer. Solving differential equations with radial basis functions: multilevel methods and smoothing. Advances in Computational Mathematics, 11(2–3):139– 159, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018919824891. Ra-

dial basis functions and their applications.

Floater:2013:SCM

[FBCR13] M. Floater, C. Beccari, T. Cashman, and L. Romani. A smoothness criterion for monotonicity-preserving subdivision. Advances in Computational Mathematics, 39 (1):193–204, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9275-y.

Fan:2019:NCR

[FBDH19] Ya-Ru Fan, Alessandro Buccini, Marco Donatelli, and Ting-Zhu Huang. A nonconvex regularization approach for compressive sensing. Advances in Computational Mathematics, 45(2): 563–588, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9627-3.

Fischer:2015:APM

[FE15] Michael Fischer and Peter Eberhard. Application of parametric model reduction with matrix interpolation for simulation of moving loads in elastic multibody systems. Advances in Computational Mathematics, 41(5):1049– 1072, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9379-7.

Farhloul:1995:MFE

[FF95] Mohamed Farhloul and Michel Fortin. A mixed finite element for the Stokes problem using quadrilateral elements. Advances in Computational Mathematics, 3(1–2): 101–113, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF02431998.

Fornberg:2005:ARB

[FF05] Bengt Fornberg and Natasha Flyer. Accuracy of radial basis function interpolation and derivative approximations on 1-D infinite grids. Advances in Computational Mathematics, 23 (1–2):5–20, July 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1812-x.

Farouki:2017:CCS

[FGG^{+}17] Rida T. Farouki, Graziano Gentili, Carlotta Giannelli, Alessandra Sestini, and Caterina Stoppato. A comprehensive characterization of the set of polynomial curves with rational rotation-minimizing frames. Advances in Computational Mathematics, 43(1): 1–24, February 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9473-0.

Farouki:2019:MPU

[FGMS19] Rida T. Farouki, Graziano Gentili, Hwan Pyo Moon, and Caterina Stoppato. Minkowski products of unit quaternion sets. Advances in Computational Mathematics, 45(3): 1607–1629, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09687-9.

Feichtinger:2014:ADG

[FGO14] Hans G. Feichtinger, Anna Grybos, and Darian M. Onchis. Approximate dual Gabor atoms via the adjoint lattice method. Advances in Computational Mathematics, 40(3):651–665, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9324-
                                        1.

Farouki:2013:ISD

[FGS13] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. An interpolation scheme for designing rational rotation-minimizing camera motions. Advances in Computational Mathematics, 38 (1):63–82, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9226-z.

Farouki:2016:LMP

[FGS16] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. Local modification of Pythagorean-hodograph quintic spline curves using the B-spline form. Advances in Computational Mathematics, 42(1):199–225, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9419-y.

Farouki:2005:SEE

[FH05] Rida T. Farouki and Chang Yong Han. Solution of elementary equations in the Minkowski geometric algebra of complex sets. Advances in Computational Mathematics, 22(4): 301–323, May 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-2605-3.

Fuhrer:2018:WPM

[FH18] Thomas Führer and Norbert Heuer. A wirebasket preconditioner for the mortar boundary element method. Advances in Computational Mathematics, 44(1): 23–49, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9534-z.

Floater:2020:CSM

[FH20] Michael S. Floater and Kaibo Hu. A characterization of supersmoothness of multivariate splines. Advances in Computational Mathematics, 46(5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09813-y.

Floater:2006:GCB

[FHK06] Michael S. Floater, Kai Hormann, and Géza Kós. A general construction of barycentric coordinates over convex polygons. Advances in Computational Mathematics, 24(1–4):311–331, January 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7611-6.

Franke:1994:LSS

[FHN94] Richard Franke, Hans Hagen, and Gregory M. Nielson. Least squares surface approximation to scattered data using multiquadratic functions. Advances in Computational Mathematics, 2(1): 81–99, January 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02519037.

Fink:2019:HOA

[Fin19] Thomas Fink. Higher order analysis of the geometry of singularities using the Taylorlet transform. Advances in Computational Mathematics, 45(2): 675–705, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9635-3.

Fasshauer:1999:MAA

[FJ99] Gregory E. Fasshauer and Joseph W. Jerome. Multistep approximation algorithms: improved convergence rates through postconditioning with smoothing kernels. Advances in Computational Mathematics, 10(1): 1–27, January 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018962112170.

Fairweather:1998:MFS

[FK98] Graeme Fairweather and Andreas Karageorghis. The method of fundamental solutions for elliptic boundary value problems. Advances in Computational Mathematics, 9(1–2):69–95, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018981221740. Nu-

merical treatment of boundary integral equations.

Floater:2010:IWM

[FK10] Michael S. Floater and Jiří Kosinka. On the injectivity of Wachspress and mean value mappings between convex polygons. Advances in Computational Mathematics, 32(2):163–174, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9098-z.

Fairweather:2005:MDM

[FKS05] Graeme Fairweather, Andreas Karageorghis, and Yiorgos-Sokratis Smyrlis. A matrix decomposition MFS algorithm for axisymmetric biharmonic problems. Advances in Computational Mathematics, 23(1–2):55– 71, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1808-6.

Fort:2019:FVM

[FKT^{+}19] J. Fořt, J. Karel, D. Trdlička, F. Benkhaldoun, I. Kissami, J.-B. Montavon, K. Hassouni, and J. Zs. Mezei. Finite volume methods for numerical simulation of the discharge motion described by different physical models. Advances in Computational Mathematics, 45(4): 2163–2189, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09706-9.

Fryklund:2020:IEB

[FKT20] Fredrik Fryklund, Mary Catherine A. Kropinski, and Anna-Karin Tornberg. An integral equation-based numerical method for the forced heat equation on complex domains. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09804-z.

Floater:2000:ACD

[FL00] Michael S. Floater and Tom Lyche. Asymptotic convergence of degree-raising. Advances in Computational Mathematics, 12(2–3):175– 187, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018913118047.

Fu:2023:DSD

[FL23] Jun-Liang Fu and Jijun Liu.

On the determination of the spatial-dependent potential coefficient in a linear pseudoparabolic equation. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10023-5.

Flinth:2018:SBD

[Fli18] Axel Flinth. Sparse blind deconvolution and demixing through ℓ_{1}, 2-minimization. Advances in Computational Mathematics, 44(1):1–21, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9533-0.

Lin:2005:EHT

[fLL05] Jia fu Lin and Qun Lin. Extrapolation of the Hood– Taylor elements for the Stokes problem. Advances in Computational Mathematics, 22(2):115–123, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-004-1089-
                                        0.

Li:2021:NMP

[fLLfDX21] Jiao fen Li, Wen Li, Xue feng Duan, and Mingqing Xiao. Newton’s method for the parameterized generalized eigenvalue problem with nonsquare matrix pencils. Advances in Computational Mathematics, 47 (2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09855-w.

Fairweather:1996:GMS

[FLM96] G. Fairweather and J. C.

López-Marcos. Galerkin methods for a semilinear parabolic problem with nonlocal boundary conditions. Advances in Computational Mathematics, 6(1):243–262, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127706.  John

Crank 80th birthday special issue.

Floater:1994:WCC

[Flo94]  Michael S. Floater.  A

weak condition for the convexity of tensor-product Bézier and B-spline surfaces. Advances in Computational Mathematics, 2(1): 67–80, January 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02519036.

Floater:1996:ACA

[Flo96a] Michael S. Floater. An analysis of cubic approximation schemes for conic sections. Advances in Computational Mathematics, 5(1):361–379, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124751.

Floater:1996:ZCB

[Flo96b]  Michael S. Floater.  On

zero curves of bivariate polynomials. Advances in Computational Mathematics, 5(1):399–415, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124753.

Freeden:2004:OZT

[FM04] W. Freeden and V. Michel. Orthogonal zonal, tesseral and sectorial wavelets on the sphere for the analysis of satellite data. Advances in Computational Mathematics, 21(1–2):181– 217, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/B:ACOM.0000016429.
                                        07498.77.

Fornberg:2014:SHB

[FM14] Bengt Fornberg and Jordan M. Martel. On spherical harmonics based numerical quadrature over the surface of a sphere. Advances in Computational Mathematics, 40(5–6):1169–1184, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9346-3.

Farouki:2018:RFM

[FM18] Rida T. Farouki and Hwan Pyo Moon. Rational frames of minimal twist along space curves under specified boundary conditions. Advances in Computational Mathematics, 44(5):1627–1650, October 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9599-3.

Farazandeh:2021:RRI

[FM21] Elham Farazandeh and Davoud Mirzaei. A rational RBF interpolation with conditionally positive definite kernels. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09900-8.

Faustmann:2022:MHM

[FMP22] Markus Faustmann, Jens Markus Melenk, and Maryam Parvizi. H-matrix approximability of inverses of FEM matrices for the time-harmonic Maxwell equations. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09965-z.

Fickus:2013:CAS

[FMPS13] Matthew Fickus, Dustin G. Mixon, Miriam J. Poteet, and Nate Strawn. Constructing all self-adjoint matrices with prescribed spectrum and diagonal. Advances in Computational Mathematics, 39(3–4):585– 609, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9298-z.

Farouki:2000:AMP

[FMR00] Rida T. Farouki, Hwan Pyo Moon, and Bahram Ravani. Algorithms for Minkowski products and implicitlydefined complex sets. Advances in Computational Mathematics, 13(3):199–229, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1018910412112.

Foucart:2023:FRP

[Fou23] Simon Foucart. Full recovery from point values: an optimal algorithm for Chebyshev approximability prior. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10063-x.

Feichtinger:2014:CAD

[FOW14] Hans G. Feichtinger, Darian M. Onchis, and Christoph Wiesmeyr. Construction of approximate dual wavelet frames. Advances in Computational Mathematics, 40(1): 273–282, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9307-2.

Floater:1998:TPM

[FP98] Michael S. Floater and J. M.

Peña. Tensor-product monotonicity preservation. Advances in Computational Mathematics, 9(3–4):353– 362, November 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018906027191.

Faul:1999:PCI

[FP99] A. C. Faul and M. J. D.

Powell. Proof of convergence of an iterative technique for thin plate spline interpolation in two dimensions. Advances in Computational Mathematics, 11(2–3):183– 192, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018923925800. Ra-

dial basis functions and their applications.

Feichtinger:2012:ARB

[FPR^{+}12] Hans G. Feichtinger, José C. Príncipe, José Luis Romero, Alexander Singh Alvarado, and Gino Angelo Velasco. Approximate reconstruction of bandlimited functions for the integrate and fire sampler. Advances in Computational Mathematics, 36(1): 67–78, January 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9180-9.

Floater:2006:CCM

[FPT06] Michael S. Floater and Valérie Pham-Trong. Convex combination maps over triangulations, tilings, and tetrahedral meshes. Advances in Computational Mathematics, 25(4):347–356, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7620-5.

Fu:2010:NPO

[FQ10] Chu-Li Fu and Zhi Qian. Numerical pseudodifferential operator and Fourier regularization. Advances in Computational Mathematics, 33 (4):449–470, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/

10.1007/s10444-009-9136-
5.

Fermo:2010:NMF

[FR10] L. Fermo and M. G. Russo.

Numerical methods for Fredholm integral equations with singular right-hand sides. Advances in Computational Mathematics, 33(3):305–330, October 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9137-4.

Franssens:1999:NNP

[Fra99]  Ghislain Franssens.  A

new non-polynomial univariate interpolation formula of Hermite type. Advances in Computational Mathematics, 10(3–4):367– 388, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018947103510.

Fornasier:2009:AJS

[FRT09] Massimo Fornasier, Ronny Ramlau, and Gerd Teschke. The application of joint sparsity and total variation minimization algorithms to a real-life art restoration problem. Advances in Computational Mathematics, 31(1–3): 157–184, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-008-9103-6.

Farouki:1994:PHS

[FS94] Rida T. Farouki and Takis Sakkalis. Pythagoreanhodograph space curves. Advances in Computational Mathematics, 2(1):41–66, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02519035.

Franke:1998:COE

[FS98] Carsten Franke and Robert Schaback. Convergence order estimates of meshless collocation methods using radial basis functions. Advances in Computational Mathematics, 8(4):381–399, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018916902176.

Feichtinger:2007:SCI

[FS07] Hans G. Feichtinger and Wenchang Sun. Sufficient conditions for irregular Gabor frames. Advances in Computational Mathematics, 26(4):403–430, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-004-7210-

Fuhr:2008:SEV

[Füh08] H. Führ. Simultaneous estimates for vector-valued Gabor frames of Hermite functions. Advances in Computational Mathematics, 29 (4):357–373, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9053-
4.

Fuhr:2016:VMC

[Füh16]  Hartmut Führ.  Vanish-

ing moment conditions for wavelet atoms in higher dimensions. Advances in Computational Mathematics, 42 (1):127–153, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9414-
3.

Fuselier:2008:EIS

[Fus08a] Edward J. Fuselier. Erratum: “Improved stability estimates and a characterization of the native space for matrix-valued RBFs” [Adv. Comput. Math. 29 (2008), no. 3, 269–290; MR2438345]. Advances in Computational Mathematics, 29(3):311–313, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1007/s10444-008-9091-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        008-9091-6.pdf. See [Fus08b].

Fuselier:2008:ISE

[Fus08b] Edward J. Fuselier. Improved stability estimates and a characterization of the native space for matrix-valued RBFs. Advances in Computational Mathematics, 29(3): 269–290, October 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9046-3.

See erratum [Fus08a].

Freeden:1996:SWT

[FW96] Willi Freeden and Ulrich Windheuser. Spherical wavelet transform and its discretization. Advances in Computational Mathematics, 5(1):51–94, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124735.

Fuselier:2015:OPD

[FW15] Edward Fuselier and Grady B. Wright. Order-preserving derivative approximation with periodic radial basis functions. Advances in Computational Mathematics, 41(1): 23–53, February 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9348-1.

Feng:2006:GFT

[FWW06] De-Jun Feng, Long Wang, and Yang Wang. Generation of finite tight frames by Householder transformations. Advances in Computational Mathematics, 24(1–4): 297–309, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7637-9.

Ford:1996:DCQ

[FXZ96] William F. Ford, Yuesheng Xu, and Yunhe Zhao. Derivative corrections for quadrature formulas. Advances in Computational Mathematics, 6(1):139–157, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127701.

Fasshauer:2013:RKS

[FY13] Gregory E. Fasshauer and Qi Ye. Reproducing kernels of Sobolev spaces via a green kernel approach with differential operators and boundary operators. Advances in Computational Mathematics, 38(4):891–921, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9264-
                                        6.

Feng:2008:AFE

[FYL08] Tao Feng, Ningning Yan, and Wenbin Liu. Adaptive finite element methods for the identification of distributed parameters in elliptic equation. Advances in Computational Mathematics, 29 (1):27–53, July 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9035-6.

Fang:2012:RIB

[FZ12] Weifu Fang and Suxing Zeng. Recovery of an interface from boundary measurement in an elliptic differential equation. Advances in Computational Mathematics, 36(2): 201–220, February 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9204-5.

Furst:2019:NSU

[FZ19] Jirí Fürst and Zdenek Zák. Numerical simulation of unsteady flows through a radial turbine. Advances in Computational Mathematics, 45(4): 1939–1952, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10444-019-09670-4.

Fu:2021:ESS

[FZLZ21] Hongfei Fu, Chen Zhu, Xueting Liang, and Bingyin Zhang. Efficient spatial second-/fourth-order finite difference ADI methods for multi-dimensional variableorder time-fractional diffusion equations. Advances in Computational Mathematics, 47(4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09881-8.

Garcia-Archilla:2020:EAF

[GAN20] Bosco García-Archilla and Julia Novo. Error analysis of fully discrete mixed finite element data assimilation schemes for the Navier– Stokes equations. Advances in Computational Mathematics, 46(4):??, August 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-020-09806-
x.

Gataric:2021:HRS

[Gat21]  Milana Gataric.  High-

resolution signal recovery via generalized sampling and functional principal component analysis. Advances in Computational Mathematics, 47(6):??, December 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09908-0.

Goyal:2024:DSH

[GDB24] Pawan Goyal, Igor Pontes Duff, and Peter Benner. Dominant subspaces of highfidelity polynomial structured parametric dynamical systems and model reduction. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10133-8.

Gomes:2021:ATT

[GDMS21] Anna Karina Fontes Gomes, Margarete Oliveira Domingues, Odim Mendes, and Kai Schneider. Adaptive two- and three-dimensional multiresolution computations of resistive magnetohydrodynamics. Advances in Computational Mathematics, 47 (2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09845-y.

Gerot:2019:EFB

[Gér19]  Cédric Gérot.  Elemen-

tary factorisation of Box spline subdivision. Advances in Computational Mathematics, 45(1):153–171, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9612-x.

Goh:2006:ITC

[GG06] Say Song Goh and Tim N. T. Goodman. Inequalities on time-concentrated or frequency-concentrated functions. Advances in Computational Mathematics, 24(1–4): 333–351, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4145-x.

Gimbutas:2016:FMM

[GG16] Zydrunas Gimbutas and Leslie Greengard. A fast multipole method for the evaluation of elastostatic fields in a half-space with zero normal stress. Advances in Computational Mathematics, 42(1): 175–198, February 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9416-1.

Galan-Garcia:2019:ECI

[GGAVGG^{+}19] José L. Galán-García, Gabriel Aguilera-Venegas, María Á. Galán-García, Pedro Rodríguez-Cielos, Iván Atencia-McKillop, Yolanda Padilla-Domínguez, and Ricardo Rodríguez-Cielos. Enhancing Cas improper integrals computations using extensions of the residue theorem. Advances in Computational Mathematics, 45(4):1825– 1841, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-09660-y.

Goh:2007:CPR

[GGL07] Say Song Goh, Tim N. T. Goodman, and S. L. Lee. Causality properties of refinable functions and sequences. Advances in Computational Mathematics, 26(1–3):231– 250, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-8007-3.

Guillen-Gonzalez:2021:CMS

[GGRBRG21] F. Guillén-González, M. A. Rodríguez-Bellido, and D. A. Rueda-Gómez. A chemorepulsion model with superlinear production: analysis of the continuous problem and two approximately positive and energy-stable schemes. Advances in Computational Mathematics, 47 (6):??, December 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-021-09907-1.

Gosea:2024:SBF

[GGW24] Ion Victor Gosea, Serkan Gugercin, and Steffen W. R. Werner. Structured barycentric forms for interpolationbased data-driven reduced modeling of second-order systems. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10118-7.

Goktas:1999:ACG

[GH99] Ünal Göktaş and Willy Hereman. Algorithmic computation of generalized symmetries of nonlinear evolution and lattice equations. Advances in Computational Mathematics, 11 (1):55–80, July 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018955405327.

Gabardo:2003:FAM

[GH03] Jean-Pierre Gabardo and Deguang Han. Frames associated with measurable spaces. Advances in Computational Mathematics, 18(2–4):127– 147, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021312429186. Frames.

Guo:2006:OCV

[GH06] Benqi Guo and Norbert Heuer. The optimal convergence of the h–p version of the boundary element method with quasiuniform meshes for elliptic problems on polygonal domains. Advances in Computational Mathematics, 24(1–4):353– 374, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7618-z.

Guglielmi:2008:CBP

[GH08] Nicola Guglielmi and Ernst Hairer. Computing breaking points in implicit delay differential equations. Advances in Computational Mathematics, 29(3):229–247, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9044-5.

Grohs:2016:SAL

[GH16] P. Grohs and S. Hosseini. ε-subgradient algorithms for locally Lipschitz functions on Riemannian manifolds. Advances in Computational Mathematics, 42 (2):333–360, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9426-z.

Grassle:2018:PRO

[GH18] Carmen Gräßle and Michael Hinze. POD reduced-order modeling for evolution equations utilizing arbitrary finite element discretizations. Advances in Computational Mathematics, 44(6):1941– 1978, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9620-x.

Garrigos:2012:DFO

[GHdN12] Gustavo Garrigós, Eugenio Hernández, and Maria de Natividade. Democracy functions and optimal embeddings for approximation spaces. Advances in Computational Mathematics, 37(2): 255–283, August 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9197-0.

Grohs:2023:STE

[GHJZ23] Philipp Grohs, Fabian Hornung, Arnulf Jentzen, and Philipp Zimmermann. Spacetime error estimates for deep neural network approximations for differential equations. Advances in Computational Mathematics, 49 (1):??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09970-2.

Garcke:2018:PFA

[GHKL18] Harald Garcke, Michael Hinze, Christian Kahle, and Kei Fong Lam. A phase field approach to shape optimization in Navier–Stokes flow with integral state constraints. Advances in Computational Mathematics, 44(5):1345–1383, October 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9586-8.

Grassle:2019:PMO

[GHLU19] Carmen Gräßle, Michael Hinze, Jens Lang, and Sebastian Ullmann. POD model order reduction with spaceadapted snapshots for incompressible flows. Advances in Computational Mathematics, 45(5-6):2401–2428, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09716-7.

Giani:2020:ACD

[Gia20] Stefano Giani. An adaptive composite discontinuous Galerkin method for elliptic problems on complicated domains with discontinuous coefficients. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09759-1.

Gross:2021:DAC

[GIKV21] Craig Gross, Mark A. Iwen, Lutz Kämmerer, and Toni Volkmer. A deterministic algorithm for constructing multiple rank-1 lattices of near-optimal size. Advances in Computational Mathematics, 47(6):??, December 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09916-0.

Gillman:2017:IET

[Gil17] Adrianna Gillman. An integral equation technique for scattering problems with mixed boundary conditions. Advances in Computational Mathematics, 43(2):351–364, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-016-9488-
6.

Gray:2019:BIA

[GJMY19] L. J. Gray, Jas Jakowski, M. N. J. Moore, and Wenjing Ye. Boundary integral analysis for non-homogeneous, incompressible Stokes flows. Advances in Computational Mathematics, 45(3):1729– 1734, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09699-5.

Giannelli:2014:SSB

[GJS14] Carlotta Giannelli, Bert Jüttler, and Hendrik Speleers. Strongly stable bases for adaptively refined multilevel spline spaces. Advances in Computational Mathematics, 40(2):459–490, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9315-
                                        2.

Greif:2022:LRE

[GJU22] Constantin Greif, Philipp Junk, and Karsten Urban. Linear/Ridge expansions: enhancing linear approximations by ridge functions. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09936-4.

Greengard:2020:AEU

[GJW20] Leslie Greengard, Shidong Jiang, and Jun Wang. On the accurate evaluation of unsteady Stokes layer potentials in moving two-dimensional geometries. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09760-8.

Ghoreishi:2019:PRC

[GK19] F. Ghoreishi and Changiz Goli Keshavarzi. Preconditioning RBF collocation method by adapted finite difference matrix. Advances in Computational Mathematics, 45(5-6):3293–3325, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09738-1.

Graf:2023:OBS

[GKKS23] Olga Graf, Felix Krahmer, and Sara Krause-Solberg. One-bit Sigma–Delta modulation on the circle. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10032-4.

Grohs:2020:AMS

[GKM^{+}20] Philipp Grohs, Gitta Kutyniok, Jackie Ma, Philipp Petersen, and Mones Raslan. Anisotropic multiscale systems on bounded domains. Advances in Computational Mathematics, 46(2):??, April 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        020-09784-0.

Geppert:2019:SPF

[GKS19] Jakob Geppert, Felix Krahmer, and Dominik Stöger. Sparse power factorization: balancing peakiness and sample complexity. Advances in Computational Mathematics, 45(3):1711– 1728, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09698-6.

Guglielmi:2021:CLR

[GKS21] Nicola Guglielmi, Daniel Kressner, and Carmela Scalone. Computing low-rank rightmost eigenpairs of a class of matrix-valued linear operators. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09895-2.

Guo:2013:ORX

[GL13] Kanghui Guo and Demetrio Labate. Optimal recovery of 3D X-ray tomographic data via shearlet decomposition. Advances in Computational Mathematics, 39(2): 227–255, August 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9276-x.

Guo:2017:MAE

[GL17] Kanghui Guo and Demetrio Labate. Microlocal analysis of edge flatness through directional multiscale representations. Advances in Computational Mathematics, 43(2): 295–318, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9486-8.

Ganesh:2023:HOA

[GL23] M. Ganesh and F. Le Louër. A high-order algorithm for time-domain scattering in three dimensions. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10033-3.

Girometti:2023:TID

[GLM23] Laura Girometti, Alessandro Lanza, and Serena Morigi. Ternary image decomposition with automatic parameter selection via auto- and crosscorrelation. Advances in Computational Mathematics, 49(1):??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-10000-4.

Gracia:2010:CSS

[GLO10] J. L. Gracia, F. J. Lisbona, and E. O’Riordan. A coupled system of singularly perturbed parabolic reaction–diffusion equations. Advances in Computational Mathematics, 32(1):43–61, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9086-3.

Gerbeau:2015:ROM

[GLS15] Jean-Frédéric Gerbeau, Damiano Lombardi, and Elisa Schenone. Reduced order model in cardiac electrophysiology with approximated Lax pairs. Advances in Computational Mathematics, 41(5):1103– 1130, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9393-9.

Goodman:1993:WBS

[GLT93] T. N. T. Goodman, S. L. Lee, and W. S. Tang. Wavelet bases for a set of commuting unitary operators. Advances in Computational Mathematics, 1(1):109–126, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02070823.

Gong:2023:PEE

[GLX23] Wei Gong, Dongdong Liang, and Xiaoping Xie. Pointwise error estimates for linear finite element approximation to elliptic Dirichlet problems in smooth domains. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10017-3.

Gillman:2014:ACS

[GM14] Adrianna Gillman and Per-Gunnar Martinsson. An O(N) algorithm for constructing the solution operator to 2D elliptic boundary value problems in the absence of body loads. Advances in Computational Mathematics, 40(4):773–796, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9326-z.

Ghelardoni:2017:MMF

[GM17] Paolo Ghelardoni and Cecilia Magherini. A matrix method for fractional Sturm–Liouville problems on bounded domain. Advances in Computational Mathematics, 43(6):1377–1401, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9529-9.

Gopal:2022:AHO

[GM22] Abinand Gopal and Per-Gunnar Martinsson. An accelerated, high-order accurate direct solver for the Lippmann–Schwinger equation for acoustic scattering in the plane. Advances in Computational Mathematics, 48(4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09963-1.

Gomez:2024:STD

[GM24] Sergio Gómez and Andrea Moiola. A space-time DG method for the Schrödinger equation with variable potential. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10444-024-10108-9.

Graves-Morris:1997:CSI

[GMB97] P. R. Graves-Morris and B. Beckermann. The compass (star) identity for vector-valued rational interpolants. Advances in Computational Mathematics, 7(3): 279–294, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018951020406.

Giesselmann:2023:PEE

[GMBS23] Jan Giesselmann, Elena Mäder-Baumdicker, and David Jakob Stonner. A posteriori error estimates for wave maps into spheres. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10051-1.

Gatica:2009:NCM

[GMM09] Gabriel N. Gatica, Antonio Márquez, and Salim Meddahi. A new coupling of mixed finite element and boundary element methods for an exterior Helmholtz problem in the plane. Advances in Computational Mathematics, 30(3): 281–301, April 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9068-5.

Goodman:1997:SFL

[GMRS97] Tim N. T. Goodman, Charles A. Micchelli, Giuseppe Rodriguez, and Sebastiano Seatzu. Spectral factorization of Laurent polynomials. Advances in Computational Mathematics, 7(4): 429–454, August 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018915407202.

Guessab:2012:CNF

[GMS12] Allal Guessab, María Moncayo, and Gerhard Schmeisser. A class of nonlinear fourpoint subdivision schemes. Advances in Computational Mathematics, 37(2):151–190, August 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9199-y.

Gribonval:2001:AWG

[GN01] R. Gribonval and M. Nielsen. Approximate weak greedy algorithms. Advances in Computational Mathematics, 14 (4):361–378, May 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1012255021470.

Gribonval:2008:BSR

[GN08] Rémi Gribonval and Morten Nielsen. Beyond sparsity: recovering structured representations by ℓ^{1} minimization and greedy algorithms. Advances in Computational Mathematics, 28(1): 23–41, January 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9009-5.

Gupta:2013:PEE

[GN13] Nupur Gupta and Neela Nataraj. A posteriori error estimates for an optimal control problem of laser surface hardening of steel. Advances in Computational Mathematics, 39 (1):69–99, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9270-8.

Gnaneshwar:2007:DKM

[Gna07]  N. Gnaneshwar.  A de-

generate kernel method for eigenvalue problems of compact integral operators. Advances in Computational Mathematics, 27(3):339–354, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9005-9.

Griebel:1995:TPT

[GO95] M. Griebel and P. Oswald. Tensor product type subspace splittings and multilevel iterative methods for anisotropic problems. Advances in Computational Mathematics, 4(1–2):171– 206, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02123478.

Goodman:2005:SPI

[GO05] T. N. T. Goodman and B. H. Ong. Shape-preserving interpolation by splines using vector subdivision. Advances in Computational Mathematics, 22(1):49–77, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-1174-z.

Gomes:1995:CEW

[Gom95]  Sônia M. Gomes.  Con-

vergence estimates for the wavelet-Galerkin method: superconvergence at the node points. Advances in Computational Mathematics, 4(1–2):261–282, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF03177516.

Gonsor:1993:SAN

[Gon93] Daniel E. Gonsor. Subdivision algorithms with nonnegative masks generally converge. Advances in Computational Mathematics, 1(2): 215–221, June 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02071386.

Guo:2003:P

[GOZ03] Benqi Guo, Hae-Soo Oh, and Zhimin Zhang. Preface. Advances in Computational Mathematics, 19(1– 3):1–2, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022844220807.

Gasca:1995:CAS

[GP95] M. Gasca and J. M. Peña.

On the characterization of almost strictly totally positive matrices. Advances in Computational Mathematics, 3(3): 239–250, April 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02432001.

Garrappa:2013:EGM

[GP13] Roberto Garrappa and Marina Popolizio. Evaluation of generalized Mittag-Leffler functions on the real line. Advances in Computational Mathematics, 39(1): 205–225, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9274-z.

Grochenig:2014:URT

[GP14] Karlheinz Gröchenig and Elmar Pauwels. Uniqueness and reconstruction theorems for pseudodifferential operators with a bandlimited Kohn–Nirenberg symbol. Advances in Computational Mathematics, 40(1): 49–63, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9297-0.

Gosea:2018:BTL

[GPAF18] Ion Victor Gosea, Mihaly Petreczky, Athanasios C. Antoulas, and Christophe Fiter. Balanced truncation for linear switched systems. Advances in Computational Mathematics, 44(6):1845– 1886, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9610-z.

Guy:2015:GMI

[GPG15] Robert D. Guy, Bobby Philip, and Boyce E. Griffith. Geometric multigrid for an implicit-time immersed boundary method. Advances in Computational Mathematics, 41(3):635–662, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-014-9380-
1.

Garcia-Puente:2010:LPP

[GPS10] Luis David Garcia-Puente and Frank Sottile. Linear precision for parametric patches. Advances in Computational Mathematics, 33(2): 191–214, August 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9126-7.

Graham:2021:AHP

[GPS21] I. G. Graham, O. R. Pembery, and E. A. Spence. Analysis of a Helmholtz preconditioning problem motivated by uncertainty quantification. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09889-0.

Gerhards:2017:PCS

[GPT17] Christian Gerhards, Sergiy Pereverzyev, Jr., and Pavlo Tkachenko. A parameter choice strategy for the inversion of multiple observations. Advances in Computational Mathematics, 43(1): 101–112, February 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9477-9.

Guichard:2022:WPN

[GQ22] Cindy Guichard and El Houssaine Quenjel. Weighted positive nonlinear finite volume method for dominated anisotropic diffusive equations. Advances in Computational Mathematics, 48(6): ??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09995-7.

Gao:2024:OCM

[GQL24] Xinyue Gao, Yi Qin, and Jian Li. Optimally convergent mixed finite element methods for the time-dependent 2D/ 3D stochastic closed-loop geothermal system with multiplicative noise. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10122-x.

Greengard:2019:AEI

[GR19] Philip Greengard and Vladimir Rokhlin. An algorithm for the evaluation of the incomplete gamma function. Advances in Computational Mathematics, 45(1):23–49, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9604-x.

Graf:2012:UAS

[Grä12] Manuel Gräf. A unified approach to scattered data approximation on S^{3} and SO(3). Advances in Computational Mathematics, 37(3): 379–392, October 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9214-3.

Gillette:2012:EEG

[GRB12] Andrew Gillette, Alexander Rand, and Chandrajit Bajaj. Error estimates for generalized barycentric interpolation. Advances in Computational Mathematics, 37(3): 417–439, October 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9218-z.

Gomes:2019:TCA

[GRdSA19] Daniele C. R. Gomes, Mauro A.

Rincon, Maria Darci G. da Silva, and Gladson O. Antunes. Theoretical and computational analysis of a nonlinear Schrödinger problem with moving boundary. Advances in Computational Mathematics, 45(2): 981–1004, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9643-3.

Grochenig:2003:LFF

[Grö03] Karlheinz Gröchenig. Localized frames are finite unions of Riesz sequences. Advances in Computational Mathematics, 18(2–4):149– 157, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1021368609918. Frames.

Grohs:2013:RFD

[Gro13] Philipp Grohs. Refinable functions for dilation families. Advances in Computational Mathematics, 38(3): 531–561, April 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9248-6.

Gemignani:2022:BLP

[GRV22] Luca Gemignani, Lucia Romani, and Alberto Viscardi. Bezout-like polynomial equations associated with dual univariate interpolating subdivision schemes. Advances in Computational Mathematics, 48(1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09912-4.

Gasca:2000:PIS

[GS00] Mariano Gasca and Thomas Sauer. Polynomial interpolation in several variables. Advances in Computational Mathematics, 12(4): 377–410, March 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018981505752. Mul-

tivariate polynomial interpolation.

Guessab:2010:CPD

[GS10] Allal Guessab and Gerhard Schmeisser. Construction of positive definite cubature formulae and approximation of functions via Voronoi tessellations. Advances in Computational Mathematics, 32(1): 25–41, January 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9080-9.

Greengard:2012:E

[GS12] Leslie Greengard and Michael Shelley. Editorial. Advances in Computational Mathematics, 36(1):1, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9267-
                                        3;   http://link.springer.
                                        com/content/pdf/10.1007/
                                        s10444-011-9267-3.pdf.

Guo:2013:LCB

[GS13] Zheng-Chu Guo and Lei Shi. Learning with coefficientbased regularization and ℓ^{1}-penalty. Advances in Computational Mathematics, 39(3–4):493–510, October 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-012-9288-6.

Gao:2018:ALG

[GS18] Huadong Gao and Weiwei Sun. Analysis of linearized Galerkin-mixed FEMs for the time-dependent Ginzburg– Landau equations of superconductivity. Advances in Computational Mathematics, 44(3):923–949, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9568-
                                        2.

Guglielmi:2019:CCR

[GS19a] Nicola Guglielmi and Carmela Scalone. Computing the closest real normal matrix and normal completion. Advances in Computational Mathematics, 45(5-6):2867–2891, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09717-6.

Guo:2019:FSC

[GS19b] Zheng-Chu Guo and Lei Shi.

Fast and strong convergence of online learning algorithms. Advances in Computational Mathematics, 45(5-6):2745– 2770, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09707-8.

Guglielmi:2020:EMN

[GS20] Nicola Guglielmi and Carmela Scalone. An efficient method for non-negative low-rank completion. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09779-x.

Gottlich:2022:CST

[GS22] Simone Göttlich and Thomas Schillinger. Control strategies for transport networks under demand uncertainty. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09993-9.

Giannelli:2022:IDS

[GSS22] Carlotta Giannelli, Lorenzo Sacco, and Alessandra Sestini. Interpolation of 3D data streams with C^{2} PH quintic splines. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09976-w.

Garcia:2014:NOV

[GT14] Galina C. García and Takéo Takahashi. Numerical observers with vanishing viscosity for the 1 d wave equation. Advances in Computational Mathematics, 40(4): 711–745, August 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9320-5.

Tang:2002:SMI

[gTpM02] Jian guo Tang and He ping Ma. Single and multi-interval Legendre τ-methods in time for parabolic equations. Advances in Computational Mathematics, 17(4):349–367, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016273820035.

Gravesen:2002:CIF

[GU02] Jens Gravesen and Michael Ungstrup. Constructing invariant fairness measures for surfaces. Advances in Computational Mathematics, 17 (1–2):67–88, July 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1015229622042. Ad-

vances in geometrical algorithms and representations.

Gu:2017:PSC

[Gu17] Zhendong Gu. Piecewise spectral collocation method for system of Volterra integral equations. Advances in Computational Mathematics, 43 (2):385–409, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9490-z.

Gu:2019:SCM

[Gu19] Zhendong Gu. Spectral collocation method for system of weakly singular Volterra integral equations. Advances in Computational Mathematics, 45(5-6):2677–2699, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-019-09703-y.

Gu:2020:SCM

[Gu20] Zhendong Gu. Spectral collocation method for nonlinear Caputo fractional differential system. Advances in Computational Mathematics, 46(5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09808-9.

Gutleb:2021:FSS

[Gut21] Timon S. Gutleb. A fast sparse spectral method for nonlinear integro-differential Volterra equations with general kernels. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09866-7.

Gonzalez-Vera:1996:SRA

[GVSLN96] P. González-Vera, J. C. Santos-León, and O. Njåstad. Some results about numerical quadrature on the unit circle. Advances in Computational Mathematics, 5(1):297–328, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124749.

Guo:2014:CMG

[GW14] Ben-Yu Guo and Zhong-Qing Wang. A collocation method for generalized nonlinear Klein–Gordon equation. Advances in Computational Mathematics, 40(2): 377–398, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9312-5.

Gao:2017:CRK

[GW17] Wenwu Gao and Zongmin Wu. Constructing radial kernels with higher-order generalized Strang–Fix conditions. Advances in Computational Mathematics, 43(6):1355– 1375, December 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9528-x.

Guo:2021:EHV

[GW21] Yuling Guo and Zhongqing Wang. An hp-version Legendre spectral collocation method for multi-order fractional differential equations. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09858-7.

Gu:2024:ESM

[GWC24] Xuelong Gu, Yushun Wang, and Wenjun Cai. Energy stable and maximum bound principle preserving schemes for the Allen–Cahn equation based on the saul’yev methods. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10142-7.

Goodman:1999:IM

[GY99] Tim N. T. Goodman and Thomas P. Y. Yu. Interpolation of medians. Advances in Computational Mathematics, 11(1):1–10, July 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018959506235.

Goh:2000:UPL

[GY00] Say Song Goh and Chee Heng Yeo. Uncertainty products of local periodic wavelets. Advances in Computational Mathematics, 13(3):319–333, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1018962428951.

Yi:2012:GJR

[gYyG12] Yong gang Yi and Ben yu Guo. Generalized Jacobi rational spectral method on the half line. Advances in Computational Mathematics, 37(1):1–37, July 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9193-4.

Guo:2017:ORL

[GYZ17] Zheng-Chu Guo, Yiming Ying, and Ding-Xuan Zhou. Online regularized learning with pairwise loss functions. Advances in Computational Mathematics, 43(1):127–150, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9479-7.

Guo:2013:CEL

[GZ13] Zheng-Chu Guo and Ding-Xuan Zhou. Concentration estimates for learning with unbounded sampling. Advances in Computational Mathematics, 38(1): 207–223, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9238-8.

Gosea:2020:PDS

[GZA20] Ion Victor Gosea, Qiang Zhang, and Athanasios C. Antoulas. Preserving the DAE structure in the Loewner model reduction and identification framework. Advances in Computational Mathematics, 46(1):??, February 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-020-09752-
8.

Gong:2018:LSO

[GZW18] Yuezheng Gong, Jia Zhao, and Qi Wang. Linear second order in time energy stable schemes for hydrodynamic models of binary mixtures based on a spatially pseudospectral approximation. Advances in Computational Mathematics, 44(5):1573– 1600, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9597-5.

Hagger:1998:ADD

[Hag98] M. J. Hagger. Automatic domain decomposition on unstructured grids (DOUG). Advances in Computational Mathematics, 9(3–4):281– 310, November 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018997725374.

Hakopian:2000:CHI

[Hak00] Hakop A. Hakopian. On a class of Hermite interpolation problems. Advances in Computational Mathematics, 12(4):303–309, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018933622590. Mul-

tivariate polynomial interpolation.

Hall:1995:NSS

[Hal95] George Hall. A new stepsize strategy for explicit Runge– Kutta codes. Advances in Computational Mathematics, 3(4):343–352, May 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03028367.

Han:1998:SOS

[Han98] Bin Han. Symmetric orthonormal scaling functions and wavelets with dilation factor 4. Advances in Computational Mathematics, 8(4): 221–247, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018948314471.

Hansen:2005:FCM

[Han05] Olaf Hansen. A fast collocation method for the radiosity equation, based on the hierarchical algorithm of Hanrahan and Salzman: the 1 d case. Advances in Computational Mathematics, 22 (4):353–375, May 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-3962-7.

Han:2006:SSS

[Han06]  Bin Han.  Solutions in

Sobolev spaces of vector refinement equations with a general dilation matrix. Advances in Computational Mathematics, 24(1–4):375– 403, January 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7615-2.

Han:2010:SOC

[Han10] Bin Han. Symmetric orthonormal complex wavelets with masks of arbitrarily high linear-phase moments and sum rules. Advances in Computational Mathematics, 32(2):209–237, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9102-7.

Hangelbroek:2012:LRA

[Han12] T. Hangelbroek. On local RBF approximation. Advances in Computational Mathematics, 37(2):285–299, August 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9212-5.

Hollig:2005:IWM

[HAS05] Klaus Höllig, Christian Apprich, and Anja Streit. Introduction to the Web-method and its applications. Advances in Computational Mathematics, 23(1–2):215– 237, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1811-y.

Hauber:1997:NTR

[Hau97] R. Hauber. Numerical treatment of retarded differential– algebraic equations by collocation methods. Advances in Computational Mathematics, 7(4):573–592, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018919508111.

Hachama:2024:FOM

[HB24] Mohammed Hachama and Fatiha Boutaous. A fractional osmosis model for image fusion. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10103-6.

Herkert:2024:DBO

[HBH24] Robin Herkert, Patrick Buchfink, and Bernard Haasdonk. Dictionary-based online-adaptive structure-preserving model order reduction for parametric Hamiltonian systems. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10102-7.

Hoefkens:2003:CVS

[HBM03] Jens Hoefkens, Martin Berz, and Kyoko Makino. Computing validated solutions of implicit differential equations. Advances in Computational Mathematics, 19(1– 3):231–253, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1022858921155. Chal-

lenges in computational mathematics (Pohang, 2001).

Hao:2014:HOA

[HBMY14] S. Hao, A. H. Barnett, P. G. Martinsson, and P. Young. High-order accurate methods for Nyström discretization of integral equations on smooth curves in the plane. Advances in Computational Mathematics, 40(1): 245–272, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9306-3.

Huang:2021:ACE

[HCF^{+}21] Jingfang Huang, Jian Cao, Fuhui Fang, Marc G. Genton, David E. Keyes, and George Turkiyyah. An O(N) algorithm for computing expectation of N-dimensional truncated multi-variate normal distribution I: fundamentals. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09888-1.

Hu:2020:CFE

[HCH20] Hanzhang Hu, Yanping Chen, and Yunqing Huang. A characteristic finite element two-grid algorithm for a compressible miscible displacement problem. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09768-0.

Hou:2020:TGR

[HCYY20] Tianliang Hou, Luoping Chen, Yueting Yang, and Yin Yang. Two-grid Raviart– Thomas mixed finite element methods combined with Crank–Nicolson scheme for a class of nonlinear parabolic equations. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09777-z.

Herzet:2020:PGV

[HD20] C. Herzet and M. Diallo. Performance guarantees for a variational “multi-space” decoder. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09746-6.

Hemker:1995:SGF

[Hem95]  P. W. Hemker.  Sparse-

grid finite-volume multigrid for 3D-problems. Advances in Computational Mathematics, 4(1–2):83–110, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02123474. ???

Hemker:1996:CSG

[Hem96] P. W. Hemker. Corrigendum: “Sparse-grid finite-volume multigrid for 3D-problems” [Adv. Comput. Math. 4 (1995), no. 1–2, 83–110; MR1338896 (96c:65193)]. Advances in Computational Mathematics, 5(1):125–126, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124737;   http:
//link.springer.com/content/
pdf/10.1007/BF02124737.

pdf. See [Hem95].

Hesse:2009:ESD

[Hes09]  Kerstin Hesse.  The s-

energy of spherical designs on S^{2}. Advances in Computational Mathematics, 30(1): 37–59, January 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9057-0.

Holman:2019:CAM

[HF19] Jirí Holman and Jirí Fürst.

Coupling the algebraic model of bypass transition with EARSM model of turbulence. Advances in Computational Mathematics, 45(4): 1977–1992, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09680-2.

Hass:2007:GCS

[HFH^{+}07] Joel Hass, Rida T. Farouki, Chang Yong Han, Xiaowen Song, and Thomas W. Sederberg. Guaranteed consistency of surface intersections and trimmed surfaces using a coupled topology resolution and domain decomposition scheme. Advances in Computational Mathematics, 27(1):1–26, July 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7539-5.

Huang:2015:ETS

[HFH15] Pengzhan Huang, Xinlong Feng, and Yinnian He. An efficient two-step algorithm for the incompressible flow problem. Advances in Computational Mathematics, 41(6):1495–1513, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9400-1.

Helsing:2015:VEK

[HH15] Johan Helsing and Anders Holst. Variants of an explicit kernel-split panelbased Nyström discretization scheme for Helmholtz boundary value problems. Advances in Computational Mathematics, 41(3):691–708, June 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/

10.1007/s10444-014-9383-
y.

He:2016:DDR

[HH16] Hongjin He and Deren Han. A distributed Douglas– Rachford splitting method for multi-block convex minimization problems. Advances in Computational Mathematics, 42(1):27–53, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9408-1.

Hric:2019:NSN

[HH19a] Vladimír Hric and Jan Halama. On the numerical solution of non-equilibrium condensation of steam in nozzles. Advances in Computational Mathematics, 45(4): 2147–2162, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09700-1.

Hu:2019:SPV

[HH19b] Qiya Hu and Shaoliang Hu. A substructuring preconditioner with vertexrelated interface solvers for elliptic-type equations in three dimensions. Advances in Computational Mathematics, 45(3):1129– 1161, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9648-y.

Hardin:2004:MVR

[HHS04] Douglas P. Hardin, Thomas A. Hogan, and Qiyu Sun. The matrix-valued Riesz lemma and local orthonormal bases in shift-invariant spaces. Advances in Computational Mathematics, 20 (4):367–384, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1027389826705.

He:2018:CPI

[HHX18a] Hongjin He, Liusheng Hou, and Hong-Kun Xu. Correction to: A partially isochronous splitting algorithm for three-block separable convex minimization problems. Advances in Computational Mathematics, 44(4):1117–1118, August 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9591-y.

See [HHX18b].

He:2018:PIS

[HHX18b] Hongjin He, Liusheng Hou, and Hong-Kun Xu. A partially isochronous splitting algorithm for threeblock separable convex minimization problems. Advances in Computational Mathematics, 44(4):1091– 1115, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9574-4.

See correction [HHX18a].

Hubbert:2023:GWF

[HJ23] Simon Hubbert and Janin Jäger. Generalised Wendland functions for the sphere. Advances in Computational Mathematics, 49(1):??, February 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
022-10005-z.

Hiptmair:2012:MTB

[HJH12] R. Hiptmair and C. Jerez-Hanckes. Multiple traces boundary integral formulation for Helmholtz transmission problems. Advances in Computational Mathematics, 37(1):39–91, July 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9194-3.

Hu:1997:ROC

[HJJV97] Min Hu, Ken Jackson, Jan Janssen, and Stefan Vandewalle. Remarks on the optimal convolution kernel for CSOR waveform relaxation. Advances in Computational Mathematics, 7(1–2): 135–156, March 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018938617680. Par-

allel methods for ODEs.

Huang:2023:LMS

[HJYZ23] Qiong-Ao Huang, Wei Jiang, Jerry Zhijian Yang, and Gengen Zhang. Linear multi-step methods and their numerical stability for solving gradient flow equations. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10043-1.

Hinze:2021:STC

[HK21] Michael Hinze and Denis Korolev. A spacetime certified reduced basis method for quasilinear parabolic partial differential equations. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09860-z.

Healy:2004:TSE

[HKR04] D. M. Healy, Jr., P. J. Kostelec, and D. Rockmore. Towards safe and effective high-order Legendre transforms with applications to FFTs for the 2sphere. Advances in Computational Mathematics, 21(1– 2):59–105, July 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016431.
03652.65.

Helsing:2021:CIE

[HKR21] Johan Helsing, Anders Karlsson, and Andreas Rosén. Comparison of integral equations for the Maxwell transmission problem with general permittivities. Advances in Computational Mathematics, 47(5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09904-4.

Haslinger:2014:DDM

[HKRS14] Jaroslav Haslinger, Radek Kučera, Julien Riton, and Taoufik Sassi. A domain decomposition method for two-body contact problems with Tresca friction. Advances in Computational Mathematics, 40(1):65–90, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9299-y.

Hansen:1996:LUC

[HL96] Lars Kai Hansen and Jan Larsen. Linear unlearning for cross-validation. Advances in Computational Mathematics, 5(1):269–280, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02124747.

Han:2005:FPA

[HL05] Weimin Han and Wing Kam Liu. Flexible piecewise approximations based on partition of unity. Advances in Computational Mathematics, 23(1–2):191– 199, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1810-z.

He:2006:MLS

[HL06] Yinnian He and Kam-Moon Liu. Multi-level spectral Galerkin method for the Navier–Stokes equations. II. Time discretization. Advances in Computational Mathematics, 25(4):403–433, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7640-1.

Hesse:2008:CES

[HL08] Kerstin Hesse and Paul Leopardi. The Coulomb energy of spherical designs on S^{2}. Advances in Computational Mathematics, 28(4): 331–354, May 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9026-7.

Hamm:2016:CIG

[HL16a] Keaton Hamm and Jeff Ledford. Cardinal interpolation with general multiquadrics. Advances in Computational Mathematics, 42(5):1149– 1186, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9457-0.

Ho:2016:QQC

[HL16b] Kin Tat Ho and Lok Ming Lui. QCMC: quasi-conformal parameterizations for multiplyconnected domains. Advances in Computational Mathematics, 42(2):279–312, April 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-015-9424-
1.

Hamm:2018:CIG

[HL18] Keaton Hamm and Jeff Ledford. Cardinal interpolation with general multiquadrics: convergence rates. Advances in Computational Mathematics, 44(4):1205– 1233, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9578-0.

Huang:2020:NAF

[HL20] Yunqing Huang and Jichun Li. Numerical analysis of a finite element method for the electromagnetic concentrator model. Advances in Computational Mathematics, 46 (6):??, December 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09817-8.

Hou:2021:TSF

[HL21] Pengyu Hou and Fang Liu. Two-scale finite element discretizations for nonlinear eigenvalue problems in quantum physics. Advances in Computational Mathematics, 47(4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09883-6.

Hintermuller:2012:SOT

[HLN12] M. Hintermüller, A. Laurain, and A. A. Novotny. Secondorder topological expansion for electrical impedance tomography. Advances in Computational Mathematics, 36 (2):235–265, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9205-
4.

Han:2011:FTA

[HLT11] Deguang Han, Pengtong Li, and Wai-Shing Tang. Frames and their associated H_{F}^{p} -subspaces. Advances in Computational Mathematics, 34(2):185–200, February 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9149-0.

Han:2003:MFR

[HM03] Bin Han and Qun Mo.

Multiwavelet frames from refinable function vectors. Advances in Computational Mathematics, 18(2–4):211– 245, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021360312348. Frames.

Hu:2021:OSA

[HMM21] Jun Hu, Limin Ma, and Rui Ma. Optimal superconvergence analysis for the Crouzeix–Raviart and the Morley elements. Advances in Computational Mathematics, 47(4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09874-7.

Hieu:2022:OSO

[HMQ22] Dang Van Hieu, Le Dung Muu, and Pham Kim Quy. One-step optimization method for equilibrium problems. Advances in Computational Mathematics, 48(3):??, June 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        022-09953-3.

Harris:2019:CIP

[HMS19] Corey Harris, Mateusz Michałek, and Emre Can Sertöz. Computing images of polynomial maps. Advances in Computational Mathematics, 45(5-6):2845–2865, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09715-8.

Hao:2022:AFE

[HMX22] Tingting Hao, Manman Ma, and Xuejun Xu. Adaptive finite element approximation for steady-state Poisson–Nernst–Planck equations. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09938-2.

Howlett:1996:F

[HN96] Jack Howlett and J. C.

Newby. Foreword. Advances in Computational Mathematics, 6(1):iii–vi, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127692.

Halla:2018:TSH

[HN18] Martin Halla and Lothar Nannen. Two scale Hardy space infinite elements for scalar waveguide problems. Advances in Computational Mathematics, 44(3):611–643, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
017-9549-5.

Hofit:2023:GHI

[HNN23] Ben-Zion Vardi Hofit, Dyn Nira, and Sharon Nir. Geometric Hermite interpolation in R^{n} by refinements. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-023-10037-z.

Hashemi:2022:REP

[HNT22] Behnam Hashemi, Yuji Nakatsukasa, and Lloyd N. Trefethen. Rectangular eigenvalue problems. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-022-09994-
                                        8.

Himpe:2015:DDC

[HO15] Christian Himpe and Mario Ohlberger. Data-driven combined state and parameter reduction for inverse problems. Advances in Computational Mathematics, 41(5):1343– 1364, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9420-5.

Hegarty:2017:PUN

[HO17] A. F. Hegarty and E. O’Riordan. Parameter-uniform numerical method for singularly perturbed convection–diffusion problem on a circular domain. Advances in Computational Mathematics, 43(5): 885–909, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10444-016-9510-z.

Hoffman:2006:GVI

[Hof06] Alan J. Hoffman. Gersgorin variations. II. On themes of Fan and Gudkov. Advances in Computational Mathematics, 25(1–3):1–6, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7624-1.

Holst:2001:ANT

[Hol01] M. Holst. Adaptive numerical treatment of elliptic systems on manifolds. Advances in Computational Mathematics, 15(1–4):139– 191, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014246117321. A

posteriori error estimation and adaptive computational methods.

Hain:2019:HPE

[HORU19] Stefan Hain, Mario Ohlberger, Mladjan Radic, and Karsten Urban. A hierarchical a posteriori error estimator for the Reduced Basis Method. Advances in Computational Mathematics, 45(5-6):2191– 2214, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-019-09675-z.

Howlett:1996:FJC

[How96] Jack Howlett. Foreword: John Crank. Advances in Computational Mathematics, 6(1):vii–viii, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). John Crank 80th birthday special issue.

Hrkac:2019:CTP

[HPP^{+}19] Gino Hrkac, Carl-Martin Pfeiler, Dirk Praetorius, Michele Ruggeri, Antonio Segatti, and Bernhard Stiftner. Convergent tangent plane integrators for the simulation of chiral magnetic skyrmion dynamics. Advances in Computational Mathematics, 45(3):1329– 1368, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09667-z.

Herzog:2019:FIS

[HPS19] Roland Herzog, John W. Pearson, and Martin Stoll. Fast iterative solvers for an optimal transport problem. Advances in Computational Mathematics, 45(2): 495–517, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9625-5.

Hieu:2020:AHM

[HQHV20] Dang Van Hieu, Pham Kim Quy, La Thi Hong, and Le Van Vy. Accelerated hybrid methods for solving pseudomonotone equilibrium problems. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09778-y.

Hess:2023:DDS

[HQR23] Martin W. Hess, Annalisa Quaini, and Gianluigi Rozza. A data-driven surrogate modeling approach for time-dependent incompressible Navier–Stokes equations with dynamic mode decomposition and manifold interpolation. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10016-4.

Heuveline:2001:PEC

[HR01] Vincent Heuveline and Rolf Rannacher. A posteriori error control for finite element approximations of elliptic eigenvalue problems. Advances in Computational Mathematics, 15(1–4):107– 138, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014291224961. A

posteriori error estimation and adaptive computational methods.

Hsiao:2002:WCM

[HR02] George C. Hsiao and Andreas Rathsfeld. Wavelet collocation methods for a first kind boundary integral equation in acoustic scattering. Advances in Computational Mathematics, 17(4):281–308, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016249415373.

Hearn:2013:EJR

[HR13] Tristan A. Hearn and Lothar Reichel. Extensions of the Justen–Ramlau blind deconvolution method. Advances in Computational Mathematics, 39(3–4):465– 491, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-012-9290-z.

Holhos:2017:APM

[HR17] Adrian Holhoş and Daniela Roşca. Area preserving maps and volume preserving maps between a class of polyhedrons and a sphere. Advances in Computational Mathematics, 43(4):677–697, August 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9502-z.

Heuer:2003:ODD

[HS03] Norbert Heuer and Ernst P.

Stephan. An overlapping domain decomposition preconditioner for high order BEM with anisotropic elements. Advances in Computational Mathematics, 19(1– 3):211–230, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022899205324. Chal-

lenges in computational mathematics (Pohang, 2001).

Hon:2008:SPD

[HS08] Y. C. Hon and Robert Schaback. Solvability of partial differential equations by meshless kernel methods. Advances in Computational Mathematics, 28(3): 283–299, April 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9023-2.

Hon:2013:SLE

[HS13] Y. C. Hon and R. Schaback. Solving the 3D Laplace equation by meshless collocation via harmonic kernels. Advances in Computational Mathematics, 38 (1):1–19, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9224-1.

Huang:2016:SCM

[HS16] Can Huang and Martin Stynes. A spectral collocation method for a weakly singular Volterra integral equation of the second kind. Advances in Computational Mathematics, 42(5):1015– 1030, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9451-6.

Huang:2020:OSC

[HS20] Chaobao Huang and Martin Stynes. Optimal H^{1} spatial convergence of a fully discrete finite element method for the time-fractional Allen– Cahn equation. Advances in Computational Mathematics, 46(4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09805-y.

Hiptmair:2018:LDS

[HSSS18] R. Hiptmair, L. Scarabosio, C. Schillings, and Ch. Schwab. Large deformation shape uncertainty quantification in acoustic scattering. Advances in Computational Mathematics, 44(5):1475– 1518, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9594-8.

Hong:2009:ELC

[HSV09] Byung-Woo Hong, Stefano Soatto, and Luminita A. Vese. Enforcing local context into shape statistics. Advances in Computational Mathematics, 31(1–3):185– 213, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9104-5.

Hu:2016:STD

[HSY16] Jun Hu, Zhongci Shi, and Xueqin Yang. Superconvergence of three dimensional Morley elements on cuboid meshes for biharmonic equations. Advances in Computational Mathematics, 42(6):1453–1471, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9470-3.

Hon:2011:DTR

[HT11] Y. C. Hon and Tomoya Takeuchi. Discretized Tikhonov regularization by reproducing kernel Hilbert space for backward heat conduction problem. Advances in Computational Mathematics, 34(2): 167–183, February 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9148-1.

He:2012:TLL

[HT12] Lianhua He and Lutz Tobiska. The two-level local projection stabilization as an enriched one-level approach. Advances in Computational Mathematics, 36 (4):503–523, May 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9188-1.

Hu:2007:RDD

[Hu07]  Qiya Hu.  A regular-

ized domain decomposition method with Lagrange multiplier. Advances in Computational Mathematics, 26 (4):367–401, May 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-4094-4.

Huang:2006:CMB

[Hua06] Min Huang. A construction of multiscale bases for Petrov– Galerkin methods for integral equations. Advances in Computational Mathematics, 25 (1–3):7–22, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-7607-7.

Hubbert:2012:CFR

[Hub12]  Simon Hubbert.  Closed

form representations for a class of compactly supported radial basis functions. Advances in Computational Mathematics, 36(1): 115–136, January 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9184-5.

Hakopian:2019:CPE

[HV19a] Hakop Hakopian and Vahagn Vardanyan. On a correction of a property of GC_{n} sets. Advances in Computational Mathematics, 45(1): 311–325, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9618-4.

Hakopian:2019:ULS

[HV19b] Hakop Hakopian and Vahagn Vardanyan. On the usage of lines in GC_{n} sets. Advances in Computational Mathematics, 45(5-6):2721– 2743, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09705-w.

Hickernell:2000:IAA

[HW00] F. J. Hickernell and H. Woźniakowski. Integration and approximation in arbitrary dimensions. Advances in Computational Mathematics, 12(1): 25–58, January 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018948631251. High

dimensional integration.

Hesse:2012:NIP

[HW12] Kerstin Hesse and Robert S. Womersley. Numerical integration with polynomial exactness over a spherical cap. Advances in Computational Mathematics, 36(3): 451–483, April 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9187-2.

Hong:2018:CRM

[HW18] Qi Hong and Jiming Wu. Coercivity results of a modified Q_{1}-finite volume element scheme for anisotropic diffusion problems. Advances in Computational Mathematics, 44(3):897–922, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-017-9567-
3.

Huang:2019:ASM

[HW19a] Can Huang and Li-Lian Wang. An accurate spectral method for the transverse magnetic mode of Maxwell equations in Cole–Cole dispersive media. Advances in Computational Mathematics, 45(2):707–734, April 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-018-9636-
2.

Huning:2019:CSS

[HW19b] Svenja Hüning and Johannes Wallner. Convergence of subdivision schemes on Riemannian manifolds with nonpositive sectional curvature. Advances in Computational Mathematics, 45(3):1689– 1709, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09693-x.

Huang:2013:AEE

[HWZ13] Jin Huang, Zhu Wang, and Rui Zhu. Asymptotic error expansions for hypersingular integrals. Advances in Computational Mathematics, 38(2):257–279, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9236-x.

He:2022:NLF

[HWZ22] Yuyu He, Xiaofeng Wang, and Ruihua Zhong. A new linearized fourth-order conservative compact difference scheme for the SRLW equations. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09951-5.

Hu:2023:DNM

[HWZ23] Shenglong Hu, Yong Wang, and Jinling Zhou. A DCA– Newton method for quartic minimization over the sphere. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10040-4.

Huang:2002:CFE

[HX02] Yunqing Huang and Weimin Xue. Convergence of finite element approximations and multilevel linearization for Ginzburg–Landau model of d-wave superconductors. Advances in Computational Mathematics, 17(4):309–330, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016293508648.

Hou:2017:FSM

[HX17] Dianming Hou and Chuanju Xu. A fractional spectral method with applications to some singular problems. Advances in Computational Mathematics, 43(5): 911–944, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9511-y.

Huang:2021:ASV

[HX21] Rong Huang and Jungong Xue. Accurate singular values of a class of parameterized negative matrices. Advances in Computational Mathematics, 47(5): ??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09898-z.

Hu:2010:NMI

[HXC10] Xiaoyi Hu, Xiang Xu, and Wenbin Chen. Numerical method for the inverse heat transfer problem in composite materials with Stefan– Boltzmann conditions. Advances in Computational Mathematics, 33(4):471–489, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9131-x.

Hu:2014:NPM

[HY14] Qiya Hu and Long Yuan. A Newton-penalty method for a simplified liquid crystal model. Advances in Computational Mathematics, 40(1): 201–244, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9305-4.

Hu:2018:PWM

[HY18] Qiya Hu and Long Yuan. A plane wave method combined with local spectral elements for nonhomogeneous Helmholtz equation and timeharmonic Maxwell equations. Advances in Computational Mathematics, 44(1):245–275, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9542-z.

Hou:2021:RRD

[HYHH21] Jiangyong Hou, Wenjing Yan, Dan Hu, and Zhengkang He. Robin-Robin domain decomposition methods for the dual-porosity-conduit system. Advances in Computational Mathematics, 47(1): ??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09828-5.

Huang:2012:SEA

[HZHL12] Jin Huang, Guang Zeng, Xiaoming He, and Zi-Cai Li. Splitting extrapolation algorithm for first kind boundary integral equations with singularities by mechanical quadrature methods. Advances in Computational Mathematics, 36(1): 79–97, January 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9181-8.

He:2020:LSP

[HZZ20] Wenming He, Zhimin Zhang, and Qingsong Zou. Local superconvergence of postprocessed high-order finite volume element solutions. Advances in Computational Mathematics, 46(4):??, August 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
020-09801-2.

Huang:2024:NLO

[HZZZ24] Xuehai Huang, Chao Zhang, Yaqian Zhou, and Yangxing Zhu. New low-order mixed finite element methods for linear elasticity. Advances in Computational Mathematics, 50(2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10112-z.

Iacono:2017:NAP

[IB17] Roberto Iacono and John P. Boyd. New approximations to the principal realvalued branch of the Lambert W -function. Advances in Computational Mathematics, 43(6):1403–1436, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9530-3.

Ivanyshyn:2010:HPI

[IKS10] Olha Ivanyshyn, Rainer Kress, and Pedro Serranho. Huygens’ principle and iterative methods in inverse obstacle scattering. Advances in Computational Mathematics, 33(4):413–429, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9135-6;
 http://link.springer.com/
content/pdf/10.1007/s10444-
009-9135-6.pdf.

Ionescu:2016:MSA

[IL16] Ioan R. Ionescu and Oana Lupaşcu. Modeling shallow avalanche onset over complex basal topography. Advances in Computational Mathematics, 42(1):5–26, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9407-2.

Idriss:2022:TSB

[ILHH22] El Mourabit Idriss, Amine Laghrib, Aissam Hadri, and Abdelilah Hakim. A theoretical study of a bilateral term with a tensor-based fourthorder PDE for image superresolution. Advances in Computational Mathematics, 48 (6):??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09996-6.

Ivanov:1999:BLA

[IMS99] Tjavdar Ivanov, Vladimir Maz’ya, and Gunther Schmidt. Boundary layer approximate approximations and cubature of potentials in domains. Advances in Computational Mathematics, 10(3– 4):311–342, May 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1023/A:1018990918531.

Ivanov:2015:FME

[IP15] Kamen Ivanov and Pencho Petrushev. Fast memory efficient evaluation of spherical polynomials at scattered points. Advances in Computational Mathematics, 41(1): 191–230, February 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9354-3.

Izzo:2023:HOC

[IRT23] Federico Izzo, Olof Runborg, and Richard Tsai. Highorder corrected trapezoidal rules for a class of singular integrals. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10060-0.

Ismail:2013:AGD

[IS13] Mourad E. H. Ismail and Plamen Simeonov. Asymptotics of generalized derangements. Advances in Computational Mathematics, 39 (1):101–127, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9271-7.

Ito:1996:NCL

[Ito96] Yoshifusa Ito. Nonlinearity creates linear independence. Advances in Computational Mathematics, 5(1):189–203, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124743.

Iapichino:2017:MPC

[IUV17] L. Iapichino, S. Ulbrich, and S. Volkwein. Multiobjective PDE-constrained optimization using the reduced-basis method. Advances in Computational Mathematics, 43(5): 945–972, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9512-x.

Iguchi:2022:WAS

[IY22] Yuga Iguchi and Toshihiro Yamada. Weak approximation of SDEs for tempered distributions and applications. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09960-4.

Javanshiri:2022:EIF

[JAA22] Hossein Javanshiri, Mohammad Abolghasemi, and Ali Akbar Arefijamaal. The essence of invertible frame multipliers in scalability. Advances in Computational Mathematics, 48(3):??, June 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
022-09940-8.

Janik:1998:PIV

[Jan98] Tadeusz J. Janik. Parallel implementation of the pversion of the finite element method for elliptic equations on a shared-memory architecture. Advances in Computational Mathematics, 8(1–2): 97–110, January 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018936011745.

Jeon:1998:NBE

[Jeo98] Youngmok Jeon. New boundary element formulas for the biharmonic equation. Advances in Computational Mathematics, 9(1–2):97–115, September 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018985322648. Nu-

merical treatment of boundary integral equations.

Juttler:2002:LSF

[JF02] Bert Jüttler and Alf Felis. Least-squares fitting of algebraic spline surfaces. Advances in Computational Mathematics, 17(1–2):135– 152, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1015200504295. Ad-

vances in geometrical algorithms and representations.

Jiang:2015:ESE

[JGW15] Shidong Jiang, Leslie Greengard, and Shaobo Wang. Efficient sum-of-exponentials approximations for the heat kernel and their applications. Advances in Computational Mathematics, 41(3): 529–551, June 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9372-1.

Jia:1995:SSS

[Jia95]  Rong-Qing Jia.  Sub-

division schemes in L_{p} spaces. Advances in Computational Mathematics, 3(4): 309–341, May 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03028366.

Jiang:2000:PCO

[Jia00] Qingtang Jiang. Parameterization of m-channel orthogonal multifilter banks.

Advances in Computational Mathematics, 12(2–3):189– 211, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018965102118.

Jiang:2003:PMT

[Jia03] Qingtang Jiang. Parameterizations of masks for tight affine frames with two symmetric/antisymmetric generators. Advances in Computational Mathematics, 18(2–4): 247–268, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1021339707805. Frames.

Jia:2009:SWI

[Jia09a]  Rong-Qing Jia.  Spline

wavelets on the interval with homogeneous boundary conditions. Advances in Computational Mathematics, 30(2): 177–200, February 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9064-9.

Jiang:2009:HTF

[Jia09b] Qingtang Jiang. Hexagonal tight frame filter banks with idealized high-pass filters. Advances in Computational Mathematics, 31(1–3): 215–236, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9085-4.

Jiang:2011:BWF

[Jia11] Qingtang Jiang. Biorthogonal wavelets with 4-fold axial symmetry for quadrilateral surface multiresolution processing. Advances in Computational Mathematics, 34(2): 127–165, February 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9144-5.

Jang:2013:CSM

[JJK13] Sumi Jang, Byeongseon Jeong, and Hong Oh Kim. Compactly supported multiwindow dual Gabor frames of rational sampling density. Advances in Computational Mathematics, 38(1): 159–186, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9234-z.

Jenkins:2014:PCG

[JJLR14] Eleanor W. Jenkins, Volker John, Alexander Linke, and Leo G. Rebholz. On the parameter choice in grad–div stabilization for the Stokes equations. Advances in Computational Mathematics, 40 (2):491–516, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9316-1.

John:2008:FEE

[JK08] Volker John and Songul Kaya. Finite element error analysis of a variational multiscale method for the Navier–Stokes equations. Advances in Computational Mathematics, 28(1): 43–61, January 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-9010-z.

Jaklic:2012:AGI

[JKK^{+}12] Gašper Jaklič, Jernej Kozak, Marjeta Krajnc, Vito Vitrih, and Emil Žagar. An approach to geometric interpolation by Pythagorean-hodograph curves. Advances in Computational Mathematics, 37 (1):123–150, July 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9209-0.

Jeong:2013:EPR

[JKLY13] Byeongseon Jeong, Hong Oh Kim, Yeon Ju Lee, and Jungho Yoon. Exponential polynomial reproducing property of nonstationary symmetric subdivision schemes and normalized exponential B-splines.

Advances in Computational Mathematics, 38(3):647–666, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-011-9253-
9.

Jiang:1997:SPM

[JL97] Qingtang Jiang and S. L. Lee.

Spectral properties of matrix continuous refinement operators. Advances in Computational Mathematics, 7(3): 383–399, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018911306293.

Jia:2006:WBH

[JL06] Rong-Qing Jia and Song-Tao Liu. Wavelet bases of Hermite cubic splines on the interval. Advances in Computational Mathematics, 25(1– 3):23–39, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-7609-5.

Jungel:2019:CIE

[JL19] Ansgar Jüngel and Oliver Leingang. Convergence of an implicit Euler Galerkin scheme for Poisson–Maxwell– Stefan systems. Advances in Computational Mathematics, 45(3):1469–1498, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-019-09674-
                                        0.

Ji:2020:SMP

[JlLZ20] Bingquan Ji, Hong lin Liao, and Luming Zhang. Simple maximum principle preserving time-stepping methods for time-fractional Allen– Cahn equation. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09782-2.

Josz:2019:SPI

[JLM19] Cédric Josz, Jean Bernard Lasserre, and Bernard Mourrain. Sparse polynomial interpolation: sparse recovery, super-resolution, or Prony? Advances in Computational Mathematics, 45(3): 1401–1437, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09672-2.

Jiang:2020:NRS

[JLW20] Daijun Jiang, Yikan Liu, and Dongling Wang. Numerical reconstruction of the spatial component in the source term of a time-fractional diffusion equation. Advances in Computational Mathematics, 46(3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09754-6.

Jia:2008:AER

[JLZ08] Shanghui Jia, Deli Li, and Shuhua Zhang. Asymptotic expansions and Richardson extrapolation of approximate solutions for integrodifferential equations by mixed finite element methods. Advances in Computational Mathematics, 29 (4):337–356, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9052-
5.

Jodar:2000:SOA

[JM00] Lucas Jódar and Marco Marletta. Solving ODEs arising from non-selfadjoint Hamiltonian eigenproblems. Advances in Computational Mathematics, 13(3):231–256, July 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/A:1018954227133.

Jeon:2003:NBE

[JM03] Youngmok Jeon and William McLean. A new boundary element method for the biharmonic equation with Dirichlet boundary conditions. Advances in Computational Mathematics, 19(4):339–354, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1024206232212.

Jackiewicz:2000:CRK

[JMO00] Z. Jackiewicz, A. Marthinsen, and B. Owren. Construction of Runge–Kutta methods of Crouch–Grossman type of high order. Advances in Computational Mathematics, 13(4):405–415, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016645730465.

Jung:2018:FSS

[JN18] Chang-Yeol Jung and Thien Binh Nguyen. Fine structures for the solutions of the twodimensional Riemann problems by high-order WENO schemes. Advances in Computational Mathematics, 44 (1):147–174, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-017-9538-
                                        8.

Johnson:2013:RIS

[Joh13] Michael J. Johnson. The role of inner summaries in the fast evaluation of thinplate splines. Advances in Computational Mathematics, 39(1):1–25, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9265-5.

Jones:2017:CUP

[Jon17] Robert Stephen Jones. Computing ultra-precise eigenvalues of the Laplacian within polygons. Advances in Computational Mathematics, 43(6):1325–1354, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9527-y.

Jaming:2023:UPR

[JR23] Philippe Jaming and Martin Rathmair. Uniqueness of phase retrieval from three measurements. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10045-z.

Junior:2015:OIF

[JRS15] D. S. Almeida Júnior, A. J. A.

Ramos, and M. L. Santos. Observability inequality for the finite-difference semi-discretization of the 1d coupled wave equations.

Advances in Computational Mathematics, 41(1):105–130, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9351-6.

Jetter:1995:GBS

[JS95] K. Jetter and J. Stöckler. A generalization of de Boor’s stability result and symmetric preconditioning. Advances in Computational Mathematics, 3(4):353–367, May 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF03028368.

Ji:1999:CSB

[JS99] Hui Ji and Zuowei Shen. Compactly supported (bi)orthogonal wavelets generated by interpolatory refinable functions. Advances in Computational Mathematics, 11 (1):81–104, July 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018999220348.

Jeon:2005:ACB

[JS05] Youngmok Jeon and Dongwoo Sheen. Analysis of a cell boundary element method. Advances in Computational Mathematics, 22 (3):201–222, April 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-7671-z.

Jang:2021:PEA

[JS21] Yongseok Jang and Simon Shaw. A priori error analysis for a finite element approximation of dynamic viscoelasticity problems involving a fractional order integro-differential constitutive law. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09857-8.

Jeon:1997:DQM

[JSSE97] Y. Jeon, I. H. Sloan, E. P.

Stephan, and J. Elschner. Discrete qualocation methods for logarithmic-kernel integral equations on a piecewise smooth boundary. Advances in Computational Mathematics, 7(4):547–571, August 1997. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018967424040.

Jiang:2020:LIS

[JSW20] Chaolong Jiang, Yongzhong Song, and Yushun Wang. A linearly implicit structurepreserving Fourier pseudospectral scheme for the damped nonlinear Schrödinger equation in three dimensions. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09781-3.

Juttler:1998:DBF

[Jüt98] Bert Jüttler. The dual basis functions for the Bernstein polynomials. Advances in Computational Mathematics, 8(4):345–352, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018912801267.

Jones:1995:ZFR

[JW95] John W. Jones and Bruno D. Welfert. Zero-free regions for a rational function with applications. Advances in Computational Mathematics, 3(3): 265–289, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02988626.

Jiang:2023:SRK

[JWJ23] Yutong Jiang, Gang Wu, and Long Jiang. A semirandomized Kaczmarz method with simple random sampling for large-scale linear systems. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10018-2.

Jiang:2022:NIT

[JY22] Nan Jiang and Huanhuan Yang. Numerical investigation of two second-order, stabilized SAV ensemble methods for the Navier–Stokes equations. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09977-9.

Jetter:2004:PRS

[JZ04] K. Jetter and G. Zimmermann. Polynomial reproduction in subdivision. Advances in Computational Mathematics, 20(1–3):67–86, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025859224071. Ap-

proximation and applications.

Jia:2010:FAT

[JZ10] Rong-Qing Jia and Hanqing Zhao. A fast algorithm for the total variation model of image denoising. Advances in Computational Mathematics, 33(2):

231–241, August 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9128-5.

Konstantinov:1997:SAD

[KA97] Mihail Konstantinov and Vera Angelova. Sensitivity analysis of the differential matrix Riccati equation based on the associated linear differential system. Advances in Computational Mathematics, 7(3): 295–301, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018903104476.

Kacewicz:2018:AMS

[Kac18] Bolesław Kacewicz. Adaptive mesh selection asymptotically guarantees a prescribed local error for systems of initial value problems. Advances in Computational Mathematics, 44(5):1325– 1344, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9584-2.

Kosinka:2015:CWC

[KB15] Jiří Kosinka and Michael Bartoň. Convergence of Wachspress coordinates: from polygons to curved domains. Advances in Computational Mathematics, 41(3): 489–505, June 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9370-3.

Karjoun:2022:SPA

[KBL22] Hasan Karjoun, Abdelaziz Beljadid, and Philippe G. LeFloch. A structurepreserving algorithm for surface water flows with transport processes. Advances in Computational Mathematics, 48(1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09918-y.

Kublik:2016:LAT

[KC16] Richard A. Kublik and David L. Chopp. A locally adaptive time stepping algorithm for the solution to reaction diffusion equations on branched structures. Advances in Computational Mathematics, 42(3): 621–649, June 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9437-9.

Kohl:2023:FAS

[KCCV23] Ryan Kohl, Eduardo Corona, Vani Cheruvu, and Shravan Veerapaneni. Fast and accurate solvers for simulating Janus particle suspensions in Stokes flow. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10046-y.

Kordy:2017:NSC

[KCW17] Michal Kordy, Elena Cherkaev, and Philip Wannamaker. Null space correction and adaptive model order reduction in multi-frequency Maxwell’s problem. Advances in Computational Mathematics, 43(1):171–193, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9482-z.

Keinert:1995:NSB

[Kei95]  Fritz Keinert.  Numeri-

cal stability of biorthogonal wavelet transforms. Advances in Computational Mathematics, 4(1–2):1–26, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02123471.

Krah:2022:WAP

[KESR22] Philipp Krah, Thomas Engels, Kai Schneider, and Julius Reiss. Wavelet adaptive proper orthogonal decomposition for large-scale flow data. Advances in Computational Mathematics, 48 (2):??, April 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09922-2.

Krivodonova:2003:EED

[KF03] Lilia Krivodonova and Joseph E.

Flaherty. Error estimation for discontinuous Galerkin solutions of two-dimensional hyperbolic problems. Advances in Computational Mathematics, 19(1–3):57– 71, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022894504834. Chal-

lenges in computational mathematics (Pohang, 2001).

Kunc:2019:GEM

[KF19] Oliver Kunc and Felix Fritzen. Generation of energy-minimizing point sets on spheres and their application in mesh-free interpolation and differentiation. Advances in Computational Mathematics, 45(5-6):3021– 3056, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09726-5.

Keck:2005:PTI

[KH05] R. Keck and D. Hietel. A projection technique for incompressible flow in the meshless finite volume particle method. Advances in Computational Mathematics, 23(1– 2):143–169, July 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1831-7.

Khleborodov:2018:FEC

[Khl18] Denis Khleborodov. Fast elliptic curve point multiplication based on binary and binary non-adjacent scalar form methods. Advances in Computational Mathematics, 44(4):1275– 1293, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9581-5.

Kriva:2016:ACC

[KHM16] Zuzana Krivá, Angela Handlovičová, and Karol Mikula. Adaptive cell-centered finite volume method for diffusion equations on a consistent quadtree grid. Advances in Computational Mathematics, 42(2): 249–277, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9423-2.

Kim:2006:GDA

[Kim06] M.-Y. Y. Kim. Global dynamics of approximate solutions to an age-structured epidemic model with diffusion. Advances in Computational Mathematics, 25 (4):451–474, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7639-
7.

Kirkup:1998:FCC

[Kir98] S. M. Kirkup. Fortran codes for computing the discrete Helmholtz integral operators. Advances in Computational Mathematics, 9(3–4):391– 409, November 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018953910353.

Kosinka:2009:HIP

[KJ09] Jiří Kosinka and Bert Jüttler.

C^{1} Hermite interpolation by Pythagorean hodograph quintics in Minkowski space. Advances in Computational Mathematics, 30(2):123–140, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9059-y.

Koras:1999:CCP

[KK99] G. D. Koras and P. D. Kaklis. Convexity conditions for parametric tensorproduct B-spline surfaces. Advances in Computational Mathematics, 10(3–4):291– 309, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018938901692.

Kuchment:2002:DOG

[KK02] P. Kuchment and L. Kunyansky. Differential operators on graphs and photonic crystals. Advances in Computational Mathematics, 16(2–3): 263–290, April 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014481629504. Mod-

eling and computation in optics and electromagnetics.

Karni:2005:LEA

[KK05] Smadar Karni and Alexander Kurganov. Local error analysis for approximate solutions of hyperbolic conservation laws. Advances in Computational Mathematics, 22 (1):79–99, January 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7099-8.

Kuehn:2020:CEE

[KK20] Christian Kuehn and Patrick Kürschner. Combined error estimates for local fluctuations of SPDEs. Advances in Computational Mathematics, 46(1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09766-2.

Kim:2003:QBF

[KKL03] Hong Oh Kim, Rae Young Kim, and Jae Kun Lim. Quasi-biorthogonal frame multiresolution analyses and wavelets. Advances in Computational Mathematics, 18(2–4):269–296, February 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/A:1021371227388.

Frames.

Keil:2022:AML

[KKL^{+}22] Tim Keil, Hendrik Kleikamp, Rolf J. Lorentzen, Micheal B. Oguntola, and Mario Ohlberger. Adaptive machine learningbased surrogate modeling to accelerate PDE-constrained optimization in enhanced oil recovery. Advances in Computational Mathematics, 48 (6):??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-022-09981-z.

Kim:2010:QIR

[KKLY10] Hong Oh Kim, Rae Young Kim, Yeon Ju Lee, and Jungho Yoon. Quasiinterpolatory refinable functions and construction of biorthogonal wavelet systems. Advances in Computational Mathematics, 33(3): 255–283, October 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9129-4.

Kevlahan:2019:CDA

[KKP19] N. K. R. Kevlahan, R. Khan, and B. Protas. On the convergence of data assimilation for the one-dimensional shallow water equations with sparse observations. Advances in Computational Mathematics, 45(5-6):3195– 3216, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09733-6.

Karner:2024:LNN

[KKP24] Clemens Karner, Vladimir Kazeev, and Philipp Christian Petersen. Limitations of neural network training due to numerical instability of backpropagation. Advances in Computational Mathematics, 50(1):

??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10106-x.

Kozak:2015:PCP

[KKV15] Jernej Kozak, Marjeta Krajnc, and Vito Vitrih. Parametric curves with Pythagorean binormal. Advances in Computational Mathematics, 41 (4):813–832, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9387-
7.

Kim:2000:MMF

[KL00] Yongdeok Kim and Sungyun Lee. Modified Mini finite element for the Stokes problem in R^{2} or R^{3}. Advances in Computational Mathematics, 12(2–3):261– 272, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018973303935.

Konovalov:2004:FKS

[KL04] V. N. Konovalov and D. Leviatan. Free-knot splines approximation of s-monotone functions. Advances in Computational Mathematics, 20 (4):347–366, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1027324000817.

Konovalov:2007:FSA

[KL07] V. N. Konovalov and D. Leviatan. Freeknot splines approximation of Sobolev-type classes of s-monotone functions. Advances in Computational Mathematics, 27(2): 211–236, August 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9032-9.

Kruschel:2015:CAR

[KL15] Christian Kruschel and Dirk A. Lorenz. Computing and analyzing recoverable supports for sparse reconstruction. Advances in Computational Mathematics, 41(6):1555– 1580, December 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9403-6.

Kosmas:2016:AHO

[KL16] Odysseas Kosmas and Sigrid Leyendecker. Analysis of higher order phase fitted variational integrators. Advances in Computational Mathematics, 42(3):605–619, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/

10.1007/s10444-015-9436-
x.

Kopteva:2017:IMN

[KL17] Natalia Kopteva and Torsten Linß. Improved maximumnorm a posteriori error estimates for linear and semilinear parabolic equations. Advances in Computational Mathematics, 43(5): 999–1022, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9514-3.

Kosmas:2019:VIO

[KL19] Odysseas Kosmas and Sigrid Leyendecker. Variational integrators for orbital problems using frequency estimation. Advances in Computational Mathematics, 45(1): 1–21, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9603-y.

Kim:2022:NID

[KL22] Yunho Kim and Dongsun Lee. Numerical investigation into the dependence of the Allen–Cahn equation on the free energy. Advances in Computational Mathematics, 48(3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-022-09955-1.

Kleefeld:2021:HSC

[Kle21]  Andreas Kleefeld.  The

hot spots conjecture can be false: some numerical examples. Advances in Computational Mathematics, 47(6): ??, December 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09911-5.

Koch:1995:CCK

[KLNS95] Per Erik Koch, Tom Lyche, Marian Neamtu, and Larry L. Schumaker. Control curves and knot insertion for trigonometric splines. Advances in Computational Mathematics, 3(4): 405–424, May 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03028369.

Knobloch:2019:EIO

[KLS19] Petr Knobloch, Petr Lukás, and Pavel Solin. On error indicators for optimizing parameters in stabilized methods. Advances in Computational Mathematics, 45(4): 1853–1862, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09662-4.

Karkulik:2019:MHM

[KM19] Michael Karkulik and Jens Markus Melenk. H-matrix approximability of inverses of discretizations of the fractional Laplacian. Advances in Computational Mathematics, 45(5-6):2893–2919, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09718-5.

Koc:2019:CER

[KMMI19] Birgul Koc, Muhammad Mohebujjaman, Changhong Mou, and Traian Iliescu. Commutation error in reduced order modeling of fluid flows. Advances in Computational Mathematics, 45(5-6):2587–2621, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09739-0.

Kobarg:2014:NEM

[KMO^{+}14] Jan Hendrik Kobarg, Peter Maass, Janina Oetjen, Oren Tropp, Eyal Hirsch, Chen Sagiv, Mohammad Golbabaee, and Pierre Vandergheynst. Numerical experiments with MALDI Imaging data. Advances in Computational Mathematics, 40(3): 667–682, June 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9325-0.

Kropinski:2014:FIE

[KN14] Mary Catherine A. Kropinski and Nilima Nigam. Fast integral equation methods for the Laplace–Beltrami equation on the sphere. Advances in Computational Mathematics, 40(2):577–596, April 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-013-9319-
                                        y.

Knibbeler:2024:CEM

[Kni24] Vincent Knibbeler. Computing equivariant matrices on homogeneous spaces for geometric deep learning and automorphic Lie algebras. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10126-7.

Knobloch:2009:CSP

[Kno09] Petr Knobloch. On the choice of the SUPG parameter at outflow boundary layers. Advances in Computational Mathematics, 31(4):369–389, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9075-6.

Kaneko:1998:SPG

[KNP98] Hideaki Kaneko, Richard D.

Noren, and Peter A. Padilla. Singularity preserving Galerkin method for Hammerstein equations with logarithmic kernel. Advances in Computational Mathematics, 9(3– 4):363–376, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018910128100.

Kropinski:2016:IEM

[KNQ16] M. C. Kropinski, N. Nigam, and B. Quaife. Integral equation methods for the Yukawa– Beltrami equation on the sphere. Advances in Computational Mathematics, 42(2): 469–488, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9431-2.

Krizek:2001:PEE

[KNV01] Michal Křížek, Jan Němec, and Tomáš Vejchodský. A posteriori error estimates for axisymmetric and nonlinear problems. Advances in Computational Mathematics, 15(1–4):219–236, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014234911830. A

posteriori error estimation and adaptive computational methods.

Kalmoun:2023:NCP

[KNV23] El Mostafa Kalmoun, Mohamed M. S. Nasser, and Matti Vuorinen. Numerical computation of a preimage domain for an infinite strip with rectilinear slits. Advances in Computational Mathematics, 49(1): ??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-10006-y.

Kazeev:2017:QFE

[KORS17] Vladimir Kazeev, Ivan Oseledets, Maxim Rakhuba, and Christoph Schwab. QTTfinite-element approximation for multiscale problems I: model problems in one dimension. Advances in Computational Mathematics, 43 (2):411–442, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9491-y.

Kostic:2015:GPE

[Kos15] Vladimir Kostić. On general principles of eigenvalue localizations via diagonal dominance. Advances in Computational Mathematics, 41(1): 55–75, February 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9349-0.

Kotnik:2008:PCF

[Kot08] Tadej Kotnik. The primecounting function and its analytic approximations: π(x) and its approximations. Advances in Computational Mathematics, 29(1):55–70, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-007-9039-
2.

Khoromskij:1995:MPR

[KP95] Boris N. Khoromskij and Siegfried Prössdorf. Multilevel preconditioning on the refined interface and optimal boundary solvers for the Laplace equation. Advances in Computational Mathematics, 4(1–2):331–355, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123480.

Khoromskij:1998:FCH

[KP98] Boris N. Khoromskij and Siegfried Prössdorf. Fast computations with the harmonic Poincaré–Steklov operators on nested refined meshes. Advances in Computational Mathematics, 8(1–2): 111–135, January 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018988028583.

Kouibia:2004:ASF

[KP04] A. Kouibia and M. Pasadas. Approximation of surfaces by fairness bicubic splines. Advances in Computational Mathematics, 20(1–3):87– 103, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025805701726. Ap-

proximation and applications.

Kundu:2018:FEA

[KP18] Sudeep Kundu and Amiya Kumar Pani. Finite element approximation to global stabilization of the Burgers’ equation by Neumann boundary feedback control law. Advances in Computational Mathematics, 44(2): 541–570, April 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9553-9.

Krishnan:2009:PDA

[KPY09] D. Krishnan, Quang Vinh Pham, and Andy M. Yip. A primal-dual active-set algorithm for bilaterally constrained total variation deblurring and piecewise constant Mumford–Shah segmentation problems. Advances in Computational Mathematics, 31(1–3):237– 266, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9101-8.

Kunsch:2019:OCM

[KR19] Robert J. Kunsch and Daniel Rudolf. Optimal confidence for Monte Carlo integration of smooth functions. Advances in Computational Mathematics, 45(5-6):3095– 3122, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09728-3.

Krasauskas:2002:TSP

[Kra02] Rimvydas Krasauskas. Toric surface patches. Advances in Computational Mathematics, 17(1–2):89–113, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1015289823859.

Advances in geometrical algorithms and representations.

Keslerova:2019:NMG

[KRP19] Radka Keslerová, Hynek Reznícek, and Tomás Padelek. Numerical modelling of generalized Newtonian fluids in bypass tube. Advances in Computational Mathematics, 45(4):2047–2063, August 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09684-y.

Kopteva:2017:ANS

[KS17] Natalia Kopteva and Martin Stynes. Analysis and numerical solution of a Riemann–Liouville fractional derivative two-point boundary value problem. Advances in Computational Mathematics, 43(1):77–99, February 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9476-x.

Kashefi:2018:FEC

[KS18] Ali Kashefi and Anne E. Staples. A finite-element coarsegrid projection method for incompressible flow simulations. Advances in Computational Mathematics, 44(4): 1063–1090, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9573-5.

Karvonen:2020:WCO

[KS20] Toni Karvonen and Simo Särkkä. Worst-case optimal approximation with increasingly flat Gaussian kernels. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09767-1.

Kaye:2023:FTD

[KS23] Jason Kaye and Hugo U. R.

Strand. A fast time domain solver for the equilibrium Dyson equation. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10067-7.

Kapl:2021:FQF

[KST21] Mario Kapl, Giancarlo Sangalli, and Thomas Takacs. A family of C^{1} quadrilateral finite elements. Advances in Computational Mathematics, 47(6):??, December 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-021-09878-3.

Krieg:2023:ETA

[KSUW23] David Krieg, Paweł Siedlecki, Mario Ullrich, and Henryk Woźniakowski. Exponential tractability of L_{2}- approximation with function values. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10021-7.

Karasozen:2018:EPM

[KU18] Bülent Karasözen and Murat Uzunca. Energy preserving model order reduction of the nonlinear Schrödinger equation. Advances in Computational Mathematics, 44(6):1769–1796, December 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9593-9.

Kumar:2023:SAS

[Kum23] Naresh Kumar. Supercloseness analysis of a stabilizerfree weak Galerkin finite element method for viscoelastic wave equations with variable coefficients. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10010-w.

Kunoth:1995:MPA

[Kun95] Angela Kunoth. Multilevel preconditioning — appending boundary conditions by Lagrange multipliers. Advances in Computational Mathematics, 4(1–2):145–170, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123477.

Kunert:2001:RPE

[Kun01]  Gerd Kunert.  Robust

a posteriori error estimation for a singularly perturbed reaction–diffusion equation on anisotropic tetrahedral meshes. Advances in Computational Mathematics, 15(1–4):237–259, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014248711347. A

posteriori error estimation and adaptive computational methods.

Kunis:2009:NSR

[Kun09] Stefan Kunis. A note on stability results for scattered data interpolation on Euclidean spheres. Advances in Computational Mathematics, 30(4):303–314, May 2009.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9069-
                                        4.

Kurschner:2018:BTM

[Kür18]  Patrick Kürschner.  Bal-

anced truncation model order reduction in limited time intervals for large systems. Advances in Computational Mathematics, 44(6):1821– 1844, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9608-6.

Kapl:2021:SIS

[KV21] Mario Kapl and Vito Vitrih. C^{s}-smooth isogeometric spline spaces over planar bilinear multi-patch parameterizations. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09868-5.

Kvasov:2014:MCI

[Kva14] Boris I. Kvasov. Monotone and convex interpolation by weighted quadratic splines. Advances in Computational Mathematics, 40(1): 91–116, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9300-9.

Kuijt:2001:LAS

[KvD01] Frans Kuijt and Ruud van Damme. A linear approach to shape preserving spline approximation. Advances in Computational Mathematics, 14(1):25–48, January 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016660513725.

Koren:1995:APM

[KvL95] B. Koren and B. van Leer.

Analysis of preconditioning and multigrid for Euler flows with low-subsonic regions. Advances in Computational Mathematics, 4(1–2):127– 144, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123476.

Kang:2024:FNI

[KXL24] Hongchao Kang, Qi Xu, and Guidong Liu. Fast numerical integration of highly oscillatory Bessel transforms with a Cauchy type singular point and exotic oscillators. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10134-7.

Kim:2004:UCM

[KXZ04] Hwanho Kim, Jinchao Xu, and Ludmil Zikatanov. Uniformly convergent multigrid methods for convection– diffusion problems without any constraint on coarse grids. Advances in Computational Mathematics, 20(4): 385–399, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1027378015262.

Kang:2001:SPE

[KY01] Tong Kang and Dehao Yu. Some a posteriori error estimates of the finite-difference streamline–diffusion method for convection-dominated diffusion equations. Advances in Computational Mathematics, 15(1–4):193– 218, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014294002686. A

posteriori error estimation and adaptive computational methods.

Levie:2022:RCF

[LA22] Ron Levie and Haim Avron. Randomized continuous frames in time-frequency analysis. Advances in Computational Mathematics, 48(3): ??, June 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09941-7.

Lamerand:2024:IPD

[LAG^{+}24] Louis Lamérand, Didier Auroux, Philippe Ghendrih, Francesca Rapetti, and Eric Serre. Inverse problem for determining free parameters of a reduced turbulent transport model for tokamak plasma. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10135-6.

Lai:2006:CMC

[Lai06] Ming-Jun Lai. Construction of multivariate compactly supported orthonormal wavelets. Advances in Computational Mathematics, 25(1–3):41–56, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7643-
y.

Lara:2018:NRS

[Lar18] F. Lara. A note on “Reguralizers for structured sparsity”. Advances in Computational Mathematics, 44(4):

1321–1323, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9583-3.

Lasserre:2016:LDA

[Las16] Jean B. Lasserre. Lebesgue decomposition in action via semidefinite relaxations. Advances in Computational Mathematics, 42(5):1129– 1148, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9456-1.

Liu:1993:EBF

[LB93] W. B. Liu and John W. Barrett. Error bounds for the finite element approximation of a degenerate quasilinear parabolic variational inequality. Advances in Computational Mathematics, 1(2): 223–239, June 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02071387.

Lu:2018:SCM

[LBCZ18] Hong Lu, Peter W. Bates, Wenping Chen, and Mingji Zhang. The spectral collocation method for efficiently solving PDEs with fractional Laplacian. Advances in Computational Mathematics, 44 (3):861–878, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9564-6.

Li:2009:NSF

[LC09] Jian Li and Zhangxin Chen.

A new stabilized finite volume method for the stationary Stokes equations. Advances in Computational Mathematics, 30(2):141–152, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9060-5.

Li:2013:SDS

[LC13] Jian Li and Zhangxin Chen.

On the semi-discrete stabilized finite volume method for the transient Navier– Stokes equations. Advances in Computational Mathematics, 38(2):281–320, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9237-9.

Leng:2018:CQO

[LC18] Haitao Leng and Yanping Chen. Convergence and quasi-optimality of an adaptive finite element method for optimal control problems with integral control constraint. Advances in Computational Mathematics, 44(2): 367–394, April 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9546-8.

Liu:2019:NVE

[LC19] Xin Liu and Zhangxin Chen. The nonconforming virtual element method for the Navier–Stokes equations. Advances in Computational Mathematics, 45(1):51–74, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9602-z.

Leng:2020:AHD

[LC20] Haitao Leng and Yanping Chen. Adaptive hybridizable discontinuous Galerkin methods for nonstationary convection diffusion problems. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09795-x.

Leng:2021:RTP

[LC21] Haitao Leng and Yanping Chen. Residual-type a posteriori error analysis of HDG methods for Neumann boundary control problems. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09864-9.

Li:2019:DAF

[LCC19] Jichun Li, Meng Chen, and Min Chen. Developing and analyzing fourth-order difference methods for the metamaterial Maxwell’s equations. Advances in Computational Mathematics, 45(1): 213–241, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9614-8.

Lipovetsky:2007:EAC

[LD07] Evgeny Lipovetsky and Nira Dyn. An efficient algorithm for the computation of the metric average of two intersecting convex polygons, with application to morphing. Advances in Computational Mathematics, 26(1–3): 269–282, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7473-6.

Li:2016:HOS

[LD16] Can Li and Weihua Deng.

High order schemes for the tempered fractional diffusion equations. Advances in Computational Mathematics, 42 (3):543–572, June 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9434-z.

Lai:2022:FSE

[LD22] J. Lai and H. Dong. A fast solver for elastic scattering from axisymmetric objects by boundary integral equations. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09935-5.

LeMehaute:2000:SAM

[Le 00]  A. Le Méhauté.  On

some aspects of multivariate polynomial interpolation. Advances in Computational Mathematics, 12(4): 311–333, March 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018985606661. Mul-

tivariate polynomial interpolation.

LeGia:2005:APP

[Le 05] Q. T. Le Gia. Approximation of parabolic PDEs on spheres using spherical basis functions. Advances in Computational Mathematics, 22 (4):377–397, May 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-3960-9.

Ledford:2015:CRF

[Led15] Jeff Ledford. On the convergence of regular families of cardinal interpolators. Advances in Computational Mathematics, 41(2): 357–371, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9361-4.

Lee:2003:ACM

[Lee03] Hyung-Chun Lee. Analysis and computational methods of Dirichlet boundary optimal control problems for 2D Boussinesq equations. Advances in Computational Mathematics, 19(1–3):255– 275, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022872602498. Chal-

lenges in computational mathematics (Pohang, 2001).

Lee:2016:TUA

[Lee16]  Jeonghun J. Lee.  To-

wards a unified analysis of mixed methods for elasticity with weakly symmetric stress. Advances in Computational Mathematics, 42(2):

361–376, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9427-y.

Leeb:2021:RES

[Lee21] William Leeb. Rapid evaluation of the spectral signal detection threshold and Stieltjes transform. Advances in Computational Mathematics, 47(4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09890-7.

Lemvig:2009:CPD

[Lem09] Jakob Lemvig. Constructing pairs of dual bandlimited framelets with desired time localization. Advances in Computational Mathematics, 30(3):231–247, April 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9066-
                                        7.

Lenze:1994:LGS

[Len94] Burkhard Lenze. On local and global sigma–pi neural networks. A common link. Advances in Computational Mathematics, 2(4):479–491, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF02521610.

Lenze:1996:HSP

[Len96] Burkhard Lenze. Hyperbolic sigma–pi neural network operators for compactly supported continuous functions. Advances in Computational Mathematics, 5(1):163–172, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124741.

Leopoldseder:2002:CSS

[Leo02] Stefan Leopoldseder. Cone spline surfaces and spatial arc splines — a sphere geometric approach. Advances in Computational Mathematics, 17 (1–2):49–66, July 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1015215802130. Ad-

vances in geometrical algorithms and representations.

Leopardi:2013:DSR

[Leo13] Paul Leopardi. Discrepancy, separation and Riesz energy of finite point sets on the unit sphere. Advances in Computational Mathematics, 39 (1):27–43, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9266-4.

Lepe:2023:IPD

[Lep23] Felipe Lepe. Interior penalty discontinuous Galerkin methods for the velocity-pressure formulation of the Stokes spectral problem. Advances in Computational Mathematics, 49(4):??, August 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-023-10062-
                                        y.

Levin:1999:ULP

[Lev99] David Levin. Using Laurent polynomial representation for the analysis of nonuniform binary subdivision schemes. Advances in Computational Mathematics, 11 (1):41–54, July 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018907522165.

Li:2022:NLP

[LFL22] Yang Li, Minfu Feng, and Yan Luo. A new local projection stabilization virtual element method for the Oseen problem on polygonal meshes. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09952-4.

Li:2013:CCN

[LG13] J. Li and O. Gonzalez.

Convergence and conditioning of a Nyström method for Stokes flow in exterior three-dimensional domains. Advances in Computational Mathematics, 39 (1):143–174, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9272-1.

Li:2020:DFV

[LGC^{+}20] Rui Li, Yali Gao, Jie Chen, Li Zhang, Xiaoming He, and Zhangxin Chen. Discontinuous finite volume element method for a coupled Navier– Stokes–Cahn–Hilliard phase field model. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09764-4.

Liu:2022:UOE

[LGN22] Ying Liu, Zhen Guan, and Yufeng Nie. Unconditionally optimal error estimates of a linearized weak Galerkin finite element method for semilinear parabolic equations. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09961-3.

Li:2024:ERP

[LGQ24] Jian Li, Jiawei Gao, and Yi Qin. An efficient rotational pressure-correction schemes for 2D/3D closedloop geothermal system. Advances in Computational Mathematics, 50(3):??, June 2024. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        024-10154-3.

Leng:2013:OPD

[LH13] Jinsong Leng and Deguang Han. Orthogonal projection decomposition of matrices and construction of fusion frames. Advances in Computational Mathematics, 38(2): 369–381, February 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9241-0.

Liu:2023:SOS

[LHH23] Shuaijun Liu, Pengzhan Huang, and Yinnian He. A second-order scheme based on blended BDF for the incompressible MHD system. Advances in Computational Mathematics, 49 (5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10073-9.

Liu:2008:NOD

[LHY08] Yang Liu, Qiya Hu, and Dehao Yu. A non-overlapping domain decomposition for low-frequency time-harmonic Maxwell’s equations in unbounded domains. Advances in Computational Mathematics, 28(4):355–382, May 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-007-9027-
6.

Li:1994:HOB

[Li94]  Xin Li.  Hankel opera-

tors and best Hankel approximation on the halfplane. Advances in Computational Mathematics, 2(3): 343–355, April 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521115.

Li:2003:IMC

[Li03] Yuesheng Li. The inversion of multiscale convolution approximation and average of distributions. Advances in Computational Mathematics, 19(1–3):293– 306, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1023/A:1022811005225. Chal-

lenges in computational mathematics (Pohang, 2001).

Li:2004:CSM

[Li04]  Song Li.  Characteriza-

tion of smoothness of multivariate refinable functions and convergence of cascade algorithms of nonhomogeneous refinement equations. Advances in Computational Mathematics, 20(4): 311–331, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1027373528712.

Li:2005:MMF

[Li05a] Jichun Li. Mixed methods for fourth-order elliptic and parabolic problems using radial basis functions. Advances in Computational Mathematics, 23(1–2):21– 30, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1807-7.

Li:2005:CMF

[Li05b] Xin Li. On convergence of the method of fundamental solutions for solving the Dirichlet problem of Poisson’s equation. Advances in Computational Mathematics, 23(3): 265–277, October 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1782-z.

Li:2008:CMF

[Li08a]  Xin Li.  Convergence

of the method of fundamental solutions for Poisson’s equation on the unit sphere. Advances in Computational Mathematics, 28(3): 269–282, April 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9022-3.

Li:2008:RCM

[Li08b] Xin Li. Rate of convergence of the method of fundamental solutions and hyperinterpolation for modified Helmholtz equations on the unit ball. Advances in Computational Mathematics, 29(4):393–413, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9056-1.

Li:2009:API

[Li09]  Xin Li.  Approximation

of potential integral by radial bases for solutions of Helmholtz equation. Advances in Computational Mathematics, 30(3):201–230, April 2009. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-008-9065-
                                        8.

Li:2010:RBAa

[Li10a] Xin Li. Radial basis approximation and its application to biharmonic equation. Advances in Computational Mathematics, 32(3): 275–302, April 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9107-2.

Li:2010:RBAb

[Li10b] Xin Li. Radial basis approximation for Newtonian potentials. Advances in Computational Mathematics, 33 (1):1–24, July 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9117-8.

Licht:2022:BCF

[Lic22]  Martin W. Licht.  On

basis constructions in finite element exterior calculus. Advances in Computational Mathematics, 48 (2):??, April 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09926-6.

Liu:1999:RKC

[Liu99] Yunkang Liu. Runge–Kuttacollocation methods for systems of functional-differential and functional equations. Advances in Computational Mathematics, 11(4):315–329, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018936328525.

Liu:2018:NMS

[LJ18] Na Liu and Wei Jiang.

A numerical method for solving the time fractional Schrödinger equation. Advances in Computational Mathematics, 44(4):1235– 1248, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9579-z.

Lu:2020:AAS

[LJL20] Jian Lu, Qingtang Jiang, and Lin Li. Analysis of adaptive synchrosqueezing transform with a time-varying parameter. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09814-x.

Ling:2005:LSP

[LK05] Leevan Ling and Edward J. Kansa. A least-squares preconditioner for radial basis functions collocation methods. Advances in Computational Mathematics, 23(1– 2):31–54, July 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1809-5.

Luan:2020:DTD

[LK20] Tran Nhat Luan and Tra Quoc Khanh. Determination of temperature distribution and thermal flux for two-dimensional inhomogeneous sideways heat equations. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09796-w.

Li:2018:RAD

[LKT18] Huamin Li, Yuval Kluger, and Mark Tygert. Randomized algorithms for distributed computation of principal component analysis and singular value decomposition. Advances in Computational Mathematics, 44(5):1651– 1672, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-018-9600-1.

Lewanowicz:2017:BFD

[LKW17] Stanisław Lewanowicz, Paweł Keller, and Paweł Woźny. Bézier form of dual bivariate Bernstein polynomials. Advances in Computational Mathematics, 43(4): 777–793, August 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9506-8.

Laghchim-Lahlou:1998:FSC

[LL98] Mohammed Laghchim-Lahlou.

The C^{r}-fundamental splines of Clough–Tocher and Powell– Sabin types for Lagrange interpolation on a three direction mesh. Advances in Computational Mathematics, 8(4): 353–366, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018904532218.

Levesley:1999:DFS

[LL99] Jeremy Levesley and Will Light. Direct form seminorms arising in the theory of interpolation by translates of a basis function. Advances in Computational Mathematics, 11(2–3):161– 182, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1023/A:1018971808961. Ra-

dial basis functions and their applications.

Lai:2004:NKT

[LL04] Ming-Jun Lai and A. LeMéhauté. A new kind of trivariate C^{1} macro-element. Advances in Computational Mathematics, 21(3–4):273– 292, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/B:ACOM.0000032047.
                                        05052.27.

Lin:2007:CAE

[LL07a] Qun Lin and Jiaquan Liu. Counterexamples to the asymptotic expansion of interpolation in finite elements. Advances in Computational Mathematics, 27(2): 167–177, August 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9030-y.

Lu:2007:FSA

[LL07b] Shujuan Lü and Qishao Lu. Fourier spectral approximation to long-time behavior of three dimensional Ginzburg–Landau type equation. Advances in Computational Mathematics, 27(3): 293–318, October 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9004-x.

Lv:2010:EEF

[LL10] Junliang Lv and Yonghai Li. L^{2} error estimate of the finite volume element methods on quadrilateral meshes. Advances in Computational Mathematics, 33(2): 129–148, August 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9121-z.

Lian:2011:GFS

[LL11] Qiao-Fang Lian and Yun-Zhang Li. Gabor frame sets for subspaces. Advances in Computational Mathematics, 34(4):391–411, May 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-010-9161-
4.

Lin:2012:AHO

[LL12a] Fu-Rong Lin and Fen Liang.

Application of high order numerical quadratures to numerical inversion of the Laplace transform. Advances in Computational Mathematics, 36(2):267–278, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-011-9202-7.

Lv:2012:EES

[LL12b] Junliang Lv and Yonghai Li. L^{2} error estimates and superconvergence of the finite volume element methods on quadrilateral meshes. Advances in Computational Mathematics, 37(3):393–416, October 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9215-2.

Li:2013:SGF

[LL13] Yun-Zhang Li and Qiao-Fang Lian. Super Gabor frames on discrete periodic sets. Advances in Computational Mathematics, 38(4): 763–799, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9259-3.

Li:2021:ROP

[LL21] Huiping Li and Song Li. Riemannian optimization for phase retrieval from masked Fourier measurements. Advances in Computational Mathematics, 47(6):??, December 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09914-2.

Liu:2016:WGF

[LLC16] Xin Liu, Jian Li, and Zhangxin Chen. A weak Galerkin finite element method for the Oseen equations. Advances in Computational Mathematics, 42(6):1473– 1490, December 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9471-2.

Lin:2023:ECA

[LLLH23] Jia-Wei Lin, Tiexiang Li, Wen-Wei Lin, and Tsung-Ming Huang. Ellipsoidal conformal and area-/volume-preserving parameterizations and associated optimal mass transportations. Advances in Computational Mathematics, 49(4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10048-w.

Li:2024:AVP

[LLR24] Hongpeng Li, Xu Li, and Hongxing Rui. Analysis of a P_{1} ⊕ RT_{0} finite element method for linear elasticity with Dirichlet and mixed boundary conditions. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-024-10107-w.

Laghchim-Lahlou:1994:TFE

[LLS94] M. Laghchim-Lahlou and P. Sablonnière. Triangular finite elements of HCT type and class C^{ρ}. Advances in Computational Mathematics, 2(1):101–122, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02519038.

Lawton:1995:CCS

[LLS95] Wayne Lawton, S. L. Lee, and Zuowei Shen. Characterization of compactly supported refinable splines. Advances in Computational Mathematics, 3(1–2):137– 145, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF03028364.

Laghchim-Lahlou:1996:FEP

[LLS96] M. Laghchim-Lahlou and P. Sablonnière. C^{r}-finite elements of Powell–Sabin type on the three direction mesh. Advances in Computational Mathematics, 6(1):191–206, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127703.

Lee:2003:LFN

[LLS03] Chang-Ock Lee, Jongwoo Lee, and Dongwoo Sheen. A locking-free nonconforming finite element method for planar linear elasticity. Advances in Computational Mathematics, 19(1–3):277– 291, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022838628615. Chal-

lenges in computational mathematics (Pohang, 2001).

Lee:2009:CNA

[LLS09] Cheng-Feng Lee, Leevan Ling, and Robert Schaback. On convergent numerical algorithms for unsymmetric collocation. Advances in Computational Mathematics, 30(4):339–354, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9071-
x.

Lee:2006:SBS

[LLY06] Byung-Gook Lee, Yeon Ju Lee, and Jungho Yoon. Stationary binary subdivision schemes using radial basis function interpolation. Advances in Computational Mathematics, 25(1–3):57– 72, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7642-z.

Luo:2012:PPC

[LLY12] Xingjun Luo, Fanchun Li, and Suhua Yang. A posteriori parameter choice strategy for fast multiscale methods solving ill-posed integral equations. Advances in Computational Mathematics, 36(2): 299–314, February 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9229-9.

Li:2010:ESS

[LLZ10] Mingxia Li, Qun Lin, and Shuhua Zhang. Extrapolation and superconvergence of the Steklov eigenvalue problem. Advances in Computational Mathematics, 33 (1):25–44, July 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9118-7.

Lamberti:2004:TQI

[LM04] P. Lamberti and C. Manni. Tensioned quasi-interpolation via geometric continuity. Advances in Computational Mathematics, 20(1–3):105– 127, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1023/A:1025823221346. Ap-

proximation and applications.

Lyche:2006:TPE

[LM06] Tom Lyche and Marie-Laurence Mazure. Total positivity and the existence of piecewise exponential Bsplines. Advances in Computational Mathematics, 25(1– 3):105–133, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7633-0.

LeGia:2014:SHE

[LM14] Quoc Thong Le Gia and William McLean. Solving the heat equation on the unit sphere via Laplace transforms and radial basis functions. Advances in Computational Mathematics, 40(2): 353–375, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9311-6.

Lemvig:2014:PTF

[LMO14] Jakob Lemvig, Christopher Miller, and Kasso A. Okoudjou. Prime tight frames. Advances in Computational Mathematics, 40(2): 315–334, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-013-9309-0.

Lepe:2023:CPV

[LMRV23a] F. Lepe, D. Mora, G. Rivera, and I. Velásquez. Correction: a posteriori virtual element method for the acoustic vibration problem. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10022-6.

Lepe:2023:PVE

[LMRV23b] F. Lepe, D. Mora, G. Rivera, and I. Velásquez. A posteriori virtual element method for the acoustic vibration problem. Advances in Computational Mathematics, 49(1): ??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-10003-1.

Liu:2001:MEE

[LMT01] Wenbin Liu, Heping Ma, and Tao Tang. On mixed error estimates for elliptic obstacle problems. Advances in Computational Mathematics, 15(1–4):261– 283, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014261013164. A

posteriori error estimation and adaptive computational methods.

Linh:2011:MEA

[LMV11] Vu Hoang Linh, Volker Mehrmann, and Erik S. Van Vleck. QR methods and error analysis for computing Lyapunov and Sacker–Sell spectral intervals for linear differentialalgebraic equations. Advances in Computational Mathematics, 35(2–4):281– 322, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9156-1.

Liu:2013:FRK

[LMWX13] Jianqiang Liu, Charles A.

Micchelli, Rui Wang, and Yuesheng Xu. Finite rank kernels for multi-task learning. Advances in Computational Mathematics, 38(2): 427–439, February 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9244-x.

Li:2019:ENS

[LMYL19] Qi Li, Liquan Mei, Xiaofeng Yang, and Yibao Li. Efficient numerical schemes with unconditional energy stabilities for the modified phase field crystal equation. Advances in Computational Mathematics, 45(3): 1551–1580, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09678-w.

Lechleiter:2014:TGM

[LN14] Armin Lechleiter and Dinh-Liem Nguyen. A trigonometric Galerkin method for volume integral equations arising in TM grating scattering. Advances in Computational Mathematics, 40(1): 1–25, February 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9295-2.

Li:2019:ICU

[LN19] Fang Li and Michael K. Ng. Image colorization by using graph bi-Laplacian. Advances in Computational Mathematics, 45(3):1521– 1549, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09677-x.

Lyche:2002:PIM

[LNW02] Tom Lyche, Trygve K. Nilssen, and Ragnar Winther. Preconditioned iterative methods for scattered data interpolation. Advances in Computational Mathematics, 17(3):237–256, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016050415412.

Li:2008:ONS

[LO08] Shidong Li and Hidemitsu Ogawa. Optimal noise suppression: a geometric nature of pseudoframes for subspaces. Advances in Computational Mathematics, 28(2): 141–155, February 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9014-3.

Lohi:2023:NDF

[Loh23] Jonni Lohi. New degrees of freedom for differential forms on cubical meshes. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10047-x.

Lopez:2018:CEB

[Lóp18]  José L. López.  Conver-

gent expansions of the Bessel functions in terms of elementary functions. Advances in Computational Mathematics, 44(1):277–294, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-017-9543-y.

Lowitzsch:2005:MVR

[Low05] Svenja Lowitzsch. Matrixvalued radial basis functions: stability estimates and applications. Advances in Computational Mathematics, 23(3): 299–315, October 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1786-8.

Lyche:2004:OSM

[LP04] T. Lyche and J. M. Peña. Optimally stable multivariate bases. Advances in Computational Mathematics, 20(1–3): 149–159, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1025863309959. Ap-

proximation and applications.

Li:2010:SSP

[LP10] Song Li and Yali Pan. Subdivision schemes with polynomially decaying masks. Advances in Computational Mathematics, 32(4):487–507, May 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-009-9116-

Lyon:2014:FAS

[LP14] M. Lyon and J. Picard.

The Fourier approximation of smooth but nonperiodic functions from unevenly spaced data. Advances in Computational Mathematics, 40(5–6):1073–1092, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9342-7.

Lasserre:2019:ECF

[LP19] Jean B. Lasserre and Edouard Pauwels. The empirical Christoffel function with applications in data analysis. Advances in Computational Mathematics, 45(3): 1439–1468, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09673-1.

Laakmann:2021:EAS

[LP21] Fabian Laakmann and Philipp Petersen. Efficient approximation of solutions of parametric linear transport equations by ReLU DNNs. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09834-7.

Lorente-Pardo:2000:CSA

[LPSSP00] J. Lorente-Pardo, P. Sablonnière, and M. C. Serrano-Pérez. On the convexity of C^{1} surfaces associated with some quadrilateral finite elements. Advances in Computational Mathematics, 13 (3):271–292, July 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018958328042.

Lammers:2010:ADF

[LPY10] Mark Lammers, Alexander M. Powell, and Özgür Yılmaz. Alternative dual frames for digital-to-analog conversion in sigma–delta quantization. Advances in Computational Mathematics, 32(1):73–102, January 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-008-9088-
                                        1.

Li:2014:HWS

[LQ14] Gang Li and Jianxian Qiu. Hybrid WENO schemes with different indicators on curvilinear grids. Advances in Computational Mathematics, 40(4):747–772, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10444-013-9322-3.

Lindsay:2018:BIE

[LQW18] Alan E. Lindsay, Bryan Quaife, and Laura Wendelberger. A boundary integral equation method for mode elimination and vibration confinement in thin plates with clamped points. Advances in Computational Mathematics, 44(4):1249– 1273, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9580-6.

Levesley:2007:RBI

[LR07] J. Levesley and D. L.

Ragozin. Radial basis interpolation on homogeneous manifolds: convergence rates. Advances in Computational Mathematics, 27(2):237–246, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9000-1.

Lehrenfeld:2019:MCR

[LR19] Christoph Lehrenfeld and Stephan Rave. Mass conservative reduced order modeling of a free boundary osmotic cell swelling problem. Advances in Computational Mathematics, 45(5-6):2215– 2239, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09691-z.

Linss:2012:NMS

[LRS12] Torsten Linß, Hans-Görg Roos, and Martin Schopf. Nitsche-mortaring for singularly perturbed convection– diffusion problems. Advances in Computational Mathematics, 36(4):581–603, May 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9195-
                                        2.

Landoulsi:2023:IOF

[LRY23] Oussama Landoulsi, Svetlana Roudenko, and Kai Yang. Interaction with an obstacle in the 2D focusing nonlinear Schrödinger equation. Advances in Computational Mathematics, 49 (5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10055-x.

Lai:1998:APB

[LS98] Ming-Jun Lai and Larry L. Schumaker. On the approximation power of bivariate splines. Advances in Computational Mathematics, 9(3– 4):251–279, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018958011262.

Lu:2005:PDD

[LS05] Fangyan Lu and Hongwei Sun. Positive definite dot product kernels in learning theory. Advances in Computational Mathematics, 22(2): 181–198, February 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-3140-6.

Lv:2018:RPR

[LS18] Fusheng Lv and Wenchang Sun. Real phase retrieval from unordered partial frame coefficients. Advances in Computational Mathematics, 44(3):879–896, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-017-9566-
4.

Liang:2019:CMG

[LS19] Hui Liang and Martin Stynes.

Collocation methods for general Riemann–Liouville twopoint boundary value problems. Advances in Computational Mathematics, 45(2): 897–928, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9645-1.

Li:2020:SEE

[LS20] Xiaoli Li and Jie Shen. Stability and error estimates of the SAV Fourier-spectral method for the phase field crystal equation. Advances in Computational Mathematics, 46(3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09789-9.

Levie:2014:ATA

[LSLS14] R. Levie, H.-G. Stark, F. Lieb, and N. Sochen. Adjoint translation, adjoint observable and uncertainty principles. Advances in Computational Mathematics, 40 (3):609–627, June 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9336-x.

LeGia:2011:SNP

[LST11] Q. T. Le Gia, E. P. Stephan, and T. Tran. Solution to the Neumann problem exterior to a prolate spheroid by radial basis functions. Advances in Computational Mathematics, 34(1):83–103, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9145-4.

LeGia:2017:ZGL

[LSW17] Q. T. Le Gia, I. H. Sloan, and H. Wendland. Zooming from global to local: a multiscale RBF approach. Advances in Computational Mathematics, 43 (3):581–606, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9498-4.

Li:2015:MSF

[LSXZ15] Qia Li, Lixin Shen, Yuesheng Xu, and Na Zhang. Multi-step fixed-point proximity algorithms for solving a class of optimization problems arising from image processing. Advances in Computational Mathematics, 41(2): 387–422, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9363-2.

Li:2022:DAT

[LSY22] Jichun Li, Chi-Wang Shu, and Wei Yang. Development and analysis of two new finite element schemes for a time-domain carpet cloak model. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-022-09948-0.

Langer:2024:RST

[LSY24] Ulrich Langer, Olaf Steinbach, and Huidong Yang. Robust space-time finite element methods for parabolic distributed optimal control problems with energy regularization. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10123-w.

Liandrat:1998:FAS

[LT98] J. Liandrat and Ph. Tchamitchian. On the fast approximation of some nonlinear operators in nonregular wavelet spaces. Advances in Computational Mathematics, 8(4): 179–192, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018992129492.

Leviatan:2006:SAG

[LT06] D. Leviatan and V. N. Temlyakov. Simultaneous approximation by greedy algorithms. Advances in Computational Mathematics, 25 (1–3):73–90, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10444-004-7613-4.

Liu:2012:SGT

[LT12] Entao Liu and Vladimir N.

Temlyakov. Super greedy type algorithms. Advances in Computational Mathematics, 37(4):493–504, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9220-5.

Lopez:2013:NSE

[LT13] José L. López and Nico M.

Temme. New series expansions of the Gauss hypergeometric function. Advances in Computational Mathematics, 39(2):349–365, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9283-y.

Luo:1999:SPD

[Luo99] Zuhua Luo. Strictly positive definiteness of Hermite interpolation on spheres. Advances in Computational Mathematics, 10(3–4):261– 270, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018982716713.

Luther:2004:ASN

[Lut04] U. Luther. Approximation spaces in the numerical analysis of operator equations. Advances in Computational Mathematics, 20(1–3):129– 147, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025806612592. Ap-

proximation and applications.

Liu:2003:UNU

[LW03] Youming Liu and Yang Wang. The uniformity of non-uniform Gabor bases. Advances in Computational Mathematics, 18(2–4):345– 355, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021350103925. Frames.

Lai:2004:SMS

[LW04] Ming-Jun Lai and Paul Wenston. L_{1} spline methods for scattered data interpolation and approximation. Advances in Computational Mathematics, 21(3–4):293– 315, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032042.
35918.c2.

LeGia:2014:DCS

[LW14] Q. T. Le Gia and H. Wendland. Data compression on the sphere using multiscale radial basis function approximation. Advances in Computational Mathematics, 40(4): 923–943, August 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9334-z.

Li:2023:CSA

[LW23] Ting Li and Bin Wang.

Continuous-stage adapted exponential methods for charged-particle dynamics with arbitrary magnetic fields. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10093-5.

Liu:2017:FDS

[LWL17] Fu-Jun Liu, Zhong-Qing Wang, and Hui-Yuan Li. A fully diagonalized spectral method using generalized Laguerre functions on the half line. Advances in Computational Mathematics, 43(6):1227–1259, December 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-017-9522-3.

Liu:2021:OEE

[LWW21] Wenjie Liu, Li-Lian Wang, and Boying Wu. Optimal error estimates for Legendre expansions of singular functions with fractional derivatives of bounded variation. Advances in Computational Mathematics, 47(6): ??, December 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09905-3.

Liu:2022:WCD

[LWWZ22] Jun Liu, Xiang-Sheng Wang, Shu-Lin Wu, and Tao Zhou. A well-conditioned direct PinT algorithm for firstand second-order evolutionary equations. Advances in Computational Mathematics, 48(3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09928-4.

Lu:2019:CIP

[LWX19] Peipei Lu, Haijun Wu, and Xuejun Xu. Continuous interior penalty finite element methods for the timeharmonic Maxwell equation with high wave number. Advances in Computational Mathematics, 45(5-6):3265– 3291, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09737-2.

Lian:2006:CLT

[LWY06] Qiaofang Lian, Yongge Wang, and Dunyan Yan. A characterization of L^{p}(R) by local trigonometric bases with 1 < p < ∞. Advances in Computational Mathematics, 25(1– 3):91–104, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7625-0.

Wang:2006:SLP

[lWyG06] Li lian Wang and Ben yu Guo. Stair Laguerre pseudospectral method for differential equations on the half line. Advances in Computational Mathematics, 25(1– 3):305–322, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-003-7608-6.

Liu:2019:NGC

[LWZ19] Zhifang Liu, Chunlin Wu, and Yanan Zhao. A new globally convergent algorithm for non-Lipschitz l_{p}–l_{q} minimization. Advances in Computational Mathematics, 45(3): 1369–1399, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09668-y.

Xu:2002:MLL

[lXyG02] Cheng long Xu and Ben yu Guo. Mixed Laguerre– Legendre spectral method for incompressible flow in an infinite strip. Advances in Computational Mathematics, 16 (1):77–96, January 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014249613222.

Liu:2001:PEE

[LY01] Wenbin Liu and Ningning Yan. A posteriori error estimates for distributed convex optimal control problems. Advances in Computational Mathematics, 15(1–4):285– 309, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014239012739. A

posteriori error estimation and adaptive computational methods.

Li:2007:CMK

[LY07] Haitao Li and Dunyan Yan. Characterizations of multi-knot piecewise linear spectral sequences. Advances in Computational Mathematics, 27(4):401–422, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9006-8.

Liu:2008:SAN

[LY08] Huipo Liu and Ningning Yan. Superconvergence analysis of the nonconforming quadrilateral linear-constant scheme for Stokes equations. Advances in Computational Mathematics, 29(4):375–392, November 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9054-3.

Li:2021:NCA

[LY21] Zhi Li and Ming Yan. New convergence analysis of a primal-dual algorithm with large stepsizes. Advances in Computational Mathematics, 47(1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09840-9.

Li:2021:SSO

[LYF^{+}21] Yibao Li, Qian Yu, Weiwei Fang, Binhu Xia, and Junseok Kim. A stable second-order BDF scheme for the three-dimensional Cahn– Hilliard–Hele–Shaw system. Advances in Computational Mathematics, 47(1):??, February 2021. CODEN ACMHEX.

ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        020-09835-6.

Li:2013:BMS

[LYY13] Youfa Li, Shouzhi Yang, and Dehui Yuan. Bessel multiwavelet sequences and dual multiframelets in Sobolev spaces. Advances in Computational Mathematics, 38(3): 491–529, April 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9246-8.

Li:2020:CSP

[LZ20a] Xin Li and Luming Zhang. A conservative sine pseudospectral-difference method for multi-dimensional coupled Gross–Pitaevskii equations. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09769-z.

Li:2020:UOC

[LZ20b] Yuan Li and Chunfang Zhai. Unconditionally optimal convergence analysis of secondorder BDF Galerkin finite element scheme for a hybrid MHD system. Advances in Computational Mathematics, 46(5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09815-w.

Liang:2021:RPT

[LZ21] Zhao-Zheng Liang and Guo-Feng Zhang. Robust preconditioning techniques for multiharmonic finite element method with application to time-periodic parabolic optimal control problems. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09887-2.

Li:2022:HOC

[LZ22] Xin Li and Luming Zhang.

High-order conservative energy quadratization schemes for the Klein–Gordon–Schrödinger equation. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09962-2.

Liu:2021:CFE

[LZCW21] Huan Liu, Xiangcheng Zheng, Chuanjun Chen, and Hong Wang. A characteristic finite element method for the time-fractional mobile/ immobile advection diffusion model. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09867-6.

Li:2023:ADF

[LZL23] Jian Li, Jiyao Zeng, and Rui Li. An adaptive discontinuous finite volume element method for the Allen–Cahn equation. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10031-5.

Zhu:2003:ASS

[lZmCRX03] You lan Zhu, Bin mu Chen, Hongliang Ren, and Hanping Xu. Application of the singularity-separating method to American exotic option pricing. Advances in Computational Mathematics, 19(1–3):147– 158, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1022835722199. Chal-

lenges in computational mathematics (Pohang, 2001).

Lin:1998:AMI

[LZY98] Qun Lin, Shuhua Zhang, and Ningning Yan. An acceleration method for integral equations by using interpolation post-processing. Advances in Computational Mathematics, 9(1–2):117– 129, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018925103993. Nu-

merical treatment of boundary integral equations.

Lin:2009:SSP

[LZZ09] Yanping Lin, Kai Zhang, and Jun Zou. Studies on some perfectly matched layers for one-dimensional timedependent systems. Advances in Computational Mathematics, 30(1):1–35, January 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9055-2.

MacLeod:1994:TBT

[Mac94] Allan J. MacLeod. Tablebased tests for Bessel function software. Advances in Computational Mathematics, 2(2):251–260, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02521111.

Maiorov:2005:LBR

[Mai05] V. Maiorov. On lower bounds in radial basis approximation. Advances in Computational Mathematics, 22(2): 103–113, February 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1090-7.

Maiorov:2006:PDE

[Mai06]  V. Maiorov.  Pseudo-

dimension and entropy of manifolds formed by affineinvariant dictionary. Advances in Computational Mathematics, 25(4):435–450, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7645-9.

Maiorov:2010:GPR

[Mai10] Vitaly Maiorov. Geometric properties of the ridge function manifold. Advances in Computational Mathematics, 32(2):239–253, February 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9106-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        008-9106-3.pdf.

Maier:2020:ARP

[Mai20] L.-B. Maier. Ambient residual penalty approximation of partial differential equations on embedded submanifolds. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09802-1.

Marletta:1994:NSE

[Mar94] Marco Marletta. Numerical solution of eigenvalue problems for Hamiltonian systems. Advances in Computational Mathematics, 2(2): 155–184, March 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521106.

Maroni:1995:ISD

[Mar95] P. Maroni. An introduction to second degree forms. Advances in Computational Mathematics, 3(1–2):59–88, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02431996.

Matthies:2009:LPS

[Mat09]  Gunar Matthies.  Local

projection stabilisation for higher order discretisations of convection–diffusion problems on Shishkin meshes. Advances in Computational Mathematics, 30(4):315–337, May 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-008-9070-
                                        y.

Mazure:1998:VTD

[Maz98] Marie-Laurence Mazure. Vandermonde type determinants and blossoming. Advances in Computational Mathematics, 8(4):291–315, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018956616288.

Mazure:1999:CSP

[Maz99] Marie-Laurence Mazure. Chebyshev spaces with polynomial blossoms. Advances in Computational Mathematics, 10(3–4):219– 238, May 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018995019439.

Mazure:2001:CSB

[Maz01] Marie-Laurence Mazure. Chebyshev splines beyond total positivity. Advances in Computational Mathematics, 14(2): 129–156, February 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016616731472.

Mazure:2004:BOB

[Maz04] Marie-Laurence Mazure. Blossoms and optimal bases. Advances in Computational Mathematics, 20(1–3):177– 203, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025855123163. Ap-

proximation and applications.

Mazure:2007:PBA

[Maz07] Marie-Laurence Mazure. Pseudoaffinity, de Boor algorithm, and blossoms. Advances in Computational Mathematics, 26(1–3):305– 322, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7450-0.

Morton:1996:SBC

[MB96] K. W. Morton and N. A.

Burgess. The stability of boundary conditions for an angled-derivative difference scheme. Advances in Computational Mathematics, 6(1):263–279, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127707.  John

Crank 80th birthday special issue.

Malik:2020:FRM

[MB20] Osman Asif Malik and Stephen Becker. Fast randomized matrix and tensor interpolative decomposition using CountSketch. Advances in Computational Mathematics, 46(6):??, December 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09816-9.

Mao:2008:QNF

[MC08] Shipeng Mao and Shaochun Chen. A quadrilateral nonconforming finite element for linear elasticity problem. Advances in Computational Mathematics, 28(1): 81–100, January 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-006-9017-0.

Milovanovic:2010:NGQ

[MC10] Gradimir V. Milovanović and Aleksandar S. Cvetković. Nonstandard Gaussian quadrature formulae based on operator values. Advances in Computational Mathematics, 32 (4):431–486, May 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9114-y.

Miao:2023:FBC

[MC23] Cun-Qiang Miao and Lan Cheng. On flexible block Chebyshev–Davidson method for solving symmetric generalized eigenvalue problems. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10078-4.

Massar:2011:FCS

[MFB^{+}11] Melody L. Massar, Matthew Fickus, Erik Bryan, Douglas T. Petkie, and Andrew J. Terzuoli, Jr. Fast computation of spectral centroids. Advances in Computational Mathematics, 35 (1):83–97, July 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9167-y.

Mezzadri:2019:NLM

[MG19] Francesco Mezzadri and Emanuele Galligani. A nonlinearity lagging method for non-steady diffusion equations with nonlinear convection terms. Advances in Computational Mathematics, 45(3):1185– 1220, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-018-9652-2.

Marchand:2022:AGH

[MGSS22] P. Marchand, J. Galkowski, E. A. Spence, and A. Spence. Applying GMRES to the Helmholtz equation with strong trapping: how does the number of iterations depend on the frequency? Advances in Computational Mathematics, 48(4):??, August 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        022-09931-9.

Mustafa:2017:FUB

[MH17] Ghulam Mustafa and Rabia Hameed. Families of univariate and bivariate subdivision schemes originated from quartic B-spline. Advances in Computational Mathematics, 43(5):1131– 1161, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9519-y.

Maierhofer:2022:CAO

[MH22] Georg Maierhofer and Daan Huybrechs. Convergence analysis of oversampled collocation boundary element methods in 2D. Advances in Computational Mathematics, 48(2):??, April 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09924-8.

Mhaskar:1993:APM

[Mha93] H. N. Mhaskar. Approximation properties of a multilayered feedforward artificial neural network. Advances in Computational Mathematics, 1(1):61–80, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02070821.

Mohaoui:2021:TCB

[MHR21] Souad Mohaoui, Abdelilah Hakim, and Said Raghay. Tensor completion via bilevel minimization with fixedpoint constraint to estimate missing elements in noisy data. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09841-8.

Miyajima:2019:VCM

[Miy19]  Shinya Miyajima.  Ver-

ified computation of the matrix exponential. Advances in Computational Mathematics, 45(1):137–152, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9609-5.

Merchel:2023:ALR

[MJM23] Sandra Merchel, Bert Jüttler, and Dominik Mokris. Adaptive and local regularization for data fitting by tensor-product spline surfaces. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10035-1.

Macon:2002:RSC

[MKS02] Charles A. Macon, Leo C. Kempel, and Stephen W. Schneider. Radiation and scattering by complex conformal antennas on a circular cylinder. Advances in Computational Mathematics, 16(2–3):191–209, April 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1014425527687.

Modeling and computation in optics and electromagnetics.

Mason:1993:E

[MM93] John C. Mason and Charles A. Micchelli. Editorial. Advances in Computational Mathematics, 1(1):i, February 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/BF02070817.

Mason:1999:CE

[MM99] John C. Mason and Charles A.

Micchelli. The current Editor. Advances in Computational Mathematics, 11(2– 3):i, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018984211687.

Maiorov:2000:NOS

[MM00] V. E. Maiorov and R. Meir.

On the near optimality of the stochastic approximation of smooth functions by neural networks. Advances in Computational Mathematics, 13 (1):79–103, April 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018993908478.

Meng:2020:MVE

[MM20] Jian Meng and Liquan Mei.

A mixed virtual element method for the vibration problem of clamped Kirchhoff plate. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09810-1.

Moseley:2023:FBP

[MMNM23] Ben Moseley, Andrew Markham, and Tarje Nissen-Meyer. Finite basis physics-informed neural networks (FBPINNs): a scalable domain decomposition approach for solving differential equations. Advances in Computational Mathematics, 49(4):??, August 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-023-10065-
9.

Micchelli:2013:RSS

[MMP13] Charles A. Micchelli, Jean M. Morales, and Massimiliano Pontil. Regularizers for structured sparsity. Advances in Computational Mathematics, 38(3): 455–489, April 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9245-9.

Morigi:2000:SRC

[MN00] Serena Morigi and Marian Neamtu. Some results for a class of generalized polynomials. Advances in Computational Mathematics, 12(2–3): 133–149, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018908917139.

Manzoni:2015:HSA

[MN15] Andrea Manzoni and Federico Negri. Heuristic strategies for the approximation of stability factors in quadratically nonlinear parametrized PDEs. Advances in Computational Mathematics, 41(5):1255– 1288, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9413-4.

Mallik:2016:CFE

[MN16] Gouranga Mallik and Neela Nataraj. Conforming finite element methods for the von Kármán equations. Advances in Computational Mathematics, 42(5):1031– 1054, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9452-5.

Mustapha:2016:CSA

[MNC16] Kassem Mustapha, Maher Nour, and Bernardo Cockburn. Convergence and superconvergence analyses of HDG methods for time fractional diffusion problems. Advances in Computational Mathematics, 42(2): 377–393, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-015-9428-x.

Mukherjee:2006:LTS

[MNPR06] Sayan Mukherjee, Partha Niyogi, Tomaso Poggio, and Ryan Rifkin. Learning theory: stability is sufficient for generalization and necessary and sufficient for consistency of empirical risk minimization. Advances in Computational Mathematics, 25(1– 3):161–193, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7634-z.

Mhaskar:2000:PFS

[MNPW00] H. N. Mhaskar, F. J. Narcowich, J. Prestin, and J. D. Ward. Polynomial frames on the sphere. Advances in Computational Mathematics, 13(4):387–403, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016639802349.

Meinardus:1996:BSA

[MNW96] G. Meinardus, G. Nürnberger, and G. Walz. Bivariate segment approximation and splines. Advances in Computational Mathematics, 6(1): 25–45, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/BF02127694.

Mhaskar:1999:APZ

[MNW99] H. N. Mhaskar, F. J. Narcowich, and J. D. Ward. Approximation properties of zonal function networks using scattered data on the sphere. Advances in Computational Mathematics, 11(2–3):121– 137, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018967708053. Ra-

dial basis functions and their applications.

Mhaskar:2004:RBD

[MNW04] H. N. Mhaskar, F. J. Narcowich, and J. D. Ward. On the representation of banddominant functions on the sphere using finitely many bits. Advances in Computational Mathematics, 21(1– 2):127–146, July 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016434.
97920.1f.

Marthinsen:2016:GIN

[MO16] Håkon Marthinsen and Brynjulf Owren. Geometric integration of non-autonomous linear Hamiltonian problems. Advances in Computational Mathematics, 42(2):

313–332, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9425-0.

Matthies:2019:APM

[MO19] Hermann G. Matthies and Roger Ohayon. Analysis of parametric models. Advances in Computational Mathematics, 45(5-6):2555– 2586, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09735-4.

Monterde:2009:CHP

[Mon09] J. Monterde. A characterization of helical polynomial curves of any degree. Advances in Computational Mathematics, 30(1): 61–78, January 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9063-x.

Moosmuller:2017:HSM

[Moo17] Caroline Moosmüller. Hermite subdivision on manifolds via parallel transport. Advances in Computational Mathematics, 43(5):1059– 1074, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.

1007/s10444-017-9516-1;
http://link.springer.com/
content/pdf/10.1007/s10444-
017-9516-1.pdf.

Mossner:2010:JSR

[Möß10]  Bernhard Mößner.  On

the joint spectral radius of matrices of order 2 with equal spectral radius. Advances in Computational Mathematics, 33(2):243–254, August 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9130-y.

Mortreux:1996:APA

[MP96] P. Mortreux and M. Prévost.

An acceleration property of the E-algorithm for alternate sequences. Advances in Computational Mathematics, 5(1):443–482, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124755.

Mhaskar:2000:DSP

[MP00] H. N. Mhaskar and J. Prestin.

On the detection of singularities of a periodic function. Advances in Computational Mathematics, 12(2–3): 95–131, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1023/A:1018921319865.

Mainar:2004:QCC

[MP04] E. Mainar and J. M. Peña. Quadratic-cycloidal curves. Advances in Computational Mathematics, 20(1–3):161– 175, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025813919473. Ap-

proximation and applications.

Mergia:2018:FSE

[MP18] Woinshet D. Mergia and Kailash C. Patidar. Fractionalstep θ -method for solving singularly perturbed problem in ecology. Advances in Computational Mathematics, 44 (3):645–671, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9554-8.

Mixon:2023:MFR

[MP23] Dustin G. Mixon and Daniel Packer. Max filtering with reflection groups. Advances in Computational Mathematics, 49(6):??, December 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10084-6.

Mainar:2022:ACM

[MPR22] E. Mainar, J. M. Peña, and B. Rubio. Accurate computations with matrices related to bases t^{i}e^{λt}. Advances in Computational Mathematics, 48(4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09954-2.

Marcellan:1996:OPL

[MPS96] Francisco Marcellán, Franz Peherstorfer, and Robert Steinbauer. Orthogonality properties of linear combinations of orthogonal polynomials. Advances in Computational Mathematics, 5(1):281–295, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124748.

Marcellan:1997:OPL

[MPS97] Francisco Marcellán, Franz Peherstorfer, and Robert Steinbauer. Orthogonality properties of linear combinations of orthogonal polynomials. II. Advances in Computational Mathematics, 7(3): 401–428, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018963323132.

Marusic:1996:CMS

[MR96] Miljenko Marušić and Mladen Rogina. A collocation method for singularly perturbed two-point boundary value problems with splines in tension. Advances in Computational Mathematics, 6(1): 65–76, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127696.

Massey:2010:MCF

[MR10] Pedro Massey and Mariano Ruiz. Minimization of convex functionals over frame operators. Advances in Computational Mathematics, 32(2): 131–153, February 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9092-5.

Madden:2015:MSG

[MR15] Niall Madden and Stephen Russell. A multiscale sparse grid finite element method for a two-dimensional singularly perturbed reaction– diffusion problem. Advances in Computational Mathematics, 41(6):1423–1450, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9395-7.

Martini:2015:RBA

[MRH15] Immanuel Martini, Gianluigi Rozza, and Bernard Haasdonk. Reduced basis approximation and a-posteriori error estimation for the coupled Stokes–Darcy system. Advances in Computational Mathematics, 41(5):1131– 1157, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9396-6.

Matache:2003:HOC

[MRS03] Ana-Maria Matache, Tomáš Roubíček, and Christoph Schwab. Higher-order convex approximations of Young measures in optimal control. Advances in Computational Mathematics, 19(1– 3):73–97, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022883930432. Chal-

lenges in computational mathematics (Pohang, 2001).

Massey:2014:OFC

[MRS14] Pedro G. Massey, Mariano A. Ruiz, and Demetrio Stojanoff. Optimal frame completions. Advances in Computational Mathematics, 40(5–6):1011–1042, December 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9339-7.

Massey:2018:FCP

[MRS18] Pedro G. Massey, Noelia B. Rios, and Demetrio Stojanoff. Frame completions with prescribed norms: local minimizers and applications. Advances in Computational Mathematics, 44(1): 51–86, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9535-y.

Marcati:2022:TRB

[MRS22] C. Marcati, M. Rakhuba, and Ch. Schwab. Tensor rank bounds for point singularities in r^{3}. Advances in Computational Mathematics, 48(3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09925-7.

Micchelli:1997:RM

[MS97] C. A. Micchelli and Thomas Sauer. Regularity of multiwavelets. Advances in Computational Mathematics, 7(4): 455–545, August 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018971524949.

Mund:1998:PGM

[MS98] Patrick Mund and Ernst P.

Stephan. The preconditioned GMRES method for systems of coupled FEM– BEM equations. Advances in Computational Mathematics, 9(1–2):131– 144, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018921003084. Nu-

merical treatment of boundary integral equations.

McLean:1999:BEP

[MS99] W. McLean and O. Steinbach. Boundary element preconditioners for a hypersingular integral equation on an interval. Advances in Computational Mathematics, 11(4):271–286, December 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018944530343.

Moller:2000:BPI

[MS00] H. Michael Möller and Thomas Sauer. H-bases for polynomial interpolation and system solving. Advances in Computational Mathematics, 12(4):335–362, March 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1023/A:1018937723499. Mul-

tivariate polynomial interpolation.

Moller:2004:MRF

[MS04] H. Michael Möller and Thomas Sauer. Multivariate refinable functions of high approximation order via quotient ideals of Laurent polynomials. Advances in Computational Mathematics, 20(1–3):205– 228, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1025889132677. Ap-

proximation and applications.

Manni:2007:QSQ

[MS07] Carla Manni and Paul Sablonnière. Quadratic spline quasi-interpolants on Powell–Sabin partitions. Advances in Computational Mathematics, 26(1–3):283– 304, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-006-9025-0.

Merrien:2012:HSS

[MS12] Jean-Louis Merrien and Tomas Sauer. From Hermite to stationary subdivision schemes in one and several variables. Advances in Computational Mathematics, 36 (4):547–579, May 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9190-7.

Munch:2016:MFD

[MS16] Arnaud Münch and Diego A.

Souza. A mixed formulation for the direct approximation of L^{2}-weighted controls for the linear heat equation. Advances in Computational Mathematics, 42(1): 85–125, February 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9412-5.

Mao:2018:SEM

[MS18] Zhiping Mao and Jie Shen.

Spectral element method with geometric mesh for twosided fractional differential equations. Advances in Computational Mathematics, 44 (3):745–771, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9561-9.

Meng:2019:CAA

[MS19] Xiangyun Meng and Martin Stynes. Convergence analysis of the Adini element on a Shishkin mesh for a singularly perturbed fourthorder problem in two dimensions. Advances in Computational Mathematics, 45(2): 1105–1128, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9646-0.

Mahata:2022:NDO

[MS22] Shantiram Mahata and Rajen Kumar Sinha. Nonsmooth data optimal error estimates by energy arguments for subdiffusion equations with memory. Advances in Computational Mathematics, 48(4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09967-x.

Mehra:2021:ASG

[MSL21] Mani Mehra, Ankita Shukla, and Günter Leugering. An adaptive spectral graph wavelet method for PDEs on networks. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09824-9.

Mazroui:2002:ECS

[MSS02] A. Mazroui, D. Sbibih, and P. Sablonnière. Existence and construction of H_{1}-splines of class C^{k} on a three directional mesh. Advances in Computational Mathematics, 17(3):167–198, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016066825676.

Maass:2014:SRU

[MSST14] Peter Maass, Chen Sagiv, Hans-Georg Stark, and Bruno Torresani. Signal representation, uncertainty principles and localization measures. Advances in Computational Mathematics, 40(3): 597–607, June 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9341-8.

Mazroui:2005:SMS

[MST05] A. Mazroui, D. Sbibih, and A. Tijini. A simple method for smoothing functions and compressing Hermite data. Advances in Computational Mathematics, 23(3):279–297, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1783-y.

Ma:2022:SCF

[MSW^{+}22] Xijun Ma, Chungen Shen, Li Wang, Lei-Hong Zhang, and Ren-Cang Li. A selfconsistent-field iteration for MAXBET with an application to multi-view feature extraction. Advances in Computational Mathematics, 48 (2):??, April 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09929-3.

Micchelli:2013:PAL

[MSXZ13] Charles A. Micchelli, Lixin Shen, Yuesheng Xu, and Xueying Zeng. Proximity algorithms for the L1/ TV image denoising model. Advances in Computational Mathematics, 38(2):401–426, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9243-y.

Mao:2006:AFE

[MSZ06] Dong Mao, Lihua Shen, and Aihui Zhou. Adaptive finite element algorithms for eigenvalue problems based on local averaging type a posteriori error estimates. Advances in Computational Mathematics, 25(1–3):135– 160, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7617-0.

Mehlum:1998:ISM

[MT98] Even Mehlum and Christian Tarrou. Invariant smoothness measures for surfaces. Advances in Computational Mathematics, 8(1–2):49–63, January 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018931910836.

Muoi:2021:EOF

[MTH21] Pham Quy Muoi, Wee Chin Tan, and Viet Ha Hoang. Essentially optimal finite elements for multiscale elliptic eigenvalue problems. Advances in Computational Mathematics, 47(6):??, December 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09903-5.

Mikula:2014:NTS

[MU14] Karol Mikula and Jozef Urbán. A new tangentially stabilized 3D curve evolution algorithm and its application in virtual colonoscopy. Advances in Computational Mathematics, 40(4):819–837, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9328-x.

Muir:1999:ODC

[Mui99] P. H. Muir. Optimal discrete and continuous monoimplicit Runge–Kutta schemes for BVODEs. Advances in Computational Mathematics, 10(2):135–167, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1018926631734.

Muradova:2008:SMN

[Mur08] Aliki D. Muradova. The spectral method and numerical continuation algorithm for the von Kármán problem with postbuckling behaviour of solutions. Advances in Computational Mathematics, 29(2):179–206, August 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9050-7.

Mason:1999:CPM

[MV99] J. C. Mason and E. Venturino. A Chebyshev polynomial method for line integrals with singularities. Advances in Computational Mathematics, 10(2):187–208, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018978615805.

Meister:2001:EPL

[MV01] Andreas Meister and Christof Vömel. Efficient preconditioning of linear systems arising from the discretization of hyperbolic conservation laws. Advances in Computational Mathematics, 14(1): 49–73, January 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016645505973.

Melenk:2001:RBP

[MW01] J. M. Melenk and B. I.

Wohlmuth. On residualbased a posteriori error estimation in hp–FEM. Advances in Computational Mathematics, 15(1–4):311– 331, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014268310921. A

posteriori error estimation and adaptive computational methods.

Micchelli:2013:AAP

[MWW13] Charles A. Micchelli, Jianzhong Wang, and Yi Wang. On an asymptotic analysis of polynomial approximation to halfband filters. Advances in Computational Mathematics, 38(3):601–622, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9251-
y.

Ma:2023:MEC

[MWZZ23] Shu Ma, Jilu Wang, Mingyan Zhang, and Zhimin Zhang. Mass- and energy-conserving Gauss collocation methods for the nonlinear Schrödinger equation with a wave operator. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10077-5.

Melenk:2013:ARS

[MXO13] J. M. Melenk, C. Xenophontos, and L. Oberbroeckling. Analytic regularity for a singularly perturbed system of reaction–diffusion equations with multiple scales. Advances in Computational Mathematics, 39(2):367–394, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9284-x.

Micchelli:2013:CHS

[MXY13] Charles A. Micchelli, Yuesheng Xu, and Bo Yu. On computing with the Hilbert spline transform. Advances in Computational Mathematics, 38 (3):623–646, April 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9252-x.

Micchelli:2008:TIO

[MXZ08] Charles A. Micchelli, Yuesheng Xu, and Haizhang Zhang. On translation invariant operators which preserve the B-spline recurrence. Advances in Computational Mathematics, 28(2):157–169, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9018-z.

Mu:2017:SFE

[MY17] Lin Mu and Xiu Ye. A simple finite element method for the Stokes equations. Advances in Computational Mathematics, 43(6):1305– 1324, December 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9526-z.

Micchelli:2013:DR

[MZ13] Charles A. Micchelli and Tao Zhao. On discrete ℓ^{1}- regularization. Advances in Computational Mathematics, 38(2):441–454, February 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-011-9247-
7.

Mao:2022:AFK

[MZ22] Tong Mao and Ding-Xuan Zhou. Approximation of functions from Korobov spaces by deep convolutional neural networks. Advances in Computational Mathematics, 48(6):??, December 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09991-x.

Nair:2012:QBC

[Nai12] M. Thamban Nair. Quadrature based collocation methods for integral equations of the first kind. Advances in Computational Mathematics, 36(2):315–329, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9196-1.

Neytcheva:2011:FEB

[NBL11] Maya Neytcheva, Erik Bängtsson, and Elisabeth Linnér. Finiteelement based sparse approximate inverses for blockfactorized preconditioners. Advances in Computational Mathematics, 35(2–4):323– 355, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9176-5.

Vu:2013:SAD

[nCV13] B. ng Công Vũ. A splitting algorithm for dual monotone inclusions involving cocoercive operators. Advances in Computational Mathematics, 38(3): 667–681, April 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9254-8.

Nees:1996:CAD

[Nee96]  Manuela Nees.  Cheby-

shev approximation by discrete superposition. application to neural networks. Advances in Computational Mathematics, 5(1):137–151, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124739.

Nestler:2016:APT

[Nes16]  Franziska Nestler.  Au-

tomated parameter tuning based on RMS errors for nonequispaced FFTs. Advances in Computational Mathematics, 42(4):889–919, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9446-8.

Newby:1996:JCH

[New96] J. C. Newby. John Crank: his association with Brunel University. Advances in Computational Mathematics, 6(1): ix–x, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). John Crank 80th birthday special issue.

Niyogi:1999:GBF

[NG99] Partha Niyogi and Federico Girosi. Generalization bounds for function approximation from scattered noisy data. Advances in Computational Mathematics, 10(1): 51–80, January 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018966213079.

Niessen:1998:ENR

[Nie98] Gero Nießen. An explicit norm representation for the analysis of multilevel methods. Advances in Computational Mathematics, 9(3– 4):311–335, November 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018949809444.

Nielsen:2007:PSS

[Nie07] Morten Nielsen. On polynomial symbols for subdivision schemes. Advances in Computational Mathematics, 27(2):195–209, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7476-3.

Nitsche:2022:CTR

[Nit22] Monika Nitsche. Corrected trapezoidal rule for nearsingular integrals in axisymmetric Stokes flow. Advances in Computational Mathematics, 48(5):??, October 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09973-z.

Nikl:2019:WBL

[NKL^{+}19] Jan Nikl, Milan Kuchařík, Jiří Limpouch, Richard Liska, and Stefan Weber. Wave-based laser absorption method for high-order transport-hydrodynamic codes. Advances in Computational Mathematics, 45(4):1953– 1976, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09671-3.

Noakes:1996:NCB

[Noa96] J. Lyle Noakes. Networks and closed balls. Advances in Computational Mathematics, 5(1):153–161, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/BF02124740.

Noakes:1998:NCC

[Noa98]  Lyle Noakes.  Nonlinear

corner-cutting. Advances in Computational Mathematics, 8(4):165–177, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018940112654.

Noakes:2002:QIS

[Noa02] Lyle Noakes. Quadratic interpolation on spheres. Advances in Computational Mathematics, 17(4):385–395, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016277023669.

Noakes:2006:DRC

[Noa06] Lyle Noakes. Duality and Riemannian cubics. Advances in Computational Mathematics, 25(1–3):195– 209, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7621-4.

Noakes:2014:MAR

[Noa14]  Lyle Noakes.  Minimum

L^{∞} accelerations in Riemannian manifolds. Advances in Computational Mathematics, 40(4):839–863, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9329-9.

Noelle:1995:CHO

[Noe95] Sebastian Noelle. Convergence of higher order finite volume schemes on irregular grids. Advances in Computational Mathematics, 3(3): 197–218, April 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02431999.

Novak:2000:HDI

[Nov00] Erich Novak. High dimensional integration. Advances in Computational Mathematics, 12(1):1–2, January 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1023/A:1018996514413.

Navas-Palencia:2018:HPC

[NP18] Guillermo Navas-Palencia.

High-precision computation of the confluent hypergeometric functions via Franklin– Friedman expansion. Advances in Computational Mathematics, 44(3):841–859, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
017-9565-5.

Nouy:2024:DBM

[NP24] Anthony Nouy and Alexandre Pasco. Dictionary-based model reduction for state estimation. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10129-4.

Nasser:2022:PDN

[NRR^{+}22] Mohamed Nasser, Oona Rainio, Antti Rasila, Matti Vuorinen, Terry Wallace, Hang Yu, and Xiaohui Zhang. Polycircular domains, numerical conformal mappings, and moduli of quadrilaterals. Advances in Computational Mathematics, 48(5): ??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09975-x.

Neamtu:2004:AOS

[NS04] Marian Neamtu and Larry L. Schumaker. On the approximation order of splines on spherical triangulations. Advances in Computational Mathematics, 21(1–2):3–20, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/B:ACOM.0000016430.
93487.ec.

Needham:2021:SGC

[NS21] Tom Needham and Clayton Shonkwiler. Symplectic geometry and connectivity of spaces of frames. Advances in Computational Mathematics, 47(1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09842-7.

Nie:2024:NMF

[NSD24] Daxin Nie, Jing Sun, and Weihua Deng. Numerical methods for forward fractional Feynman–Kac equation. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10152-5.

Nouisser:2004:PBS

[NSS04] O. Nouisser, D. Sbibih, and P. Sablonnière. Pairs of Bsplines with small support on the four-directional mesh generating a partition of unity. Advances in Computational Mathematics, 21(3–4): 317–355, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1023/B:ACOM.0000032048.
                                        57043.dc.

Narcowich:1998:SRS

[NSW98] F. J. Narcowich, N. Sivakumar, and J. D. Ward. Stability results for scattered-data interpolation on Euclidean spheres. Advances in Computational Mathematics, 8(4): 137–163, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018996230401.

Narcowich:2007:APR

[NSW07] Francis J. Narcowich, Xinping Sun, and Joseph D. Ward. Approximation power of RBFs and their associated SBFs: a connection. Advances in Computational Mathematics, 27(1): 107–124, July 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7506-1.

Nuyens:2016:RLR

[NSW16] Dirk Nuyens, Gowri Suryanarayana, and Markus Weimar. Rank-1 lattice rules for multivariate integration in spaces of permutation-invariant functions: error bounds and tractability. Advances in Computational Mathematics, 42(1):55–84, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9411-
6.

Nurnberger:2004:LIC

[NSZ04] Günther Nürnberger, Larry L.

Schumaker, and Frank Zeilfelder. Lagrange interpolation by C^{1} cubic splines on triangulated quadrangulations. Advances in Computational Mathematics, 21(3–4): 357–380, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032044.
49282.8a.

Ndjansi:2023:ECI

[NT23] Lionel Ouya Ndjansi and Laurent Tchoualag. An efficient computation of the inverse of the single layer matrix for the resolution of the linear elasticity problem in BEM. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10039-x.

Novak:2017:COI

[NUWZ17] Erich Novak, Mario Ullrich, Henryk Woźniakowski, and Shun Zhang. Complexity of oscillatory integrals on the real line. Advances in Computational Mathematics, 43 (3):537–553, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9496-6.

Nicaise:2010:QED

[NV10] Serge Nicaise and Julie Valein. Quasi exponential decay of a finite difference space discretization of the 1-d wave equation by pointwise interior stabilization. Advances in Computational Mathematics, 32(3): 303–334, April 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9108-1.

Narcowich:2004:PCA

[NW04] Francis J. Narcowich and Joseph D. Ward. Preface [Computational and approximation on the sphere]. Advances in Computational Mathematics, 21(1–2):1–2, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/B:ACOM.0000016666.
                                        68274.5d.

Nannen:2022:CSI

[NW22] Lothar Nannen and Markus Wess. Complex-scaled infinite elements for resonance problems in heterogeneous open systems. Advances in Computational Mathematics, 48(2):??, April 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09923-1.

Nurnberger:2004:LIB

[NZ04] Günther Nürnberger and Frank Zeilfelder. Lagrange interpolation by bivariate C^{1}-splines with optimal approximation order. Advances in Computational Mathematics, 21(3–4):381– 419, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032043.
07621.62.

Noakes:2022:FGJ

[NZ22] Lyle Noakes and Erchuan Zhang. Finding geodesics joining given points. Advances in Computational Mathematics, 48(4):??, August 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
022-09966-y.

Ober-Blobaum:2015:CAH

[OBS15] Sina Ober-Blöbaum and Nils Saake. Construction and analysis of higher order Galerkin variational integrators. Advances in Computational Mathematics, 41(6):1391–1422, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9394-8.

Ockendon:1996:RCG

[Ock96] J. R. Ockendon. The role of the Crank–Gupta model in the theory of free and moving boundary problems. Advances in Computational Mathematics, 6(1):281–293, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127708.  John

Crank 80th birthday special issue.

ONeil:2018:SKI

[O’N18a] Michael O’Neil. Second-kind integral equations for the Laplace–Beltrami problem on surfaces in three dimensions. Advances in Computational Mathematics, 44(5):1385– 1409, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9587-7.

Ottosen:2018:NAN

[ON18b] Emil Solsbæk Ottosen and Morten Nielsen. Nonlinear approximation with nonstationary Gabor frames. Advances in Computational Mathematics, 44(4):1183– 1203, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9577-1.

Onchis:2014:OFM

[Onc14] Darian M. Onchis. Optimized frames and multi-dimensional challenges in time-frequency analysis. Advances in Computational Mathematics, 40 (3):703–709, June 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9332-1.

ORiordan:2011:PUN

[OP11] E. O’Riordan and M. L.

Pickett. A parameteruniform numerical method for a singularly perturbed two parameter elliptic problem. Advances in Computational Mathematics, 35 (1):57–82, July 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9164-1.

Opfer:2006:MK

[Opf06] Roland Opfer. Multiscale kernels. Advances in Computational Mathematics, 25 (4):357–380, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-7622-
3.

ORiordan:2009:NAS

[OS09] Eugene O’Riordan and Martin Stynes. Numerical analysis of a strongly coupled system of two singularly perturbed convection– diffusion problems. Advances in Computational Mathematics, 30(2):101–121, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9058-z.

Oswald:2004:SNM

[Osw04]  Peter Oswald.  Smooth-

ness of nonlinear medianinterpolation subdivision. Advances in Computational Mathematics, 20(4):401–423, May 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1023/A:1027315032100.

Pan:1995:AAA

[Pan95] Victor Y. Pan. An algebraic approach to approximate evaluation of a polynomial on a set of real points. Advances in Computational Mathematics, 3(1–2): 41–58, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF02431995.

Park:2020:ODD

[Par20] Jongho Park. An overlapping domain decomposition framework without dual formulation for variational imaging problems. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09799-7.

Park:2022:FGM

[Par22] Jongho Park. Fast gradient methods for uniformly convex and weakly smooth problems. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09943-5.

Paulsen:2019:TCB

[Pau19]  William Paulsen.  Tetra-

tion for complex bases. Advances in Computational Mathematics, 45(1):243–267, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9615-7.

Paulsen:2017:SCP

[PC17] William Paulsen and Samuel Cowgill. Solving F (z + 1) = b^{F (z)} in the complex plane. Advances in Computational Mathematics, 43(6):1261– 1282, December 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9524-1.

Pericleous:1996:FSN

[PCM^{+}96] K. A. Pericleous, M. Cross, G. Moran, P. Chow, and K. S. Chan. Free surface Navier–Stokes flows with simultaneous heat transfer and solidification/melting. Advances in Computational Mathematics, 6(1):295–308, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127709.  John

Crank 80th birthday special issue.

Przybilla:2024:SAD

[PDB24] Jennifer Przybilla, Igor Pontes Duff, and Peter Benner. Semi-active damping optimization of vibrational systems using the reduced basis method. Advances in Computational Mathematics, 50(3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-024-10141-8.

Peixin:2005:CCI

[Pei05] Ye Peixin. Computational complexity of the integration problem for anisotropic classes. Advances in Computational Mathematics, 23 (4):375–392, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-004-1830-
8.

Penzl:1998:NSG

[Pen98] Thilo Penzl. Numerical solution of generalized Lyapunov equations. Advances in Computational Mathematics, 8(1– 2):33–48, January 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018979826766.

Pena:2011:DDS

[Peñ11] Juan Manuel Peña. Diagonal dominance, Schur complements and some classes of H-matrices and P -matrices. Advances in Computational Mathematics, 35(2–4):357– 373, November 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9160-5.

Peralta:2022:EEM

[Per22] Gilbert Peralta. Error estimates for mixed and hybrid FEM for elliptic optimal control problems with penalizations. Advances in Computational Mathematics, 48 (6):??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09980-0.

Peters:1996:IRC

[Pet96] Jörg Peters. Interpolation regions for convex cubic curve segments. Advances in Computational Mathematics, 6(1): 87–96, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127698.

Petras:2000:SCE

[Pet00] Knut Petras. On the Smolyak cubature error for analytic functions. Advances in Computational Mathematics, 12 (1):71–93, January 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018904816230. High

dimensional integration.

Petzoldt:2002:PEE

[Pet02] Martin Petzoldt. A posteriori error estimators for elliptic equations with discontinuous coefficients. Advances in Computational Mathematics, 16(1):47–75, January 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014221125034.

Pu:2023:NSF

[PF23] Tianyi Pu and Marco Fasondini. The numerical solution of fractional integral equations via orthogonal polynomials in fractional powers. Advances in Computational Mathematics, 49(1): ??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-10009-9.

Pelosi:2005:GHI

[PFMS05] Francesca Pelosi, Rida T.

Farouki, Carla Manni, and Alessandra Sestini. Geometric Hermite interpolation by spatial Pythagoreanhodograph cubics. Advances in Computational Mathematics, 22(4):325–352, May 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-2599-
x.

Pottmann:2010:EOM

[PGB10] Helmut Pottmann, Philipp Grohs, and Bernhard Blaschitz. Edge offset meshes in Laguerre geometry. Advances in Computational Mathematics, 33(1):45–73, July 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-009-9119-
6.

Peherstorfer:2015:RMS

[PGB15] Benjamin Peherstorfer, Pablo Gómez, and Hans-Joachim Bungartz. Reduced models for sparse grid discretizations of the multiasset Black–Scholes equation. Advances in Computational Mathematics, 41(5):1365– 1389, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9421-4.

Pottmann:2009:LMS

[PGM09] Helmut Pottmann, Philipp Grohs, and Niloy J. Mitra. Laguerre minimal surfaces, isotropic geometry and linear elasticity. Advances in Computational Mathematics, 31(4):391–419, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9076-5.

Pinkus:2004:SPD

[Pin04]  Allan Pinkus.  Strictly

positive definite functions on a real inner product space. Advances in Computational Mathematics, 20(4): 263–271, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1027362918283.

Pitolli:2016:BSN

[Pit16]  Francesca Pitolli.  Bell-

shaped nonstationary refinable ripplets. Advances in Computational Mathematics, 42(6):1427–1451, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9469-9.

PazCalvo:2001:SEM

[PIZ01] Mari Paz Calvo, Arieh Iserles, and Antonella Zanna. Semiexplicit methods for isospectral flows. Advances in Computational Mathematics, 14 (1):1–24, January 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016635812817.

Puligilla:2020:AEE

[PJ20] Shivakanth Chary Puligilla and Balaji Jayaraman. Assessment of end-to-end and sequential data-driven learning for non-intrusive modeling of fluid flows. Advances in Computational Mathematics, 46(4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09753-7.

Prautzsch:1994:CSD

[PK94] Hartmut Prautzsch and Leif Kobbelt. Convergence of subdivision and degree elevation. Advances in Computational Mathematics, 2(1): 143–154, January 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02519040.

Piquemal:2005:NWP

[PL05] Anne-Sophie Piquemal and Jacques Liandrat. A new wavelet preconditioner for finite difference operators. Advances in Computational Mathematics, 22(2):125–163, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1111-6.

Plantevin:1995:WIM

[Pla95] Frédérique Plantevin. Wavelets on irregular meshes. Advances in Computational Mathematics, 4(1–2):293– 329, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123479.

Plaskota:2000:EDS

[Pla00] Leszek Plaskota. The exponent of discrepancy of sparse grids is at least 2.1933. Advances in Computational Mathematics, 12 (1):3–24, January 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018900715321. High

dimensional integration.

Plato:2012:RPC

[Pla12] Robert Plato. The regularizing properties of the composite trapezoidal method for weakly singular Volterra integral equations of the first kind. Advances in Computational Mathematics, 36(2): 331–351, February 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9217-0.

Plonka:1995:TSS

[Plo95]  Gerlind Plonka.  Two-

scale symbol and autocorrelation symbol for B-splines with multiple knots. Advances in Computational Mathematics, 3(1–2):1–22, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02431993.

Perez-Lopez:2022:NAC

[PLRGVR22] Jhean E. Pérez-López, Diego A. Rueda-Gómez, and Élder J. Villamizar-Roa. Numerical analysis of a chemotaxis model for tumor invasion. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09949-z.

Pandey:2017:HAS

[PM17] Rishi Kumar Pandey and Hradyesh Kumar Mishra. Homotopy analysis Sumudu transform method for timefractional third order dispersive partial differential equation. Advances in Computational Mathematics, 43(2): 365–383, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9489-5.

Peherstorfer:2019:TBM

[PM19] Benjamin Peherstorfer and Youssef Marzouk. A transportbased multifidelity preconditioner for Markov chain Monte Carlo. Advances in Computational Mathematics, 45(5-6):2321–2348, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09711-y.

Pottmann:1995:CDR

[Pot95] Helmut Pottmann. Curve design with rational Pythagoreanhodograph curves. Advances in Computational Mathematics, 3(1–2):147– 170, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03028365.

Potts:2004:ASD

[Pot04] Daniel Potts. Approximation of scattered data by trigonometric polynomials on the torus and the 2sphere. Advances in Computational Mathematics, 21 (1–2):21–36, July 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016432.
97333.e8.

Pouly:1997:AFE

[PP97] L. Pouly and J. Pousin.

Adaptive finite element for semi-linear convection–diffusion problems. Advances in Computational Mathematics, 7(3):

235–259, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018946919497.

Poveda:2023:PEE

[PP23] Leonardo A. Poveda and Pedro Peixoto. On pointwise error estimates for Voronoïbased finite volume methods for the Poisson equation on the sphere. Advances in Computational Mathematics, 49(3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10041-3.

Pintore:2021:ECB

[PPH^{+}21] Moreno Pintore, Federico Pichi, Martin Hess, Gianluigi Rozza, and Claudio Canuto. Efficient computation of bifurcation diagrams with a deflated approach to reduced basis spectral element method. Advances in Computational Mathematics, 47(1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09827-6.

Prautzsch:1999:DEP

[PR99] Hartmut Prautzsch and Ulrich Reif. Degree estimates for C^{k}-piecewise polynomial subdivision surfaces.

Advances in Computational Mathematics, 10(2):209–217, February 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018922530826.

Pilarczyk:2015:CCH

[PR15] Paweł Pilarczyk and Pedro Real. Computation of cubical homology, cohomology, and (co)homological operations via chain contraction. Advances in Computational Mathematics, 41(1): 253–275, February 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9356-1.

Petersen:2019:APH

[PR19] Philipp Petersen and Mones Raslan. Approximation properties of hybrid shearletwavelet frames for Sobolev spaces. Advances in Computational Mathematics, 45(3): 1581–1606, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09679-9.

Prautzsch:1998:SSS

[Pra98] Hartmut Prautzsch. Smoothness of subdivision surfaces at extraordinary points. Advances in Computational Mathematics, 9(3–4):377– 389, November 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018945708536.

Pradovera:2023:TCG

[Pra23] Davide Pradovera. Toward a certified greedy Loewner framework with minimal sampling. Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10091-7.

Prevost:1994:APA

[Pré94]  Marc Prévost.  Accelera-

tion property for the Ealgorithm and an application to the summation of series. Advances in Computational Mathematics, 2(3): 319–341, April 1994. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521114.

Peiris:2024:ANN

[PRS24] Vinesha Peiris, Vera Roshchina, and Nadezda Sukhorukova. Artificial neural networks with uniform norm-based loss functions. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10124-9.

Prossdorf:1998:MAM

[PS98] Siegfried Prössdorf and Jörg Schult. Multiwavelet approximation methods for pseudodifferential equations on curves. Stability and convergence analysis. Advances in Computational Mathematics, 9(1–2):145– 171, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018977120831. Nu-

merical treatment of boundary integral equations.

Pousin:2005:PEE

[PS05] J. Pousin and T. Sassi. A posteriori error estimates and domain decomposition with nonmatching grids. Advances in Computational Mathematics, 23(3):241–263, October 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1779-7.

Pena:2007:EPR

[PS07] Juan Manuel Peña and Tomas Sauer. Efficient polynomial reduction. Advances in Computational Mathematics, 26(1–3):323– 336, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7208-0.

Patel:2023:RSS

[PS23] Dhiraj Patel and S. Sivananthan. Random sampling of signals concentrated on compact set in localized reproducing kernel subspace of L^{p}(R^{n}). Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10075-7.

Pradhan:2011:RTN

[PSNP11] Debasish Pradhan, Baskar Shalini, Neela Nataraj, and Amiya K. Pani. A Robintype non-overlapping domain decomposition procedure for second order elliptic problems. Advances in Computational Mathematics, 34(4): 339–368, May 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9157-0.

Plonka:1995:CWB

[PST95] Gerlind Plonka, Kathi Selig, and Manfred Tasche. On the construction of wavelets on a bounded interval. Advances in Computational Mathematics, 4(1–2):357–388, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123481.

Plonka:2008:NSB

[PST08] Gerlind Plonka, Hagen Schumacher, and Manfred Tasche. Numerical stability of biorthogonal wavelet transforms. Advances in Computational Mathematics, 29(1):1–25, July 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-007-9034-
7.

Pietschmann:2024:NIA

[PSW24] Jan-Frederik Pietschmann, Ailyn Stötzner, and Max Winkler. Numerical investigation of agent-controlled pedestrian dynamics using a structure-preserving finite volume scheme. Advances in Computational Mathematics, 50(1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10098-0.

Paluszynski:2003:TFW

[PŠWX03] Maciej Paluszyński, Hrvoje Šikić, Guido Weiss, and Shaoliang Xiao. Tight frame wavelets, their dimension functions, MRA tight frame wavelets and connectivity properties. Advances in Computational Mathematics, 18(2–4):297– 327, February 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021312110549. Frames.

Potts:2021:CWF

[PT21] Daniel Potts and Manfred Tasche. Continuous window functions for NFFT. Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09873-8.

Pham:2012:OAS

[PTC12] Duong Pham, Thanh Tran, and Simon Crothers. An overlapping additive Schwarz preconditioner for the Laplace– Beltrami equation using spherical splines. Advances in Computational Mathematics, 37(1):93–121, July 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9200-9.

Pottmann:1994:HSE

[PW94] Helmut Pottmann and Michael G. Wagner. Helix splines as an example of affine Tchebycheffian splines. Advances in Computational Mathematics, 2(1):123–142, January 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02519039.

Pottmann:2008:FGC

[PW08] Helmut Pottmann and Johannes Wallner. The focal geometry of circular and conical meshes. Advances in Computational Mathematics, 29(3): 249–268, October 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9045-4.

Potts:2024:VTC

[PW24] Daniel Potts and Laura Weidensager. Variable transformations in combination with wavelets and ANOVA for high-dimensional approximation. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10147-2.

Pang:2020:CSP

[PWL20] Tongyao Pang, Chunlin Wu, and Zhifang Liu. A cubic spline penalty for sparse approximation under tight frame balanced model. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09786-y.

Plonka:2016:RBT

[PZ16] Gerlind Plonka and Yi Zheng. Relation between total variation and persistence distance and its application in signal processing. Advances in Computational Mathematics, 42(3):651–674, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9438-
8.

Qin:2012:ISP

[QC12] Hai-Hua Qin and David Colton. The inverse scattering problem for cavities with impedance boundary condition. Advances in Computational Mathematics, 36(2): 157–174, February 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9179-2.

Qiu:2023:OEE

[Qiu23] Wenlin Qiu. Optimal error estimate of an accurate secondorder scheme for Volterra integrodifferential equations with tempered multi-term kernels. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10050-2.

Qian:2011:AFS

[QW11] Tao Qian and Yan-Bo Wang.

Adaptive Fourier series — a variation of greedy algorithm. Advances in Computational Mathematics, 34(3): 279–293, April 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9153-4.

Qu:2024:MPU

[QWS24] Wei Qu, Yanbo Wang, and Xiaoyun Sun. Matching pursuit with unbounded parameter domains. Advances in Computational Mathematics, 50(1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10097-1.

Qian:2010:OBN

[QWXZ10] Tao Qian, Rui Wang, Yuesheng Xu, and Haizhang Zhang. Orthonormal bases with nonlinear phases. Advances in Computational Mathematics, 33(1):75–95, July 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9120-0.

Qiao:2021:FAO

[QX21] Leijie Qiao and Da Xu. A fast ADI orthogonal spline collocation method with graded meshes for the two-dimensional fractional integro-differential equation. Advances in Computational Mathematics, 47(5):??, October 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09884-5.

Qi:2023:EAS

[QZX23] Xiao Qi, Yanrong Zhang, and Chuanju Xu. An efficient approximation to the stochastic Allen–Cahn equation with random diffusion coefficient field and multiplicative noise. Advances in Computational Mathematics, 49(5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10072-w.

Romero:2009:GAP

[RAB^{+}09] Juan R. Romero, Simon K. Alexander, Shikha Baid, Saurabh Jain, and Manos Papadakis. The geometry and the analytic properties of isotropic multiresolution analysis. Advances in Computational Mathematics, 31(1–3):283–328, October 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9111-6.

Ramaswami:1995:PCS

[Ram95] G. Ramaswami. Perturbed collocation and symplectic RKN methods. Advances in Computational Mathematics, 3(1–2):23–40, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02431994.

Ramos:2018:UHO

[Ram18] Alberto Gil C. P. Ramos.

Uniform and high-order discretization schemes for Sturm–Liouville problems via Fer streamers. Advances in Computational Mathematics, 44(2):395–421, April 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-017-9547-
7.

Rauhut:2005:BTL

[Rau05] Holger Rauhut. Best time localized trigonometric polynomials and wavelets. Advances in Computational Mathematics, 22(1):1–20, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7411-7.

Rodriguez:2018:ECC

[RBGS18] Ana Alonso Rodríguez, Enrico Bertolazzi, Riccardo Ghiloni, and Ruben Specogna. Efficient construction of 2chains representing a basis of H_{2}(\overline{Ω}, ∂Ω; Z). Advances in Computational Mathematics, 44(5):1411– 1440, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9588-6.

Ramos:2023:NED

[RCJ^{+}23] A. J. A. Ramos, A. D. S. Campelo, D. S. Almeida Júnior, M. M. Freitas, and R. C. Barbosa. Numerical exponential decay of thermoelastic waves connected in parallel. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10027-1.

Reiter:2024:BTE

[RDEG24] Sean Reiter, Tobias Damm, Mark Embree, and Serkan Gugercin. On the balanced truncation error bound and sign parameters from arrowhead realizations. Advances in Computational Mathematics, 50(1):??, February 2024. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-024-10105-
y.

Reichel:1993:CPO

[Rei93] Lothar Reichel. Construction of polynomials that are orthogonal with respect to a discrete bilinear form. Advances in Computational Mathematics, 1(2): 241–258, June 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02071388.

Remogna:2012:BCS

[Rem12]  Sara Remogna.  Bivari-

ate C^{2} cubic spline quasiinterpolants on uniform Powell–Sabin triangulations of a rectangular domain. Advances in Computational Mathematics, 36(1):39–65, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9178-3.

Ren:2009:CWT

[Ren09] Jian Ren. A cryptographic watermarking technique for multimedia signals. Advances in Computational Mathematics, 31(1–3):267– 281, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9096-1.

Rabin:2019:TDL

[RF19] Neta Rabin and Dalia Fishelov. Two directional Laplacian pyramids with application to data imputation. Advances in Computational Mathematics, 45(4):2123– 2146, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09697-7.

Rand:2013:IEE

[RGB13] Alexander Rand, Andrew Gillette, and Chandrajit Bajaj. Interpolation error estimates for mean value coordinates over convex polygons. Advances in Computational Mathematics, 39(2): 327–347, August 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-012-9282-z.

Rodriguez:2018:IMR

[RGB18] Andrea Carracedo Rodriguez, Serkan Gugercin, and Jeff Borggaard. Interpolatory model reduction of parameterized bilinear dynamical systems. Advances in Computational Mathematics, 44(6):1887–1916, December 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9611-y.

Raus:2012:QOR

[RH12] Toomas Raus and Uno Hämarik. On the quasioptimal rules for the choice of the regularization parameter in case of a noisy operator. Advances in Computational Mathematics, 36(2): 221–233, February 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9203-6.

Redeker:2015:PER

[RH15] Magnus Redeker and Bernard Haasdonk. A POD–EIM reduced two-scale model for crystal growth. Advances in Computational Mathematics, 41(5):987–1013, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-014-9367-y.

Rong:2017:NAS

[RHZ17] Yao Rong, Yanren Hou, and Yuhong Zhang. Numerical analysis of a second order algorithm for simplified magnetohydrodynamic flows. Advances in Computational Mathematics, 43(4): 823–848, August 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9508-6.

Rippa:1999:ASG

[Rip99] Shmuel Rippa. An algorithm for selecting a good value for the parameter c in radial basis function interpolation. Advances in Computational Mathematics, 11(2–3):193– 210, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018975909870. Ra-

dial basis functions and their applications.

Reddy:2000:ADO

[RJ00] Muni V. Reddy and Stephen Joe. An average discrepancy for optimal vertexmodified number-theoretic rules. Advances in Computational Mathematics, 12(1): 59–69, January 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018952732160. High

dimensional integration.

Rolania:2007:ABS

[RL07] D. Barrios Rolanía and G. López Lagomasino. Asymptotic behavior of solutions of general three term recurrence relations. Advances in Computational Mathematics, 26(1–3):9–37, January 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-4137-x.

Roanes-Lozano:2019:PFA

[RLGGAV19] Eugenio Roanes-Lozano, José Luis Galán-García, and Gabriel Aguilera-Venegas. A prototype of a functional approach to personalized menus generation using set operations. Advances in Computational Mathematics, 45(4): 1881–1895, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09665-1.

Romani:2019:ATP

[RM19] Lucia Romani and Francesca Montagner. Algebraic– trigonometric Pythagorean– hodograph space curves. Advances in Computational Mathematics, 45(1):75–98, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9606-8.

Rodriguez:2019:MAO

[RMCCG^{+}19] C. Destenave Rodriguez, Javier Morales-Castillo, José Roberto Cantú-González, F-Javier Almaguer, and J. M. Martinez. Mathematical analysis to optimize crystal growth. Advances in Computational Mathematics, 45(4):2083– 2096, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09686-w.

Roberts:1998:VA

[Rob98] D. E. Roberts. On a vector q–d algorithm. Advances in Computational Mathematics, 8(4):193–219, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018944213562.

Riemenschneider:1999:GRB

[RS99] S. D. Riemenschneider and N. Sivakumar. Gaussian radial-basis functions: cardinal interpolation of l^{p} and power-growth data. Advances in Computational Mathematics, 11(2–3):229– 251, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018980110778. Ra-

dial basis functions and their applications.

Reif:2001:CIS

[RS01] Ulrich Reif and Peter Schröder.

Curvature integrability of subdivision surfaces. Advances in Computational Mathematics, 14(2):157–174, February 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016685104156.

Rekatsinas:2019:OAT

[RS19] Nikolaos Rekatsinas and Rob Stevenson. An optimal adaptive tensor product wavelet solver of a spacetime FOSLS formulation of parabolic evolution problems. Advances in Computational Mathematics, 45(2): 1031–1066, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9644-2.

Romani:2014:ATP

[RSA14] Lucia Romani, Laura Saini, and Gudrun Albrecht. Algebraic– trigonometric Pythagoreanhodograph curves and their use for Hermite interpolation.

Advances in Computational Mathematics, 40(5–6):977– 1010, December 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9338-8.

Ricaud:2014:OCG

[RST^{+}14] Benjamin Ricaud, Guillaume Stempfel, Bruno Torrésani, Christoph Wiesmeyr, Hélène Lachambre, and Darian Onchis. An optimally concentrated Gabor transform for localized time-frequency components. Advances in Computational Mathematics, 40(3):683–702, June 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-013-9337-
9.

Rohwedder:2011:PPI

[RSZ11] Thorsten Rohwedder, Reinhold Schneider, and Andreas Zeiser. Perturbed preconditioned inverse iteration for operator eigenvalue problems with applications to adaptive wavelet discretization. Advances in Computational Mathematics, 34(1): 43–66, January 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9141-8.

Ricaud:2014:SUP

[RT14] Benjamin Ricaud and Bruno Torrésani. A survey of uncertainty principles and some signal processing applications. Advances in Computational Mathematics, 40(3): 629–650, June 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9323-2.

Rapun:2021:ASM

[RTV21] M.-L. Rapún, F. Terragni, and J. M. Vega. Adaptive sampling and modal expansions in pattern-forming systems. Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09870-x.

Reif:2021:CFG

[RW21] Ulrich Reif and Andreas Weinmann. Clothoid fitting and geometric Hermite subdivision. Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09876-5.

Rong:2013:AWC

[RWT13] Zhijian Rong, Li-Lian Wang, and Xue-Cheng Tai. Adaptive wavelet collocation methods for image segmentation using TV–Allen–Cahn type models. Advances in Computational Mathematics, 38(1): 101–131, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9227-y.

Rim:2013:GPR

[RY13] Kyung Soo Rim and Beong In Yun. Gibbs phenomenon removal by adding Heaviside functions. Advances in Computational Mathematics, 38 (4):683–699, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9255-7.

Richardson:2016:MIR

[RY16] Casey L. Richardson and Laurent Younes. Metamorphosis of images in reproducing kernel Hilbert spaces. Advances in Computational Mathematics, 42 (3):573–603, June 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9435-y.

Rydel:2019:NIO

[Ryd19] Marek Rydel. New integerorder approximations of discrete-time non-commensurate fractional-order systems using the cross Gramian. Advances in Computational Mathematics, 45(2):631–653, April 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
018-9633-5.

Rynne:2000:CGM

[Ryn00] Bryan P. Rynne. Convergence of Galerkin method solutions of the integral equation for thin wire antennas. Advances in Computational Mathematics, 12(2–3):251– 259, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018956800300.

Rude:1998:MPE

[RZ98] Ulrich Rüde and Aihui Zhou.

Multi-parameter extrapolation methods for boundary integral equations. Advances in Computational Mathematics, 9(1–2):173– 190, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018989523557. Nu-

merical treatment of boundary integral equations.

Rieger:2010:SII

[RZ10] Christian Rieger and Barbara Zwicknagl. Sampling inequalities for infinitely smooth functions, with applications to interpolation and machine learning. Advances in Computational Mathematics, 32(1): 103–129, January 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9089-0;
 http://link.springer.com/
content/pdf/10.1007/s10444-
008-9089-0.pdf.

Sablonniere:2004:BTB

[Sab04] Paul Sablonnière. Bernsteintype bases and corner cutting algorithms for C^{1} Merrien’s curves. Advances in Computational Mathematics, 20(1–3):229–246, January 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025870005361. Ap-

proximation and applications.

Salam:1998:FVO

[Sal98]  A. Salam.  Formal vec-

tor orthogonal polynomials. Advances in Computational Mathematics, 8(4): 267–289, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018952415379.

Sangawi:2015:SSM

[San15] Ali W. K. Sangawi. Straight slits map and its inverse of bounded multiply connected regions. Advances in Computational Mathematics, 41(2): 439–455, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9368-x.

Sauer:1995:CAM

[Sau95]  Thomas Sauer.  Compu-

tational aspects of multivariate polynomial interpolation. Advances in Computational Mathematics, 3(3): 219–237, April 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02432000.

Sauer:2006:DMR

[Sau06] Thomas Sauer. Differentiability of multivariate refinable functions and factorization. Advances in Computational Mathematics, 25(1– 3):211–235, July 2006. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7635-y.

Stein:2022:QFS

[SB22] David B. Stein and Alex H.

Barnett. Quadrature by fundamental solutions: kernelindependent layer potential evaluation for large collections of simple objects. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09971-1.

Sorgente:2022:RMQ

[SBMS22] T. Sorgente, S. Biasotti, G. Manzini, and M. Spagnuolo. The role of mesh quality and mesh quality indicators in the virtual element method. Advances in Computational Mathematics, 48 (1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09913-3.

Stabile:2019:ROV

[SBZR19] Giovanni Stabile, Francesco Ballarin, Giacomo Zuccarino, and Gianluigi Rozza. A reduced order variational multiscale approach for turbulent flows. Advances in Computational Mathematics, 45(5-6):2349–2368, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09712-x.

Shih:2008:EPG

[SC08] Shagi-Di Shih and Shue-Sum Chow. Equivalence of n-point Gauss–Chebyshev rule and 4n-point midpoint rule in computing the period of a Lotka–Volterra system. Advances in Computational Mathematics, 28(1): 63–79, January 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9013-4.

Schaback:1995:EEC

[Sch95] Robert Schaback. Error estimates and condition numbers for radial basis function interpolation. Advances in Computational Mathematics, 3(3): 251–264, April 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02432002.

Schumaker:2004:PMS

[Sch04] Larry L. Schumaker. Preface [Multivariate splines]. Advances in Computational Mathematics, 21(3–4):221– 222, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032130.
41067.34.

Schaback:2009:SLE

[Sch09]  Robert Schaback.  Solv-

ing the Laplace equation by meshless collocation using harmonic kernels. Advances in Computational Mathematics, 31(4):457–470, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9078-3;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        008-9078-3.pdf.

Schaback:2011:MWF

[Sch11a]  Robert Schaback.  The

missing Wendland functions. Advances in Computational Mathematics, 34(1):67–81, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9142-7;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        009-9142-7.pdf.

Scheuerer:2011:APS

[Sch11b] Michael Scheuerer. An alternative procedure for selecting a good value for the parameter c in RBF-interpolation. Advances in Computational Mathematics, 34(1):105–126, January 2011. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/s10444-010-9146-3;
http://link.springer.com/
content/pdf/10.1007/s10444-
010-9146-3.pdf.

Schaback:2015:CTC

[Sch15] Robert Schaback. A computational tool for comparing all linear PDE solvers: erroroptimal methods are meshless. Advances in Computational Mathematics, 41(2): 333–355, April 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9360-5.

Schlegel:2021:WTR

[Sch21] Kevin Schlegel. When is there a representer theorem? Advances in Computational Mathematics, 47 (4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09877-4.

Schaback:2023:SEI

[Sch23] Robert Schaback. Small errors imply large evaluation instabilities. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10026-2.

Chang:2013:BHM

[sCLC13] Shih sen Chang, H. W. Joseph Lee, and Chi Kin Chan. A block hybrid method for solving generalized equilibrium problems and convex feasibility problem. Advances in Computational Mathematics, 38(3):563–580, April 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9249-
5.

Song:1996:MMB

[SEE96] J. H. Song, M. J. Edirisinghe, and J. R. G. Evans. Mathematical modelling of binder removal from plasticallyformed engineering ceramics. Advances in Computational Mathematics, 6(1): 325–331, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

Smith:2014:DAF

[SF14] Lindsay N. Smith and Matthew Fickus. Determining angular frequency from images of rotating objects via a generalized Fast Fourier Transform. Advances in Computational Mathematics, 40(1):27–47, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-013-9296-1.

Sommer:2015:FCP

[SFDE15] Alexander Sommer, Ortwin Farle, and Romanus Dyczij-Edlinger. A fast certified parametric nearfield-to-far-field transformation technique for electrically large antenna arrays. Advances in Computational Mathematics, 41(5):1015– 1034, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9373-0.

Siyu:2023:CDD

[SH23] Bai Siyu and Huang Hongying. Coupling of direct discontinuous Galerkin method and natural boundary element method for exterior interface problems with curved elements. Advances in Computational Mathematics, 49 (1):??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-10007-x.

Shekhtman:2008:BIP

[She08] Boris Shekhtman. Bivariate ideal projectors and their perturbations. Advances in Computational Mathematics, 29(3):207–228, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9040-9.

Shen:2009:BDE

[She09] Jianhong (Jackie) Shen. Beamlets are densely embedded in H^{−1}. Advances in Computational Mathematics, 31(1–3):329– 348, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9094-3.

Sherratt:2013:NCB

[She13] Jonathan A. Sherratt. Numerical continuation of boundaries in parameter space between stable and unstable periodic travelling wave (wavetrain) solutions of partial differential equations. Advances in Computational Mathematics, 39(1): 175–192, July 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9273-0.

Shi:2014:QRR

[SHTS14] Lei Shi, Xiaolin Huang, Zheng Tian, and Johan A. K. Suykens. Quantile regression with ℓ_{1}- regularization and Gaussian kernels. Advances in Computational Mathematics, 40(2):

517–551, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9317-0.

San:2015:SPO

[SI15] Omer San and Traian Iliescu. A stabilized proper orthogonal decomposition reduced-order model for large scale quasigeostrophic ocean circulation. Advances in Computational Mathematics, 41(5):1289–1319, October 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9417-0.

Sidi:2017:MPR

[Sid17] Avram Sidi. Minimal polynomial and reduced rank extrapolation methods are related. Advances in Computational Mathematics, 43(1): 151–170, February 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9481-0.

Sidi:2020:VVP

[Sid20] Avram Sidi. Vector versions of Prony’s algorithm and vector-valued rational approximations. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09751-9.

Silbermann:1999:TPD

[Sil99] Bernd Silbermann. Tribute to Professor Dr. rer. nat. habil. Siegfried Prössdorf 02.01.1939—19.07.1998. Advances in Computational Mathematics, 11(1):i–ii, July 1999. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1018903421256.

Sun:2015:NIS

[SJJL15] Zhiyuan Sun, Yuling Jiao, Bangti Jin, and Xiliang Lu. Numerical identification of a sparse Robin coefficient. Advances in Computational Mathematics, 41(1):131–148, February 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9352-5.

Sui:2021:SOA

[SJJY21] Yubing Sui, Jingzhou Jiang, Guigen Jin, and Xiaofeng Yang. Second-order accurate and energy stable numerical scheme for an immiscible binary mixture of nematic liquid crystals and viscous fluids with strong anchoring potentials. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09865-8.

Sun:2005:GFT

[SL05] Jiachang Sun and Huiyuan Li. Generalized Fourier transform on an arbitrary triangular domain. Advances in Computational Mathematics, 22(3):223–248, April 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-003-7667-
8.

Shen:2012:RIP

[SL12] Yi Shen and Song Li. Restricted p-isometry property and its application for nonconvex compressive sensing. Advances in Computational Mathematics, 37(3): 441–452, October 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9219-y.

Schauer:2015:RBM

[SL15] Volker Schauer and Christian Linder. The reduced basis method in all-electron calculations with finite elements. Advances in Computational Mathematics, 41(5):1035– 1047, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9374-z.

Sticko:2020:HOC

[SLK20] Simon Sticko, Gustav Ludvigsson, and Gunilla Kreiss. High-order cut finite elements for the elastic wave equation. Advances in Computational Mathematics, 46 (3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09785-z.

Sommer:2014:OGE

[SLN14] S. Sommer, F. Lauze, and M. Nielsen. Optimization over geodesics for exact principal geodesic analysis. Advances in Computational Mathematics, 40(2): 283–313, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9308-1.

Sun:2021:AAO

[SLT21] Xue Sun, Zhongli Liu, and Hongjiong Tian. An alternative approach for order conditions of Runge–Kutta– Nyström methods. Advances in Computational Mathematics, 47(5):??, October 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10444-021-09894-3.

Sun:1999:SPD

[SM99] Xingping Sun and V. A.

Menegatto. Strictly positive definite functions on the complex Hilbert sphere. Advances in Computational Mathematics, 11(2–3):105– 119, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018915723982. Ra-

dial basis functions and their applications.

Salah:2000:CBS

[SM00] I. Ben Salah and P. Maroni.

The connection between selfassociated two-dimensional vector functionals and third degree forms. Advances in Computational Mathematics, 13(1):51–77, April 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018941924408.

San:2018:NNC

[SM18] Omer San and Romit Maulik.

Neural network closures for nonlinear model order reduction. Advances in Computational Mathematics, 44(6):1717–1750, December 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9590-z.

Saibaba:2022:ERT

[SMK22] Arvind K. Saibaba, Rachel Minster, and Misha E. Kilmer. Efficient randomized tensor-based algorithms for function approximation and low-rank kernel interactions. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09979-7.

Smoch:2007:SBG

[Smo07] Laurent Smoch. Spectral behaviour of GMRES applied to singular systems. Advances in Computational Mathematics, 27(2):151–166, August 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9029-4.

Sushnikova:2020:SNE

[SO20] Daria A. Sushnikova and Ivan V. Oseledets. Simple non-extensive sparsification of the hierarchical matrices. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-020-09794-y.

Sondergaard:2007:GFS

[Søn07] Peter L. Søndergaard. Gabor frames by sampling and periodization. Advances in Computational Mathematics, 27(4):355–373, November 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9003-y.

Sorokina:2018:BBT

[Sor18] Tatyana Sorokina. Bernstein– Bézier techniques for divergence of polynomial spline vector fields in R^{n}. Advances in Computational Mathematics, 44(1):227–244, February 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9541-0.

Sousa:2022:CCI

[Sou22] Ercília Sousa. Convergence of consistent and inconsistent schemes for fractional diffusion problems with boundaries. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09984-w.

Speleers:2017:HSS

[Spe17] Hendrik Speleers. Hierarchical spline spaces: quasiinterpolants and local approximation estimates. Advances in Computational Mathematics, 43(2):235–255, April 2017. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-016-9483-
y.

Spence:2023:SPE

[Spe23]  E. A. Spence.  A sim-

ple proof that the hp-FEM does not suffer from the pollution effect for the constant-coefficient full-space Helmholtz equation. Advances in Computational Mathematics, 49(2):??, April 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
023-10025-3.

Sharma:2018:EMF

[SPS18] Nisha Sharma, Amiya K. Pani, and Kapil K. Sharma. Expanded mixed FEM with lowest order RT elements for nonlinear and nonlocal parabolic problems. Advances in Computational Mathematics, 44(5):1537– 1571, October 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-018-9596-6.

Schober:2021:EDT

[SPS21] Kevin Schober, Jürgen Prestin, and Serhii A. Stasyuk. Edge detection with trigonometric polynomial shearlets. Advances in Computational Mathematics, 47(1): ??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09838-3.

Sprott:2002:KGR

[SR02] Kenneth Sprott and Bahram Ravani. Kinematic generation of ruled surfaces. Advances in Computational Mathematics, 17(1–2):115– 133, July 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1015211729988. Ad-

vances in geometrical algorithms and representations.

Song:2012:MII

[SRFH12] Guohui Song, John Riddle, Gregory E. Fasshauer, and Fred J. Hickernell. Multivariate interpolation with increasingly flat radial basis functions of finite smoothness. Advances in Computational Mathematics, 36(3): 485–501, April 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9192-5.

Sommer:1996:CSW

[SS96] Manfred Sommer and Hans Strauss. A condition of Schoenberg–Whitney type for multivariate spline interpolation. Advances in Computational Mathematics, 5 (1):381–397, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02124752.

Schumaker:2004:QST

[SS04] Larry L. Schumaker and Tatyana Sorokina. C^{1} quintic splines on type-4 tetrahedral partitions. Advances in Computational Mathematics, 21(3–4):421– 444, October 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000032045.
03775.8d.

Sloan:2008:ASU

[SS08] Ian H. Sloan and Alvise Sommariva. Approximation on the sphere using radial basis functions plus polynomials. Advances in Computational Mathematics, 29(2): 147–177, August 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9048-1.

Sun:2013:IFW

[SS13] Xudong Sun and Wenchang Sun. Inversion formula for the windowed Fourier transform, II. Advances in Computational Mathematics, 38(1): 21–34, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9223-2.

Schumaker:2014:CPS

[SS14] Larry L. Schumaker and Hendrik Speleers. Convexity preserving C^{0} splines. Advances in Computational Mathematics, 40(1):117–135, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9301-8.

Son:2015:MOR

[SS15] Nguyen Thanh Son and Tatjana Stykel. Model order reduction of parameterized circuit equations based on interpolation. Advances in Computational Mathematics, 41(5):1321– 1342, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9418-z.

Santin:2016:AEK

[SS16] G. Santin and R. Schaback. Approximation of eigenfunctions in kernel-based spaces. Advances in Computational Mathematics, 42(4):973–993, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9449-5.

Strossner:2024:AEF

[SSK24] Christoph Strössner, Bonan Sun, and Daniel Kressner. Approximation in the extended functional tensor train format. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10140-9.

Sorgentone:2023:EQE

[ST23] Chiara Sorgentone and Anna-Karin Tornberg. Estimation of quadrature errors for layer potentials evaluated near surfaces with spherical topology. Advances in Computational Mathematics, 49(6): ??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10083-7.

Steidl:1998:NFF

[Ste98] Gabriele Steidl. A note on Fast Fourier Transforms for nonequispaced grids. Advances in Computational Mathematics, 9(3–4):337– 352, November 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018901926283.

Stonag:1996:CPL

[Sto96] Eduardo D. Stonag. Critical points for least-squares problems involving certain analytic functions, with applications to sigmoidal nets. Advances in Computational Mathematics, 5(1):245–268, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124746.

Stolk:2017:ISD

[Sto17] Christiaan C. Stolk. An improved sweeping domain decomposition preconditioner for the Helmholtz equation. Advances in Computational Mathematics, 43(1): 45–76, February 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9475-y;
http://link.springer.com/

content/pdf/10.1007/s10444-
016-9475-y.pdf.

Strom:1995:BSH

[Str95]  Kyrre Strøm.  B-splines

with homogenized knots. Advances in Computational Mathematics, 3(3):291–308, January 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02988627.

Struble:2013:CWA

[Str13]  Dale W. Struble.  The

construction of wavelets adapted to compact domains. Advances in Computational Mathematics, 38(3): 581–600, April 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9250-z.

Sun:2012:LIF

[SU12] Qiyu Sun and Michael Unser. Left-inverses of fractional Laplacian and sparse stochastic processes. Advances in Computational Mathematics, 36(3):399–441, April 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9183-
6;   http://link.springer.
com/content/pdf/10.1007/
s10444-011-9183-6.pdf.

Sukhorukova:2022:GVP

[SU22] Nadezda Sukhorukova and Julien Ugon. A generalisation of de la Vallée–Poussin procedure to multivariate approximations. Advances in Computational Mathematics, 48(1):??, February 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09919-x.

Sun:2008:FSF

[Sun08] Qiyu Sun. Frames in spaces with finite rate of innovation. Advances in Computational Mathematics, 28(4): 301–329, May 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9021-4.

Sun:2010:LRS

[Sun10] Qiyu Sun. Local reconstruction for sampling in shiftinvariant spaces. Advances in Computational Mathematics, 32(3):335–352, April 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-008-9109-
0.

Sun:2014:LNF

[Sun14] Qiyu Sun. Localized nonlinear functional equations and two sampling problems in signal processing. Advances in Computational Mathematics, 40(2): 415–458, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9314-3.

Svacek:2019:IAF

[Svá19] Petr Svácek. On implementation aspects of finite element method and its application. Advances in Computational Mathematics, 45(4): 2065–2081, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09685-x.

Stevenson:2022:WTF

[SvVW22] Rob Stevenson, Raymond van Venetië, and Jan Westerdiep. A wavelet-in-time, finite element-in-space adaptive method for parabolic evolution equations. Advances in Computational Mathematics, 48(3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09930-w.

Shaw:1996:TAF

[SW96] Simon Shaw and J. R. Whiteman. Towards adaptive finite element schemes for partial differential Volterra equation solvers. Advances in Computational Mathematics, 6(1):309–323, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127710.  John

Crank 80th birthday special issue.

Steinbach:1998:CSE

[SW98] O. Steinbach and W. L.

Wendland. The construction of some efficient preconditioners in the boundary element method. Advances in Computational Mathematics, 9(1–2):191– 216, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018937506719. Nu-

merical treatment of boundary integral equations.

Sloan:2004:ESP

[SW04] Ian H. Sloan and Robert S.

Womersley. Extremal systems of points and numerical integration on the sphere. Advances in Computational Mathematics, 21(1–2):107– 125, July 2004. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016428.
25905.da.

Schaback:2006:LCR

[SW06] R. Schaback and J. Werner. Linearly constrained reconstruction of functions by kernels with applications to machine learning. Advances in Computational Mathematics, 25(1–3):237– 258, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7616-1.

Sorokina:2008:MPS

[SW08] T. Sorokina and A. J. Worsey. A multivariate Powell–Sabin interpolant. Advances in Computational Mathematics, 29(1):71–89, July 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-007-9041-8.

Sun:2010:RLS

[SW10] Hongwei Sun and Qiang Wu. Regularized least square regression with dependent samples. Advances in Computational Mathematics, 32(2): 175–189, February 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9099-y.

Swietlicka:2019:RTS

[’Sw19] Aleksandra ’Swietlicka. Regularization theory in the study of generalization ability of a biological neural network model. Advances in Computational Mathematics, 45(4): 1793–1805, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-09658-6.

Schmidt:2020:REP

[SWH20] Andreas Schmidt, Dominik Wittwar, and Bernard Haasdonk. Rigorous and effective a-posteriori error bounds for nonlinear problems-application to RB methods. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09741-x.

Sauer:1995:MHI

[SX95] Thomas Sauer and Yuan Xu. On multivariate Hermite interpolation. Advances in Computational Mathematics, 4(1–2):207–259, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03177515.

Shi:2007:COP

[SX07] Ying Guang Shi and Guoliang Xu. Construction of σorthogonal polynomials and Gaussian quadrature formulas. Advances in Computational Mathematics, 27(1): 79–94, July 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9033-8.

Shen:2009:PMM

[SX09] Lixin Shen and Yuesheng Xu. Preface [Mathematical methods for image processing]. Advances in Computational Mathematics, 31(1– 3):1–4, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9113-z;
 http://link.springer.com/
content/pdf/10.1007/s10444-
009-9113-z.pdf.

Sun:2022:OVS

[SXWG22] Yafei Sun, Yingxiang Xu, Shuangbin Wang, and Shan Gao. Optimized Ventcel– Schwarz methods for the Cahn–Hilliard equation discretized by the stabilized linear Crank–Nicolson scheme. Advances in Computational Mathematics, 48(5):??, October 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09982-y.

Shu:2006:AMM

[SXYY06] Shi Shu, Jinchao Xu, Ying Yang, and Haiyuan Yu. An algebraic multigrid method for finite element systems on criss-cross grids. Advances in Computational Mathematics, 25(1–3):287– 304, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7627-y.

Schneider:2006:ADG

[SXZ06] Reinhold Schneider, Yuesheng Xu, and Aihui Zhou. An analysis of discontinuous Galerkin methods for elliptic problems. Advances in Computational Mathematics, 25(1–3):259– 286, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7619-y.

Shi:2019:UOE

[SY19] Dongyang Shi and Huaijun Yang. Unconditionally optimal error estimates of a new mixed FEM for nonlinear Schrödinger equations. Advances in Computational Mathematics, 45(5-6):3173– 3194, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-019-09732-7.

Sun:2018:ADM

[SYCJ18] Tao Sun, Penghang Yin, Lizhi Cheng, and Hao Jiang. Alternating direction method of multipliers with difference of convex functions. Advances in Computational Mathematics, 44(3): 723–744, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9559-3.

Sun:2009:CLS

[SZ09] Wenchang Sun and Xingwei Zhou. Characterization of local sampling sequences for spline subspaces. Advances in Computational Mathematics, 30(2):153–175, February 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-008-9062-y.

Song:2021:VBI

[SZJ21] Xiaoyan Song, Guang-Hui Zheng, and Lijian Jiang. Variational Bayesian inversion for the reaction coefficient in space-time nonlocal diffusion equations. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09850-1.

Taddei:2019:OOP

[Tad19] Tommaso Taddei. An offline/ online procedure for dual norm calculations of parameterized functionals: empirical quadrature and empirical test spaces. Advances in Computational Mathematics, 45(5-6):2429–2462, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09721-w.

Talaska:2019:PAF

[Tal19] Tomasz Tala’ska. Parallel, asynchronous, fuzzy logic systems realized in CMOS technology. Advances in Computational Mathematics, 45(4): 1807–1823, August 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-09659-5.

Tang:2017:SID

[Tan17] Sui Tang. System identification in dynamical sampling. Advances in Computational Mathematics, 43(3): 555–580, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9497-5.

Tausch:2019:NMB

[Tau19] Johannes Tausch. Nyström method for BEM of the heat equation with moving boundaries. Advances in Computational Mathematics, 45(5-6):2953–2968, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09720-x.

Turner:2016:TDC

[TB16] M. R. Turner and T. J. Bridges. Time-dependent conformal mapping of doublyconnected regions. Advances in Computational Mathematics, 42(4):947–972, August 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9448-6;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        015-9448-6.pdf.

Twizell:1994:NME

[TBD94] E. H. Twizell, A. Boutayeb, and K. Djidjeli. Numerical methods for eighth- , tenth- and twelfth-order eigenvalue problems arising in thermal instability. Advances in Computational Mathematics, 2(4):407–436, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521607.

Tomljanovic:2018:DOP

[TBG18] Zoran Tomljanović, Christopher Beattie, and Serkan Gugercin. Damping optimization of parameter dependent mechanical systems by rational interpolation. Advances in Computational Mathematics, 44(6):1797– 1820, December 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9605-9.

Toshniwal:2021:CDS

[TD21] Deepesh Toshniwal and Michael DiPasquale. Counting the dimension of splines of mixed smoothness. Advances in Computational Mathematics, 47(1):??, February 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-020-09830-
x.

Tian:2016:ACS

[TDG16] Xiaochuan Tian, Qiang Du, and Max Gunzburger. Asymptotically compatible schemes for the approximation of fractional Laplacian and related nonlocal diffusion problems on bounded domains. Advances in Computational Mathematics, 42(6):1363–1380, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9466-z.

Temlyakov:1998:BTA

[Tem98] V. N. Temlyakov. The best m-term approximation and greedy algorithms. Advances in Computational Mathematics, 8(4):249–265, June 1998. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1018900431309.

Temlyakov:2000:WGA

[Tem00] V. N. Temlyakov. Weak greedy algorithms. Advances in Computational Mathematics, 12(2–3):213– 227, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018917218956.

Temlyakov:2001:GAB

[Tem01] V. N. Temlyakov. Greedy algorithms in Banach spaces. Advances in Computational Mathematics, 14(3):277–292, April 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1016657209416.

Temlyakov:2002:CCW

[Tem02] V. N. Temlyakov. A criterion for convergence of weak greedy algorithms. Advances in Computational Mathematics, 17(3):269–280, October 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016061804993.

Temlyakov:2007:GEB

[Tem07] V. N. Temlyakov. Greedy expansions in Banach spaces. Advances in Computational Mathematics, 26(4):431–449, May 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-005-7452-
y.

Twizell:1996:SOS

[TGA96] E. H. Twizell, A. B. Gumel, and M. A. Arigu. Secondorder, L_{0}-stable methods for the heat equation with timedependent boundary conditions. Advances in Computational Mathematics, 6 (1):333–352, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/BF02127712. John

Crank 80th birthday special issue.

Tan:2019:HDF

[TH19] Wee Chin Tan and Viet Ha Hoang. High dimensional finite elements for time-space multiscale parabolic equations. Advances in Computational Mathematics, 45(3): 1291–1327, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-09657-7.

Tonnon:2024:SLF

[TH24] Wouter Tonnon and Ralf Hiptmair. Semi-Lagrangian finite element exterior calculus for incompressible flows. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10092-6.

Thanh:2019:WBN

[Tha19] Mai Duc Thanh. A wellbalanced numerical scheme for a model of two-phase flows with treatment of nonconservative terms. Advances in Computational Mathematics, 45(5-6):2701–2719, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09704-x.

Timme:2021:MPP

[Tim21] Sascha Timme. Mixed precision path tracking for polynomial homotopy continuation. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09899-y.

Takhirov:2018:CEM

[TL18] Aziz Takhirov and Alexander Lozovskiy. Computationally efficient modular nonlinear filter stabilization for high Reynolds number flows. Advances in Computational Mathematics, 44(1): 295–325, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9544-x.

Taddele:2022:LAM

[TLG^{+}22] Guash Haile Taddele, Yuan Li, Aviv Gibali, Poom Kumam, and Jing Zhao. Linear approximation method for solving split inverse problems and its applications. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09959-x.

Toshniwal:2021:PSS

[TMH21] Deepesh Toshniwal, Bernard Mourrain, and Thomas J. R. Hughes. Polynomial spline spaces of non-uniform bidegree on T-meshes: combinatorial bounds on the dimension. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09829-4.

Tommasini:1996:CRL

[Tom96] T. Tommasini. Complexity reduction of least squares problems involving special Vandermonde matrices. Advances in Computational Mathematics, 6(1):77–86, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127697.

Tornberg:2016:ESS

[Tor16] Anna-Karin Tornberg. The Ewald sums for singly, doubly and triply periodic electrostatic systems. Advances in Computational Mathematics, 42(1):227–248, February 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9422-3.

Tao:2017:DDM

[TQ17] Zhanjing Tao and Jianxian Qiu. Dimension-by-dimension moment-based central Hermite WENO schemes for directly solving Hamilton–Jacobi equations. Advances in Computational Mathematics, 43(5):1023– 1058, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9515-2.

Tsukanov:2005:MMA

[TS05] I. Tsukanov and V. Shapiro.

Meshfree modeling and analysis of physical fields in heterogeneous media. Advances in Computational Mathematics, 23(1–2):95– 124, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1835-3.

Thao:2021:CCA

[TSV21] Nguyen Hieu Thao, Oleg Soloviev, and Michel Verhaegen. Convex combination of alternating projection and Douglas–Rachford operators for phase retrieval. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

springer.com/article/10.
1007/s10444-021-09861-y.

Tan:2010:ROB

[TSY10] Lihui Tan, Lixin Shen, and Lihua Yang. Rational orthogonal bases satisfying the Bedrosian identity. Advances in Computational Mathematics, 33(3): 285–303, October 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9133-8.

Tran:1998:WBP

[TSZ98] Thanh Tran, Ernst P. Stephan, and Stefan Zaprianov. Wavelet-based preconditioners for boundary integral equations. Advances in Computational Mathematics, 9(1–2):233– 249, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018993624466. Nu-

merical treatment of boundary integral equations.

Thanh:2021:WBH

[TTC21] Nguyen Xuan Thanh, Mai Duc Thanh, and Dao Huy Cuong. A well-balanced high-order scheme on van Leer-type for the shallow water equations with temperature gradient and variable bottom topography. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09832-9.

Tausch:1998:SKI

[TW98] Johannes Tausch and Jacob White. Second-kind integral formulations of the capacitance problem. Advances in Computational Mathematics, 9(1–2):217– 232, September 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018973019922. Nu-

merical treatment of boundary integral equations.

Twizell:1996:P

[Twi96]  E. H. Twizell.  Preface.

Advances in Computational Mathematics, 6(1):i–ii, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02127691.

Tang:2019:AOA

[TX19] Xiao Tang and Aiguo Xiao.

Asymptotically optimal approximation of some stochastic integrals and its applications to the strong secondorder methods. Advances in Computational Mathematics, 45(2):813–846, April 2019.

CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-018-9638-
                                        0.

Tao:2018:CAD

[TY18] Min Tao and Xiaoming Yuan. Convergence analysis of the direct extension of ADMM for multiple-block separable convex minimization. Advances in Computational Mathematics, 44(3): 773–813, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9560-x.

Tygert:2023:CVC

[Tyg23] Mark Tygert. Calibration of P -values for calibration and for deviation of a subpopulation from the full population. Advances in Computational Mathematics, 49 (5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10068-6.

Temlyakov:2011:GAR

[TYY11] V. N. Temlyakov, Mingrui Yang, and Peixin Ye. Greedy approximation with regard to non-greedy bases. Advances in Computational Mathematics, 34(3):

319–337, April 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9155-2.

Tasche:2003:REA

[TZ03] Manfred Tasche and Hansmartin Zeuner. Roundoff error analysis of the recursive moving window discrete Fourier transform. Advances in Computational Mathematics, 18(1):65–78, January 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1021254709641.

Tan:2005:FSR

[TZ05] Jieqing Tan and Ping Zhou.

On the finite sum representations of the Lauricella functions F_{D}. Advances in Computational Mathematics, 23(4):333–351, November 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-1838-0.

Tebou:2007:UBS

[TZ07] Louis T. Tebou and Enrique Zuazua. Uniform boundary stabilization of the finite difference space discretization of the 1-d wave equation. Advances in Computational Mathematics, 26(1–3):

337–365, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-7629-9.

Teng:2016:BCD

[TZL16] Zhongming Teng, Yunkai Zhou, and Ren-Cang Li. A block Chebyshev–Davidson method for linear response eigenvalue problems. Advances in Computational Mathematics, 42(5):1103– 1128, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9455-2.

Unger:2019:KWL

[UG19] Benjamin Unger and Serkan Gugercin. Kolmogorov nwidths for linear dynamical systems. Advances in Computational Mathematics, 45(5-6):2273–2286, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09701-0.

Urban:1995:DFW

[Urb95] Karsten Urban. On divergencefree wavelets. Advances in Computational Mathematics, 4(1–2):51–81, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123473.

Varas:1996:RSM

[Var96] María Leonor S. Varas. Reconstruction of signals from multiscale edges. Advances in Computational Mathematics, 5(1):31–49, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124734.

Venturino:2000:CLI

[VC00] E. Venturino and G. A. Chandler. Calculation of line integrals in boundary integral methods. Advances in Computational Mathematics, 13 (3):293–318, July 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018914513021.

vanderHouwen:1993:PII

[vdH93] P. J. van der Houwen.

Preconditioning in implicit initial-value problem methods on parallel computers. Advances in Computational Mathematics, 1(1): 39–60, February 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02070820.

vanderHouwen:1997:PLS

[vdHdS97] P. J. van der Houwen and J. J. B. de Swart. Parallel linear system solvers for Runge–Kutta methods. Advances in Computational Mathematics, 7(1–2):157– 181, March 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018990601750. Par-

allel methods for ODEs.

vanderHouwen:2000:DIR

[vdHS00] P. J. van der Houwen and B. P. Sommeijer. Diagonally implicit Runge–Kutta methods for 3D shallow water applications. Advances in Computational Mathematics, 12(2–3):229– 250, February 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018969203026.

vanderHouwen:1997:WRM

[vdHv97] P. J. van der Houwen and W. A. van der Veen. Waveform relaxation methods for implicit differential equations. Advances in Computational Mathematics, 7(1–2): 183–197, March 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1023/A:1018942718589. Par-

allel methods for ODEs.

vanderMee:2003:SEI

[vdMNS03] Cornelis V. M. van der Mee, M. Z. Nashed, and Sebastiano Seatzu. Sampling expansions and interpolation in unitarily translation invariant reproducing kernel Hilbert spaces. Advances in Computational Mathematics, 19(4):355–372, November 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1024233232215.

Venturino:1994:LIA

[Ven94]  Ezio Venturino.  A la-

cunary interpolation algorithm on arbitrary points. Advances in Computational Mathematics, 2(2):223–233, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521109.

Voss:1996:TSA

[VK96] D. A. Voss and A. Q. M.

Khaliq. Time-stepping algorithms for semidiscretized linear parabolic PDEs based on rational approximants with distinct real poles. Advances in Computational Mathematics, 6(1):353–363, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02127713.  John

Crank 80th birthday special issue.

Voronin:2017:EAC

[VM17] Sergey Voronin and Per-Gunnar Martinsson. Efficient algorithms for CUR and interpolative matrix decompositions. Advances in Computational Mathematics, 43 (3):495–516, June 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9494-8.

Vogeli:2018:FDG

[VNS18] Urs Vögeli, Khadijeh Nedaiasl, and Stefan A. Sauter. A fully discrete Galerkin method for Abel-type integral equations. Advances in Computational Mathematics, 44(5):1601–1626, October 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9598-4.

Van:2002:ATD

[VW02] Tri Van and Aihua Wood. Analysis of time-domain Maxwell’s equations for 3-D cavities. Advances in Computational Mathematics, 16(2– 3):211–228, April 2002. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014489832230. Mod-

eling and computation in optics and electromagnetics.

Van:2005:FEA

[VW05] Tri Van and Aihua Wood.

Finite element analysis of transient electromagnetic scattering problems. Advances in Computational Mathematics, 22(1):21–48, January 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-1176-x.

Walz:1995:UAB

[Wal95] Guido Walz. A unified approach to B-spline recursions and knot insertion, with application to new recursion formulas. Advances in Computational Mathematics, 3(1–2): 89–100, January 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02431997.

Wang:2008:SDQ

[Wan08] Yang Wang. Sigma–delta quantization errors and the traveling salesman problem. Advances in Computational Mathematics, 28(2):101–118, February 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-006-9016-1.

Wang:2012:PNM

[Wan12]  Y. F. Wang.  Precondi-

tioning non-monotone gradient methods for retrieval of seismic reflection signals. Advances in Computational Mathematics, 36(2):353–376, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9207-2.

Wang:2019:OCO

[Wan19]  Wansheng Wang.  Opti-

mal convergence orders of fully geometric mesh one-leg methods for neutral differential equations with vanishing variable delay. Advances in Computational Mathematics, 45(3):1631– 1655, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09688-8.

Wang:2022:AFV

[Wan22] Zhoufeng Wang. An adaptive finite volume method for the diffraction grating problem with the truncated DtN boundary condition. Advances in Computational Mathematics, 48 (4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09969-9.

Wang:2023:CMS

[Wan23a]  Xiaojie Wang.  Correc-

tion: Mean-square convergence rates of implicit Milstein type methods for SDEs with non-Lipschitz coefficients. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10056-w.

Wang:2023:MSC

[Wan23b] Xiaojie Wang. Mean-square convergence rates of implicit Milstein type methods for SDEs with non-Lipschitz coefficients. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10034-2.

Warren:1996:BCC

[War96] Joe Warren. Barycentric coordinates for convex polytopes. Advances in Computational Mathematics, 6(1): 97–108, December 1996. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/BF02127699.

Ward:2013:EEA

[War13] John Paul Ward. L^{p} error estimates for approximation by Sobolev splines and Wendland functions on R^{d}. Advances in Computational Mathematics, 38(4): 873–889, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9263-7.

Watson:1994:SOP

[Wat94] G. A. Watson. The solution of orthogonal Procrustes problems for a family of orthogonally invariant norms. Advances in Computational Mathematics, 2(4):393–405, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521606.

Watanabe:1996:SML

[Wat96] Sumio Watanabe. Solvable models of layered neural networks based on their differential structure. Advances in Computational Mathematics, 5(1):205–231, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124744.

Wang:2024:RPC

[WaZL24] Meiting Wang, Guang an Zou, and Jian Li. A rotational pressure-correction discontinuous Galerkin scheme for the Cahn–Hilliard-Darcy–Stokes system. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10151-6.

Wurth:2022:DDU

[WBGG22] Alexandra Würth, Mickaël Binois, Paola Goatin, and Simone Göttlich. Data-driven uncertainty quantification in macroscopic traffic flow models. Advances in Computational Mathematics, 48(6): ??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09989-5.

Wang:2024:AFE

[WBY24] Shixi Wang, Hai Bi, and Yidu Yang. An adaptive FEM for the elastic transmission eigenvalue problem with different elastic tensors and different mass densities. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-023-10099-z.

Wendland:1995:PPP

[Wen95] Holger Wendland. Piecewise polynomial, positive definite and compactly supported radial functions of minimal degree. Advances in Computational Mathematics, 4(1–2):389–396, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02123482.

Wu:2002:TMP

[WFL02] Wei Wu, Guorui Feng, and Xin Li. Training multilayer perceptrons via minimization of sum of ridge functions. Advances in Computational Mathematics, 17(4):331–347, November 2002. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1016249727555.

Wang:2019:HAN

[WG19] Jun Wang and Leslie Greengard. Hybrid asymptotic/ numerical methods for the evaluation of layer heat potentials in two dimensions. Advances in Computational Mathematics, 45(2): 847–867, April 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.

springer.com/article/10.
1007/s10444-018-9641-5.

Wei:2007:NDR

[WH07] T. Wei and Y. C. Hon. Numerical differentiation by radial basis functions approximation. Advances in Computational Mathematics, 27(3): 247–272, October 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-9001-0.

Wang:2018:NMS

[WHS18] Liqun Wang, Songming Hou, and Liwei Shi. A numerical method for solving threedimensional elliptic interface problems with triple junction points. Advances in Computational Mathematics, 44(1): 175–193, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9539-7.

Wang:2018:MWX

[WHTZ18] Wenqing Wang, Xuehai Huang, Kai Tang, and Ruiyue Zhou. Morley– Wang–Xu element methods with penalty for a fourth order elliptic singular perturbation problem. Advances in Computational Mathematics, 44(4):1041– 1061, August 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9572-6.

Wierse:1997:NTM

[Wie97] Monika Wierse. A new theoretically motivated higher order upwind scheme on unstructured grids of simplices. Advances in Computational Mathematics, 7(3): 303–335, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018955121314.

Wieland:2015:IPM

[Wie15]  Bernhard Wieland.  Im-

plicit partitioning methods for unknown parameter sets: in the context of the reduced basis method. Advances in Computational Mathematics, 41(5):1159– 1186, October 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9404-5.

Wille:1998:ESC

[Wil98] David R. Willé. Experiments in stepsize control for Adams linear multistep methods. Advances in Computational Mathematics, 8(4): 335–344, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1023/A:1018960717197.

Williams:1993:LSC

[WK93] Jack Williams and Z. Kalogiratou. Least squares and Chebyshev fitting for parameter estimation in ODEs. Advances in Computational Mathematics, 1(3):357–366, October 1993. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02072016.

Wendland:2020:SPD

[WK20] Holger Wendland and Jens Künemund. Solving partial differential equations on (evolving) surfaces with radial basis functions. Advances in Computational Mathematics, 46(4):??, August 2020. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-020-09803-
0.

Wu:2005:GSF

[WL05] Zong Min Wu and Jian Ping Liu. Generalized Strang– Fix condition for scattered data quasi-interpolation. Advances in Computational Mathematics, 23(1–2):201– 214, July 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-004-1832-6.

Wang:2012:RDN

[WL12] Haibing Wang and Jijun Liu. Recovering the Dirichlet-to-Neumann map in inverse scattering problems using integral equation methods. Advances in Computational Mathematics, 36(2):279–297, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9191-6.

Wang:2022:IGI

[WLB22] Tingyu Wang, Simon K. Layton, and Lorena A. Barba. Inexact GMRES iterations and relaxation strategies with fast-multipole boundary element method. Advances in Computational Mathematics, 48(3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09932-8.

Witkowski:2015:SCN

[WLPV15] T. Witkowski, S. Ling, S. Praetorius, and A. Voigt. Software concepts and numerical algorithms for a scalable adaptive parallel finite element method. Advances in Computational Mathematics, 41(6):1581–1613, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9405-4.

Wang:2016:MAL

[WLW16] Chuanlong Wang, Chao Li, and Jin Wang. A modified augmented Lagrange multiplier algorithm for Toeplitz matrix completion. Advances in Computational Mathematics, 42(5):1209– 1224, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9459-y.

Wen:2023:SBP

[WLW23] Jin Wen, Zhi-Yuan Li, and Yong-Ping Wang. Solving the backward problem for timefractional wave equations by the quasi-reversibility regularization method. Advances in Computational Mathematics, 49(6):??, December 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10080-w.

Wu:2021:ZCN

[WM21] Bowei Wu and Per-Gunnar Martinsson. Zeta correction: a new approach to constructing corrected trapezoidal quadrature rules for singular integral operators. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09872-9.

Wang:2013:PCS

[WMB13] Chuan-Long Wang, Guo-Yan Meng, and Yan-Hong Bai. Practical convergent splittings and acceleration methods for non-Hermitian positive definite linear systems. Advances in Computational Mathematics, 39(2): 257–271, August 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9278-8.

Wang:2021:SEE

[WMW21] Wansheng Wang, Mengli Mao, and Zheng Wang. Stability and error estimates for the variable step-size BDF2 method for linear and semilinear parabolic equations. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09839-2.

Wang:2013:MPM

[WMY13] Chuan-Long Wang, Guo-Yan Meng, and Xue-Rong Yong.

Modified parallel multisplitting iterative methods for non-Hermitian positive definite systems. Advances in Computational Mathematics, 38(4):859–872, May 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-011-9262-
8.

Wenger:2017:CAH

[WOBL17] Theresa Wenger, Sina Ober-Blöbaum, and Sigrid Leyendecker. Construction and analysis of higher order variational integrators for dynamical systems with holonomic constraints. Advances in Computational Mathematics, 43(5):1163– 1195, October 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-017-9520-5.

Wu:2023:USL

[WQP23] Xiaoxin Wu, Weifeng Qiu, and Kejia Pan. An unconditionally stable and L^{2} optimal quadratic finite volume scheme over triangular meshes for anisotropic elliptic equations. Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-023-10085-5.

Womersley:2001:HGC

[WS01] Robert S. Womersley and Ian H. Sloan. How good can polynomial interpolation on the sphere be? Advances in Computational Mathematics, 14(3):195–226, April 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1023/A:1016630227163.

Warren:2007:BCC

[WSHD07] Joe Warren, Scott Schaefer, Anil N. Hirani, and Mathieu Desbrun. Barycentric coordinates for convex sets. Advances in Computational Mathematics, 27(3): 319–338, October 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-9008-6.

Wang:2019:FPE

[WSL^{+}19] Xiao Wang, Wenpeng Shang, Xiaofan Li, Jinqiao Duan, and Yanghong Huang. Fokker– Planck equation driven by asymmetric Lévy motion. Advances in Computational Mathematics, 45(2):787–811, April 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        018-9642-4.

Wu:2013:SSD

[WSM13] Zongmin Wu, Xingping Sun, and Limin Ma. Sampling scattered data with Bernstein polynomials: stochastic and deterministic error estimates. Advances in Computational Mathematics, 38(1): 187–205, January 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9233-0.

Wu:1995:CSP

[Wu95] Zongmin Wu. Compactly supported positive definite radial functions. Advances in Computational Mathematics, 4(1–2):283–292, December 1995. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF03177517.

Wang:2012:SPI

[WW12] Yang Wang and Qiang Wu.

Sparse PCA by iterative elimination algorithm. Advances in Computational Mathematics, 36(1):137–151, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9186-3.

Wang:2019:LTM

[WW19] Bin Wang and Xinyuan Wu. Long-time momentum and actions behaviour of energy-preserving methods for semi-linear wave equations via spatial spectral semi-discretisations. Advances in Computational Mathematics, 45(5-6):2921– 2952, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09719-4.

Wang:2021:DFW

[WWH21] Gang Wang, Feng Wang, and Yinnian He. A divergencefree weak virtual element method for the Navier– Stokes equation on polygonal meshes. Advances in Computational Mathematics, 47 (6):??, December 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09909-z.

Wang:2018:ACS

[WWX18] Bo Wang, Li-Lian Wang, and Ziqing Xie. Accurate calculation of spherical and vector spherical harmonic expansions via spectral element grids. Advances in Computational Mathematics, 44(3): 951–985, June 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9569-1.

Wang:2019:PPE

[WXWL19] Liang Wang, Chunguang Xiong, Huibin Wu, and Fusheng Luo. A priori and a posteriori error analysis for discontinuous Galerkin finite element approximations of biharmonic eigenvalue problems. Advances in Computational Mathematics, 45(5-6):2623–2646, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09689-7.

Wang:2016:WGF

[WY16] Junping Wang and Xiu Ye. A weak Galerkin finite element method for the Stokes equations. Advances in Computational Mathematics, 42(1): 155–174, February 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9415-2.

Wang:2017:HPV

[WY17] Lina Wang and Lijun Yi. An h–p version of the continuous Petrov–Galerkin method for Volterra delay-integro-differential equations. Advances in Computational Mathematics, 43(6):1437– 1467, December 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-017-9531-2.

Wei:2020:EHV

[WY20] Yichen Wei and Lijun Yi. An hp-version of the C^{0}- continuous Petrov–Galerkin time stepping method for nonlinear second-order initial value problems. Advances in Computational Mathematics, 46(4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09800-3.

Wallner:2011:CSA

[WYW11] Johannes Wallner, Esfandiar Nava Yazdani, and Andreas Weinmann. Convergence and smoothness analysis of subdivision rules in Riemannian and symmetric spaces. Advances in Computational Mathematics, 34(2): 201–218, February 2011. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-010-9150-7.

Wei:2010:CAC

[WZ10] T. Wei and D. Y. Zhou. Convergence analysis for the Cauchy problem of Laplace’s equation by a regularized method of fundamental solutions. Advances in Computational Mathematics, 33(4):491–510, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-009-9134-7.

Wang:2019:ANS

[WZ19] Hong Wang and Xiangcheng Zheng. Analysis and numerical solution of a nonlinear variable-order fractional differential equation. Advances in Computational Mathematics, 45(5-6):2647– 2675, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09690-0.

Wang:2020:AMG

[WZ20a] Haiyong Wang and Lun Zhang. Analysis of multivariate Gegenbauer approximation in the hypercube. Advances in Computational Mathematics, 46 (4):??, August 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09792-0.

Wu:2020:PCO

[WZ20b] Zongmin Wu and Xuan Zhou.

Polynomial convergence order of stochastic Bernstein approximation. Advances in Computational Mathematics, 46(1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09742-w.

Wang:2018:UOH

[WZJ18] Tingchun Wang, Xiaofei Zhao, and Jiaping Jiang. Unconditional and optimal h^{2} - error estimates of two linear and conservative finite difference schemes for the Klein– Gordon-Schrödinger equation in high dimensions. Advances in Computational Mathematics, 44(2):477–503, April 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

                                        https://link.springer.com/
                                        article/10.1007/s10444-
                                        017-9557-5.

Wang:2017:PGT

[WZL17] Tongke Wang, Zhiyue Zhang, and Zhifang Liu. The practical Gauss type rules for Hadamard finite-part integrals using Puiseux expansions. Advances in Computational Mathematics, 43(2): 319–350, April 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9487-7.

Wu:2021:AOP

[WZM21] Cong Wu, Xiaochen Zeng, and Na Mi. Adaptive and optimal pointwise deconvolution density estimations by wavelets. Advances in Computational Mathematics, 47 (1):??, February 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09844-z.

Wu:2016:FHA

[WZtX16] Gang Wu, Lu Zhang, and Ting ting Xu. A framework of the harmonic Arnoldi method for evaluating φ-functions with applications to exponential integrators. Advances in Computational Mathematics, 42(3):505–541, June 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-015-9433-
0.

Wang:2021:PEE

[WZzS21] Xuping Wang, Qifeng Zhang, and Zhi zhong Sun. The pointwise error estimates of two energy-preserving fourthorder compact schemes for viscous Burgers’ equation. Advances in Computational Mathematics, 47(2):??, April 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
021-09848-9.

Xiao:2005:MSA

[XB05] S. P. Xiao and T. Belytschko. Material stability analysis of particle methods. Advances in Computational Mathematics, 23(1– 2):171–190, July 2005. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-1817-5.

Xiao:2013:MCC

[XH13] Junmin Xiao and Qiya Hu. Multilevel correction for collocation solutions of Volterra integral equations with proportional delays. Advances in Computational Mathematics, 39(3–4):611– 644, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9294-3.

Xiang:2015:EBF

[XHC15] Shuhuang Xiang, Guo He, and Yeol Je Cho. On error bounds of Filon– Clenshaw–Curtis quadrature for highly oscillatory integrals. Advances in Computational Mathematics, 41(3): 573–597, June 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-014-9377-9.

Xiang:2013:CQV

[Xia13] Dao-Hong Xiang. Conditional quantiles with varying Gaussians. Advances in Computational Mathematics, 38 (4):723–735, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9257-5.

Xie:2008:ENE

[Xie08] Hehu Xie. Extrapolation of the Nédélec element for the Maxwell equations by the mixed finite element method. Advances in Computational Mathematics, 29(2): 135–145, August 2008. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9047-2.

Xiong:2015:IES

[XKY15] Xiaogang Xiong, Ryo Kikuuwe, and Motoji Yamamoto. Implicit Euler simulation of one-dimensional Burridge– Knopoff model of earthquakes with set-valued friction laws. Advances in Computational Mathematics, 41(6):1475–1493, December 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9398-4.

Xu:2010:SND

[XL10] Huilin Xu and Jijun Liu. Stable numerical differentiation for the second order derivatives. Advances in Computational Mathematics, 33 (4):431–447, November 2010. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-009-9132-
                                        9.

Xiao:2010:LNI

[XP10] Quan-Wu Xiao and Zhi-Wei Pan. Learning from nonidentical sampling for classification. Advances in Computational Mathematics, 33 (1):97–112, July 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-009-9123-x.

Xu:1998:MCF

[Xu98] Yuan Xu. Minimal cubature formulae for a family of radial weight functions. Advances in Computational Mathematics, 8(4):367–380, June 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1018964818105.

Xu:2000:PIS

[Xu00] Yuan Xu. Polynomial interpolation in several variables, cubature formulae, and ideals. Advances in Computational Mathematics, 12(4): 363–376, March 2000. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.

1023/A:1018989707569. Mul-

tivariate polynomial interpolation.

Xu:2004:AMH

[Xu04a] Yuan Xu. Approximation by means of h-harmonic polynomials on the unit sphere. Advances in Computational Mathematics, 21 (1–2):37–58, July 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B:ACOM.0000016433.
27278.2a.

Xu:2004:PIU

[Xu04b] Yuan Xu. Polynomial interpolation on the unit sphere and on the unit ball. Advances in Computational Mathematics, 20(1–3):247– 260, January 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1025851005416. Ap-

proximation and applications.

Xu:2014:LTE

[Xu14] Da Xu. The long time error analysis in the second order difference type method of an evolutionary integral equation with completely monotonic kernel. Advances in Computational Mathematics, 40(4): 881–922, August 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9331-2.

Xiao:2013:SLA

[XWL13] Yunhai Xiao, Soon-Yi Wu, and Dong-Hui Li. Splitting and linearizing augmented Lagrangian algorithm for subspace recovery from corrupted observations. Advances in Computational Mathematics, 38(4): 837–858, May 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9261-9.

Xu:2012:PSI

[XY12] Yuesheng Xu and Hongqi Yang. Preface [Special issue: Numerical methods for ill-posed problems]. Advances in Computational Mathematics, 36(2):153–156, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9269-1;
                                         http://link.springer.com/
                                        content/pdf/10.1007/s10444-
                                        011-9269-1.pdf.

Xie:2015:ASF

[XY15] Hehu Xie and Xiaobo Yin. Acceleration of stabilized finite element discretizations for the Stokes eigenvalue problem. Advances in Computational Mathematics, 41 (4):799–812, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-014-9386-
8.

Xie:2020:CCF

[XY20] Shusen Xie and Su-Cheol Yi.

A conservative compact finite difference scheme for the coupled Schrödinger–KdV equations. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09758-2.

Xu:2021:MCA

[XYZ21] Fei Xu, Meiling Yue, and Bin Zheng. Multilevel correction adaptive finite element method for solving nonsymmetric eigenvalue problems. Advances in Computational Mathematics, 47 (2):??, April 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09849-8.

Xu:2001:LPF

[XZ01] Jinchao Xu and Aihui Zhou.

Local and parallel finite element algorithms based on two-grid discretizations for nonlinear problems. Advances in Computational Mathematics, 14(4):293–327, May 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1023/A:1012284322811.

Xu:2003:AWM

[XZ03] Yuesheng Xu and Qingsong Zou. Adaptive wavelet methods for elliptic operator equations with nonlinear terms. Advances in Computational Mathematics, 19(1– 3):99–146, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1022816511874. Chal-

lenges in computational mathematics (Pohang, 2001).

Yan:2001:PEE

[Yan01]  Ningning Yan.  A pos-

teriori error estimators of gradient recovery type for elliptic obstacle problems. Advances in Computational Mathematics, 15(1–4):333– 361, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1014284306804. A

posteriori error estimation and adaptive computational methods.

Yan:2003:PEE

[Yan03] Ningning Yan. A posteriori error estimators of gradient recovery type for FEM of a model optimal control problem. Advances in Computational Mathematics, 19(1– 3):323–336, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1022800401298. Chal-

lenges in computational mathematics (Pohang, 2001).

Yano:2019:DGR

[Yan19] Masayuki Yano. Discontinuous Galerkin reduced basis empirical quadrature procedure for model reduction of parametrized nonlinear conservation laws. Advances in Computational Mathematics, 45(5-6):2287–2320, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09710-z.

Yao:2016:CAM

[Yao16] Xudong Yao. Convergence analysis of a minimax method for finding multiple solutions of hemivariational inequality in Hilbert space. Advances in Computational Mathematics, 42(6):1331– 1362, December 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9465-0.

Yao:2019:LSM

[Yao19] Xudong Yao. A Ljusternik– Schnirelman minimax algorithm for finding equality constrained saddle points and its application for solving eigen problems: part I. Algorithm and global convergence. Advances in Computational Mathematics, 45(1): 269–310, February 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-018-9616-6.

Yan:2021:KAF

[YB21] Wen Yan and Robert Blackwell. Kernel aggregated fast multipole method. Advances in Computational Mathematics, 47(5):??, October 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09896-1.

Yang:2024:SAF

[YCQZ24] Fang Yang, Jiecheng Chen, Tao Qian, and Jiman Zhao. A sparse approximation for fractional Fourier transform. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10127-6.

Yang:1997:EAM

[YDF97] Geng Yang, Michel C.

Delfour, and Michel Fortin. Error analysis of mixed finite elements for cylindrical shells. Advances in Computational Mathematics, 7(3): 261–277, May 1997. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018998903567.

Ye:2000:SFE

[Ye00] Xiu Ye. Stabilized finite element approximations for the Reissner–Mindlin plate. Advances in Computational Mathematics, 13(4):375–386, December 2000. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1016693613626.

Ye:2004:DSE

[Ye04] Xiu Ye. Discontinuous stable elements for the incompressible flow. Advances in Computational Mathematics, 20 (4):333–345, May 2004. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1027363218427.

Yemm:2022:DAE

[Yem22] Liam Yemm. Design and analysis of the Extended Hybrid High-Order method for the Poisson problem. Advances in Computational Mathematics, 48(4):??, August 2022. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
022-09958-y.

Yang:2023:HQA

[YF23] Yuning Yang and Yunlong Feng. Half-quadratic alternating direction method of multipliers for robust orthogonal tensor approximation. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10014-6.

Guo:2010:CLL

[yGjW10] Ben yu Guo and Tian jun Wang. Composite Laguerre–Legendre spectral method for exterior problems. Advances in Computational Mathematics, 32(4): 393–429, May 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9112-5.

Guo:1998:AMA

[yGmW98] Ben yu Guo and Yuan ming Wang. An almost monotone approximation for a nonlinear two-point boundary value problem. Advances in Computational Mathematics, 8(1– 2):65–96, January 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018983927675.

Guo:2009:LGC

[yGqW09] Ben yu Guo and Zhong qing Wang. Legendre– Gauss collocation methods for ordinary differential equations. Advances in Computational Mathematics, 30(3): 249–280, April 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9067-6.

Guo:2008:IAT

[yGS08] Ben yu Guo and Jie Shen. Irrational approximations and their applications to partial differential equations in exterior domains. Advances in Computational Mathematics, 28(3):237–267, April 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/
10.1007/s10444-006-9020-
5.

Guo:2003:SPA

[yGSlX03] Ben yu Guo, Jie Shen, and Cheng long Xu. Spectral and pseudospectral approximations using Hermite functions: application to the Dirac equation. Advances in Computational Mathematics, 19(1–3):35– 55, July 2003. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1022892132249. Chal-

lenges in computational mathematics (Pohang, 2001).

Guo:2007:EAS

[yGW07] Ben yu Guo and Li-Lian Wang. Error analysis of spectral method on a triangle. Advances in Computational Mathematics, 26(4): 473–496, May 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-7471-8.

Ying:2007:CAO

[Yin07] Yiming Ying. Convergence analysis of online algorithms. Advances in Computational Mathematics, 27(3):273–291, October 2007. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-005-9002-z.

Yashtini:2019:EAM

[YKZ19] Maryam Yashtini, Sung Ha Kang, and Wei Zhu. Efficient alternating minimization methods for variational edge-weighted colorization models. Advances in Computational Mathematics, 45(3): 1735–1767, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09702-z.

Yang:2024:USA

[YL24a] Huaijun Yang and Meng Li. Unconditional superconvergence analysis of a structure-preserving finite element method for the Poisson–Nernst–Planck equations. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10145-4.

Yu:2024:BRS

[YL24b] Yajie Yu and Hanyu Li.

A block-randomized stochastic method with importance sampling for CP tensor decomposition. Advances in Computational Mathematics, 50(2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10119-6.

Yang:2012:EAB

[YLBL12] Yidu Yang, Qun Lin, Hai Bi, and Qin Li. Eigenvalue approximations from below using Morley elements. Advances in Computational Mathematics, 36(3): 443–450, April 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9185-4.

Yoon:2001:CAA

[Yoo01] Jungho Yoon. Computational aspects of approximation to scattered data by using “shifted” thin-plate splines. Advances in Computational Mathematics, 14 (4):329–359, May 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1012205804632.

Yang:2023:USE

[YS23] Huaijun Yang and Dongyang Shi. Unconditionally superconvergent error estimates of a linearized Galerkin finite element method for the nonlinear thermistor problem. Advances in Computational Mathematics, 49 (3):??, June 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10038-y.

Yang:2020:SGR

[YTL^{+}20] Ying Yang, Ming Tang, Chun Liu, Benzhuo Lu, and Liuqiang Zhong. Superconvergent gradient recovery for nonlinear Poisson–Nernst– Planck equations with applications to the ion channel problem. Advances in Computational Mathematics, 46 (6):??, December 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09819-6.

Ying:2012:LCG

[YWC12] Yiming Ying, Qiang Wu, and Colin Campbell. Learning the coordinate gradients. Advances in Computational Mathematics, 37(3): 355–378, October 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-011-9211-6.

Yang:2024:UFE

[YX24a] Fanyi Yang and Xiaoping Xie. An unfitted finite element method with direct extension stabilization for time-harmonic Maxwell problems on smooth domains. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10148-1.

Yang:2024:LBF

[YX24b] Shunfeng Yang and Shuhuang Xiang. Local behaviors of Fourier expansions for functions of limited regularities. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10136-5.

Yin:2023:SCA

[YXM23] Daopeng Yin, Yingying Xie, and Liquan Mei. The stability and convergence analysis of finite difference methods for the fractional neutron diffusion equation. Advances in Computational Mathematics, 49(5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10070-y.

Yue:2022:MNM

[YXX22] Meiling Yue, Fei Xu, and Manting Xie. A multilevel Newton’s method for the Steklov eigenvalue problem. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09934-6.

Ye:2008:LAG

[YZ08] Gui-Bo Ye and Ding-Xuan Zhou. Learning and approximation by Gaussians on Riemannian manifolds. Advances in Computational Mathematics, 29(3):291–310, October 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-007-9049-0.

Ye:2013:FSI

[YZ13] Weicai Ye and Yongdong Zhang. A fast solver for integral equations with convolution-type kernel. Advances in Computational Mathematics, 39(1):45–67, July 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-011-9268-
2.

Ye:2021:SFP

[YZ21] Xiu Ye and Shangyou Zhang.

A stabilizer-free pressurerobust finite element method for the Stokes equations. Advances in Computational Mathematics, 47(2):??, April 2021. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
021-09856-9.

Yu:2023:DLT

[YZ23] Zhan Yu and Ding-Xuan Zhou. Deep learning theory of distribution regression with CNNs. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10054-y.

Ye:2024:OTS

[YZ24] Xiu Ye and Shangyou Zhang. Order two superconvergence of the CDG finite elements for non-self adjoint and indefinite elliptic equations. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10100-9.

Yang:2020:NCC

[YZB20] Yidu Yang, Yu Zhang, and Hai Bi. Non-conforming Crouzeix–Raviart element approximation for Stekloff eigenvalues in inverse scattering. Advances in Computational Mathematics, 46 (6):??, December 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09818-7.

Ye:2024:NAO

[YZLL24] Xiao Ye, Xiangcheng Zheng, Jun Liu, and Yue Liu. Numerical analysis for optimal quadratic spline collocation method in two space dimensions with application to nonlinear time-fractional diffusion equation. Advances in Computational Mathematics, 50(2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10116-9.

Zheludev:2010:CIS

[ZA10] Valery A. Zheludev and Amir Z. Averbuch. Computation of interpolatory splines via triadic subdivision. Advances in Computational Mathematics, 32(1): 63–72, January 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9087-2.

Zala:1999:WAP

[ZB99] Cedric A. Zala and Ian Barrodale. Warping aerial photographs to orthomaps using thin plate splines. Advances in Computational Mathematics, 11(2–3):211– 227, November 1999. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.

                                        1023/A:1018928026708. Ra-

dial basis functions and their applications.

Zhang:2012:NIG

[ZB12] Qinghui Zhang and Uday Banerjee. Numerical integration in Galerkin meshless methods, applied to elliptic Neumann problem with nonconstant coefficients. Advances in Computational Mathematics, 37(4):453–492, November 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-011-9216-1.

Zampini:2022:HDM

[ZBT^{+}22] Stefano Zampini, Wajih Boukaram, George Turkiyyah, Omar Knio, and David Keyes. H2Opus: a distributedmemory multi-GPU software package for non-local operators. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-09942-6.

Zhao:2014:RPE

[ZC14] Jikun Zhao and Shaochun Chen. Robust a posteriori error estimates for conforming discretizations of a singularly perturbed reaction–diffusion problem on anisotropic meshes. Advances in Computational Mathematics, 40(4):797–818, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9327-y.

Zhang:2021:QFV

[ZCB21] Yuanyuan Zhang, Chuanjun Chen, and Chunjia Bi. A quadratic finite volume method for nonlinear elliptic problems. Advances in Computational Mathematics, 47(3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09853-y.

zuCastell:2007:PIU

[zCFX07] Wolfgang zu Castell, Noemí Laín Fernández, and Yuan Xu. Polynomial interpolation on the unit sphere. II. Advances in Computational Mathematics, 26(1–3):155– 171, January 2007. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-005-7510-5.

Zhang:2020:EBL

[ZCL20] Fan Zhang, Jian Cheng, and Tiegang Liu. Exponential boundary-layer approximation space for solving the compressible laminar Navier– Stokes equations. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09776-0.

Chen:2009:DAM

[zCpST09] Hua zhu Chen, Jin ping Song, and Xue-Cheng Tai. A dual algorithm for minimization of the LLT model. Advances in Computational Mathematics, 31(1–3):115– 130, October 2009. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9097-0.

Zhao:2020:IAM

[ZCY20] Fei Zhao, Xia Cui, and Guangwei Yuan. Iterative acceleration methods with second-order time accuracy for nonlinear diffusion equations. Advances in Computational Mathematics, 46 (1):??, February 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09756-4.

Zhao:2014:JPC

[ZD14] Lijing Zhao and Weihua Deng. Jacobian-predictor– corrector approach for fractional differential equations.

Advances in Computational Mathematics, 40(1):137–165, February 2014. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-013-9302-7.

Zhao:2016:HOF

[ZD16] Lijing Zhao and Weihua Deng. High order finite difference methods on non-uniform meshes for space fractional operators. Advances in Computational Mathematics, 42 (2):425–468, April 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-015-9430-3.

Zhang:2017:NAF

[ZD17] Zhijiang Zhang and Weihua Deng. Numerical approaches to the functional distribution of anomalous diffusion with both traps and flights. Advances in Computational Mathematics, 43(4): 699–732, August 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9503-y.

Zeng:2018:DTS

[ZD18] Chao Zeng and Jiansong Deng. On the dimension of trivariate spline spaces with the highest order smoothness on 3D tmeshes. Advances in Computational Mathematics, 44(2): 423–451, April 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-017-9551-y.

Zeng:2019:LHV

[Zen19] Xianyi Zeng. Linear hybridvariable methods for advection equations. Advances in Computational Mathematics, 45(2):929–980, April 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-018-9647-
                                        z.

Zhu:2016:ARM

[ZG16] Hongqiang Zhu and Zhen Gao. An h-adaptive RKDG method with troubled-cell indicator for one-dimensional detonation wave simulations. Advances in Computational Mathematics, 42(5):1081– 1102, October 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9454-3.

Zhang:2012:NAS

[ZGGW12] Deyue Zhang, Yukun Guo, Chengchun Gong, and Guan Wang. Numerical analysis for the scattering by obstacles in a homogeneous chiral environment. Advances in Computational Mathematics, 36 (1):3–20, January 2012. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-010-9169-9.

Zhang:2022:FDS

[ZGV22] Yabin Zhang, Adrianna Gillman, and Shravan Veerapaneni. A fast direct solver for integral equations on locally refined boundary discretizations and its application to multiphase flow simulations. Advances in Computational Mathematics, 48 (5):??, October 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09974-y.

Zhang:2001:PEE

[Zha01] Zhimin Zhang. A posteriori error estimates on irregular grids based on gradient recovery. Advances in Computational Mathematics, 15(1–4):363–374, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014221409940. A

posteriori error estimation and adaptive computational methods.

Zhang:2005:PWF

[Zha05]  Zhihua Zhang.  Periodic

wavelet frames. Advances in Computational Mathematics, 22(2):165–180, February 2005. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

                                        //link.springer.com/article/
                                        10.1007/s10444-004-3139-
                                        z.

Zhang:2014:SFV

[Zha14] Tie Zhang. Superconvergence of finite volume element method for elliptic problems. Advances in Computational Mathematics, 40 (2):399–413, April 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9313-4.

Zheludev:2006:ISS

[Zhe06] Valery A. Zheludev. Interpolatory subdivision schemes with infinite masks originated from splines. Advances in Computational Mathematics, 25(4):475–506, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-004-4149-6.

Zheng:2023:MMC

[Zhe23]  Chao Zheng.  Multilevel

Monte Carlo simulation for the Heston stochastic volatility model. Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10076-6.

Zhu:2015:SLC

[Zhu15] Wei Zhu. Simulation of liquid crystal elastomers using Chebyshev spectral method with a new preconditioner. Advances in Computational Mathematics, 41(4):853–879, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9389-5.

Zhu:2019:FOI

[Zhu19] Wei Zhu. A first-order image denoising model for staircase reduction. Advances in Computational Mathematics, 45(5-6):3217–3239, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-019-09734-5.

Zhi:2016:HWM

[ZhXpZ16] Shi Zhi, Yan hua Xu, and Jun ping Zhao. Haar wavelets method for solving Poisson equations with jump conditions in irregular domain. Advances in Computational Mathematics, 42(4): 995–1012, August 2016. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-015-9450-z.

Zhang:2019:SCR

[ZHY19] Yidan Zhang, Yunqing Huang, and Nianyu Yi. Superconvergence of the Crouzeix– Raviart element for elliptic equation. Advances in Computational Mathematics, 45(5-6):2833–2844, December 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09714-9.

Zhang:2021:PEE

[ZHzSR21] Qifeng Zhang, Jan S. Hesthaven, Zhi zhong Sun, and Yunzhu Ren. Pointwise error estimate in difference setting for the two-dimensional nonlinear fractional complex Ginzburg–Landau equation. Advances in Computational Mathematics, 47 (3):??, June 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09862-x.

Zhou:2006:APK

[ZJ06] Ding-Xuan Zhou and Kurt Jetter. Approximation with polynomial kernels and SVM classifiers. Advances in Computational Mathematics, 25(1–3):323– 344, July 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7206-2.

Zhang:2023:AHO

[ZJ23] Gengen Zhang and Chaolong Jiang. Arbitrary high-order structure-preserving methods for the quantum Zakharov system. Advances in Computational Mathematics, 49 (6):??, December 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10074-8.

Zhang:2021:CSM

[ZJZ21] Hui Zhang, Xiaoyun Jiang, and Fanhai Zeng. An H^{1} convergence of the spectral method for the timefractional non-linear diffusion equations. Advances in Computational Mathematics, 47 (5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09892-5.

Zheng:2016:IFA

[ZL16] Xiaoming Zheng and John Lowengrub. An interfacefitted adaptive mesh method for elliptic problems and its application in free interface problems with surface tension. Advances in Computational Mathematics, 42(5):1225–1257, October 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9460-5.

Zhang:2017:SSS

[ZL17] Jin Zhang and Xiaowei Liu. Supercloseness of the SDFEM on Shishkin triangular meshes for problems with exponential layers. Advances in Computational Mathematics, 43(4): 759–775, August 2017. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-016-9505-9.

Zheng:2020:SBE

[ZL20] Bing Zheng and Peng Lv. Structured backward error analysis for generalized saddle point problems. Advances in Computational Mathematics, 46(2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09787-x.

Zheng:2021:EAS

[ZL21] Nan Zheng and Xiaoli Li. Error analysis of the SAV Fourier-spectral method for the Cahn–Hilliard–Hele– Shaw system. Advances in Computational Mathematics, 47(5):??, October 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-021-09897-0.

Zhan:2024:ITE

[ZL24] Weijun Zhan and Yuyuan Li. The improvement of the truncated Euler–Maruyama method for non-Lipschitz stochastic differential equations. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-024-10131-w.

Zhang:2023:RBF

[ZLPX23] Yongle Zhang, Guoyin Li, Ting Kei Pong, and Shiqi Xu. Retraction-based firstorder feasible methods for difference-of-convex programs with smooth inequality and simple geometric constraints. Advances in Computational Mathematics, 49(1):??, February 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL

https://link.springer.com/
article/10.1007/s10444-
022-10002-2.

Zappon:2023:ECS

[ZMQ23] Elena Zappon, Andrea Manzoni, and Alfio Quarteroni. Efficient and certified solution of parametrized one-way coupled problems through DEIM-based data projection across non-conforming interfaces. Advances in Computational Mathematics, 49 (2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-10008-w.

Zhang:2018:LLR

[ZN18] Erchuan Zhang and Lyle Noakes. Left Lie reduction for curves in homogeneous spaces. Advances in Computational Mathematics, 44(5):1673–1686, October 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9601-0.

Zozulya:2015:RDI

[Zoz15a] V. V. Zozulya. Regularization of divergent integrals: a comparison of the classical and generalizedfunctions approaches. Advances in Computational Mathematics, 41(4):727–780, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9399-3.

Zozulya:2015:RAR

[Zoz15b]  V. V. Zozulya.  RE-

TRACTED ARTICLE: Regularization of divergent integrals: An approach using generalized functions. Advances in Computational Mathematics, 41(4):953, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-013-9321-
4;   http://link.springer.
com/content/pdf/10.1007/
s10444-013-9321-4.pdf.

Zampieri:2006:ESO

[ZP06] Elena Zampieri and Luca F.

Pavarino. An explicit second order spectral element method for acoustic waves. Advances in Computational Mathematics, 25(4):381–401, November 2006. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-004-7626-z.

Zhang:2019:OAE

[ZP19] Ran Zhang and Gerlind Plonka. Optimal approximation with exponential sums by a maximum likelihood modification of Prony’s method. Advances in Computational Mathematics, 45(3): 1657–1687, June 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10444-019-09692-y.

Zampieri:2024:CSP

[ZP24] Elena Zampieri and Luca F. Pavarino. Conditioning and spectral properties of isogeometric collocation matrices for acoustic wave problems. Advances in Computational Mathematics, 50 (2):??, April 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10113-y.

Zhang:2003:EMU

[ZPR03] Lei Zhang, Haesun Park, and J. Ben Rosen. Exponential modeling with unknown model order using structured nonlinear total least norm. Advances in Computational Mathematics, 19(1– 3):307–322, July 2003. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A:1022881907141. Chal-

lenges in computational mathematics (Pohang, 2001).

Zhang:2015:LTS

[ZPY15] Tong Zhang, Damazio Pedro, and JinYun Yuan. A large time stepping viscositysplitting finite element method for the viscoelastic flow problem. Advances in Computational Mathematics, 41(1):

149–190, February 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9353-4.

Zhu:2013:ARM

[ZQ13] Hongqiang Zhu and Jianxian Qiu. An h-adaptive RKDG method with troubled-cell indicator for two-dimensional hyperbolic conservation laws. Advances in Computational Mathematics, 39(3–4):445– 463, October 2013. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-012-9287-7.

Zheng:2022:SCS

[ZQS22] Bo Zheng, Jin Qin, and Yueqiang Shang. Stability and convergence of some parallel iterative subgrid stabilized algorithms for the steady Navier–Stokes equations. Advances in Computational Mathematics, 48 (3):??, June 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09950-6.

Zampa:2023:UFF

[ZRR23] Enrico Zampa, Ana Alonso Rodríguez, and Francesca Rapetti. Using the FES framework to derive new physical degrees of freedom for finite element spaces of differential forms. Advances in Computational Mathematics, 49(2):??, April 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-022-10001-3.

Zhu:2021:CTB

[ZS21] Xiaojing Zhu and Hiroyuki Sato. Cayley-transform-based gradient and conjugate gradient algorithms on Grassmann manifolds. Advances in Computational Mathematics, 47(4):??, August 2021. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-021-09880-9.

Zheng:2024:TSD

[ZS24] Bo Zheng and Yueqiang Shang. A three-step defectcorrection stabilized algorithm for incompressible flows with non-homogeneous Dirichlet boundary conditions. Advances in Computational Mathematics, 50 (1):??, February 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10101-8.

Zhang:2001:SFE

[ZSB01a] L. Zhang, T. Strouboulis, and I. Babuška. η%- superconvergence of finite element solutions and error estimators. Advances in Computational Mathematics, 15(1–4):393–404, November 2001. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014267804986. A

posteriori error estimation and adaptive computational methods.

Zhang:2001:PEF

[ZSB01b] L. Zhang, T. Strouboulis, and I. Babuška. A posteriori estimators for the FEM: analysis of the robustness of the estimators for the Poisson equation. Advances in Computational Mathematics, 15(1–4):375– 392, November 2001. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1014232321825. A

posteriori error estimation and adaptive computational methods.

Zheng:2015:PUP

[ZSHZ15] Haibiao Zheng, Lina Song, Yanren Hou, and Yuhong Zhang. The partition of unity parallel finite element algorithm. Advances in Computational Mathematics, 41(4): 937–951, August 2015. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

                                        springer.com/article/10.
                                        1007/s10444-014-9392-x.

Zeng:2019:FVD

[ZST19] Xianyi Zeng, Mashriq Ahmed Saleh, and Jianjun Paul Tian. On finite volume discretization of infiltration dynamics in tumor growth models. Advances in Computational Mathematics, 45(5-6):3057– 3094, December 2019. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-019-09727-4.

Zak:2014:NSC

[ZT14] Mohammad Khorsand Zak and Faezeh Toutounian. Nested splitting CG-like iterative method for solving the continuous Sylvester equation and preconditioning. Advances in Computational Mathematics, 40(4): 865–880, August 2014. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10444-013-9330-3.

Zhang:2015:SFV

[ZT15] Tie Zhang and Lixin Tang. A stabilized finite volume method for Stokes equations using the lowest order P_{1}–P_{0} element pair. Advances in Computational Mathematics, 41(4):781–798, August 2015. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-014-9385-9.

Zheng:2022:ACE

[ZTHW22] Chaowen Zheng, Zhuochao Tang, Jingfang Huang, and Yichao Wu. An O(N) algorithm for computing expectation of N-dimensional truncated multi-variate normal distribution II: computing moments and sparse grid acceleration. Advances in Computational Mathematics, 48 (6):??, December 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09988-6.

Zhang:2008:SCE

[ZV08] Chengjian Zhang and Stefan Vandewalle. Stability criteria for exact and discrete solutions of neutral multidelayintegro-differential equations. Advances in Computational Mathematics, 28(4):383–399, May 2008. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http:

//link.springer.com/article/
10.1007/s10444-007-9037-
4.

Zeng:1995:NSI

[ZVW95] S. Zeng, C. Vuik, and P. Wesseling. Numerical solution of the incompressible Navier– Stokes equations by Krylov subspace and multigrid methods. Advances in Computational Mathematics, 4(1–2): 27–49, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF02123472.

Zheng:2012:NRM

[ZW12] G. H. Zheng and T. Wei. A new regularization method for a Cauchy problem of the time fractional diffusion equation. Advances in Computational Mathematics, 36 (2):377–398, February 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s10444-011-9206-
                                        3.

Zeng:2019:DIR

[ZW19] Chao Zeng and Chunlin Wu. On the discontinuity of images recovered by noncovex nonsmooth regularized isotropic models with box constraints. Advances in Computational Mathematics, 45(2):589–610, April 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s10444-018-9629-
                                        1.

Zhou:2020:UAC

[ZW20] Yanhui Zhou and Jiming Wu. A unified analysis of a class of quadratic finite volume element schemes on triangular meshes. Advances in Computational Mathematics, 46 (5):??, October 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09809-8.

Zhang:2022:UEB

[ZW22] Teng Zhang and Tingchun Wang. Uniform error bound of a conservative fourthorder compact finite difference scheme for the Zakharov system in the subsonic regime. Advances in Computational Mathematics, 48(4):??, August 2022. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-09944-4.

Zhang:2023:UCA

[ZW23a] Yuqing Zhang and Xiang Wang. Unified construction and L^{2} analysis for the finite volume element method over tensorial meshes. Advances in Computational Mathematics, 49(1):??, February 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-022-10004-0.

Zhou:2023:AFE

[ZW23b] Zhaojie Zhou and Qiming Wang. Adaptive finite element approximation of optimal control problems with the integral fractional Laplacian. Advances in Computational Mathematics, 49 (4):??, August 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10064-w.

See correction [ZW23c].

Zhou:2023:CAF

[ZW23c] Zhaojie Zhou and Qiming Wang. Correction to: Adaptive finite element approximation of optimal control problems with the integral fractional Laplacian. Advances in Computational Mathematics, 49 (5):??, October 2023. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-023-10071-x.

See [ZW23b].

Zeng:2016:DSS

[ZWDD16] Chao Zeng, Meng Wu, Fang Deng, and Jiansong Deng. Dimensions of spline spaces over non-rectangular T-meshes. Advances in Computational Mathematics, 42(6):1259–1286, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-016-9461-4.

Zwick:1994:PMA

[Zwi94] D. Zwick. A planar minimax algorithm for analysis of coordinate measurements. Advances in Computational Mathematics, 2(4):375–391, September 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02521605.

Zhou:2019:TNM

[ZWZ19] Xuanxuan Zhou, Tingchun Wang, and Luming Zhang. Two numerical methods for the Zakharov–Rubenchik equations. Advances in Computational Mathematics, 45 (3):1163–1184, June 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://

link.springer.com/article/
10.1007/s10444-018-9651-
3.

Zou:2012:GBE

[ZXC12] Bin Zou, Zongben Xu, and Xiangyu Chang. Generalization bounds of ERM algorithm with V -geometrically ergodic Markov chains. Advances in Computational Mathematics, 36(1):99–114, January 2012. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.

springer.com/article/10.
1007/s10444-011-9182-7.

Zhang:2020:NIU

[ZY20] Jun Zhang and Xiaofeng Yang. Non-iterative, unconditionally energy stable and large time-stepping method for the Cahn–Hilliard phasefield model with Flory– Huggins–de Gennes free energy. Advances in Computational Mathematics, 46 (3):??, June 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-020-09793-z.

Zhang:2018:SOU

[ZYB18] Guo-Dong Zhang, Jinjin Yang, and Chunjia Bi. Second order unconditionally convergent and energy stable linearized scheme for MHD equations. Advances in Computational Mathematics, 44 (2):505–540, April 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-017-9552-x.

Zhang:2019:HWF

[ZYC19] Yuanyuan Zhang, Min Yang, and Chuanjun Chen. The hybrid Wilson finite volume method for elliptic problems on quadrilateral meshes. Advances in Computational Mathematics, 45(1):429–452, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10444-018-9623-7.

Zhang:2016:OCS

[ZYY16] Hui Zhang, Ming Yan, and Wotao Yin. One condition for solution uniqueness and robustness of both ℓ_{1}-synthesis and ℓ_{1}-analysis minimizations. Advances in Computational Mathematics, 42(6):1381–1399, December 2016. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-016-9467-y.

Zhou:2009:HOP

[ZZ09] Xiang-Jun Zhou and Ding-Xuan Zhou. High order Parzen windows and randomized sampling. Advances in Computational Mathematics, 31(4):349–368, November 2009. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10444-008-9073-8.

Zhang:2020:NVE

[ZZC20] Bei Zhang, Jikun Zhao, and Shaochun Chen. The nonconforming virtual element method for fourth-order singular perturbation problem. Advances in Computational Mathematics, 46 (2):??, April 2020. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-020-09743-9.

Zhang:2024:MPH

[ZZL^{+}24] Hong Zhang, Gengen Zhang, Ziyuan Liu, Xu Qian, and Songhe Song. On the maximum principle and highorder, delay-free integrators for the viscous Cahn–Hilliard equation. Advances in Computational Mathematics, 50 (3):??, June 2024. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-024-10143-6.

Zhan:2023:ITG

[ZZP23] Jiajun Zhan, Liuqiang Zhong, and Jie Peng. Iterative two-grid methods for discontinuous Galerkin finite element approximations of semilinear elliptic problem. Advances in Computational Mathematics, 49(6):??, December 2023. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10444-023-10082-8.